A Complete Bibliography of Publications in Biometrics: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

21 March 2023 Version 1.04

Title word cross-reference

2 [1171]. 2 × 2 [1527]. $99.95 [842]. ^{+} [581]. ^{2} [1001]. C [1198, 1396]. D [1198]. G [1369, 1649]. γ [1585]. J × K [906]. k [499]. L^{2} [750]. l_{1} [536]. Λ [280]. m [644]. M_{t,α} [800]. N [1347, 1501, 836, 1104, 644, 1535]. O − E [565]. p [559, 194, 1535, 1030, 1330]. t [226, 860, 534, 1271]. τ [148]. u [354].

-divergence [1585]. -eQTLs [1612]. -estimation [1369, 1649]. -FWER [194]. -index [1396]. -mixture [1347, 1501, 836, 1104]. -optimal [1198]. -Out-of- [644]. -small- [1535]. -Statistic [354, 860]. -Statistics [534]. -value [194]. -Values [559, 1030, 1330]. -Year [226, 1271]. -ZIG [499].

035 [775]. 0420.2011.01689.x [621]. 05 [1063]. 091037PR4 [417].

10.1111/j.0006 [622]. 10.1111/j.0006-341X.1999.00603.x [622]. 10.1111/j.1541 [621]. 10.1111/j.1541-0420.2011.01689.x [621].

1424-1432 [1547]. 1999 [622, 810].

2005 [515]. 2011 [383]. 2012 [474, 621, 655]. 2013 [690]. 2014 [691]. 2015 [1672]. 2017 [1256, 1547, 1675]. 2019 [1673, 1674, 1667]. 2nd [1403, 1139, 1299, 1409, 1034, 1074].

341X.1999.00603.x [622]. 3rd [659, 656].

55 [622, 810]. 5th [657].

61 [515]. 65 [127]. 67 [383]. 68 [621, 655]. 69 [690].

70 [691]. 71 [1672]. 73 [1256, 1547, 1675]. 75 [1673, 1674].

978-1-4398-1917-3 [842].

A. [41, 554, 44, 470, 46, 431]. A.-M. [46]. Aalen [1435]. Abbreviated [395]. Aberrations [218, 94]. Abhaya [656]. abilities [1063]. ability [1022]. Absence [574, 545, 111, 460]. Absolute [464, 1629]. Absorbing [122]. Abundance [621, 240, 26, 454, 680, 1111, 1606]. abundances [1031]. Abuse [205]. Accelerated [564, 712, 416, 421, 1292, 756, 165, 668, 56, 601, 291, 964, 1432, 1270, 1573, 1189, 1433]. accelerometers [1018]. Acceptance [651]. access [933]. access-points [933]. accommodate [1474, 983]. Accommodating [521]. account [1543]. Accounting [1087, 1607, 361, 475, 479, 888, 1572, 922, 1431, 1626, 1514]. accounts [808]. accuracies [807]. Accuracy [545, 334, 133, 460, 546, 442, 1176, 822, 808, 732, 873, 863]. Acknowledgements [1123, 1306, 1508, 1671]. Acoustic [260]. acquired [964, 708]. Acquisition [660]. across [1528, 1626]. Action [548]. activation [1420, 785]. active [1174]. activity [1500]. acute [709, 829]. Acyclic [613, 973, 1052, 1515]. adaptation [1099]. Adaptive [644, 386, 1513, 1202, 357, 992, 423, 1375, 677, 525, 375, 1192, 448, 337, 9, 1444, 1241, 1336, 864, 1150, 866, 1446, 841, 754, 1029, 1249, 1638, 1157, 1614, 1155, 1265, 1242, 925, 1315, 1530, 1008, 950, 993]. Adaptively [640]. Addition [531]. Additive [290, 392, 890, 447, 1522, 1073, 824, 1530, 819]. Address [593]. Addressing [1062]. Adelaide [722]. Adjust [219]. Adjusted [565, 303, 1011, 1000, 1674, 1382, 1117, 1290, 1315, 1608, 1281]. Adjusting [933, 145, 576, 894]. Adjustment [444, 194, 407, 78, 568, 478, 475, 479, 18, 424, 1465, 909, 1445, 1194, 1441, 910]. Administrative [296]. admission [1378]. Admixture [281, 216]. Adrian [1033]. Advanced [743]. advancement [1631]. Advances [660]. Advancing [1301]. Adverse [398, 1624]. aerial [1543]. Affymetrix [173]. African [766]. AFT [1014]. after [714, 1572, 1020]. Against [583, 1070]. Age [463, 73, 1298, 427]. Age-Dependent [73]. Age-Related [463]. agent [831]. Agents [106]. Ages [134]. Aggarwal [1301]. aggregate [919]. Aggregated [508, 823]. Agreement [669, 1319, 1663]. Agresti [659, 1145]. ahead [801]. Ahn [1349]. AIDS [1591, 1034, 150]. Air [495, 158, 532, 1662, 737, 1201, 709, 1609]. Aisner [1505]. al [513, 514, 512, 840]. al. [914, 1415, 915]. Alan [1145, 659]. Albert [622]. Alessandra [950]. Alessandro [950]. Alessio [1185]. Alex [1352]. Alexandre [1595]. Alfredo [1595]. Algorithm [531, 576, 540, 727, 1208, 755]. Algorithms [1502, 642, 403, 991, 989, 990]. alignment [1663, 1337]. Alignments [95]. Alive [1001]. all-comers [814, 1182]. all-or-nothing [1047]. Allen [1504]. Allison [380]. Allocation [1507, 367, 733, 1604, 950]. Allometric [201]. Allowing [1557, 1176, 1556, 1559, 1558]. Along [159]. Also [689]. Alston [556]. Alternating [1181]. alternation [1641]. Alternative [209, 35, 1193, 235, 1179, 1429, 826, 1546, 1545]. Alternatives [404, 1016]. alzheimer [885, 1113]. Amanda [1122]. AMMI [762]. Among [581, 515, 4, 679, 1490, 1486, 1488, 1487, 1489, 1396, 1201, 596]. amphibian [938]. Analyses [476, 662, 148, 363, 157, 1190, 1495, 1029, 1254, 1570, 1462, 1289, 715, 623]. Analysis [1279, 1032, 355, 211, 564, 780, 379, 380, 1547, 383, 1144, 404, 610, 634, 160, 329, 171, 681, 335, 207, 338, 544, 1141, 429, 79, 545, 281, 50, 99, 26, 373, 1416, 331, 1303, 487, 585, 1505, 265, 310, 678, 146, 264, 10, 144, 1409, 189, 602, 951, 1473, 197, 718, 1380, 635, 674, 181, 530, 667, 25, 91, 233, 230, 1305, 115, 154, 51, 371, 582, 1187, 45, 541, 617, 494, 120, 122, 1140, 136, 543, 533, 534, 1595, 377, 318, 228, 324, 52, 315, 492, 647, 86, 60, 525, 372, 46, 376, 1300, 156, 1518, 1076, 555, 842, 273, 351, 537]. Analysis [438, 1665, 188, 268, 607, 686, 97, 540, 183, 94, 57, 247, 178, 1408, 285, 1256, 1673, 1053, 1336, 701, 359, 1332, 1632, 1620, 812, 944, 1544, 874, 1601, 1009, 945, 1440, 702, 1469, 1637, 1523, 751, 416, 421, 854, 1138, 1450, 1294, 1251, 1328, 1092, 1589, 1023, 1051, 1414, 1377, 1659, 1497, 791, 1295, 1478, 1529, 1639, 1057, 1193, 1498, 1413, 1417, 1670, 746, 873, 985, 1600, 1237, 1619, 1344, 869, 697, 984, 1439, 912, 1278, 1581, 794, 1056, 880, 905, 1005, 1155, 1340, 1484, 1232, 1607, 1331, 1415, 1067, 1280, 837, 1618, 908, 695, 923, 835, 1275, 1655, 1617, 965, 1266, 1091, 1463, 1641]. analysis [1613, 729, 1470, 1230, 1118, 1262, 1297, 1365, 787, 980, 1022, 1605, 975, 1622, 819, 1371, 1116, 1077, 1184, 1302, 953, 1074, 1215, 330, 130, 659, 556, 470, 657, 259, 744, 327]. Analytic [488, 813, 1172, 806]. Analytics [1301]. Analyze [588, 589]. Analyzing [498, 490, 54, 168, 19, 1273, 1020, 924, 1392]. Anand [1503]. ancestry [1326]. anchoring [1568]. anchovies [833]. Ando [387]. Andrew [720, 1557, 1667, 1558]. Animal [681, 118, 260, 102, 934, 1363, 998, 1607, 1543]. Animals [621, 454]. ANOVA [41, 289, 566]. Anticipation [319]. Antidiabetic [462]. Antiviral [202]. Antognini [950]. Antonio [655]. Application [608, 237, 1657, 495, 610, 463, 394, 161, 865, 137, 679, 429, 1339, 281, 198, 194, 585, 428, 162, 1393, 578, 440, 598, 631, 1396, 969, 1201, 635, 201, 447, 486, 490, 1341, 683, 611, 63, 1207, 708, 1093, 1019, 632, 687, 236, 149, 76, 540, 19, 424, 481, 775, 796, 964, 1324, 1059, 1131, 862, 1674, 1174, 852, 1490, 1486, 898, 1368, 817, 1488, 1487, 1519, 1378, 1542, 1489, 802, 1063, 1240, 1113, 1020, 885, 851, 1106, 1249, 715, 1484, 1334, 1067, 1280, 870, 902, 1530, 1261, 1230, 1118, 825, 1608, 1605, 1533, 1611, 877, 1632]. Applications [41, 95, 64, 919, 73, 425, 489, 809, 688, 534, 670, 1595, 1034, 658, 572, 1074, 448, 83, 65, 437, 1033, 804, 1027, 890, 735, 1198, 752, 821, 832, 1422, 1157, 1107, 929, 960, 899, 1141, 1121, 551, 555]. Applied [138, 1354, 842, 1077, 1216, 912, 433, 333]. Applying [45]. Approach [227, 139, 209, 28, 439, 317, 634, 350, 359, 207, 338, 449, 206, 306, 287, 356, 428, 164, 162, 260, 203, 328, 496, 578, 40, 602, 68, 305, 300, 181, 145, 257, 639, 685, 88, 561, 280, 318, 450, 46, 322, 93, 510, 451, 311, 247, 505, 285, 504, 1040, 820, 1540, 1435, 1043, 1000, 1531, 862, 1173, 1332, 1575, 1398, 696, 776, 1576, 1097, 702, 973, 1041, 808, 752, 1013, 1464, 795, 1178, 1165, 763, 826, 1600, 872, 833, 1204, 947, 1331, 1083, 726, 1617, 1335, 736, 921, 1470, 1343, 1436, 1428, 1477, 705, 1022, 975, 1215, 1670]. Approaches [154, 168, 759, 1616, 1198, 1631, 728]. Approximate [199, 1114, 1621, 1449, 832, 984]. approximated [1048]. Approximation [653, 654, 1501, 741, 1254, 1330, 1599]. Approximations [73, 1262]. Aquitaine [1207]. Arbitrary [82, 855, 1103]. Ardo [1405]. Area [145, 568, 238, 23]. Arm [676, 575, 1282, 1443, 925, 1315]. armed [1614]. Arnason [1028]. Arnason-Schwarz [1028]. Array [491, 80, 218, 276]. arrival [938]. ART [646]. artery [1539]. arthritis [901]. article [47]. Artificial [425]. ASA-CRC [1593]. Asbestos [77]. aspects [836, 993]. assay [1069]. Assays [275, 241, 1538, 1625]. Assemblages [305]. Assess [164, 1006, 1022]. Assessing [20, 359, 360, 480, 1663, 123, 1164, 205, 774, 485, 736, 664, 942, 1418, 1512]. Assessment [810, 1211, 92, 231, 548, 162, 68, 668, 80, 2, 321, 1481, 1674, 1250, 1386, 963, 1058, 1608]. Assessments [279]. assigned [923]. Assignment [261, 864, 1151, 1373, 1254, 728]. assistance [959]. assisted [1576, 1462]. Associated [317, 987, 1482, 1062]. associating [1231]. Association [608, 463, 98, 428, 190, 144, 35, 119, 37, 576, 670, 134, 532, 607, 451, 333, 1408, 135, 27, 1239, 1450, 745, 883, 1328, 904, 1003, 1400, 1163, 974, 1327, 1253, 1638, 694, 924, 1275, 700, 875, 1297, 1399, 792, 1366]. Associations [579, 882, 1394]. assumption [1180, 1471, 1212]. Assumptions [323, 518, 1665]. Assurance [1578]. Asymptotic [404, 217, 169, 1501]. Asymptotically [588, 589]. attempt [944]. attraction [857]. attraction-repulsion [857]. Attributable [67, 1238]. AUC [239]. Audit [362, 361]. auditory [1252]. Augmentation [254]. Augmented [395, 961, 366, 1369, 112, 1271, 1320, 1533]. Augmenting [261]. August [593]. author [166]. Authors [956, 996, 1038, 1080, 1126, 1148, 1219, 1259, 128, 210, 325, 126]. autism [1533]. autism-related [1533]. Auto [348]. Auto-probit [348]. Autocorrelation [508, 736]. Autologistic [1658]. Automated [1537, 1304].

Autoregressive [493, 737]. Auxiliary [53, 238, 352, 66, 18, 1540, 865, 699, 795]. Auxiliary-Dependent [66]. Availability [643, 934]. Average [70, 632, 1547, 1191, 1294, 723, 1429, 888]. averaged [1154, 945, 1441]. Averaging [317, 391, 1492]. Aylin [1352].

B [1673, 329, 950, 258, 951, 553, 375, 842]. B. [742]. backward [1318]. Baddeley [1033]. Bagging [1587]. Baio [719]. Baker [515, 514]. Balakrishnan [172]. Balance [467]. Balanced [455, 651]. band [909, 910]. Bands [563, 565, 1616, 1017]. Banerjee [842]. Banks [1301]. Baptiste [1350]. BAS [651]. Base [92, 816]. Base-Calling [92]. Based [47, 6, 28, 70, 642, 92, 116, 633, 524, 562, 224, 105, 669, 69, 30, 239, 328, 578, 40, 631, 145, 458, 472, 560, 81, 8, 280, 606, 488, 397, 342, 466, 437, 451, 135, 285, 1040, 1673, 853, 714, 1324, 1006, 1283, 712, 1620, 1577, 1674, 1027, 917, 1319, 1584, 946, 986, 970, 1662, 1132, 1224, 1191, 487, 1150, 883, 353, 1102, 1054, 1573, 1542, 1377, 1663, 348, 1437, 791, 1222, 1226, 1029, 788, 1512, 1024, 1438, 1223, 1344, 1447, 697, 1225, 1162, 502, 1475, 924, 1389, 1367, 799, 773, 1290, 1645, 1062, 1333, 929, 1421, 15, 333, 705, 968, 1608]. based [1599, 1238, 1583, 975, 1371]. Baseline [21, 401, 917, 1289]. basis [1530, 1262]. basket [1242]. Basso [41]. batch [1285]. Batschelet [408]. Bayes [369, 939, 706, 318, 1025, 65, 364, 1170, 977]. BayesCAT [1337]. Bayesian [622, 386, 845, 846, 1362, 1360, 556, 847, 1361, 848, 849, 255, 673, 987, 613, 439, 610, 1574, 964, 463, 317, 307, 319, 106, 1444, 161, 1336, 1195, 58, 701, 903, 186, 1531, 1485, 862, 1173, 370, 804, 1026, 1172, 1575, 1174, 519, 522, 699, 449, 545, 99, 199, 675, 1001, 945, 972, 898, 1, 527, 702, 344, 831, 461, 1150, 146, 1359, 33, 162, 803, 30, 203, 506, 1542, 462, 1101, 578, 278, 1586, 802, 766, 1153, 1528, 1386, 36, 1656, 737, 832, 1201, 123, 577, 68, 13, 674, 1529, 530, 1480, 732, 1253, 978, 985, 1024, 61, 576, 1105, 244]. Bayesian [639, 1296, 879, 801, 688, 163, 523, 1344, 1288, 120, 711, 205, 1196, 881, 611, 560, 1075, 880, 833, 103, 341, 638, 1387, 423, 612, 988, 1467, 228, 844, 786, 1083, 616, 1443, 1337, 60, 406, 102, 93, 1621, 1350, 1335, 410, 834, 1242, 1609, 34, 475, 479, 77, 700, 1530, 1343, 1068, 1436, 1262, 191, 17, 405, 346, 738, 1623, 575, 59, 424, 505, 1008, 1022, 938, 1622, 62, 507, 504, 586, 1281, 899, 591, 329, 911, 1668, 254, 553, 719, 387, 1630, 375]. beamforming [705]. Becker [1353, 952]. Bécue [1669]. Behavior [681, 581]. behavioral [971]. Benchmark [548, 734, 945]. Benefit [67, 645, 1651, 1571]. Bent [201]. Berger [1507, 512]. Bergsma [44]. Bernstein [712, 782]. Bernstein-Schoenberg [782]. Berry [375]. Bertalanffy [682]. Bertaut [1669]. Beta [360]. Beta-Mixture [360]. Betensky [653]. Between [550, 7, 81, 125, 813, 1176, 1208, 549, 1102, 1663, 440, 1193, 1394, 532, 805, 1333]. Between-Group [81]. between-study [1193]. between-subject [1333]. bi [896]. bi-level [896]. biallelic [1240]. Bias [643, 359, 71, 1416, 345, 208, 143, 584, 667, 145, 343, 45, 183, 1414, 926, 1413, 1417, 1462, 976, 1415, 473]. Bias-Corrected [345, 143]. Biased [22, 165, 1112, 293, 54, 677, 759, 1523, 1380, 1318, 1517]. Biclustering [142, 1129]. Big [1143, 919, 1406]. Bilder [953]. Bill [1409]. Binary [422, 28, 620, 335, 260, 246, 323, 37, 391, 122, 52, 315, 672, 365, 193, 401, 247, 1547, 1097, 1294, 1345, 802, 1064, 784, 1044, 1512, 1462, 1379, 988, 1443, 753, 1658, 748]. Binomial [493, 796, 1529]. Binomials [618]. Bioassay [366, 702]. Biodiversity [26]. Bioinformatics [244, 553]. biologic [1068]. biological [1483, 1340]. Biologics [780]. Biology [1409, 326, 432]. Biomarker [182, 1302, 300, 238, 443, 66, 862, 831, 793, 713, 1061]. Biomarker-Calibrated [443]. Biomarkers [301, 538, 1518, 11, 442, 1176, 1287, 1085, 867, 767, 772, 768, 889, 1288, 1084, 863, 1652, 1335, 771, 1008, 770, 942]. biomass [833]. Biomedical [1034, 1119, 437, 1336]. Biometric [1221, 721, 593]. Biometrics [127, 514, 1221, 513, 324]. Biosimilars [780]. biospecimens [1069]. Biostatistical [1186, 258]. Biostatistics [1142, 656, 1592]. Birth [398, 532, 427, 870, 929]. birth-death-shift [929]. bisulfite [1532]. Bivariate [172, 207, 365, 537, 964, 696, 1051, 969, 1163, 1317, 694]. Blackwell [1094]. bleeding [804]. blocked [926]. blocks [884]. blue [1106]. Blunders [1593]. BM [278]. BM-Map [278]. Board [957, 997, 1039, 1081, 1127, 1149, 1220, 1260, 1309, 1358, 1412, 1461, 1511, 1554, 1598, 1635]. Boca [1032, 1403, 1628, 1144, 950, 1186, 1141, 1402, 1406, 1185, 1142, 1299, 1629, 1214, 1505, 1143, 1668, 1405, 1355, 993, 1631, 1401, 1352, 1353, 1670, 1305, 1301, 1187, 1349, 1140, 952, 953, 1506, 1594, 1075, 1667, 1407, 1595, 1669, 1122, 1300, 1630, 1350, 1213, 842, 1183, 1593, 811, 1033]. Bolus [320]. Book [1032, 211, 780, 1077, 1035, 1216, 1455, 1403, 1628, 379, 381, 380, 382, 433, 434, 435, 690, 591, 1453, 1502, 1549, 1144, 41, 1348, 1504, 329, 172, 1184, 171, 950, 1186, 1507, 1632, 911, 552, 252, 1591, 1451, 1141, 1139, 430, 384, 386, 471, 84, 85, 551, 1402, 330, 258, 1406, 1185, 554, 1142, 256, 1452, 373, 1299, 1120, 212, 1629, 331, 129, 42, 1303, 1214, 1505, 374, 215, 1454, 130, 1590, 1143, 1188, 328, 1409, 1668, 1405, 1355, 1215, 1302, 557, 1457, 993, 1145, 1631, 659, 720, 951, 992, 1401, 90, 1352, 1353, 1121, 556, 778, 718, 878]. Book [214, 43, 660, 1670, 91, 44, 1305, 661, 326, 169, 253, 1592, 1548, 1301, 257, 1187, 1404, 254, 45, 1349, 170, 385, 742, 1458, 553, 1550, 470, 472, 657, 719, 88, 779, 1140, 656, 744, 952, 953, 877, 912, 1506, 1594, 1075, 473, 1667, 387, 1407, 1595, 332, 1503, 132, 213, 377, 1034, 174, 173, 592, 1669, 1304, 1551, 87, 1119, 658, 86, 625, 167, 1456, 255, 1122, 46, 376, 1074, 131, 1300, 1630, 375, 1350, 1298, 1076, 1213, 555, 624, 842, 83, 1183, 168, 378, 327, 259, 432, 431, 743, 1593, 811, 89, 333, 1408, 590, 623, 1033]. Boomerang [1072]. Bootstrap [644, 124, 1435, 762, 1644]. bootstrapping [1324, 986]. Borrowing [459, 1425]. Both [180]. Bottle [1352]. bound [765]. Boundaries [539]. Boundary [51, 1098]. Bounds [503, 1561, 663, 748, 1562, 1560, 1563]. Bové [654]. boxplot [1499]. Bradley [1298]. brain [1420, 1026, 706, 1230]. Branching [73]. Breast [515, 4, 817]. Breeding [680]. Bretz [384, 131]. Bridging [620, 1580]. British [368]. Broad [397, 1663]. Broemeling [1668, 1075]. Broman [85]. Brooks [329, 661]. Buckland [1121]. Buckley [293]. Buehlmann [1143]. Bun [1300]. burden [1063]. Burgess [1213].

Burt [1303]. Business [1184]. Buyse [1140].

C [211, 1673, 552, 1402, 1224, 1121, 556, 1301, 1223, 657, 1225, 332, 213, 131, 1427, 707]. C-learning [1427]. C. [172, 326, 169, 385, 744, 87]. C.-D. [172]. CACE [1622]. Cadmium [368]. Cai [653]. calculation [1496, 807, 1477]. Calculations [618, 1349]. Calibrated [443]. Calibrating [662, 805]. Calibration [227, 735, 527, 164, 33, 181, 56, 247, 424, 960]. Calling [92]. Cambridge [1120, 1188, 1409, 1404, 1503, 1119, 1298, 1351]. Campagnoli [170]. can [1517]. Cancer [515, 4, 242, 106, 778, 539, 580, 66, 299, 862, 1173, 1205, 1174, 1208, 817, 1180, 1190, 1063, 851, 746, 1342, 1175, 1526, 902, 1477]. Cancers [1034]. Candidate [119]. cannot [1346]. Canonical [144, 189, 1279, 1600, 1340, 1484]. Canopy [23]. Capitalizing [1031]. Capture [82, 453, 683, 114, 680, 682, 322, 261, 781, 1043, 1363, 970, 971, 1661, 1338, 1110, 1385, 1543, 895, 800, 728, 1333]. capture-mark-resight [1363]. capture-recapture [781, 1043, 971, 1661, 1338, 1110, 1385, 1543, 895, 800, 728]. carcinoma [1334, 1335]. Cardiovascular [776, 462, 617, 1611]. CARDS [1011]. Carini [1407]. Carlin [375, 842]. Carlo [213, 1372, 540, 661]. Carpenter [877]. carried [1289]. Carryover [21]. Casals [1595]. Case [237, 810, 439, 319, 498, 1398, 110, 29, 35, 197, 457, 674, 53, 538, 576, 617, 120, 136, 195, 464, 816, 52, 488, 671, 572, 156, 147, 2, 178, 135, 27, 1547, 892, 1000, 823, 1173, 961, 827, 1294, 1189, 1494, 1473, 788, 1610, 1527, 1439, 1004, 708, 1025, 943, 924, 1275, 1091, 1255, 1238, 942, 977, 556]. Case-base [816]. Case-Cohort [538, 178, 1189, 1494, 1473, 1439, 924, 1255]. Case-Control [237, 810, 439, 498, 110, 35, 53, 576, 120, 136, 195, 52, 156, 2, 135, 27, 892, 823, 1189, 1494, 788, 1610, 1527, 708, 1025, 943, 1275, 1238, 942]. Case-Crossover [674]. Case-Only [147, 1398, 961]. Case-Parent [488]. Casella [213]. cases [1610]. Categorical [1032, 508, 1188, 44, 953, 612, 109, 659]. Categories [64, 123]. category [1256, 1138]. Catherine [1301]. Causal [355, 209, 482, 850, 812, 519, 1572, 1589, 1409, 1603, 458, 1604, 1213, 298, 664, 466, 401, 481, 597, 504, 813, 1040, 1491, 1588, 723, 1393, 1007, 1153, 1311, 1429, 724, 1369, 1467, 1265, 748, 773, 888, 1436, 1119, 1590]. causal-inference [813]. Causation [473]. Cause [180, 922, 1409]. Cause-Specific [180]. causes [869, 1323]. Cavity [198]. CCor [1102]. CD4 [1207]. CDF [409]. Cell [611, 424, 299, 1020, 1107, 1636]. cell-cycle [1107]. cells [1525]. Censored [622, 3, 58, 666, 1, 1405, 399, 294, 316, 119, 225, 233, 234, 293, 54, 638, 510, 295, 601, 483, 1568, 1012, 1272, 1270, 1469, 1523, 1014, 1466, 1524, 693, 1046, 1386, 1063, 1164, 1237, 1474, 1065, 1093, 965, 1318, 1200, 982, 981, 1091, 901, 1137, 1045, 1271, 1316]. Censoring [180, 599, 264, 10, 584, 425, 148, 296, 179, 266, 1566, 822, 1073, 1092, 1292, 1091, 1434]. Census [539]. Censuses [113, 730]. center [1339]. Cessation [288, 1386]. CGH [218]. Chaeryon [771]. Chain [661, 493, 1628]. Chains [614]. Chakraborti [381]. Challenge [647]. Challenge-Response [647]. Challenges [425, 1214]. Chambers [1183]. Chan [1591]. Chandan [1301].

Chang [1256, 1407]. Change [93, 987, 1235, 831, 885, 809]. Change-Point [93, 885]. changes [885]. Changing [539]. Chao [684]. Chapman [950, 993, 1631, 951, 952, 953, 842, 1593, 811, 1194]. Chapple [1557, 1558]. Characteristic [75, 342]. Characteristics [436, 834]. Characterize [645]. characterizing [1208]. Charu [1301]. Chasan [1142]. Chasan-Taber [1142]. Chaudhuri [554]. checking [892]. chemicals [716]. Chen [1673, 1488, 1487, 1489, 912, 811]. Chenguang [383]. Cheung [1215, 1300]. Chi [58, 1234]. chi-square [1234]. Chi-Squared [58]. Chichester [1590]. Child [161, 498]. Children [635, 340]. China [907]. ChIP [163, 192]. ChIP-chip [163]. ChIP-seq [192]. Choice [394, 441, 866, 832, 1130]. Choices [321]. cholera [774]. Chough [744]. Chow [780, 1591, 87]. Chris [1035]. Christian [1304]. Christiane [1628]. Christophe [1506]. Christopher [953, 1667]. chromatin [1133]. Chronic [570]. Chu [1416]. Chul [1349]. Chung [1591, 780]. Cichosz [1502]. Circular [408, 1107, 1499]. cis [1612, 1384]. cis-eQTL [1384]. Cities [317]. Clarice [810]. clarifications [740]. Class [673, 529, 185, 344, 7, 205, 520, 460, 1518, 1055, 970, 1064, 1093]. Classes [7]. classical [1540]. Classification [853, 143, 391, 284, 62, 542, 1466, 808, 1101, 752, 860, 1113, 1500, 753, 902, 1419, 1427, 1170]. Classifiers [460]. Claude [90]. Claudio [1407]. clearance [898, 707]. Cliburn [1591]. climate [987]. Clinical [564, 1403, 379, 434, 1139, 1214, 1505, 182, 357, 646, 993, 778, 1557, 123, 68, 300, 43, 1404, 639, 541, 1349, 1140, 86, 876, 375, 431, 107, 597, 299, 1657, 1059, 1556, 1559, 804, 1674, 814, 865, 961, 1624, 1282, 831, 1180, 1524, 1496, 866, 841, 1631, 1182, 754, 1047, 1060, 1090, 1592, 1341, 1288, 806, 1443, 925, 1315, 1608, 1622, 1558, 1407, 386, 88]. clinically [867, 815]. CLIP [738]. clonal [1342]. Closed [781, 322, 1110, 1385, 895]. Closed-population [781]. Cluster [1011, 104, 336, 246, 1543, 157, 55, 1674, 1010, 1150, 1525, 1495, 1619, 1604, 1087, 740, 1431, 805, 1243, 1608]. cluster-adjusted [1674, 1608]. cluster-based [1150]. cluster-period [1525]. cluster-randomized [1087]. Clustered [290, 524, 37, 44, 177, 468, 176, 1349, 248, 446, 178, 764, 1027, 1235, 1010, 1014, 1619, 1539, 1234]. Clustering [562, 105, 1210, 221, 280, 492, 141, 1381, 1531, 1131, 1252, 1233, 1581, 1166, 378]. clusters [1103, 939, 805]. Co [1504, 207, 1393, 707, 1337, 785]. co-activation [785]. co-estimation [1337]. Co-expression [207]. co-infected [707]. co-infection [1393]. Co. [1034]. coagulation [1623]. Coarsened [236, 1092]. coast [833]. Coefficient [116, 353, 103, 776, 966, 1272, 1492, 1108, 1157, 1611]. Coefficients [600, 445, 223, 874, 1240, 760, 1246]. Coexpression [449]. Cognitive [38, 340, 885]. Cohort [394, 538, 1207, 178, 859, 1189, 1494, 1473, 730, 1439, 1004, 1084, 1005, 707, 924, 709, 1255]. Cohorts [315, 1493]. Coin [677]. cold [1343]. Collazo [1628]. Collection [1304]. Collective [242]. Colorectal [515, 4]. Columbia [368]. Combination [69, 1176, 1198, 1376]. Combinations [106, 239, 793]. combinatorial [1446]. combine [835]. Combined [47, 6, 759]. Combining [827, 739, 1250, 282, 767, 768, 402, 327, 771, 135, 823, 1061, 1099, 770, 772].

Combs [1301]. comers [814, 1182]. Commenges [1186]. Commensurate [357, 711]. Comment [513, 689, 653, 654, 127, 512]. Comments [514]. Committees [1403, 43]. Common [1641, 1212, 434]. Communication [118]. Communities [138, 1626]. community [1450, 1297, 1533]. community-level [1450, 1297]. Companion [212]. Comparative [435, 1274, 1589, 1496, 1060, 1447, 856]. compare [1049]. Comparing [1131, 1085, 550, 549, 301, 959, 582, 363, 822, 807]. Comparison [383, 160, 962, 36, 1136, 683, 273, 269, 801]. Comparisons [384, 81, 1199, 473]. compartmental [1000]. Competing [224, 951, 225, 176, 685, 133, 134, 442, 251, 1256, 822, 757, 1205, 1339, 890, 891, 1138, 1051, 693, 698, 1292, 1497, 1237, 869, 694, 1093, 1050]. completer [1289]. Complex [118, 734, 1114, 1292, 906, 1117, 1087, 700]. Compliance [520, 1047]. Complier [520]. Component [310, 678, 509, 484, 757, 854, 1648, 1057, 1498, 715, 1232, 837]. Components [116, 101, 563, 349, 246, 670, 484, 528, 1251, 829]. Composite [428, 1583, 764, 1648]. Compositional [686, 1534]. Comprehensive [42]. Computation [199, 1326, 832, 254]. Computational [836, 168, 432, 253]. Computationally [468, 611, 981]. Computed [263]. Computer [495, 33, 1298]. computerization [913, 916, 914, 915]. Computing [1402]. Concealment [1507]. Concentration [320, 74]. Concentrations [368]. Concepts [623]. concerning [1674, 1608]. Concordance [116, 353, 1062]. Conditional [227, 620, 1241, 1336, 272, 207, 441, 206, 599, 356, 7, 217, 465, 683, 1176, 970, 1191, 1094, 737, 1470]. conditionally [786]. conditions [1131]. Conference [593]. Confidence [219, 1561, 874, 1616, 563, 243, 982, 1109, 714, 909, 827, 1651, 1562, 1560, 1563, 1090, 826, 893, 927, 1665, 910]. Confirmatory [398]. Confounder [297, 789, 1359, 1362, 1360, 1361, 888, 1441]. Confounders [146]. confounding [1465, 1572]. Conjunction [541]. Connecting [935]. connectivity [1420, 1026, 1230]. Connie [1488, 1487, 1489]. Consider [689]. Considerations [992, 1407, 157]. Consistent [627, 628, 629, 448, 626]. Constrained [115, 243, 735, 1015]. constraints [1575, 1480, 1534, 1098]. construct [918]. Constructing [1651, 427]. Construction [714, 1161, 1419]. Consumption [321]. content [1525]. Contents [474]. Contingency [510, 906]. Continual [312]. Continuous [160, 137, 84, 164, 598, 37, 606, 1385, 908, 320, 298, 196, 1176, 1329, 1158, 1589, 1007, 1442, 969, 784, 998, 1621, 172]. Continuous-time [1385, 1589, 1442, 1621]. contrast [1019, 1192]. Contribution [417]. Control [237, 810, 439, 498, 79, 633, 194, 110, 29, 35, 349, 53, 576, 120, 136, 195, 52, 156, 2, 135, 27, 892, 823, 1189, 1494, 788, 1610, 1527, 1425, 708, 1614, 1025, 806, 943, 1275, 1290, 834, 1238, 942, 977, 1353, 952]. control-based [1290]. Controlled [671, 572, 1060]. Controlling [64, 1264, 632]. Controls [571]. Convex [1481, 1129, 1581, 817]. Conway [1027]. coordinate [1344]. coordinate-based [1344]. Copula [1645, 1437]. Copula-based [1645, 1437]. Copulas [227, 1566, 1205]. copy [1482, 1054]. coronary [1539]. Corrected [563, 345, 143, 567, 1094]. Correcting [642, 1324]. Correction [515, 474, 621, 622, 690, 655, 810, 1256, 1547, 1672, 1673, 383, 47, 1674, 1675, 155, 691, 1289, 1618, 787]. Correlated [403, 457, 450, 1575, 1624, 1196, 794, 1636, 902]. Correlation [116, 353, 144, 1409, 349, 466, 55, 1279, 874, 777, 1437, 1382, 1600, 1340, 1484, 1234, 1247, 1116]. Correlation-Based [466]. Correlation-Selected [349]. correlations [736]. Cosmatos [87]. Cost [352, 808, 1448]. Costs [483]. Count [245, 1032, 116, 304, 269, 581, 1347, 1195, 1432, 1027, 1646, 1001, 1496, 1293, 1529, 1071, 1470, 1599, 819, 1371]. counterfactuals [1603]. countermatched [1005]. Countermeasures [32]. counting [1273]. Countries [593]. Counts [240, 499, 77, 1043, 1052, 1196, 1346]. couple [1542]. couple-based [1542]. Course [255, 1343, 731]. Covariance [610, 115, 1131, 1283, 1577, 696, 904, 1615, 1665, 705]. Covariate [684, 444, 110, 1311, 78, 177, 343, 568, 1382, 155, 677, 1315, 183, 57, 787, 567, 1426, 1012, 909, 1041, 1445, 889, 1418, 1069, 1474, 695, 805, 1602, 910, 1564, 1281]. Covariate-Adaptive [677]. Covariate-adjusted [1382, 1315, 1281]. Covariate-Dependent [177, 1069]. Covariate-Specific [343, 889]. Covariates [383, 160, 335, 366, 16, 662, 10, 203, 336, 506, 271, 635, 181, 667, 53, 663, 352, 612, 228, 525, 322, 273, 268, 505, 820, 1244, 798, 917, 699, 1364, 1322, 1110, 1295, 795, 963, 1467, 894, 726, 921]. Covariates-Application [635]. Cover [1307, 1356, 1410, 1459, 1509, 1552, 1596, 1633]. Cowpertwait [214]. Cox [1567, 1089, 16, 1236, 1322, 930, 1321, 600, 667, 53, 1370, 54, 1395, 760, 1048, 1320, 445, 1564]. crab [1106]. CRC [1032, 1403, 1628, 1144, 1186, 1507, 1591, 1141, 1402, 1406, 1185, 1142, 1299, 1629, 1214, 1505, 1143, 1668, 1405, 1355, 1401, 1352, 1353, 1670, 1305, 1301, 1187, 1349, 1140, 912, 1506, 1594, 1075, 1667, 1407, 1595, 1669, 1122, 1300, 1630, 1350, 1213, 842, 1183, 1593, 1033, 950, 993, 1631, 951, 952, 953, 811]. Creating [531]. credible [1444, 1083, 789]. creel [933]. Cressie [552, 1667]. Criteria [358, 1673, 1620, 1048]. Criterion [397, 297, 773]. Critical [532]. Croon [44]. Cross [583, 394, 395, 672, 112, 1626, 868, 940, 976, 1468]. cross-cut [976]. Cross-Dependent [672]. Cross-Sectional [583, 395, 112, 1626, 868, 940, 1468]. Cross-validation [394]. Crossover [674, 21, 122]. CT [902]. Cubic [339]. CUDA [1402]. Cumulative [689, 104, 444, 180, 584, 179, 1199, 892, 891, 1268]. Cure [484, 1567, 865, 1115, 1162, 1137]. Current [1631, 364, 268, 57, 1291, 1317, 1472, 729, 1470]. current-status [1317]. Curve [145, 343, 568, 238, 75, 682, 311, 1006, 1423, 809, 980]. Curves [84, 99, 342, 1211, 889]. CUSUM [565]. cut [714, 976]. cut-off [714]. Cycle [469, 978, 1107]. Cystic [76].

D [379, 381, 622, 41, 84, 1668, 1301, 553, 1075, 1503, 87, 327, 432, 1566, 1171, 1437]. D-vine [1566, 1437]. D. [380, 172, 91, 472, 255]. daily [852]. Dalrymple [1488, 1487, 1489]. Damla [1488, 1487, 1489]. Dan [1076]. Danh [1488, 1487, 1489]. Daniel [1216, 1186, 1408]. Daniels [383]. Daphnia [903]. Data [245, 1032, 237, 673, 125, 1035, 622, 383, 1144, 495, 404, 634, 620, 394, 92, 3, 290, 226, 526, 58, 171, 116, 456, 370, 335, 498, 604, 482, 529, 347, 574, 522, 429, 79, 121, 548, 666, 50, 547, 289, 306, 1402, 330, 1406, 1185, 633, 98, 1, 373, 524, 110, 563, 224, 1303, 487, 585, 64, 1359, 1362, 428, 579, 33, 508, 130, 264, 10, 304, 506, 1302, 578, 278, 399, 40, 277, 646, 440, 408, 1360, 659, 631, 271, 217, 123, 577, 13, 305, 262, 300, 674, 660, 37, 25, 91, 233, 177, 468, 1305, 176, 605, 115, 154, 603, 371]. Data [582, 1301, 254, 639, 541, 688, 683, 163, 523, 205, 293, 339, 648, 63, 136, 953, 292, 54, 533, 534, 221, 464, 638, 561, 1361, 334, 174, 619, 228, 324, 96, 52, 315, 606, 682, 616, 488, 1304, 672, 361, 632, 60, 365, 525, 520, 372, 436, 267, 376, 389, 80, 510, 270, 509, 236, 295, 168, 65, 193, 302, 273, 364, 351, 409, 438, 188, 268, 261, 259, 337, 686, 97, 55, 76, 401, 9, 284, 183, 19, 191, 446, 601, 17, 266, 178, 269, 222, 483, 135, 251, 285, 507, 542, 1291, 796, 1391, 1256, 1547, 1673, 853, 764, 1397, 1347, 1540]. data [1329, 1574, 1011, 1499, 1269, 1195, 1420, 823, 1485, 1173, 804, 750, 727, 725, 920, 1026, 1426, 1332, 1620, 1577, 759, 1432, 1423, 757, 1027, 1174, 761, 1285, 1239, 696, 1235, 1646, 1010, 1449, 735, 1662, 1206, 1114, 1272, 1132, 1270, 1469, 1637, 1523, 1014, 1368, 1642, 717, 1031, 1191, 1138, 1537, 1466, 1052, 838, 1294, 1251, 1328, 1524, 919, 1536, 1023, 1651, 693, 1338, 1496, 1276, 802, 1494, 1135, 1396, 1013, 1190, 1386, 1659, 1063, 821, 832, 1113, 1020, 1292, 1497, 1380, 784, 791, 885, 1295, 1163, 1478, 1529, 758, 1267, 1229, 1498, 1670, 906, 756, 746, 1178, 1317, 1106, 763, 788, 1472, 1512, 1534, 985]. data [1433, 1370, 1237, 1638, 1606, 809, 1296, 879, 1619, 1042, 1344, 711, 697, 984, 749, 1210, 1581, 794, 1056, 1157, 884, 905, 963, 1017, 1387, 1395, 1379, 1093, 1107, 1340, 1484, 1669, 943, 1636, 1175, 1280, 1234, 1658, 1532, 835, 876, 1209, 965, 1318, 709, 1645, 1061, 1071, 828, 895, 982, 800, 1002, 981, 1091, 1613, 1647, 729, 1470, 1493, 1516, 1137, 1320, 1255, 1262, 1434, 775, 705, 1365, 787, 1533, 975, 1316, 1627, 819, 1371, 1116, 899, 1032, 1403, 1502, 1348, 1184, 1120, 1143, 1188, 1405, 1355, 951, 1076, 842, 591, 552, 84, 44, 45, 779, 744, 43]. Data-driven [1359, 1362, 1360, 1361]. Data-Grouped [292]. Database [31]. Databases [671]. Datasets [1503]. Datta [1076]. David [1032, 810]. Davide [1561, 1562]. Davidian [130]. DDP [497]. Deal [643]. Dealing [208]. Death [97, 401, 38, 1589, 1156, 929, 1471, 856]. Debashish [1120]. December [515, 655, 1547, 691]. Decision [634, 1192, 918]. decision-making [1192]. Decision-Theory [634]. Decisions [64, 476, 917]. Deck [314]. Decomposition [142, 217, 518, 1642]. Deductive [913, 916, 914, 915]. deep [1587]. Defined [99]. Degeneration [463]. delayed [1249, 1477]. Deletion [531]. Deletion/Substitution/Addition [531]. Delivered [320]. delta [1290]. delta-adjusted [1290]. Dementia [38, 1093]. Demographic [113, 1606]. Demography [658]. Demonstrating [32]. Denis [1350]. Density [260, 80, 243, 1363, 1364]. depend [1603]. Dependence [227, 1566, 121, 489, 572, 1176, 1131, 1201, 1253, 1050]. dependency [1087]. Dependent [599, 264, 10, 584, 73, 635, 225, 44, 177, 339, 133, 179, 315, 606, 672, 66, 266, 222, 11, 340, 196, 1073, 1565, 1322, 889, 1106, 1069, 694, 863, 923, 965, 1238, 1316]. depends [1008]. depression [1059, 864]. Depth [508, 305]. deregulated [1531]. derivation [913, 916, 914, 915]. Design [780, 1504, 106, 370, 574, 386, 675, 1282, 1505, 208, 595, 462, 992, 1557, 660, 538, 120, 122, 423, 354, 571, 52, 406, 156, 151, 66, 196, 1556, 1559, 804, 1578, 1625, 1150, 1524, 1376, 1446, 807, 1480, 1151, 1069, 1447, 1084, 1443, 1314, 1061, 1612, 968, 1558, 1564, 1140, 1408, 211, 379, 331, 129, 472, 86]. Designs [108, 242, 275, 950, 231, 500, 309, 403, 323, 674, 21, 393, 555, 710, 1390, 1244, 1241, 961, 944, 1198, 841, 1473, 1240, 754, 1088, 697, 1375, 835, 834, 1242, 1243, 555]. Despite [588, 589]. Detect [206, 1054]. Detectability [641]. Detecting [867, 1445, 489, 283, 187, 451, 94, 1212, 308, 1132, 1586, 939, 968]. Detection [39, 28, 104, 547, 194, 1483, 967, 1636, 1666, 1171, 928, 1624, 1103, 999, 902, 1533]. determinantal [1068]. Determinants [194]. Determination [32, 1390, 1495]. Developing [593]. Development [88, 1631, 1300, 742]. Developments [1035]. Deviation [467, 1130]. device [1397, 711]. Devices [370]. Dey [553, 1141]. Diabetes [462]. Diagnosing [1134]. diagnosis [1162]. Diagnostic [161, 239, 111, 440, 8, 460, 546, 1176, 1424, 807, 873]. Diagnostics [405]. Diagonal [1536, 143]. Dialysis [617, 776, 1490, 1486, 1488, 1487, 1489, 1396]. Dibben [1035]. DIC [801]. Dichotomization [298]. Dichotomized [154]. Dichotomous [575]. Dietary [443, 1674, 1618, 1231, 1608]. Difference [197, 124, 1211, 1521, 815, 1593]. Differences [441, 266, 516, 1445]. Differential [286, 449, 99, 304, 489, 218, 437, 1053, 946, 887, 825, 1246]. Differentially [279, 191, 903, 1132, 1586, 1636, 1532]. differentiated [766]. Diffusion [285]. Diggle [252]. Digital [1301]. Dilution [275, 1625]. Dimension [82, 562, 144, 1165, 234, 341, 1535, 753, 1200]. Dimensional [610, 634, 429, 105, 143, 217, 233, 491, 1506, 1518, 438, 188, 284, 1465, 1283, 1485, 1601, 1576, 973, 1251, 1364, 1536, 1007, 1542, 1095, 1639, 1643, 828, 1530, 731, 1602, 1421, 1365]. Ding [912]. Ding-Geng [912]. Dipak [1141]. Direct [596, 1249, 869, 518, 863, 893]. Directed [613, 973, 1052, 1515, 1277]. Directional [64, 1664]. Dirichlet [185, 1105, 221, 1231, 1381]. Dirichlet-tree [1231]. Discontinuation [410, 1428]. Discovered [361]. Discoveries [64, 1264]. Discovery [39, 655, 559, 79, 535, 534, 1531, 1030, 1341]. Discrepancies [615]. Discrepancy [427]. Discrete [1348, 371, 324, 372, 509, 820, 1329, 1584, 1368, 1497, 1030, 1636, 900, 1032]. discrete-continuous [1329]. discretely [1621, 929]. Discretization [437]. Discretization-Based [437]. Discretized [508]. Discriminant [143, 189, 1639]. discrimination [1224, 1222, 1226, 1223, 1225, 1022]. discriminative [1500]. Discussion [1561, 627, 845, 1416, 1224, 1488, 846, 1487, 1414, 628, 1489, 1360, 418, 768, 1562, 1557, 914, 847, 478, 1223, 417, 420, 1225, 629, 1361, 848, 1415, 915, 769, 771, 770, 1558]. Discussions [419, 916, 412, 414, 1362, 1563, 413, 411, 477, 1226]. Disease [237, 823, 282, 496, 29, 440, 1353, 582, 465, 952, 464, 134, 378, 76, 308, 1501, 1287, 1482, 1339, 1206, 808, 821, 1113, 885, 939, 707, 1658, 901, 907, 938, 1611]. disease-associated [1482]. Diseases [570, 485, 540]. dispersed [1496]. Dispersion [392, 469]. Dispersions [573]. displaced [1011]. Display [577, 1303]. Dissertation [1142]. Dissimilarity [573, 280]. Distance [1171, 40, 81, 750, 904, 1121]. Distance-Based [81]. Distributed [619, 1424, 1245, 1675, 919, 802]. Distribution [219, 456, 1568, 522, 185, 547, 407, 203, 534, 587, 229, 664, 1565, 1027, 1134, 984, 1428]. Distribution-free [1568]. Distributions [665, 26, 408, 316, 569, 132, 100, 936, 1368, 1049, 1168, 730, 1381, 172]. Dive [508]. divergence [1585]. Diverging [609]. diverse [1279]. Diversity [550, 549, 932]. Diving [621, 454]. Dmitrienko [131]. DNA [95, 31, 1054, 489, 1201, 1002]. Do [460]. DOI [621, 622]. domain [1368]. Dominic [1304]. Donald [1119]. Dorfman [426]. dormancy [1204]. Dose [70, 679, 928, 548, 162, 403, 1557, 640, 311, 861, 1556, 1559, 734, 874, 1541, 945, 1197, 1198, 716, 703, 1558]. Dose-Finding [403, 1541]. Dose-Response [679, 162, 311, 861, 874, 1197, 1198, 716, 703]. doses [976]. Dot [241]. Double [934, 24, 1539, 179, 516, 1627, 1154, 1058]. Double-Observer [24, 934]. Double-Robust [516]. Double-wavelet [1627]. Doubly [1672, 1465, 665, 335, 399, 1190, 1267, 236, 886, 725, 1482, 1521, 1572, 723, 1311, 1370, 1369, 1395]. Doubly-Robust [665, 1672, 886]. Douglas [657]. downscaling [797]. DPOAE [1211]. DPP [1068]. Drawing [1576]. drawn [781]. Drineas [1143]. Driven [633, 1273, 1359, 1362, 838, 1360, 1361]. drop [820, 1135, 1331, 1517]. drop-out [820, 1331, 1517]. drop-outs [1135]. Dropout [521, 236, 1673, 1620, 1646, 1469, 1617]. Dropouts [673, 1267]. Drug [742, 320, 861, 1376, 872, 1539, 908]. drug-eluting [1539]. Duan [1351]. Due [401]. duration [730, 985]. during [803, 923]. Dutilleul [623]. Dyadic [673, 749]. Dylan [1547]. Dynamic [117, 644, 207, 153, 277, 202, 230, 391, 685, 1425, 1207, 334, 1654, 1391, 1672, 822, 1059, 1490, 1486, 1488, 1466, 1487, 1589, 1489, 698, 1498, 1500, 1652, 886, 1058, 1649, 1428, 1427, 170]. dynamical [825, 1186]. Dynamics [641, 182, 1351, 648, 48, 776].

E. [253, 86]. Early [513, 514, 1653]. Ebola [1000]. Ecological [256, 40, 963, 1386]. Ecology [587, 167, 1187, 329]. Economics [719]. Ed [1592, 1409, 1034, 1074, 957, 997, 1039, 1081, 1127, 1149, 1220, 1260, 1309, 1358, 1412, 1461, 1511, 1554, 1598, 1635]. Eden [1184]. Edge [51]. edited [252, 330, 129, 130, 553, 87, 86, 46, 131, 555, 168, 993]. Edition [842, 1403, 1139, 1299, 659, 657, 656]. Editor [514, 512, 689, 127, 324]. Editors [388, 558, 692, 843, 958, 1128, 1310, 1555, 5, 175]. Eds [1141, 1120, 1140, 1035, 1407]. Education [593, 1301]. EEG [937, 1251]. Effect [673, 20, 209, 422, 290, 526, 544, 519, 599, 596, 300, 283, 279, 296, 645, 520, 518, 664, 216, 475, 479, 340, 504, 1199, 861, 1491, 1191, 1197, 723, 1393, 1007, 1464, 1047, 1429, 1249, 1438, 1384, 1653, 1133, 1087, 1207, 1067, 888, 1477, 960, 1583]. effective [703]. effectiveness [1589, 775]. Effects [108, 404, 140, 394, 338, 498, 527, 1299, 287, 562, 203, 506, 232, 316, 158, 452, 21, 667, 250, 230, 663, 393, 204, 469, 458, 249, 228, 179, 96, 518, 528, 401, 451, 597, 1391, 853, 1666, 1244, 1424, 920, 1287, 814, 1398, 1134, 831, 1236, 1648, 1377, 1153, 1311, 1182, 737, 1201, 1418, 1088, 1194, 1268, 724, 1369, 1570, 749, 774, 1117, 905, 963, 1083, 1636, 709, 941, 829, 888, 1441, 805, 1343, 887, 1612, 968]. Efficacy [583, 69, 598, 32, 1070, 755, 1008]. Efficiency [810, 500, 115, 1392, 509, 2, 865, 1094, 733, 1255]. Efficient [710, 764, 666, 1624, 524, 403, 1189, 596, 660, 468, 988, 216, 1262, 913, 916, 914, 1069, 915, 981]. Efron [1298]. Ege [1033]. Eggermont [89]. Eigenvalue [1366, 1484]. Elashoff [1355]. elastic [1513, 1442]. electromagnetic [706]. electronic [867, 1537, 1519, 821]. Elegant [259]. Elements [174, 929]. eligible [1588]. Elizabeth [657]. eluting [1539]. EM-algorithm [755]. Emerging [187]. Emery [1184]. emigration [1514]. Empirical [1673, 1620, 449, 444, 1368, 369, 635, 939, 318, 19, 909, 1425, 884, 1025, 1170, 910]. Empirical-likelihood-based [1673, 1620]. empirically [1000]. Empty [409]. end [1339, 707]. end-stage [1339, 707]. Endpoint [423, 151, 149, 504, 1268, 1314, 86]. Endpoints [422, 301, 646, 122, 813, 1040, 1674, 961, 1541, 1180, 1443, 1062, 900, 1608]. Enhanced [1009, 753]. enhancement [1443]. Enrichment [410, 701, 1060, 1151]. envelope [1278]. Environment [282, 120, 450, 488, 974, 1029, 1388, 1212, 1583]. Environmental [498, 882, 716, 256]. Epidemics [150]. Epidemiologic [45]. Epidemiology [1142, 60, 969, 1075]. epigenome [1253]. epigenome-wide [1253]. epigenomic [1666]. eQTL [610, 429, 1384, 389]. eQTLs [1612]. equality [1536]. equate [1250]. Equation [286, 461, 1215, 437, 759, 825, 1246]. Equations [690, 99, 336, 1409, 521, 438, 188, 946, 704, 1286, 557]. Equilibrium [34]. Equivalence [904, 674, 587]. Erica [1672]. Error [642, 495, 347, 356, 181, 250, 617, 249, 155, 235, 60, 303, 183, 567, 1540, 1324, 1632, 1322, 1041, 969, 1372, 1425, 1614, 1618, 1323, 1434, 787, 1564]. Error-Prone [181]. Errors [547, 203, 304, 506, 578, 361, 48, 424, 1643, 786, 695, 1593]. errors-in-variables [1643]. EST [191]. Estate [658]. Estes [1488, 1487, 1489]. estimate [973, 833, 703, 1427]. Estimated [484, 1379]. Estimates [307, 335, 148, 1517, 931]. Estimating [108, 125, 690, 583, 47, 6, 642, 219, 1491, 317, 220, 1059, 456, 1430, 482, 395, 441, 240, 574, 206, 306, 717, 1191, 838, 184, 260, 336, 584, 1338, 557, 29, 602, 225, 300, 1047, 667, 568, 582, 279, 1342, 1277, 648, 749, 352, 150, 464, 96, 521, 406, 709, 112, 438, 1002, 188, 932, 266, 546, 517, 775, 1246, 759, 946, 940, 737, 1606, 1087, 1207, 704, 840, 1286, 888, 921, 1441, 755, 1622]. Estimation [1100, 39, 621, 82, 209, 559, 1176, 161, 1312, 286, 814, 392, 113, 666, 444, 199, 845, 309, 524, 180, 599, 585, 345, 263, 310, 846, 1052, 369, 595, 138, 1252, 425, 730, 1240, 202, 453, 1110, 596, 889, 145, 343, 238, 490, 847, 1070, 1643, 694, 454, 611, 1384, 400, 1475, 1374, 534, 248, 179, 488, 443, 1175, 229, 93, 236, 216, 475, 409, 479, 77, 437, 805, 23, 311, 346, 89, 291, 504, 586, 1672, 1520, 764, 714, 1582, 1540, 1501, 1241, 1000, 1522, 1089, 1154, 823, 727, 725, 850, 1432, 757, 1568, 696, 1521, 1363, 913, 916, 1270, 1368, 1572]. estimation [831, 777, 922, 1104, 1094, 1573, 1189, 766, 1311, 1003, 1400, 1182, 1030, 1267, 1229, 795, 1515, 1249, 914, 826, 985, 1105, 1268, 1112, 1369, 1615, 848, 1539, 988, 844, 849, 707, 1337, 915, 1048, 1318, 1061, 941, 800, 886, 1058, 1649, 1098, 1626, 1255, 900, 1612, 1399, 1045, 1640, 1015, 1238, 1271, 1281, 856, 1468, 1213, 720]. Estimations [1565]. Estimator [313, 655, 684, 67, 544, 535, 637, 635, 293, 235, 147, 516, 1435, 1190, 1429, 1194, 1290, 1320]. Estimators [435, 665, 441, 649, 188, 1465, 853, 972, 723]. Ethics [242]. ethnic [1067]. Evaluate [340, 863, 923]. Evaluating [1339, 595, 1313, 462, 538, 747, 358, 1539, 72, 295, 466, 442, 1249, 1062, 1571]. Evaluation [31, 732, 597, 1040, 991, 754, 989, 990, 1374, 715, 1389]. Event [125, 622, 591, 1348, 289, 98, 1, 265, 263, 10, 7, 1355, 225, 468, 176, 334, 363, 86, 572, 267, 1518, 295, 314, 1628, 1256, 1566, 820, 822, 1235, 1646, 1449, 776, 1138, 1064, 1396, 1013, 1063, 979, 1394, 1433, 1268, 1136, 1159, 1443, 923, 1392, 1434, 819]. event-history [820]. Events [313, 264, 602, 156, 223, 1269, 804, 1568, 1624, 1092, 1473, 1063, 756, 979, 963, 1093, 1654, 1118]. Evolution [199, 1063, 929]. Evolutionary [494]. Exact [614, 1569, 323, 1570, 488, 927, 1118]. examination [1432]. Examining [34]. Examples [1402, 1350, 911, 434]. exceedances [1227]. exchangeable [711]. Exercises [258]. Existing [315, 363]. exp [817]. Expansion [948, 949]. Expected [441, 1373]. Expenditure [205]. expenditures [1046]. Expensive [611]. Experimental [1504, 679, 675, 1376, 462, 1625]. Experimentation [42]. Experiments [307, 681, 449, 331, 279, 472, 647, 555, 937, 1418]. Experts [546]. Explained [292, 267, 1502]. Exploiting [546, 1201]. Exploration [579]. Exploring [368]. Exponential [1478, 141, 896, 936, 1180, 1021, 1516]. Exposed [518, 903, 1603]. Exposure [237, 810, 439, 617, 532, 77, 2, 247, 340, 1073, 1662, 882, 1268, 1117, 1388, 1441, 775, 1611]. exposure-dependent [1073]. exposure-weighted [882]. Exposures [450, 443, 298, 987]. Expressed [279, 191, 903, 1132, 1636]. Expression [64, 447, 490, 173, 318, 448, 1053, 207, 791, 1484]. Extended [392, 1285, 851, 760]. extending [1437, 1299, 1183]. extension [701]. extent [1644]. external [1618]. Extraction [321]. Extrapolation [422, 1643]. Extreme [1141]. extremes [1609].

F [41, 1139, 384, 330, 1557, 131, 1558]. Factor [398, 599, 523, 494, 566, 561, 1544, 1479, 1295, 1344, 1196, 1109, 652, 1421]. Factorial [42, 1088]. Factorization [321, 1579]. factors [882]. Fahrmeir [591]. Failure [564, 290, 416, 421, 177, 165, 668, 55, 56, 446, 601, 1269, 1565, 961, 1270, 922, 1573, 1189, 1292, 1670, 1369, 869, 694, 1091, 1255, 787, 1316]. Fairclough [379]. falls [796]. False [39, 689, 559, 79, 64, 584, 1030, 534, 1264]. False-Positive [584]. Familial [134, 1239]. Families [282, 712, 1205, 1326, 132]. Family [189, 1478, 351, 451, 135, 941, 1516, 1612]. Family-Based [451, 135]. Familywise [1614]. Fang [324, 372]. Faraway [1630, 1299]. Farcomeni [1185]. Fast [1656, 1330, 704, 1599, 1450, 1318, 1297]. Fatality [464]. Favaro [655]. FDR [633, 632]. feature [1537, 752, 731]. Features [159, 1639, 1421]. Fecundability [117]. fecundity [978]. Feedback [586]. Fellner [1261]. Fellner-Schall [1261]. Feng [1256]. Feng-Chang [1256]. Ferrari [1561, 1562]. Ferraty [330]. few [1570]. fewer [933]. FFCD [1063]. Fiber [77]. Fibrosis [76]. Field [102, 858, 745, 897]. Fields [186]. Filipino [635]. Filtered [347]. filtering [1029]. Finding [70, 95, 403, 734, 1541, 1288]. Fine [1256, 180, 1438, 467, 829]. Fine-Gray [1438]. Finite [82, 609, 169, 1181, 1166]. Fink [45]. Fire [511, 512]. Firing [36]. First [255]. Fish [682]. Fisher [721, 722]. Fit [614, 58, 616, 405, 1161, 1577, 788]. fitness [1611]. fitting [1377, 981]. Fitzmaurice [130, 1077]. Fixed [140, 232]. FL [1032, 950, 993, 1670, 952, 953, 1075, 1667, 1033, 811]. Flandre [420]. FLCRM [1321]. Flexibility [500]. Flexible [1540, 231, 1616, 1096, 971, 499, 1044, 270, 1043, 1014, 988, 1175, 981, 1050]. Florida [842]. fly [733]. fMRI [1420, 121, 678, 578, 715]. Follmann [622]. Follow [780, 395, 1426, 866, 923, 965, 1571]. Follow-on [780]. Follow-up [395, 1426, 866, 923, 965, 1571]. followed [777]. Following [486, 52]. follows [128, 166, 210, 325, 126]. Food [321]. foot [907]. Forecasting [74, 1659]. Forest [101, 138, 1042]. formula [765]. Formulations [369]. forward [1289, 704, 925, 1315]. Foster [1406]. Foulkes [333]. Foundations [1145]. founder [858]. Fox [45]. FPCA [1333]. FPCA-based [1333]. fraction [1115, 1524, 1651, 1238]. Frailty [529, 289, 224, 288, 177, 446, 1291, 1646, 890, 1236, 756, 979, 1204, 899, 470]. Frailty-Based [224]. Framework [480, 440, 389, 466, 964, 1026, 1274, 1245, 1675, 1153, 1418, 1056, 1266, 1118, 1427, 792]. Frangakis [914, 915]. Frank [1074]. Frauke [1406]. Free [28, 362, 1568, 1095, 1334, 390]. frequency [765, 1106, 932]. Frequentist [39, 386, 834]. Friedman [174]. Friendly [1032]. Frying [511, 512]. Fuentes [252]. Fully [699, 287, 292, 1387]. Function [348, 646, 602, 528, 340, 854, 1496, 879, 870, 783, 921, 691, 1262, 1428, 900]. function-on-function [879, 783, 691]. function-on-scalar [854]. Functional [501, 338, 231, 330, 562, 563, 1321, 1478, 605, 523, 497, 12, 492, 436, 1323, 74, 62, 285, 542, 701, 750, 1577, 917, 751, 1251, 1663, 1659, 1113, 784, 1057, 1498, 1108, 879, 1210, 1056, 1017, 1263, 1232, 837, 1167, 902, 1516, 785, 1262, 1605, 1627, 171]. functional-response [1263]. Functions [482, 392, 1171, 962, 1319, 966, 1272, 1044, 988, 1019, 1045]. Fundamentals [385]. Fungus [137]. Fused [1099, 1657]. fusion [495]. future [987]. FWER [194].

G [434, 515, 171, 1591, 1675, 130, 425, 1353, 1557, 661, 254, 170, 657, 952, 877, 1213, 1558]. G-Estimation [425]. G. [386, 660, 213, 432]. G.-L [254]. Gadda [1186]. Gaetan [332]. Gail [1629]. gamma [890]. Gang [1355]. Gap [148, 603, 1051, 1020, 1019]. Gaps [314]. Garica [1595]. Garica-Hiernaux [1595]. Garrett [1077]. Garrote [283]. Gatekeeping [633, 1314]. Gaussian [613, 1211, 1424, 186, 727, 428, 30, 1656, 1044, 1229, 393, 560, 102, 536, 1262, 284, 1166]. Gaussian-process [1424]. Gaussian-Wishart [1262]. GEE [1495, 157, 636]. Geelbek [1122]. Gel [1122]. Gelfand [252, 842]. Gelman [661]. Gene [64, 282, 447, 490, 283, 120, 560, 280, 318, 450, 488, 448, 1131, 862, 967, 1276, 791, 1029, 1253, 1341, 872, 881, 1277, 1484, 1388, 1212, 1583, 1533, 173]. gene-drug [872]. Gene-Environment [282, 120, 450, 488, 1029, 1388, 1212, 1583]. Gene-Gene [283, 967]. Genealogies [560]. General [1426, 529, 430, 306, 310, 1198, 569, 451, 1274, 970, 1085, 1056, 1266, 1468]. Generalization [684, 144, 616]. Generalizations [1401]. Generalized [608, 355, 690, 67, 116, 604, 392, 430, 1299, 854, 118, 336, 1145, 860, 250, 1534, 1433, 204, 249, 1157, 150, 12, 677, 521, 438, 188, 1390, 712, 1111, 776, 1202, 1473, 1267, 795, 1484, 1247, 704, 1286, 888, 1261, 875, 1602, 1611, 557]. Generalizing [1588]. generated [1117]. Generation [92, 278, 277, 1076, 1132, 1276]. generative [1467]. Genes [119, 279, 191, 903, 1102, 1636]. Genetic [608, 515, 4, 319, 350, 220, 498, 281, 194, 360, 550, 549, 428, 35, 119, 75, 670, 1213, 168, 216, 607, 27, 882, 745, 883, 939, 1254, 1278, 924, 1275, 941, 700, 875, 929, 792, 1170, 1583, 1366]. genetical [1637]. Genetics [326, 221, 1034, 327, 333, 718, 385]. Geng [912]. Genome [219, 1408, 1102, 766, 1327, 1253, 1638]. genome- [1253]. genome-differentiated [766]. Genome-Scale [219]. Genome-Wide [1408, 1327, 1638]. Genomic [631, 217, 533, 193, 94, 1329, 1173, 1024, 1404]. genomics [1637, 792, 327]. Gentle [253]. Geoadditive [245]. Geoffrey [657]. Geographic [104, 547]. George [1120]. Geostatistical [496]. Gerhard [1348, 1188]. Germany [803]. Geskus [951]. Gestational [427]. GGE [762]. ggplot2 [259]. Ghani [1406]. Ghosh [1120, 417, 553, 418]. Gianluca [719]. Gibbons [381]. Gimenez [329]. Giovagnoli [950]. Giraud [1506]. Girolami [432]. Gittins [925, 1315]. Given [7]. GLM [1367]. GLM-based [1367]. Global [1065, 1331, 72, 736]. Glymour [1590]. Goergen [1628]. Göhlmann [173]. Gold [545, 111, 568, 460, 546, 873]. Goldstein [1035, 327]. Good [972, 765]. Good-Turing [972, 765]. Goodman [1592]. Goodness [614, 58, 616, 405, 1577, 788]. Goodness-of-Fit [614, 58, 405, 1577, 788]. gorillas [766]. Gouskova [1256]. Governance [274, 594]. governing [1394]. Grace [1632]. Gradients [605, 880]. Granger [1605]. Grant [1142]. graph [1230, 1015]. graph-constrained [1015]. Graphical [613, 536, 624, 295, 1656, 1229, 1383, 785]. Graphics [259, 1305, 376, 373]. graphs [973, 1052, 1515, 1628]. Graves [171]. Gray [180, 1438]. Great [369]. Greco [1185]. Greedy [1152]. Group [608, 70, 817, 310, 457, 81, 151, 687, 546, 396, 1199, 1657, 1538, 896, 1010, 1513, 1177, 1524, 991, 1445, 1311, 860, 855, 1580, 989, 1438, 990, 1296, 908, 835, 1247, 1314, 1061, 875, 1571]. group- [835]. Group-Randomized [396]. group-sparse [1580]. Grouped [292, 1286, 652]. Groups [679, 531, 396, 1066]. Groupwise [1278]. Growth [101, 99, 184, 49, 682, 1300, 1423, 1568, 1106]. Gruber [435]. guarantee [1374]. Guerra [327]. guess [946]. Guide [85, 1406, 42, 1409, 1304, 1302, 380]. guided [1161, 747, 1389]. Guido [1119]. Gumbel [407]. gut [1231]. Guttorp [252]. Guyon [332]. GWAS [1067].

H [1077, 435, 430, 258, 1629, 878, 473, 173, 167, 743]. H. [328, 167, 168, 259]. H4 [803]. HAART [486, 1207]. Haenszel [1155]. Hagan [1302]. Hagelloch [487]. Hagenaars [44]. Häggström [1360, 1361]. Hall [842, 1593, 950, 993, 1631, 951, 952, 953, 811]. Hall/ASA [1593]. Hall/ASA-CRC [1593]. Hall/CRC [950, 993, 1631, 951, 952, 953, 811]. Halloran [211]. Hand [84, 907]. Handbook [330, 46, 1143, 1187, 252, 661, 168]. Handling [779, 943, 1493]. Hans [621]. Haplotype [96, 1212]. haplotype-environment [1212]. Haplotypes [463]. Harbor [508]. Hard [206]. Hardcover [842]. Hardin [557]. Hardy [34]. Hardy-Weinberg [34]. Harm [645]. harnessing [1531]. Harrell [1074]. Harron [1035]. Harvey [1035]. Hasegawa [1547]. Hastie [1401, 174, 1298]. Hazard [653, 3, 654, 666, 444, 598, 443, 291, 962, 1319, 1521, 922, 851, 1090]. Hazards [529, 179, 178, 1522, 1179, 1073, 712, 1235, 1616, 909, 824, 1047, 788, 1472, 1048, 981, 910]. Heagerty [1224, 1223, 1225]. Health [570, 158, 635, 205, 719, 1174, 867, 1250, 868, 737, 1592, 709, 829]. Healthcare [1352, 1301, 1301]. Heat [317]. Heiberger [1303]. Heidelberg [1216, 1348, 1303, 1121, 1074, 1076, 1408]. Held [654]. Helene [1186]. helices [1422]. hematopoietic [1020]. Henderson [235]. Henning [743]. Heo [1349]. hepatitis [707]. hepatocellular [1334, 1335]. Herson [1403, 43]. Heterogeneity [319, 95, 641, 114, 645, 623, 1053, 1283, 1193, 1194, 1247, 1626]. Heterogeneous [347, 137, 1383, 436, 322, 859, 1584, 1168, 1378, 1165, 1515, 1070, 1419]. Heteroscedastic [429, 1372, 786]. Hidden [643, 1285, 1209, 1666, 1586, 871, 1462, 905, 1532, 897, 738, 374, 1594]. Hierarchical [106, 220, 186, 678, 357, 1177, 784, 51, 244, 533, 561, 687, 716, 191, 405, 62, 1390, 1211, 1026, 1174, 898, 1101, 1233, 732, 1341, 1262, 1622, 842]. Hiernaux [1595]. High [610, 634, 449, 429, 1601, 562, 105, 143, 217, 78, 233, 163, 1506, 1518, 80, 438, 188, 284, 1465, 1283, 1485, 1576, 973, 838, 1536, 1007, 1418, 1448, 1643, 1066, 828, 1530, 716, 1602, 1421, 1365]. high-cost [1448]. High-Density [80]. High-Dimensional [610, 634, 429, 105, 143, 217, 233, 1506, 1518, 438, 188, 284, 1485, 1576, 973, 1536, 1007, 1643, 828, 1602, 1421, 1365]. High-Order [163]. High-Throughput [449, 1418, 1066, 1530, 716]. higher [733]. Highest [243]. Hilbe [471, 557, 720]. Hinkelmann [555]. Hiroshi [621]. histogram [1581]. histogram-valued [1581]. histopathology [752]. Historical [422, 370, 357, 711, 806]. Histories [314]. History [591, 439, 267, 76, 820]. HIV [47, 1591, 6, 394, 161, 395, 739, 838, 1393, 940, 316, 202, 486, 502, 1207, 150, 707, 229, 112, 48, 775, 581]. HIV-Infected [502]. HIV-Positive [486]. HIV/AIDS [1591, 150]. HIV/TB [1393]. HLA [901]. Hoboken [1077, 1502, 1139, 1302, 1145, 1301, 877, 1304]. Hoff [255]. Holland [1303]. Holly [771]. Holm [515]. Holmgren [742]. Homogeneity [573, 197, 305, 509]. homogeneous [859, 1437, 1030]. Hooker [171]. hormones [1394]. Horton [1305, 376]. Hospice [51]. hospital [964, 1490, 1486, 1488, 1487, 1378, 1489, 708]. hospitalizations [1396]. Hot [314]. Hot-Deck [314]. Hothorn [384]. Hout [1405]. HPTN [775]. Hu [1561, 1562]. Hua [1214]. Huang [771]. hub [1528]. Hubbard [689]. Hudgens [1591]. Human [326, 683, 1300, 168, 1018, 1626]. hurdle [1636]. Hybrid [673, 681, 1557, 540, 1556, 1559, 835, 703, 1602, 1558]. Hydrological [619]. hypergeometric [701]. Hypotheses [633, 72, 1016, 1648, 1546, 1545, 1655]. Hypothesis [345, 1007, 588, 617, 195, 589, 760, 46, 1167, 1230, 575, 1283].

I. [215, 473]. I/II [69]. Iain [1594]. Ian [1406]. IBS [274, 594, 593]. ICU [19]. Identifiability [111, 149, 401]. Identification [859, 277, 457, 596, 447, 724, 566, 280, 218, 903, 1274, 1543, 887]. identifier [710]. Identify [101, 78, 1287, 1060, 738]. Identifying [383, 463, 1482, 449, 937, 793, 1082, 273, 532, 785, 1398, 1083, 1532, 1533]. Ignorable [673, 343, 699, 1493]. Igor [655]. II [1559, 778, 1557, 250, 89, 1558, 1556, 500, 69, 249, 1443]. II. [114]. II/III [1559, 1557, 1558, 1556]. III [1559, 1557, 1558, 1556]. Illness [97, 856]. Illness-Death [97, 856]. illustrated [700]. illustration [867]. Image [542, 732]. Images [241, 12, 752]. imaging [1278, 1387, 902, 1605, 1627, 590]. imbalance [805]. Imbens [1119]. IML [212]. Immune [48]. immunoassays [1101]. immunodeficiency [1626]. immunotherapy [1477]. Impact [570, 486, 205, 571, 485, 27, 1095, 1193, 707]. Imperfect [641, 710, 1176]. Implementation [992]. Implications [229]. Importance [150]. important [815]. Imprinting [451, 968]. Improve [810, 2]. Improved [1391, 1666, 369, 1606, 534, 236, 9, 504, 765, 1181]. improvement [864, 1289]. Improving [865, 500, 1094, 35, 1255]. Imputation [359, 272, 154, 541, 877, 8, 402, 303, 314, 302, 1250, 1494, 1117, 943, 1290, 728, 1137, 983]. Incidence [47, 6, 395, 180, 229, 112, 891, 940, 1626]. Incident [456, 1426, 1610]. Incomplete [383, 604, 522, 646, 779, 365, 273, 19, 729]. inconsistency [1377]. Incorporating [608, 242, 1657, 439, 1312, 858, 1295, 63, 533, 55, 727, 1637, 1496, 875, 775]. Incorporation [357, 1340]. Incorrectly [588, 589]. Incremental [150, 295, 942]. Independence [24, 7, 1450, 906, 1297, 1212]. Independent [310, 678, 316, 618, 715]. index [714, 1426, 1111, 776, 1396, 1095, 1108, 1389, 925, 1315, 1602, 1421, 1281]. index-based [714]. index/Cox [1567]. Indexes [490, 427]. indicators [1323]. Indirect [554, 518]. Individual [242, 761, 550, 549, 288, 682, 580, 835]. individual-level [835]. Individualized [1312, 1430, 1650, 1587, 1374]. Individually [110]. individuals [1588, 890, 1326]. Indrayan [656]. Induced [110, 1627]. Infected [502, 707]. infecting [1440]. Infection [641, 440, 486, 617, 49, 229, 1393]. Infection-Cardiovascular [617]. Infections [687, 964, 857, 1177, 1020, 1070]. Infectious [1353, 952, 569, 464, 485, 540]. Inference [209, 439, 620, 644, 95, 519, 71, 741, 1409, 40, 926, 878, 123, 577, 13, 674, 1610, 617, 296, 611, 560, 174, 612, 1119, 625, 298, 83, 66, 351, 1156, 432, 48, 183, 192, 481, 396, 483, 196, 796, 813, 1040, 1203, 1569, 1011, 1423, 699, 827, 1158, 1572, 1590, 693, 1668, 1046, 1153, 1528, 871, 1656, 1088, 1429, 809, 1570, 1603, 869, 1604, 1056, 1467, 740, 1265, 1621, 1644, 1248, 929, 1068, 1436, 1599, 1298, 381]. Inferences [243, 1588, 1576, 1234, 1118]. inferential [754]. inferiority [960]. Inferring [1329, 576]. Infertility [687, 1542]. Inflated [499, 1027, 1646, 1368, 1293, 979, 1196, 1071]. Influence [507, 1263]. Info [956, 996, 1038, 1080, 1126, 1148, 1219, 1259]. Inform [476, 1353, 952]. Information [957, 997, 1039, 1081, 956, 996, 1038, 1080, 955, 995, 1037, 1079, 954, 994, 1036, 1078, 1127, 1149, 1220, 1260, 1126, 1148, 1219, 1259, 1125, 1147, 1218, 1258, 1124, 1146, 1217, 1257, 1309, 1358, 1412, 1461, 1308, 1357, 1411, 1460, 1307, 1356, 1410, 1459, 1511, 1554, 1598, 1635, 1510, 1553, 1597, 1634, 1509, 1552, 1596, 1633, 439, 684, 604, 357, 1525, 533, 327, 18, 1040, 1657, 823, 865, 858, 1524, 1340, 806, 1048, 728, 1471]. information-theoretic [1040]. Informative [1538, 526, 10, 336, 426, 491, 820, 1011, 1432, 1646, 1010, 1469, 1135, 821, 1112, 1619, 1331, 740, 1431, 1617, 1091, 1434]. informed [1638]. Infusion [320]. Ingram [1122]. Inhomogeneous [346]. initial [946]. Initiating [486]. initiation [1393, 1589]. INLA [1630]. insertion [1249]. Insights [313]. Instability [193]. Instrumental [1522, 1073, 637, 1573, 1013, 824, 1268, 749, 1266, 1463]. Insurance [658]. Integer [1659]. Integer-valued [1659]. integral [893]. integrate [728]. Integrated [1364, 1569, 1295]. Integrating [489, 882, 1120]. integration [920]. Integrative [1637, 1580, 746, 1340, 1170, 862, 1173, 1642, 1484, 792]. Intensities [36]. Intensity [110, 465, 346, 756]. inter [1087]. inter-subject [1087]. Interaction [287, 282, 120, 566, 1029, 894, 1388, 738, 1212, 1583]. Interactions [283, 450, 488, 298, 967, 974, 872, 1087, 748]. Intercepts [231]. Intercepts-Functional [231]. interdependent [998]. Interfaces [624]. interference [774]. interfering [1604]. Intergene [533]. interim [1314]. Intermediate [951, 300]. Intermittent [49]. Intermittently [109]. internal [900, 1564]. International [593, 1221, 721]. Interpretable [634, 1057, 918]. Interpretation [1408, 1070]. interrogation [1172]. Interval [3, 30, 1405, 1568, 1270, 1469, 1523, 1014, 1651, 693, 1386, 1480, 826, 1237, 1056, 1093, 981, 1091, 901, 1045, 1316]. Interval-Censored [3, 1405, 1270, 1469, 1523, 1014, 1386, 1237, 1093, 981, 1091, 901, 1045, 1316]. intervals [1444, 874, 1090, 982, 893, 927, 1665]. Intervention [570, 458, 1674, 1618, 1608]. Interventions [458]. Intra [550, 549]. Intra-Individual [550, 549]. intraclass [874]. Intractable [675, 1001]. Introducing [213]. Introduction [374, 1351, 472, 1594, 83, 1506, 1119, 382, 430, 657]. Introductory [214]. Invasive [650, 465]. Inverse [408, 148, 1370, 179, 402]. Inverse-Probability [402]. Inverse-Weighted [179]. invertebrates [1111]. Investigating [440, 872, 829]. Involving [422]. irregular [965]. ISBN [842]. ISBN-13 [842]. ischemia [1343]. Ising [798, 639, 163]. Isotonized [409]. Issue [474, 957, 997, 1039, 1081, 956, 996, 1038, 1080, 955, 995, 1037, 1079, 954, 994, 1036, 1078, 1127, 1149, 1220, 1260, 1126, 1148, 1219, 1259, 1125, 1147, 1218, 1258, 1124, 1146, 1217, 1257, 1309, 1358, 1412, 1461, 1308, 1357, 1411, 1460, 1307, 1356, 1410, 1459, 1511, 1554, 1598, 1635, 1510, 1553, 1597, 1634, 1509, 1552, 1596, 1633]. Issues [1214, 1062]. Italian [194]. item [139, 1250]. iterative [1600].

J [381, 621, 383, 329, 171, 252, 471, 84, 1299, 212, 1224, 557, 878, 660, 44, 1305, 169, 253, 1223, 1225, 877, 1630, 375]. J. [211, 434, 435, 84, 373, 212, 331, 90, 43, 257, 174, 624]. J.-H [878]. j.0006-341X.1999.00603.x [622]. j.1541-0420.2011.01689.x [621]. Jaakko [515]. JABES [1221]. Jacqmin [1186]. Jacqmin-Gadda [1186]. Jae [779]. James [1077, 293]. James-Type [293]. Janes [771]. Jang [690]. Japan [593]. Jarmin [1406]. Jasmine [1120]. Jason [1256, 1224, 1488, 1487, 1489, 1223, 1225]. Jay [1403]. Jean [1350]. Jean-Baptiste [1350]. Jeffrey [1503]. Jehua [811]. Jenny [1360]. Jeong [878]. Jerez [1595]. Jersey [877]. Jewell [1590]. Jialiang [718]. Jim [1628]. Jinqiao [1351]. Joe [1505]. Johansen [1435]. John [1590, 1301, 1183]. Johnson [1354]. Joint [610, 140, 1269, 441, 548, 182, 203, 506, 7, 868, 440, 1293, 1135, 1063, 979, 1515, 51, 371, 469, 49, 8, 334, 1093, 318, 324, 372, 1518, 580, 640, 243, 48, 607, 38, 597, 1365, 507, 1391, 822, 1565, 1208, 1449, 745, 830, 1094, 1064, 821, 1229, 1178, 978, 1517, 1644, 921, 1002, 1470, 1612, 1583, 1355]. Joint-Action [548]. Jolly [114]. Jolly-Seber [114]. Jones [382, 661]. Jose [1595, 992]. Joseph [1505, 1302, 720]. Jr [211]. Judea [1590]. Judgment [473, 409]. Julia [1406]. Julian [1299, 1630]. Julious [431]. Jun [1141, 779]. June [621, 622]. Jung [778]. Jure [1503].

K. [433, 552, 45, 553, 377, 555]. Kalantar [1488, 1487, 1489]. Kalantar-Zadeh [1488, 1487, 1489]. Kamyar [1488, 1487, 1489]. Kane [1143]. Kang [771]. Kaprio [515]. Karl [912]. Kass [1184]. Katie [1035]. Keen [373]. Ken [1305]. Kendall [148]. Kenward [877]. Kernel [350, 283, 1004, 668, 920, 1450, 904, 793, 1003, 1400, 818, 1297, 1399, 1583]. Khmaladze [658]. Khuri [215]. kidney [862, 1173, 1019]. Kim [779]. Kinesin [241]. Kinetic [33]. kinetics [1208]. King [329]. Kingdom [1409]. Kirk [1187]. Kishiro [621]. Kjell [1354]. Kleinman [376, 1305]. Kneib [591]. Knowledge [489]. Kobe [593]. Kolaczyk [91]. Kreuter [1406]. Kriging [347]. Kruskal [608, 198, 875]. Kruskal-Wallis [608, 198, 875]. Krzanowski [84]. Kuhn [1354]. Kullback [441]. Kupper [258]. Kuznetsova [992]. Kwang [779].

L [1673, 591, 41, 258, 1215, 1302, 556, 1670, 661, 254, 45, 1594, 1122, 624]. L. [379, 374, 623]. L.L [1674]. Laan [1143, 592]. Label [1113]. Label-noise [1113]. labor [809]. Lachin [1139]. Lacking [574]. Lag [619, 898, 1245, 1675]. Lagged [579]. Lahiri [1631]. Lai [172]. Laird [1077, 385]. Lakes [369]. Lamprey [369]. Land [201]. Lane [1406]. Lange [433, 385, 377]. Langrock [1594]. Large [496, 999, 534, 561, 671, 55, 1131, 1584, 1535, 1656, 1030, 1527, 704]. large- [1535]. Large-Scale [496, 671, 999, 1584, 1030, 1527]. Laricca [89]. Lash [45]. Lasso [461, 1401, 853, 1582, 1657, 896, 1311, 855, 1265, 1099, 463, 1264, 1479]. lasso-type [853]. Latent [245, 673, 564, 526, 137, 185, 344, 428, 7, 440, 25, 523, 205, 569, 450, 460, 1518, 652, 1055, 820, 1574, 1089, 1174, 1601, 1479, 1542, 1394, 1344, 1196, 1159, 1093, 935, 1367, 1209, 1068, 1421, 1599]. Latent-Class [673]. Lawrence [624, 432]. Lawson [331]. lead [1517]. Leadership [1122]. Learning [1276, 1401, 174, 193, 299, 920, 1642, 1650, 1587, 1192, 1427, 1152, 592, 432]. Least [1165, 601, 1672, 1600, 886]. Least-Squares [601]. Lee [375]. Left [180, 262, 1063, 982, 1320]. left-censored [1063]. Left-Truncated [262, 982, 1320]. Lei [590]. Leibler [441]. Lemey [46]. Length [469, 293, 54, 682, 759, 1523, 1380, 1178, 1106, 1249, 978, 1318]. Length-Biased [293, 54, 759, 1523, 1380, 1318]. length-frequency [1106]. Leon [744]. Leskovec [1503]. Letter [689, 127, 514, 512, 324]. Level [604, 42, 336, 157, 896, 1085, 1450, 1101, 1253, 1604, 835, 1086, 1297]. Levels [219, 24, 1109, 1002]. Levene [1247]. Leveraging [1652]. Li [1674, 1561, 1355, 718, 1562]. Liang [1224, 1223, 1225]. Lichtenstein [515]. Lies [1593]. Life [379, 658, 1498, 878, 132]. lifespan [1203]. Lifetime [664, 266, 964, 1046]. Lifetimes [516, 982]. Lijoi [655]. Likelihood [28, 220, 681, 207, 498, 281, 50, 444, 585, 1197, 428, 260, 202, 635, 246, 145, 697, 322, 929, 451, 19, 89, 1673, 764, 1332, 1620, 757, 1284, 909, 1097, 1661, 1104, 1536, 693, 795, 763, 869, 872, 884, 707, 924, 773, 800, 1320, 968, 1599, 1583, 910, 1371, 1564]. Likelihood-Based [28, 145, 697, 929, 924, 1599, 1371]. Likelihoods [675, 1001]. Lima [169]. Lin [1256, 1416, 1415, 168]. Line [621, 643, 24, 646, 201, 454, 934]. Linear [140, 394, 116, 101, 604, 430, 79, 1299, 212, 118, 277, 1145, 316, 452, 447, 600, 250, 159, 204, 170, 249, 472, 657, 543, 670, 12, 677, 361, 1074, 528, 783, 691, 15, 340, 1390, 1244, 1424, 741, 1576, 1245, 1675, 1437, 1135, 1321, 855, 1639, 530, 1267, 795, 1472, 1534, 1384, 1374, 695, 799, 1167, 888, 700, 1441, 215]. Link [406, 1199, 1044, 988]. Link-Tracing [406]. Linkage [28, 281, 1035]. Linked [1579]. Liquid [190]. Lisa [1142]. list [1194]. listed [1339]. lists [918]. Liu [1674, 625]. liver [707]. Loads [316]. Local [613, 407, 919, 636, 308, 1568, 1319, 736, 1647]. Local-aggregate [919]. localization [705]. localized [737]. Locally [216]. location [1662, 893, 1261]. Loci [30, 390, 1172]. Log [1582, 530, 817, 1049, 1384, 1375]. log-exp-sum [817]. Log-linear [530]. log-rank [1375]. log-rank-type [1049]. Log-ratio [1582]. Logistic [463, 391, 616, 1074, 686, 1533, 710, 1324, 1476, 1325, 1113, 1069, 984, 1181, 1640, 471]. Logrank [9]. Logrank-Type [9]. Long [621, 666, 578, 202, 288, 454, 1115]. Long-Diving [621, 454]. Long-Term [666, 202, 288, 1115]. Longini [211]. Longitudinal [108, 673, 125, 1077, 591, 439, 526, 335, 604, 231, 98, 579, 7, 1355, 866, 440, 635, 13, 44, 230, 115, 154, 371, 541, 1349, 523, 205, 14, 136, 8, 334, 155, 228, 324, 52, 315, 606, 672, 671, 365, 525, 229, 372, 1518, 1517, 236, 109, 314, 438, 337, 183, 17, 59, 507, 581, 1391, 1673, 853, 822, 1269, 920, 1332, 1620, 1423, 761, 696, 1449, 937, 883, 830, 1445, 868, 1064, 1135, 1063, 885, 1267, 1498, 1136, 1112, 809, 1619, 697, 984, 749, 1157, 884, 905, 1093, 1618, 1209, 965, 1335, 921, 1333, 1365, 130]. looking [925, 1315]. looks [1314]. Looney [1302]. Lorien [1488, 1487, 1489]. lost [1203]. Lot [1578]. Loughin [953]. Low [78, 561, 1001, 967, 1478, 1580, 1526]. low-count [1001]. Low-Rank [561, 967, 1478, 1580]. low-risk [1526]. Lower [503, 765]. LQAS [1578]. Ltd [1590]. Ltd. [1504]. Luca [1185]. Lung [77, 299]. Luo [1561, 1562]. Lyle [1668, 1075]. Lynch [319, 1205]. Lysate [276]. LZIP [1196].

M [211, 1077, 435, 810, 1672, 1673, 252, 1139, 1629, 1488, 1303, 1487, 130, 1489, 720, 992, 44, 169, 1301, 1404, 254, 45, 385, 472, 953, 877, 1407, 592, 167, 46, 375, 624, 83, 1183, 432]. M. [471, 258, 557, 46]. Ma [718]. MacDonald [374]. Machine [350, 668, 904, 1003, 1400, 1004, 818, 1399, 1583]. Machines [283]. Macroparasite [199]. Macular [463]. Madsen [430, 324, 372]. magna [903]. magnetic [1605, 1627]. Making [40, 1192]. malaria [898]. Mallick [553]. Mammal [201]. Management [1305, 376]. Mangion [1667]. Mantel [1155]. many [977]. Map [278]. Mapping [85, 30, 496, 278, 539, 390, 389, 1169, 706, 811]. MAR [383, 273]. Marc [1140]. March [1256]. Marco [1350]. Marginal [1011, 620, 1027, 845, 846, 1092, 468, 176, 847, 1527, 1619, 683, 525, 1071, 941, 938, 1673, 1520, 1620, 717, 1670, 984, 848, 844, 849, 1431, 44]. marginalization [1345]. marginalized [1381]. Maria [1590]. Mark [160, 307, 650, 1143, 598, 453, 25, 200, 1407, 1363, 985, 1606]. Mark-Recapture [307, 650, 25, 200]. Mark-Recapture-Recovery [160]. mark-resight [985, 1606]. Mark-Specific [598]. Marked [345, 1042]. Marker [98, 537, 466, 822, 814, 1003, 1400, 1182, 747, 1653, 1389, 1399]. marker-guided [747, 1389]. marker-set [1003, 1400, 1399]. Markers [117, 350, 480, 595, 229, 59, 1398, 1313, 1663, 1240, 1082, 901]. market [711, 908]. Markov [374, 1362, 1360, 661, 1594, 1361, 1666, 820, 614, 643, 934, 101, 1285, 1359, 1442, 1586, 1028, 1386, 871, 905, 1532, 897, 1209, 102, 100, 270, 931, 738, 856].

Markov-modulated [934]. Marks [650, 598, 1666]. Marques [1121]. Marshall [132]. Martin [1401]. martingale [892]. Masking [1507]. Mass [201, 735]. Massive [1503]. Masthead [955, 995, 1037, 1079, 1125, 1147, 1218, 1258, 1308, 1357, 1411, 1460, 1510, 1553, 1597, 1634]. Matched [237, 110, 323, 354, 571, 396, 1547, 1294, 708, 1025, 943, 805]. matched-pair [805]. Matched-Pairs [323]. Matching [733, 107, 676, 467, 1465, 923, 1086, 1664]. Maternal [451, 968]. Mateu [660]. mating [1240]. Matlab [171]. Matloff [1402]. Matrices [184, 1131, 1263]. Matrix [353, 321, 777, 1377, 1579, 828, 1230]. Matrix-based [353, 1377]. Matsui [1140]. matter [829]. Matthias [1348]. Max [1354, 1609]. max-stable [1609]. MAXENT [587]. Maximin [1243]. maximize [1373]. Maximum [498, 757, 585, 1104, 202, 632, 800, 1661, 707, 89]. Maxwell [1027]. McDonald [1594]. McNemar [152]. MDR [802]. MDR-TB [802]. Mean [20, 392, 197, 1498, 352, 266, 516, 1520, 1432, 852, 1536, 1380, 959, 1431, 773, 1392, 1071, 828, 980]. meaningful [867]. Means [406, 1283, 868, 1248]. Measles [487, 1659]. Measure [267, 1319, 1224, 1102, 1222, 1226, 1223, 1225, 1050]. measured [1589]. Measurement [347, 203, 304, 506, 617, 155, 616, 1618, 60, 303, 48, 183, 424, 1540, 1324, 1632, 1041, 969, 1372, 786, 695, 1323, 1434, 787]. Measurements [125, 21, 228, 537, 247, 867, 1437, 1064, 784, 1099]. Measures [627, 573, 669, 353, 628, 629, 280, 626, 507, 1199, 1657, 1193, 1331]. mechanism [727, 1517]. Mechanistic [182, 1042, 502, 647, 1208]. Media [1184]. Median [530, 510, 483, 1114, 982]. Mediation [355, 519, 1023, 812, 1601, 1007, 1648, 1153, 787, 1605]. mediators [812, 1007]. Medical [370, 205, 565, 590, 1537, 815, 1519, 1046, 821, 711, 656]. Medicare [51]. Medicine [1142, 718, 87, 1150, 1419, 1404, 1140, 1407]. Mee [42]. MEG [705]. Meissner [1304]. Melody [1592]. Memory [578]. Mendelian [1213]. Meng [661]. Mengersen [556]. Meningococcal [465]. Menon [1407]. Menstrual [469, 978]. Mental [205]. Merging [921]. Meta [359, 544, 545, 1416, 416, 421, 585, 462, 197, 791, 233, 543, 1118, 813, 874, 1414, 1377, 1193, 1413, 1417, 873, 1570, 1344, 912, 1415, 806, 1067, 1622, 1215, 327]. meta-analyses [1570]. Meta-Analysis [544, 545, 1416, 585, 197, 233, 543, 359, 416, 421, 791, 1118, 874, 1414, 1377, 1193, 1413, 1417, 873, 1344, 912, 1415, 1067, 1622, 1215, 327]. meta-analytic [813]. meta-analytic-predictive [806]. Meta-Experimental [462]. metabolomic [1340]. Metapopulation [240]. metastatic [902]. Metcalfe [214]. Method [220, 544, 110, 455, 184, 303, 312, 276, 601, 291, 517, 567, 1055, 1632, 1072, 1132, 793, 1377, 1013, 1448, 1254, 884, 797, 863, 1367, 840, 893, 981, 703, 1261, 1333, 1612, 983, 1371, 1564]. methodological [1631, 1035]. Methodology [1033, 215]. Methods [39, 434, 621, 160, 24, 498, 1141, 551, 1406, 1185, 310, 328, 1352, 1121, 271, 78, 246, 91, 169, 1249, 51, 582, 719, 454, 779, 1075, 54, 1595, 213, 740, 658, 375, 1213, 34, 378, 302, 437, 811, 19, 1434, 308, 1591, 1624, 762, 1236, 1094, 940, 1020, 1606, 1372, 1162, 1500, 1387, 816, 1025, 924, 1019, 1389, 829, 1493, 1571, 960, 787, 1505, 386, 87, 255, 743, 720]. methylated [1586, 1532]. Methylation [489, 1201, 1529, 1484, 1002]. Metric [466]. Metrics [80]. Meyer [1032]. Michael [1032, 1672, 383, 1591, 1143]. Microarray [79, 490, 279, 63, 534, 318, 1101, 1002]. Microarrays [93, 173]. Microbiome [686, 1450, 1534, 1367, 1613, 1297]. microorganisms [1231]. Miglioretti [689]. Miguel [1595]. MILFM [1421]. Minamikawa [621]. Minimal [467]. Minimization [367]. Minimize [367]. Minimum [815, 703]. Mining [1502, 174, 1503, 1304]. miRNA [1531, 862]. miRNA-target [862]. miRNAs [277]. Misaligned [561]. Misalignment [496, 763]. misclassification [1632, 991, 989, 990]. Misclassified [57, 1014, 922]. Misidentification [25, 800]. mislabel [1325]. Mismeasured [146]. Missing [237, 335, 272, 98, 366, 16, 110, 506, 271, 568, 238, 154, 254, 541, 520, 109, 302, 268, 401, 183, 17, 505, 1256, 727, 725, 1138, 1494, 1110, 889, 1433, 724, 869, 1087, 1467, 943, 726, 876, 1266, 921, 1248, 1493, 1144]. Missing-Data-Induced [110]. Missingness [521, 365, 18, 1673, 1620, 699, 1156, 1493]. Misspecification [250, 204, 249, 1134, 1388]. misspecifications [1496]. Misuse [948, 949]. Mitchell [1629]. Mixed [673, 404, 140, 394, 290, 116, 101, 338, 79, 99, 1299, 287, 562, 203, 506, 232, 316, 453, 452, 13, 21, 250, 230, 469, 249, 14, 744, 543, 670, 963, 351, 528, 542, 1100, 1390, 853, 1666, 820, 1329, 1244, 1287, 1134, 717, 1345, 1437, 784, 795, 905, 783, 691, 1326, 819, 1371]. Mixed-Effect [673]. Mixed-Effects [140, 394, 562, 203, 506, 316, 230, 853, 1666]. Mixing [185]. Mixture [564, 383, 360, 241, 609, 1106, 649, 195, 150, 221, 341, 520, 895, 273, 484, 427, 311, 1567, 1347, 1501, 1481, 757, 944, 836, 702, 1368, 1104, 1115, 1529, 1341, 1162, 1290, 728]. Mixtures [82, 31, 936, 141, 1053, 1195, 868, 1105, 1166]. mobile [929]. Model [237, 313, 564, 383, 689, 28, 495, 317, 92, 1324, 117, 290, 526, 1154, 186, 338, 1561, 231, 522, 137, 71, 79, 289, 366, 527, 1158, 360, 1299, 16, 180, 224, 182, 105, 310, 678, 33, 215, 1442, 584, 241, 7, 40, 277, 1046, 598, 720, 271, 832, 1562, 1560, 1563, 21, 667, 145, 177, 393, 165, 649, 391, 200, 204, 469, 465, 163, 1462, 205, 387, 292, 114, 569, 566, 54, 195, 533, 221, 521, 1334, 1431, 572, 406, 390, 1518, 773, 74, 1058, 1649, 273, 484, 528, 427, 1441, 686, 445, 15, 601, 59, 291, 597, 340, 62, 285, 581, 586, 796]. model [987, 1567, 1673, 853, 1397, 892, 1501, 903, 1420, 1161, 712, 1287, 1620, 798, 1482, 1432, 1423, 1174, 865, 946, 986, 1235, 1541, 928, 1001, 1449, 891, 776, 945, 1270, 1014, 857, 922, 745, 1169, 1054, 1007, 1573, 1542, 1586, 802, 1189, 926, 1437, 1135, 969, 1321, 1063, 821, 737, 885, 851, 1047, 1095, 1478, 1492, 756, 732, 788, 959, 1472, 978, 1108, 985, 1433, 1422, 1438, 1369, 1570, 1227, 711, 706, 1196, 881, 1384, 1133, 1162, 1157, 1395, 1093, 988, 947, 760, 1636, 1066, 1067, 1280, 1658, 1532, 726, 799, 800, 981, 1231, 1665, 1641, 700, 728, 1230, 1130, 1421, 1419, 1623, 1640, 938, 1622, 1611]. model [856, 1468]. Model-averaged [1441, 945]. Model-Based [92, 105, 40, 285, 1324, 946]. Model-free [1334, 390, 1095]. Model-robust [1158]. Modeling [245, 1032, 1273, 495, 1053, 319, 501, 398, 220, 95, 1205, 1141, 641, 121, 416, 421, 499, 182, 428, 369, 745, 508, 190, 118, 1393, 1355, 1215, 348, 646, 440, 36, 158, 635, 1394, 244, 809, 163, 49, 998, 553, 341, 587, 1354, 707, 682, 1228, 134, 668, 436, 1074, 536, 842, 532, 1428, 38, 191, 222, 581, 781, 1566, 964, 1179, 1269, 1026, 971, 1206, 966, 1272, 1191, 919, 1325, 1101, 868, 1293, 1021, 1670, 1178, 1024, 1136, 1344, 1117, 880, 1159, 870, 1630, 1609, 1002, 1530, 716, 977, 1348, 1353, 952, 387, 332]. Modelling [556, 520, 1569]. Models [108, 673, 20, 613, 383, 404, 140, 643, 394, 58, 286, 116, 101, 272, 207, 604, 529, 441, 392, 240, 281, 185, 99, 675, 845, 287, 344, 416, 421, 562, 461, 846, 374, 357, 282, 118, 203, 304, 506, 232, 1405, 578, 111, 609, 588, 316, 202, 453, 452, 13, 288, 447, 600, 37, 530, 250, 44, 230, 468, 176, 582, 847, 204, 170, 688, 683, 249, 14, 470, 88, 611, 1594, 473, 358, 543, 114, 680, 502, 670, 12, 561, 587, 8, 334, 1034, 589, 619, 677, 460, 102, 1074, 322, 580, 303, 100, 270, 72, 509, 109, 273, 364]. Models [351, 493, 261, 437, 445, 56, 446, 405, 223, 11, 507, 542, 567, 1291, 1199, 1390, 1391, 861, 1520, 1666, 1347, 820, 1582, 822, 1244, 934, 1424, 1000, 1522, 1089, 1006, 1476, 1073, 1426, 759, 850, 757, 1027, 1285, 944, 836, 1208, 1646, 1616, 1134, 1001, 852, 890, 945, 1096, 970, 1576, 1158, 762, 702, 1245, 1675, 1224, 717, 1236, 1197, 1104, 830, 1345, 1115, 1094, 1023, 1479, 1377, 930, 1437, 1064, 1028, 1386, 871, 1656, 1110, 1292, 824, 1164, 1529, 1580, 1222, 1226, 758, 1267, 1229, 795, 873, 979, 1317, 1106, 1512, 1534, 1268, 1341, 1223, 801, 1042, 1643, 1383, 1225, 1278, 1162]. models [1207, 905, 1204, 963, 848, 1467, 844, 849, 1025, 1175, 695, 1209, 1048, 1167, 1617, 1652, 773, 1621, 1290, 1645, 1323, 1517, 931, 1071, 895, 941, 783, 888, 1062, 691, 729, 728, 1530, 1098, 1261, 652, 1602, 785, 738, 825, 1246, 1599, 1271, 1514, 819, 899, 1186, 1299, 1145, 430, 471, 212, 91, 472]. moderate [1464]. Moderation [20]. Modern [993, 257, 385]. modification [1464]. Modified [448, 765, 1315, 1564]. modifier [888]. modifying [754]. modulated [934]. Modules [280]. Molecular [326]. Molenberghs [130]. Moloney [1187]. Moment [1117, 303]. Moment-Adjusted [303, 1117]. momentary [1386, 963]. Moments [544, 73, 1377]. Mónica [1669]. Monitoring [1403, 622, 1, 43, 502, 565, 1282, 1031, 907, 1352]. Monographs [842]. monotone [726, 703, 755, 1045]. Monotonely [236]. Monotonicity [323, 363, 1471]. Monte [661, 213, 1372, 540]. Montgomery [657]. Moodie [1672]. Moonseong [1349]. Morgan [329]. Morphometrics [90]. Morris [472]. mortal [1493]. Mortality [317, 78, 987, 730, 1611]. Mother [161, 498]. Mother-Child [498]. Mother-to-Child [161]. Motif [95]. Motivating [1444]. Motor [241]. Mouse [198]. mouth [907]. Movement [485, 1018, 998, 1607]. moving [884]. Muller [375, 660]. Multi [139, 1179, 1544, 678, 1341, 676, 1420, 1173, 865, 1642, 1282, 1483, 1251, 1376, 1542, 1446, 807, 1295, 1580, 1614, 1621, 1192, 931, 925, 1315, 728, 1086, 942, 899, 1468, 1405]. Multi-Arm [676, 1282, 925, 1315]. multi-armed [1614]. multi-dimensional [1251, 1542]. multi-drug [1376]. Multi-item [139]. multi-level [1086]. multi-omics [1483]. Multi-parameter [1179]. multi-phase [942]. multi-platform [1173]. multi-reader [807]. multi-resolution [1420]. multi-source [1446]. multi-stage [1282, 1192]. multi-state [865, 1621, 931, 728, 899, 1468, 1405]. Multi-study [1544]. Multi-subgroup [1341]. Multi-Subject [678]. multi-test [807]. multi-treatment [1192]. multi-view [1642, 1295, 1580]. Multicategory [1650]. Multiclass [1639]. Multidimensional [64, 904]. multidrug [802]. multidrug-resistant [802]. Multilevel [185, 369, 119, 492, 672, 870, 581, 1390, 854, 879]. Multinomial [209, 25, 636, 686, 1199, 1168, 935, 1280, 1231, 1640]. Multiphase [208, 1313]. Multiple [245, 1520, 650, 350, 226, 359, 95, 272, 920, 194, 1168, 33, 111, 1494, 489, 1327, 154, 490, 61, 582, 541, 8, 571, 318, 450, 402, 361, 897, 671, 156, 390, 93, 687, 65, 314, 218, 302, 528, 152, 1137, 1421, 1100, 1279, 861, 1016, 1531, 962, 1674, 864, 1174, 1584, 812, 1578, 1449, 776, 1483, 1250, 1197, 1519, 1177, 1377, 1473, 1528, 855, 1103, 1480, 746, 1151, 1515, 1373, 1638, 1341, 1065, 1539, 943, 1275, 799, 1314, 1290, 1335, 1323, 921, 1002, 1099, 1571, 1608, 1605, 983, 877, 384, 131]. Multiple-Imputation-Based [8]. multiple-index [776]. multiple-objective [1578]. Multiple-Probe [490]. Multiple-Sample [218]. multiplex [1538]. Multiplicative [424, 1616, 762]. Multiplicity [407]. Multireads [278]. Multiscale [525, 94]. multistage [1059, 967]. Multistate [200, 1386, 821]. Multistep [276]. multisubject [1627]. Multitreatment [543]. Multitype [73]. multivariable [986]. Multivariate [1216, 117, 186, 544, 679, 573, 224, 428, 1328, 144, 930, 598, 1201, 855, 1021, 905, 103, 302, 55, 15, 581, 1100, 1574, 761, 936, 1052, 1648, 1377, 802, 1064, 1670, 1178, 1433, 1210, 1389, 799, 1209, 1609, 1441, 1255, 1434, 328]. munition [903]. Munzert [1304]. mutation [1328]. Mutational [319]. MW [1215]. MW-L [1215]. myelodyplastic [804]. myeloma [1341].

N [172, 1184, 552, 385, 432]. N. [556, 376, 89]. Naive [617]. Nan [1077]. Nason [127]. Natalia [1256]. Natural [637, 596, 651, 518]. Neelon [258]. negative [796, 1529]. Nested [20, 123, 53, 136, 861, 892, 1189, 1494, 788, 959, 1369, 708, 1609, 942, 1381]. net [1049, 1513, 1442, 1190]. Nettleton [1076]. Network [137, 487, 348, 631, 91, 63, 485, 1329, 1637, 1102, 1377, 1658, 1641, 1533]. Network-Based [631, 487, 348, 1102]. Networks [137, 1350, 862, 911, 741, 1483, 1359, 1362, 1376, 1276, 1360, 1528, 881, 1277, 1361]. Neural [1184]. neuroimages [1657]. Neuroimaging [688, 525, 1026, 919, 897, 1652, 785]. Neuron [36]. neuroscience [1252]. Newly [423]. Next [278, 277, 1076, 1132, 1276]. Next-Generation [278, 277, 1076, 1132]. Ng [254, 1214]. Nguyen [1488, 1487, 1489]. Nicholas [1590, 1305]. Nie [513]. Niels [515, 1353, 952]. Nigeria [245]. Ning [1355]. NMF [321]. no [809]. nodes [1528]. Noel [1667]. noise [741, 930, 1113]. Non [673, 613, 650, 186, 1329, 1244, 1584, 1245, 1675, 763, 711, 726, 931, 1493, 703, 755, 960, 1166, 1170, 1116, 856]. non-exchangeable [711]. Non-Gaussian [186, 1166]. Non-Ignorable [673, 1493]. non-inferiority [960]. Non-Invasive [650]. non-linear [1244, 1245, 1675]. Non-Local [613]. non-Markov [931, 856]. non-monotone [726, 703, 755]. non-normal [1329, 763, 1116]. non-parametric [1170]. non-randomized [1584]. nonadditive [1096]. Noncentrality [534]. Noncompliance [296, 1622]. Nondestructive [23]. Nonexchangeable [459]. Nonhomogeneous [270]. Nonidentifiable [72]. Nonignorable [145, 365, 109, 17, 724]. Noninferiority [370, 1214]. Noniterative [254]. Nonlinear [404, 501, 527, 506, 316, 284, 567, 1287, 874, 1208, 717, 1197, 881, 1117, 1652]. Nonmonotone [365]. Nonnegative [321]. Nonparametric [227, 1256, 1203, 503, 1299, 287, 1138, 1051, 328, 36, 871, 1020, 305, 560, 132, 155, 228, 397, 484, 1571, 900, 269, 483, 977, 1281, 856, 1468, 1195, 765, 1096, 702, 831, 802, 1153, 1044, 1372, 1467, 736, 1436, 1045, 938, 381]. nonparametrics [972]. Nonproportional [529, 179]. Nonsmall [299]. nonstationarity [1228]. nonstationary [1336]. Nordic [515, 4]. Normal [207, 203, 316, 686, 1329, 763, 1640, 1116]. Normal/Independent [316]. Normalcy [427]. Norman [1402]. Note [383, 550, 549, 588, 323, 250, 249, 589, 273, 909, 1345, 991, 1400, 989, 990, 1399, 910]. nothing [1047]. Novel [106, 576, 1531, 1204, 1067]. nowcasting [803]. Nuclear [282]. Nuisance [356]. Null [219, 1368, 1648, 1655]. Number [609, 534, 1000, 1482, 1010, 1440, 1054, 704]. Numerical [437, 377]. nutrients [1231]. Nutritional [60, 969]. Nyhuis [1304].

O [329, 171, 993, 1408]. O104 [803]. Obesity [467]. Objective [613, 1578, 1130]. O’Brien [258]. Observation [526, 49, 57, 269, 1013, 821, 1289]. Observational [665, 129, 662, 354, 571, 671, 266, 516, 1191, 1464, 1207, 976, 1155, 1175, 1086, 1664]. Observations [403, 225, 685, 1539, 1392, 1045]. Observed [219, 662, 605, 270, 963, 740, 1621, 929]. observed-cluster [740]. Observer [24, 934]. Obstetric [481]. Occupancy [574, 206]. Occurrence [465]. Odds [272, 271, 618, 636, 364, 891, 927, 1611]. Oedekoven [1121]. off [714, 833]. Oja [328]. Okamura [621]. older [776]. Olga [992]. Olkin [1122, 132]. Omics [1120, 1483]. Omitted [667]. Omnibus [668, 448, 1602]. on-the-fly [733]. Oncogenic [193]. Oncology [500, 1631]. One [197, 302, 781, 827, 801, 1375]. One-Parameter [197]. one-sample [827, 1375]. One-Sided [302]. one-step-ahead [801]. Only [547, 363, 147, 961, 1398]. Onset [617, 134]. Ontario [539]. Open [240, 574, 1661, 114, 1338]. Operating [75, 342]. opportunistic [1031]. Optimal [642, 67, 1312, 644, 1578, 1466, 153, 403, 239, 1373, 75, 632, 467, 517, 1008, 714, 1059, 1198, 793, 1480, 1341, 1587, 1475, 1175, 840, 1109, 1243, 1333, 1664, 1427, 1152]. Optimally [750]. Optimization [69, 1557, 1556, 919, 1261, 1558]. Optimize [640, 767, 772, 768, 771, 770]. Optimizing [320]. Optimum [1244, 968]. options [1136]. O’Quigley [420]. Oracle [1098]. Order [71, 679, 633, 265, 1479, 124, 163, 1248, 1575, 1284]. Order-Restricted [71, 1248]. Ordered [613, 679, 64, 508, 1479, 976]. Ordering [41, 312, 1099]. Ordinal [1199, 69, 271, 123, 1074, 1430, 1009, 1663, 1619, 1447, 984, 794, 818]. Ordinary [286, 637, 1165, 437, 946, 825]. ordination [1367]. Organ [664, 1491]. origin [1612]. orthogonalized [794]. Other [1034, 460]. Our [594, 274]. outbreak [803]. Outbreaks [187, 308, 1000]. Outcome [209, 422, 7, 440, 315, 606, 1265, 1518, 66, 607, 340, 196, 1316, 1572, 831, 1650, 1373, 1069, 724, 1136, 697, 965, 1266, 729, 887, 1152]. outcome- [1069]. Outcome-adaptive [1265]. Outcome-Dependent [315, 340, 196, 1316, 965]. outcome-related [697]. Outcomes [398, 104, 480, 294, 288, 393, 1349, 8, 363, 565, 467, 15, 575, 1100, 1520, 920, 1319, 1339, 1012, 1009, 1449, 830, 1345, 1364, 1519, 1589, 1377, 1663, 868, 1293, 698, 1201, 1095, 713, 1447, 1462, 1087, 1065, 818, 748, 870, 1275, 652]. outlying [1193]. Output [495]. outs [1135]. over-dispersed [1496]. Overall [583, 295]. Overdispersed [116, 1291]. overdispersion [1053]. overflights [933]. overidentified [1318]. Overstate [460]. Oxford [330]. Oysters [368]. ozone [987].

P [380, 382, 622, 1672, 252, 129, 1488, 1487, 1489, 214, 169, 45, 213, 255, 46, 375, 842, 743, 89, 623, 1391, 307, 871]. P-Splines [307, 1391, 871]. P. [382, 329, 252, 430, 384, 720, 170, 375]. Pacific [368]. Pair [498, 1547, 1294, 805]. paired [1449, 1280]. paired-multinomial [1280]. Pairs [323, 396, 1342]. Pairwise [681, 105, 764, 795, 1320]. Pan [511, 512]. Panel [270, 269, 1432, 1293, 1061, 819, 1371]. panel-count [819, 1371]. panels [1389]. Paper [418]. paradigm [1631]. paradigms [813]. Parallel [1402]. Parameter [197, 534, 488, 1179, 1059, 727, 947, 1617, 1099, 1261, 1255]. Parameterization [636]. Parameters [113, 199, 356, 609, 243, 946, 1345, 1606, 773, 893]. Parametric [1291, 762, 966, 1272, 693, 132, 1232, 509, 1540, 964, 1522, 712, 1341, 1331, 726, 1493, 1098, 1170]. Parametrically [397]. Parasite [641]. parasites [898]. Parasitic [369]. Parent [488, 1612]. parent-of-origin [1612]. Parity [205]. Parsimonious [1012, 918]. Part [205]. Partial [50, 777, 322, 312, 340, 1576, 1267, 968, 1246, 1564]. Partial-Likelihood [50]. Partially [604, 447, 600, 568, 15, 1135, 1472, 963]. Participants [581]. Participation [208]. particulate [829]. partition [1623]. partitioned [1165]. Partitioning [531, 1330]. partitions [1535]. Partly [599, 1270, 1191]. Partnering [245]. Path [1409, 294, 569, 899]. Path-Specific [569]. pathogen [1206, 832]. pathogen-specific [1206]. Pathogens [583, 1440]. Pathway [975, 755]. Pathways [193, 1533]. Patient [1312, 1301, 767, 772, 768, 1083, 771, 770]. Patients [486, 617, 502, 76, 1208, 1339, 776, 1490, 1486, 1488, 1487, 1489, 1396, 1448, 1342, 708, 1526]. Patrick [1224, 1223, 1225]. Pattern [383, 944, 1187, 273, 1319, 1290]. Patterns [1033, 785]. Paul [515, 1352]. Pawel [1502]. PCAN [1116]. PCB [340]. PCR [717]. Peace [912, 86]. Pearl [1590]. Pearson [1234]. Peck [657]. Pedigrees [451]. Pedroso [169]. peer [749]. PEG [1249]. Penalization [536]. Penalized [3, 207, 338, 1287, 1097, 631, 63, 438, 1647, 437, 901, 284, 89, 727, 1482, 1245, 1675, 763, 1600, 872, 947, 789]. penalties [782, 1664]. penalty [817, 777, 947]. PenPC [973]. peptide [1101]. Percentile [448]. Performance [1352, 1006, 1339, 1374, 1061]. period [1525]. Peritoneal [198]. Permutation [559, 452, 81, 947, 1330, 41]. permutations [1330]. Perrett [212]. personal [1662]. Personalized [1407, 1526, 1024, 1082, 1152]. perspective [1466, 999]. perturbations [1483]. Perú [833]. Pesarin [41]. Peter [1143, 1557, 1304, 1558]. Petersen [1194]. Petris [170]. Petrone [170]. Petros [1143]. Pettitt [556]. Pfeiffer [1629]. Phage [577]. Pharmaceutical [131]. Pharmacogenetic [404, 872]. Pharmacokinetics [403]. Pharmacophores [244]. Pharmacovigilance [39]. Phase [778, 1557, 612, 100, 1559, 1210, 1084, 942, 1558, 242, 106, 1444, 1556, 500, 69, 1443]. Phase-Type [100]. phases [898]. phenology [1607]. phenotype [904, 1327]. Phenotypes [118, 1279]. phenotyping [1519]. phosphorylation [1343]. Phylogenetic [46]. phylogeography [766]. Physical [1500]. PICS [192]. piecewise [1021, 1107]. Pinheiro [992]. Pivotal [455]. Placebo [597]. Planning [315]. plans [1578]. Plant [113, 1240]. platform [1173, 1446]. Pleiotropy [1638]. Pleuritis [545]. Plot [275]. Plots [8]. Pneumonia [19, 708]. pneumoniae [648]. Point [615, 50, 345, 159, 1187, 587, 93, 346, 1033, 714, 1235, 1284, 857, 930, 885, 809, 1042, 1068]. Point-Pattern [1187]. Point-Process [50]. points [933]. Pointwise [1263]. Poisson [653, 654, 934, 1027, 649, 587, 785]. policies [959, 1249]. Policy [476]. Pollen [74]. pollutants [903]. Pollution [158, 532, 1662, 737, 1201, 709, 1609]. Polya [222]. polynomial [1647]. polynomials [712]. polytomous [1476]. Polyvalent [583]. Pooled [810, 2, 247, 859]. Pooling [393, 1069, 713]. Population [329, 307, 456, 604, 574, 199, 360, 349, 985, 576, 649, 458, 560, 221, 406, 216, 261, 333, 135, 27, 710, 781, 1569, 1269, 823, 1363, 1661, 1338, 766, 1515, 1105, 1288, 1604, 894, 895, 1018, 1238]. Population-Based [135, 333, 1238]. Population-Level [604]. Populations [550, 549, 119, 322, 838, 1378, 1385, 1543, 938]. positional [1531]. Positive [689, 584, 177, 486, 572, 1083]. Possible [283]. Possibly [633, 1065]. post [1172, 711, 908]. post-analytic [1172]. post-market [711, 908]. Posterior [522, 674, 1444]. Postmarketing [572]. Postrandomization [363]. Poststratification [409]. Potential [480, 1060]. potentially [1011]. Power [71, 1541, 357, 1496, 35, 807, 618, 279, 248, 632, 141, 936, 733, 1425, 1463]. powered [1364]. Powerful [283, 396, 1464, 792]. pp [842]. Practical [380, 1406, 1302, 993, 88, 1304, 46, 834, 992]. Practice [1144, 1139]. pre [754]. pre-specified [754]. precision [1009, 1150, 1665, 1419]. preclinical [885]. predator [825]. predator-prey [825]. predict [1287, 831, 1652]. Predictable [366]. Predicting [422, 547, 920, 1398, 1448]. Prediction [350, 226, 186, 348, 288, 447, 204, 685, 88, 358, 235, 342, 1613, 1657, 1006, 1174, 761, 1378, 1164, 1492, 1498, 1136, 809, 1004, 1084, 708, 988, 1334, 1062, 1018, 1170, 174]. Prediction-Based [342]. Predictions [164, 334, 580, 1391]. Predictive [394, 522, 1063, 538, 75, 1404, 1140, 561, 1354, 133, 902, 11, 822, 1161, 814, 1182, 1060, 1024, 806, 863, 1082]. Predictiveness [504, 1006]. Predictor [63, 704]. Predictors [118, 12, 295, 62, 1537, 868, 1165, 1200, 977]. Preferences [1312]. Pregnancy [469]. Prenatal [340]. Prentice [1670]. Presence [319, 604, 574, 547, 599, 343, 238, 283, 133, 450, 134, 60, 216, 48, 38, 822, 1244, 1205, 699, 1646, 898, 1313, 991, 1292, 889, 979, 989, 724, 990, 694, 774, 1159, 786, 1247, 1091]. Presence-Absence [574]. Presence-Only [547]. Present [246]. Presidential [593]. Press [1667, 842, 1032, 1403, 1628, 1144, 1186, 1507, 1591, 1141, 1402, 1406, 1185, 1142, 1299, 1120, 1629, 1214, 1505, 1143, 1188, 1409, 1668, 1405, 1355, 1401, 1352, 1353, 1670, 1305, 1301, 1187, 1404, 1349, 1351, 1140, 912, 1506, 1594, 1075, 1407, 1595, 1503, 1669, 1119, 1122, 1300, 1630, 1350, 1298, 1213, 1183, 1033]. Press/Chapman [842]. Preterm [532]. Prevalence [47, 6, 29, 570, 648, 112, 1501]. Prevalent [456, 482, 1426, 850, 1610]. Prevention [687, 775]. Preventive [1142]. prey [825]. Primary [151, 1314]. Primer [167, 1590]. Principal [563, 301, 595, 349, 68, 663, 504, 854, 1251, 1313, 1057, 1498, 1232, 837, 1365]. Principle [356]. Principles [1505]. Prior [357, 489, 639, 1657, 1202, 1069]. Priors [613, 501, 459, 1044, 1425, 806]. Probabilistic [310, 678, 164, 192, 1116]. Probabilities [422, 970, 1206, 1325, 856, 1468]. Probability [655, 535, 148, 296, 402, 753, 842, 77, 1199, 1382, 1370, 1342, 931, 433]. Probability-enhanced [753]. probability-scale [1382]. Probands [29]. Probe [490]. Probit [366, 1023, 348, 1317]. Problem [149, 424, 1535, 763, 1484]. Problems [527, 425, 254, 131, 1252, 832, 1471]. Procedure [125, 314, 1584, 1585, 1060, 1341, 1056, 1314, 1644]. Procedures [453, 295, 1476, 827, 754]. Process [50, 30, 560, 221, 561, 587, 284, 1273, 934, 1424, 857, 930, 1044, 1105, 1042, 1117, 1379, 775, 1381]. Process-Based [560]. Processes [382, 615, 345, 265, 263, 669, 73, 49, 493, 97, 346, 269, 1211, 1284, 821, 1394, 1112, 1042, 880, 1609, 929, 1068, 1262, 1668]. product [930, 1530]. product-shot-noise [930]. Products [94]. Profile [225, 1332]. Profiles [548, 497, 1011, 1002]. Profiling [1343, 1490, 1486, 1488, 1487, 1489]. prognosis [746, 1419]. Prognostic [441, 11, 442, 1224, 1222, 1226, 1223, 1447, 1225, 923]. program [864]. Programming [624]. programs [1604]. Progression [362, 182, 582, 76, 1287, 1159, 1658, 1652, 901]. Progression-Free [362]. progressive [1468]. Project [687]. Projecting [150]. Projections [570]. Prone [181]. Propensity [108, 482, 478, 107, 1086, 586]. Properties [220, 611, 102, 276, 717, 1181]. proportion [1603]. Proportional [1235, 271, 364, 178, 1179, 891, 922, 1047, 788, 1472, 1048, 981]. Proposals [1142]. Proschan [127]. Prospective [136, 334, 451]. Prostate [515, 4, 580, 1174, 1208, 851, 1175, 1526]. Protein [348, 276, 1368, 1343, 738]. Proteomics [497]. Prünster [655]. PSA [1208]. Pseudo [220, 633, 685, 1470, 505, 1013, 763, 801]. Pseudo-Bayesian [505]. pseudo-DIC [801]. Pseudo-Gatekeeping [633]. Pseudo-Likelihood [220]. pseudo-observation [1013]. Pseudo-Observations [685]. pseudo-penalized [763]. Pseudolikelihood [365, 1275]. Pseudovalues [176]. psoriatic [901]. Psychologists [1504]. Pte [1504]. Public [570, 1592]. Publication [1416, 1414, 1413, 1417, 1415]. Publishing [1504, 1034]. pulsing [1394].

Q [1628, 1674]. Qian [256]. Qin [1561, 1562]. QTL [85, 811, 85]. Quadrat [306]. Qualification [466]. Quality [379, 495, 92, 1578, 80, 864]. Quantal [548, 366, 945, 702]. quantal-response [945]. Quantification [276]. Quantifying [822, 1662, 1416, 1414, 1413, 1417, 1415, 987]. Quantile [398, 399, 201, 262, 885, 1163, 603, 638, 965, 17, 1203, 1423, 966, 1272, 1003, 1400, 1164, 870, 1137, 1399]. Quantiles [481, 1436]. Quantitative [548, 30, 119, 45, 216, 542, 1481, 717, 1591]. Quantum [241]. quasi [1284, 1097, 763, 773]. quasi-likelihood [1284, 1097, 763, 773]. Questioning [554]. questionnaires [1250].

R [1032, 810, 329, 171, 1402, 1299, 129, 42, 328, 1305, 744, 953, 877, 174, 1304, 1122, 376, 1630, 327, 333, 623, 1216, 1502, 852, 1409, 953, 912, 1350, 1183, 1033, 41, 384, 85, 256, 373, 374, 90, 214, 257, 170, 1594, 1667, 213, 1669, 167, 83, 911]. R-vine [852]. R. [327]. R/qtl [85]. Radiation [32]. RAGHUNATHAN [1144]. Raiden [1547]. Rajaraman [1503]. Ramsay [171]. Random [140, 186, 544, 231, 137, 527, 16, 287, 232, 452, 250, 393, 204, 249, 228, 102, 1256, 1424, 920, 1134, 1138, 745, 1535, 1377, 821, 1110, 1178, 809, 1570, 1133, 1117, 1467, 1636, 1067, 897, 1266, 1517, 941, 1623]. random-effects [1134, 1570, 941]. random-partitions [1535]. Randomization [677, 1213, 676, 396, 864, 926, 1151, 925, 1315, 1507, 1139]. randomizations [1157]. Randomized [564, 993, 841, 588, 296, 589, 397, 157, 410, 107, 396, 864, 814, 1588, 961, 1584, 1009, 1651, 1182, 1047, 1060, 1495, 1090, 1341, 1288, 1447, 1087, 1475, 1156, 1665, 805, 1243, 1622, 554, 778]. Rank [26, 524, 561, 448, 1674, 1010, 1049, 967, 1478, 1580, 1382, 1065, 1025, 1375, 1234, 1318, 1608]. Rank-Based [524, 1674, 1608]. rank-sum [1010]. Ranking [78]. Ranks [328]. Rao [1094]. Rao-Blackwell [1094]. Rapid [488]. Rare [463, 220, 206, 974, 1118, 1212]. Rate [39, 559, 642, 79, 206, 356, 534, 645, 484, 962, 1190, 1030, 1137, 1611, 819]. Raters [546]. Rates [622, 1, 78, 1568, 898]. Ratio [272, 281, 444, 598, 618, 443, 1582, 1043, 909, 1197, 1536, 1090, 826, 1546, 1545, 982, 893, 927, 1583, 910]. Ratios [666, 464, 636, 932]. Raton [1032, 1403, 1628, 1144, 950, 1186, 1141, 1402, 1406, 1185, 1142, 1299, 1629, 1214, 1505, 1143, 1668, 1405, 1355, 993, 1631, 1401, 1352, 1353, 1670, 1305, 1301, 1187, 1349, 1140, 952, 953, 1506, 1594, 1075, 1667, 1407, 1595, 1669, 1122, 1300, 1630, 1350, 1213, 842, 1183, 1593, 811, 1033]. Rattray [432]. Rayid [1406]. Re [1557, 1556, 777, 1558]. re-estimation [777]. Re-Optimization [1557, 1556, 1558]. Reaction [1399, 839, 1073, 741, 1450, 991, 1182, 989, 990, 840, 875]. Reader [1073, 1450, 991, 1182, 989, 990, 840, 875, 1399, 839, 807]. readmission [1490, 1486, 1488, 1487, 1489]. Real [580, 717]. Real-Time [580, 717]. Reassessment [312]. rebound [739]. Recapture [82, 160, 307, 650, 453, 25, 200, 683, 114, 680, 682, 322, 261, 781, 1043, 970, 971, 1661, 1338, 1110, 1385, 1543, 895, 800, 728]. Receiver [75, 342]. Receiving [209]. reciprocal [1383]. reclassification [1072]. recommendation [768]. recommendations [767, 771, 770, 772]. reconstruction [1133, 1117]. records [867, 1537, 1519, 821]. Recovering [605, 1096]. Recovery [160]. Recurrence [580, 1433, 1175, 1318]. Recurrent [313, 583, 289, 265, 263, 264, 10, 602, 468, 603, 49, 314, 223, 1566, 1269, 804, 1092, 1396, 1063, 1020, 756, 979, 1433, 1070, 963, 923, 1654, 1434, 819]. recurrent-event [819]. Recurrent-Events [313]. recursive [1072]. Reddy [1301]. Rediscovery [972]. reduced [1025]. Reducing [35, 1193, 1641]. Reduction [144, 1165, 234, 341, 1535, 753, 1200, 1185]. Reference [545, 1176]. reference-test [1176]. Referenced [620]. refinement [1623]. Regime [67, 1650, 1175, 1428]. regimen [1672, 886, 1058]. regimens [1059]. Regimes [644, 153, 640, 517, 1491, 1466, 698, 840, 1649, 918, 1427]. Region [51, 243]. Regions [78, 714, 761, 1586, 789]. Registration [1516, 732]. Registry [76, 1190]. Regression [108, 237, 591, 398, 3, 654, 22, 429, 471, 524, 1299, 16, 164, 153, 264, 304, 578, 399, 609, 631, 588, 201, 262, 53, 603, 371, 257, 391, 541, 713, 523, 657, 63, 1439, 794, 103, 638, 589, 155, 228, 324, 616, 361, 372, 1074, 1091, 427, 337, 607, 686, 284, 15, 56, 17, 18, 57, 247, 178, 505, 251, 819, 567, 796, 710, 1574, 1011, 1043, 1324, 1179, 1481, 1476, 1426, 1482, 1423, 1027, 986, 1576, 1097, 898, 1158, 1114, 966, 1272, 1469, 1523, 1572, 854, 1197, 1345, 1202, 1092, 904, 1322, 967, 1325, 1041, 782, 1003, 1321, 1400, 1113, 855, 885, 1044, 1164]. regression [1418, 1529, 1580, 1498, 1512, 1069, 1370, 1237, 1372, 1296, 879, 1344, 869, 1474, 1263, 1395, 1107, 786, 1025, 818, 1431, 726, 1654, 965, 1630, 1266, 895, 783, 1231, 1647, 691, 1441, 652, 901, 1137, 1399, 1045, 1583, 1533, 1271, 1371, 1564, 653, 1188, 1074, 625, 89, 435]. Regressions [234, 1165, 1181, 1166]. Regret [153]. Regret-Regression [153]. Regularity [1379]. Regularization [609, 567, 1236, 1640]. Regularized [1442, 1518, 188, 887]. regulation [1100, 1277]. regulatory [862, 1276, 881]. Reilly [326]. Reinforcement [299]. Rejoinder [1559, 909, 1490, 916, 421, 550, 1362, 991, 772, 1400, 1563, 1226, 1417, 250, 1546, 990, 589, 849, 372, 415, 630, 479, 949]. Related [463, 890, 1606, 697, 1533]. Relatedness [119, 1240, 1342]. Relationship [125, 440, 813, 1208]. relationships [703]. Relative [317, 220, 400, 464, 1398, 1031, 1160, 927]. Relatives [29]. Relaxing [363]. Reliability [139, 1347, 725]. Removal [1514]. removing [1585]. renal [1339]. renewal [1379]. reoptimization [1559]. reparameterized [945]. repeat [1019]. Repeated [240, 353, 676, 944, 867, 1437, 730, 1331, 1099]. Replacement [503]. Replicated [494]. replied [128, 166, 210, 325, 126]. Reply [654, 514, 512]. Report [388, 558, 692, 843, 958, 1128, 1310, 1555, 5, 175]. reproducibility [1418]. reproducible [1130]. reproductive [1000]. repulsion [857]. Resampling [642, 986, 1616]. Resampling-Based [642, 986]. Research [434, 1301, 437, 1349, 1591, 1592]. Residual [878, 1162, 664, 1006, 1498, 982]. Residual-based [1162, 1006]. Residuals [104, 8, 892, 1382, 794]. resight [1363, 985, 1606]. resistant [802, 1113]. resolution [1420]. resonance [1605, 1627]. Resources [651]. respiratory [857]. respondent [1273, 838]. respondent-driven [1273, 838]. responders [1060]. Response [839, 335, 679, 554, 162, 202, 271, 52, 315, 606, 647, 640, 925, 337, 311, 861, 725, 874, 928, 945, 1114, 1250, 1197, 1198, 1101, 841, 1439, 1263, 1614, 1242, 1315, 1109, 716, 703]. Response-Adaptive [337, 925, 841, 1614, 1242, 1315]. Responses [506, 457, 246, 37, 154, 636, 109, 183, 1108]. resting [715]. resting-state [715]. Restoration [294]. Restricted [71, 266, 516, 1380, 1392, 1248, 980]. Restrictions [383, 273]. retest [1211]. Retrospective [924]. Return [680]. Review [1032, 211, 780, 1077, 1035, 1216, 1455, 1403, 1628, 379, 381, 380, 382, 433, 434, 435, 690, 591, 1453, 1502, 1549, 1144, 41, 1348, 1504, 329, 172, 1184, 171, 950, 1186, 1507, 1632, 911, 552, 252, 1591, 1451, 1141, 1139, 430, 384, 386, 471, 84, 85, 551, 1402, 330, 258, 1406, 1185, 554, 1142, 256, 1452, 373, 1299, 1120, 212, 1629, 331, 129, 42, 1303, 1214, 1505, 374, 215, 1454, 130, 1590, 1143, 1188, 328, 1409, 1668, 1405, 1355, 1215, 1302, 557, 1457, 993, 1145, 1631, 659, 720, 951, 992, 1401, 90, 1352, 1353, 1121, 556, 778, 718, 878]. Review [214, 43, 660, 1670, 91, 44, 1305, 661, 326, 169, 253, 1592, 1548, 257, 1187, 1404, 254, 45, 1349, 170, 385, 742, 1458, 553, 1550, 470, 472, 657, 719, 88, 779, 1140, 656, 744, 952, 953, 877, 912, 1506, 1594, 1075, 473, 1667, 387, 1407, 1595, 332, 1503, 132, 213, 377, 1034, 174, 173, 592, 1669, 1304, 1551, 87, 1119, 658, 86, 625, 167, 1456, 255, 1122, 46, 376, 1074, 131, 1300, 1630, 375, 1350, 1298, 1076, 1213, 555, 624, 842, 83, 1183, 168, 378, 327, 259, 432, 431, 743, 1593, 811, 89, 333, 1408, 590, 623, 1033, 740]. Reviews [1301]. Revisit [618]. Rexstad [1121]. Rhee [1488, 1487, 1489]. Richard [1303, 1140, 1404]. Richness [503, 306, 138, 765]. ridge [777]. Right [666, 180, 293, 54, 601, 1566, 1318, 982]. Right-Censored [666, 293, 54, 601, 1318, 982]. right-censoring [1566]. Rights [683]. Risk [689, 70, 317, 350, 1141, 548, 584, 1378, 462, 225, 181, 124, 531, 617, 539, 358, 134, 668, 565, 19, 38, 442, 1481, 823, 1398, 776, 1224, 882, 838, 808, 1542, 1222, 1226, 939, 1223, 694, 1225, 1004, 1084, 1334, 1526, 927, 1170, 1629]. Risk-Adjusted [565]. risk-based [1224, 1222, 1226, 1223, 1225]. Risk-Group-Specific [70]. Risks [220, 441, 416, 421, 224, 951, 262, 176, 685, 133, 97, 251, 1256, 822, 757, 1205, 1339, 890, 891, 1138, 1051, 693, 1292, 1497, 1237, 869, 1050]. RNA [1132, 1636, 632, 389, 738, 975]. RNA-Seq [632, 975, 1132, 389]. Robert [1184, 1355, 1401, 213]. Robinson [720]. Robotic [275]. Robust [796, 313, 665, 226, 286, 335, 392, 574, 527, 1325, 238, 75, 541, 685, 1084, 235, 806, 536, 236, 1644, 141, 516, 396, 517, 1045, 542, 1672, 1465, 1520, 1476, 1154, 725, 1239, 1521, 1158, 1572, 723, 793, 1585, 1311, 1190, 1267, 1429, 1369, 1539, 840, 886, 1477, 1185]. Robustification [637]. Robustifying [1195]. robustness [1058]. ROC [75, 714, 1575, 889, 145, 343, 568, 238, 537, 84]. Rodrigo [1628]. Roland [1594]. Role [593, 21]. Rolf [1033]. Romain [330]. Ron [1406]. Ronald [721, 951]. Rory [1504]. Rose [592]. Rosenbaum [129]. Rosenberger [1139]. Ross [1670]. Rotation [799]. Rotation-based [799]. Routledge [1592]. Roychoudhury [1631]. Rubak [1033]. Rubba [1304]. Rubin [1119]. rule [925, 1315]. Rules [1312, 143, 1430, 1589, 754, 1587, 1374, 1389, 1152]. Ruth [1629].

S [622, 329, 258, 1406, 256, 1488, 1487, 1489, 1121, 214, 661, 1592, 257, 170, 553, 375, 842, 168, 431, 743, 1593]. S. [381, 171, 85, 256, 592, 87, 743, 333]. S.-C. [87]. Sabanés [654]. Saddlepoint [73, 1254, 1599]. SAEM [1208]. Safety [1403, 585, 43, 639, 816, 908]. Salemi [46]. Salmaso [41]. Salsburg [1593]. Sample [1390, 198, 366, 550, 549, 138, 32, 940, 1495, 618, 169, 279, 248, 157, 218, 1444, 1541, 827, 1450, 1496, 754, 906, 1181, 1375, 1247, 1392, 1477, 1297, 1349, 431]. Sample-Size [279]. Samples [607, 781, 761, 1005, 1367]. Sampling [665, 503, 22, 1578, 306, 455, 40, 1121, 165, 159, 150, 651, 52, 315, 606, 406, 340, 196, 1273, 1171, 850, 838, 1313, 1189, 697, 1333, 983, 1316, 1468]. Sandeep [1407]. Sandwich [188]. Sanguinetti [432]. Sargent [417, 418]. SAS [380, 434, 212, 1305, 331, 376]. SAS/IML [212]. Satrajit [1631]. Scalable [1582, 1485, 1640, 1420]. scalar [917, 854]. Scale [219, 496, 671, 1398, 1584, 1030, 1382, 1527, 999, 1247, 1261]. Scan [407, 187, 1103, 939]. Scans [1408]. Schall [1261]. Schedule [640]. Schedule-Dose [640]. schedules [1526, 1333]. Scheme [644, 57]. Schemper [235]. Schemper-Henderson [235]. Schmid [1348]. Schoenberg [782]. Schwarz [1028]. Science [1184, 1402, 1406, 1669]. Sciences [1119]. Scientific [1504, 1034]. Sclerosis [194, 582]. Score [70, 98, 164, 7, 200, 670, 509, 107, 790, 586, 567, 882, 1115, 1094, 808, 923, 1470, 1086, 975, 1564]. score-based [975]. Scores [108, 482, 478, 1326]. scoring [1274, 1334]. Scraping [1304]. Screening [689, 219, 1364, 584, 426, 1166, 1095, 1341, 1527, 1335, 716, 790]. screens [1066]. Scutari [1350]. Sea [369]. Seals [508]. Search [613, 31]. seasonal [1111]. Seber [114]. Second [92, 1284, 265, 646, 124, 842]. Second-Generation [92]. Second-Line [646]. Second-Order [265, 124, 1284]. Secondary [151, 1439, 1275, 1314]. Sectional [583, 395, 112, 868, 940, 1626, 1468]. Segmentation [93, 1054]. SEIR [569]. select [1333]. Selected [349, 363]. selecting [947]. Selection [689, 610, 140, 359, 1561, 522, 79, 16, 344, 1236, 105, 480, 232, 189, 600, 233, 200, 14, 387, 566, 103, 521, 390, 297, 342, 901, 445, 191, 62, 1673, 896, 1485, 862, 1172, 1620, 986, 1012, 1001, 1096, 1576, 1097, 817, 1537, 1359, 1362, 830, 1054, 793, 1479, 1046, 752, 1360, 926, 1311, 1386, 1201, 1562, 791, 1560, 1563, 1024, 747, 1643, 881, 708, 1615, 1361, 947, 1082, 1431, 1265, 704, 773, 829, 1649, 888, 700, 789, 731, 1130, 825]. selection-assisted [1576]. selection-based [1054]. Self [21, 671, 572]. Self-Controlled [671, 572]. Semi [101, 1519, 1028, 100, 1493, 1522, 1292, 1341, 1093, 1331, 1050]. semi-competing [1292, 1093, 1050]. Semi-Markov [101, 1028, 100]. Semi-parametric [1493, 1522, 1341, 1331]. Semi-supervised [1519]. Semicompeting [416, 421, 262, 97, 758]. semicontinuous [969]. Semiparametric [564, 439, 1574, 22, 272, 498, 1432, 529, 71, 666, 1646, 1270, 1469, 1523, 110, 265, 461, 922, 30, 10, 506, 584, 271, 577, 13, 530, 758, 343, 795, 1317, 61, 576, 1237, 711, 14, 533, 1159, 132, 616, 1019, 60, 695, 726, 147, 65, 364, 729, 56, 446, 311, 346, 18, 59, 291, 516, 223, 11, 1279, 853, 1575, 759, 850, 757, 696, 913, 916, 1169, 1023, 1573, 1041, 1003, 1400, 1492, 914, 1448, 1157, 786, 1431, 915, 1645, 1644, 1602, 1399, 1022, 983, 1371]. Sen [85, 169]. sense [1663]. Sensitivity [1547, 383, 161, 359, 146, 1294, 662, 667, 354, 571, 363, 72, 273, 1463, 869, 1462, 976, 1155, 1331, 1618, 1617]. Senturk [1488, 1487, 1489]. September [690, 810, 1672, 383, 1675, 474]. Seq [632, 975, 1132, 389, 192]. Sequence [1239, 550, 549, 1529]. Sequences [447]. Sequencing [92, 278, 277, 1076, 1276, 974, 1636, 1532, 1613]. Sequential [559, 950, 864, 1054, 733, 1151, 136, 151, 1524, 1445, 1446, 1438, 908, 1314, 1061, 1571, 1623]. Serial [275, 808]. Series [374, 214, 674, 617, 494, 497, 1594, 1595, 671, 572, 1336, 1001, 852, 1206, 1394, 1605, 1593]. seriously [1517]. Seroadaptation [581]. Serological [229]. Service [51]. Set [634, 883, 698, 181, 752, 1003, 1400, 1029, 1275, 755, 1399]. Set-based [883, 1029]. Set-valued [698]. sets [832, 1109, 1602, 1116]. Setting [19, 340, 1339]. severity [1293]. Sexual [245, 581]. Shanshan [1670]. Shao [779]. Shape [436, 1261]. shapes [1103]. Shared [503, 1059, 1617]. shared-parameter [1059]. Sharpening [663]. Sheather [257]. Shein [1591, 780]. Shein-Chung [1591, 780]. Shen [1673]. shift [929]. shifts [894]. Shigeyuki [1140]. Shih [1505]. Shingo [621]. Shipley [1409]. Short [666, 158, 1343, 1571]. Short-Term [666, 158, 1571]. Shostak [434]. shot [930]. Shrinkage [635, 300, 61, 505]. Shuangge [718]. Sided [302]. Sieve [1089, 1440]. Signal [39, 928, 1624, 999]. signaling [1376]. Signals [118]. Significance [350, 1233, 736, 1015, 1366]. Signs [328]. Simes [1660]. similarity [1102]. Simon [1404, 1140, 1304, 1213]. Simple [125, 734, 453, 1318, 1289]. Simplified [146]. Simulation [1283, 1643]. Simulation-based [1283]. Simulation-selection-extrapolation [1643]. Simultaneous [830, 1088, 183, 696, 625]. Simultaneously [581]. Sin [778]. Sin-Ho [778]. Singapore [1504, 1034]. Singer [169]. Single [263, 36, 1108, 1389, 1567, 1426, 1636, 1443, 1602]. single-arm [1443]. single-cell [1636]. Single-index [1108, 1567, 1426, 1602]. single-index/Cox [1567]. Single-Neuron [36]. Singular [142, 217]. Sir [721]. Site [574, 51]. Site-Edge [51]. sites [1532, 738]. Size [307, 22, 198, 366, 618, 649, 279, 1349, 560, 248, 157, 55, 1390, 1011, 1541, 838, 1338, 1496, 940, 754, 1495, 985, 1105, 1619, 740, 1431, 1477]. Size-Biased [22]. Sizes [336, 32, 279, 431, 1444, 1384]. Skaug [621]. skeleton [1515]. skeletons [973]. Skew [203, 1227]. skew- [1227]. Skew-Normal [203]. skewed [1016, 1574, 1114, 713]. Skewness [506]. Slopes [231]. Small [1547, 366, 138, 761, 1450, 1535, 1330, 1346, 1297]. Small-Sample [138, 1450, 1297]. SMART [864]. smartphone [1662]. smartphone-based [1662]. SMC [1001]. Smith [1628, 382]. Smoking [288, 1386, 963]. Smoothing [591, 307, 289, 339, 437, 1577, 1576, 735, 1261, 551]. smoothing-based [1577]. SNP [428, 1275, 93]. SNPs [488]. SNVs [1132]. Social [1406, 1119]. Society [1221, 721]. Software [1595, 168]. Sojourn [100]. Sokolowski [1301]. Solari [41]. solution [834]. Solutions [258]. somatic [1328]. Some [404]. Somnath [1076]. Song [1349]. Sonia [1595]. Sons [1590, 1139]. Sotoca [1595]. Soumi [1631]. Source [705, 1446]. Sources [33, 1174]. Southwestern [539]. Space [495, 35, 230, 465, 1595, 187, 714, 801, 1227, 894, 1228]. Space-Time [495, 465, 187, 1227, 894]. Sparse [142, 1484, 818, 1048, 1582, 1657, 798, 696, 1052, 1338, 967, 855, 1580, 1600, 1340, 1166]. Sparsely [605, 1589]. Sparsity [1401, 1096, 1479]. Spatial [591, 186, 615, 347, 252, 104, 121, 1363, 368, 309, 407, 345, 328, 1041, 158, 851, 1103, 1187, 688, 1344, 49, 221, 842, 829, 532, 346, 764, 1284, 852, 1097, 1661, 1229, 732, 763, 1133, 797, 988, 1644, 736, 1647, 1033, 332]. spatial-clustered [764]. Spatial-Temporal [532]. Spatially [620, 206, 455, 103, 561, 436, 222, 802]. Spatio [552, 50, 1420, 1042, 660, 1667, 623]. Spatio-Temporal [552, 50, 1420, 1042, 660, 623, 1667]. Spatiotemporal [485, 308, 706, 880, 1645]. Spearman [1382]. Special [555]. Species [503, 547, 206, 306, 138, 305, 587, 765, 1031]. Specific [583, 665, 422, 70, 180, 598, 225, 343, 569, 682, 295, 597, 1206, 889, 921, 899]. Specified [588, 292, 589, 754]. spectral [1336, 797, 1387]. Spectrogram [118]. spectrometry [735]. Speed [201]. Spline [3, 338, 289, 484, 437]. Splines [307, 551, 339, 1391, 782, 871]. Split [275]. Split-Plot [275]. Spread [485]. Springer [1184, 992, 878, 1216, 1348, 1632, 1303, 1121, 1354, 1074, 1076, 1408]. Springer-Verlag [1632]. square [1234]. Squared [58]. Squares [1165, 601, 1672, 1600, 886]. Stable [177, 1609]. Stage [125, 113, 687, 196, 1397, 1241, 1339, 1625, 1282, 1466, 730, 707, 1443, 1192]. Stage-Structured [113]. Stagewise [1286]. stand [1042]. Standard [545, 111, 568, 460, 546, 873]. Standardized [197, 1160]. Start [486]. State [230, 865, 801, 715, 1621, 931, 728, 899, 1468, 1405, 1595]. State-Space [230, 801]. states [1174, 951]. Stationarity [615]. Statistic [407, 354, 187, 396, 860, 826, 976, 1062]. Statistical [381, 434, 1423, 993, 720, 1233, 556, 1305, 385, 387, 1407, 54, 174, 87, 255, 376, 389, 80, 1298, 72, 83, 351, 149, 743, 97, 276, 333, 907, 196, 1476, 1274, 1132, 752, 1670, 1422, 1612, 792, 1303, 1505, 1401, 1352, 878, 1300, 1076, 811, 1632, 1631, 91, 779, 658, 378]. Statisticians [377]. Statistics [1216, 256, 373, 263, 534, 332, 606, 1119, 842, 1590, 1103, 1327, 939, 1667, 1504, 1506, 1122, 252, 169, 253, 131, 552, 326, 590]. Status [440, 364, 268, 57, 1291, 1317, 1472, 729, 1470]. STEC [803]. Stefano [655]. Stem [611, 424]. stents [1539]. Step [74, 973, 801, 983]. Stephen [1302, 1213]. stepped [1525, 1495]. Stereological [23]. Stereotype [237]. Stevens [167]. Steyerberg [88]. Stochastic [41, 679, 99, 33, 1351, 458, 285, 1000, 1575, 1668, 1034, 382]. stopover [985]. Strain [583, 648]. Strain-Specific [583]. Stram [1408]. Strata [409]. Strategies [87, 1074, 410, 299, 1446, 747]. Strategy [1632, 362, 208, 611, 1069]. Stratification [349, 68]. Stratified [665, 307, 723, 538, 52, 606, 607, 251, 1004, 1084]. Strength [459, 1161]. Streptococcus [648]. stroke [708]. Struchiner [211]. Structural [20, 845, 1642, 461, 846, 1409, 1215, 847, 1369, 566, 1520, 959, 1268, 848, 844, 849, 773]. Structure [360, 447, 600, 576, 132, 1329, 1131, 1437, 1201, 855, 1133, 1068]. Structured [113, 119, 1165, 837, 342, 1485]. structures [1089, 1637]. Stuart [515]. Studies [608, 379, 810, 665, 439, 226, 395, 679, 110, 678, 162, 403, 662, 189, 35, 158, 556, 53, 154, 120, 14, 670, 173, 354, 571, 612, 318, 52, 156, 448, 236, 218, 2, 607, 451, 17, 178, 516, 135, 27, 892, 859, 734, 1541, 702, 739, 883, 815, 1198, 904, 1376, 1445, 1494, 1464, 1060, 1088, 974, 1193, 873, 1327, 1253, 1610, 939, 1112, 1527, 1570, 872, 1004, 1207, 976, 1155, 1025, 1331, 943, 924, 1618, 1275, 941, 921, 1099, 1463, 700, 875, 1333, 785, 1255, 1086, 1664, 1238, 942, 977, 211, 129, 333]. Study [237, 583, 319, 394, 198, 201, 136, 1207, 450, 315, 66, 467, 19, 424, 581, 1000, 1173, 1544, 817, 802, 1003, 1400, 1386, 1113, 1464, 885, 1193, 1638, 1439, 708, 1334, 1658, 1618, 709, 921, 1399, 968, 856, 435]. study-specific [921]. studying [816]. Subclassification [237, 1086]. Subcluster [336]. Subcluster-Level [336]. Subgroup [564, 1274, 1150, 1341, 887]. subgroups [1447, 1083]. Subject [547, 678, 225, 295, 19, 266, 1566, 1011, 1101, 713, 1087, 1333]. subject-level [1101]. Subject-Specific [225, 295]. subjects [1010, 1603]. Subpopulations [459]. subsample [900]. subsampled [1206]. Subsampling [986, 345, 66, 1431]. Subset [363]. subspace [1641]. Substance [205]. Substitution [531]. Substructure [27]. Subtype [1419]. subtypes [746]. successive [1205]. Sufficient [1535, 1165, 234, 341, 1200, 753]. Sum [448, 1010, 817]. Summary [606, 1327, 1638]. sums [892]. sup [1115]. sup-score [1115]. Superior [354]. superiority [804]. Superpopulation [261, 985]. supervised [1519]. Suppress [369]. Suppression [48, 739]. Surface [23, 1109]. surfaces [1575, 1530]. Surgical [467]. Surrogacy [416, 421, 68]. Surrogate [422, 627, 301, 595, 628, 629, 423, 626, 466, 149, 504, 813, 1040, 1519, 1313, 1653, 1475]. Surrogates [627, 628, 629, 626, 1085]. Surveillance [464, 572, 1174, 711, 908, 1526]. survey [1114, 906]. Surveys [643, 309, 933, 934, 940, 1543, 554]. Survival [665, 422, 226, 1161, 482, 529, 362, 224, 669, 1405, 1095, 181, 530, 531, 582, 296, 470, 400, 292, 8, 96, 1074, 510, 9, 59, 178, 222, 507, 1391, 1179, 1426, 759, 1319, 1541, 1208, 1224, 1191, 1049, 1466, 1180, 830, 1364, 1524, 1393, 1023, 1473, 1190, 1380, 851, 1021, 1163, 1222, 1226, 758, 1088, 1178, 1268, 1223, 711, 1225, 1160, 1066, 1019, 1655, 1392, 1062, 1571, 1428, 1471, 900, 980, 1022, 1271, 899, 380, 718]. survivor [1269, 817]. survivors [1115]. Susceptibility [515, 4, 231, 1172]. Sverdlov [993]. swimmer [1106]. Swiss [394, 1207]. Switch [646]. Switching [101, 646]. Symmetry [397]. Symptom [76]. syncytial [857]. Syndrome [319, 804, 1205]. Synergy [162]. system [1334]. Systematic [309]. Systems [230, 432, 321, 946, 700, 1034].

T [329, 1184, 384, 1121, 254, 45, 174, 1227]. T. [591, 387, 378, 590]. Taber [1142]. Table [474, 510]. tables [906, 1527]. tail [898]. Talloen [173]. Tamhane [131]. Tan [254, 1034]. Tang [1674]. Tango [378]. target [862]. Targeted [585, 831, 707, 592]. targets [902]. task [1627]. task-induced [1627]. taxonomic [1280]. Taylor [417, 949, 418, 948]. TB [802, 1393]. Techniques [1032, 554]. Tell [1593]. temperature [852]. Temporal [552, 121, 50, 368, 508, 496, 440, 532, 1011, 1420, 1229, 1042, 660, 623, 1667]. temporary [1514]. tensor [1530]. Term [666, 202, 158, 1115, 288, 1571]. Terminal [602, 223, 1063, 979]. terms [762]. Test [608, 161, 58, 725, 545, 198, 98, 282, 584, 7, 453, 974, 248, 616, 632, 448, 72, 396, 1211, 1176, 701, 1577, 1674, 1239, 1010, 1049, 1450, 882, 1115, 1536, 1007, 807, 1029, 1438, 1638, 1155, 1375, 1234, 1247, 1314, 875, 1602, 1297, 1608, 790, 1015, 1583]. test-retest [1211]. Testing [861, 47, 6, 1016, 350, 679, 281, 287, 573, 1214, 345, 428, 1180, 111, 35, 1003, 489, 197, 457, 906, 618, 788, 491, 617, 339, 195, 488, 131, 1388, 151, 687, 828, 112, 65, 528, 575, 1100, 1538, 1531, 1283, 1584, 762, 1513, 1168, 745, 1177, 991, 1585, 1252, 1446, 1400, 989, 990, 1296, 1653, 1004, 1056, 760, 897, 799, 1167, 1061, 1530, 1230, 1399, 792, 1583, 1366, 46]. Tests [689, 41, 404, 614, 615, 550, 549, 239, 588, 452, 119, 323, 305, 124, 75, 283, 576, 200, 670, 81, 589, 677, 397, 510, 509, 302, 152, 9, 27, 1584, 1197, 883, 1648, 1660, 1327, 1065, 894, 1330, 1392]. Text [1304]. Textual [1669]. Thall [1557, 1558]. theorem [1094]. theoretic [1040]. Theory [41, 634, 1139, 1595, 1250, 258, 742]. Therapeutic [106, 1446]. Therapies [462]. Therapy [69, 1060]. there [641]. Thomas [953]. Thorsten [1187]. threatened [938]. Three [157, 1066]. three-groups [1066]. Three-Level [157]. Threshold [1474, 1615, 427, 19, 1158, 1227]. thresholding [777, 705]. Throughput [449, 1418, 1066, 1530, 716]. Tian [254]. Tibshirani [1401, 174]. Tie [1214]. Tie-Hua [1214]. Tiling [80]. Time [125, 20, 564, 422, 495, 1348, 307, 161, 290, 526, 1490, 1486, 1206, 416, 421, 1488, 599, 374, 1487, 10, 7, 1355, 646, 1489, 148, 635, 214, 225, 674, 181, 177, 176, 486, 603, 165, 582, 469, 465, 617, 494, 497, 1594, 569, 1595, 334, 133, 363, 86, 668, 525, 923, 1518, 187, 580, 295, 55, 445, 56, 446, 601, 223, 11, 1391, 781, 820, 822, 1336, 1269, 1565, 961, 1001, 1449, 852, 776, 1270, 1224, 717, 1393, 1589, 1051, 1442, 1573, 866, 1189, 1013, 1020, 1292, 1497, 1380, 889, 1222, 1226, 1670, 1394, 1106, 1433, 1268, 1136, 809, 1369, 1223, 1227]. time [1277, 694, 998, 1225, 963, 1159, 760, 894, 1385, 1228, 863, 1443, 1019, 1318, 1621, 1517, 1392, 1091, 1343, 731, 1255, 1428, 787, 980, 1238, 1605, 938, 1316, 1611]. time-course [1343]. Time-Dependent [599, 10, 635, 133, 11, 923, 889, 1106, 694, 863, 1238]. Time-dynamic [1490, 1486, 1488, 1487, 1489]. Time-Series [674, 497]. Time-Stratified [307]. time-to-drop-out [1517]. Time-to-Event [125, 1348, 1355, 176, 334, 363, 86, 1518, 822, 1449, 1013, 1268, 1136, 1159, 1443]. Time-to-Pregnancy [469]. Time-Varying [20, 161, 526, 181, 617, 525, 445, 223, 1391, 1224, 1222, 1226, 1223, 1277, 1225, 963, 760, 1611]. time-zero [809]. Times [526, 646, 680, 229, 640, 1566, 1432, 739, 1092, 821, 694, 901]. TOC [954, 994, 1036, 1078, 1124, 1146, 1217, 1257, 1307, 1356, 1410, 1459, 1509, 1552, 1596, 1633]. Tolerance [1017]. tool [1006]. toolkit [1424]. Tools [1406]. Toshiya [621]. Toxicant [369]. Toxicities [123]. Toxicity [70, 69, 640, 1480, 1530, 716]. Tracing [406]. training [1061]. Trait [28, 30, 119, 216]. Trait-Model-Free [28]. traits [1638]. Trajectories [560]. Trajectory [38]. trans [1067]. trans-ethnic [1067]. Transcription [280]. transcriptomic [1340]. Transect [621, 643, 24, 159, 454, 934]. Transects [260]. transform [1346, 1627]. Transformation [594, 529, 688, 223, 759, 850, 758, 695, 729]. Transforming [274]. transfusion [804]. Transition [184, 582, 931, 856, 1468]. Translational [87]. Transmission [161, 648, 540, 832]. transplant [1020]. Transplantation [664, 1491, 1019]. transposable [828]. Treat [462]. Treatment [108, 665, 209, 422, 67, 1312, 950, 644, 917, 153, 480, 300, 667, 486, 296, 502, 179, 645, 517, 597, 1672, 1491, 1274, 1430, 864, 814, 1398, 1191, 1466, 1393, 1589, 793, 1651, 1650, 1445, 767, 772, 1182, 768, 698, 1047, 1088, 1429, 1373, 959, 1249, 1024, 1438, 747, 1587, 1604, 1653, 1087, 1157, 1374, 976, 1083, 1082, 1175, 1389, 923, 1192, 840, 886, 1058, 1649, 805, 887, 1428, 1477, 771, 918, 960, 1008, 1427, 770, 1152]. Treatment-Selection [480]. Treatment-Specific [665, 597]. treatments [1430, 1373, 1539, 1086]. Tree [101, 137, 15, 222, 1280, 1337, 1231, 1152]. Tree-based [15]. Tree-Fungus [137]. trees [766]. Trend [27, 1365]. Trends [150, 937]. Trevor [1298, 1401]. Trial [564, 242, 386, 992, 1557, 123, 639, 423, 1059, 1556, 1559, 1674, 864, 1588, 1085, 1524, 1651, 1063, 754, 1341, 1288, 1475, 1608, 1558]. trial-eligible [1588]. Trials [422, 106, 370, 1139, 500, 1214, 1505, 357, 69, 595, 993, 588, 68, 288, 300, 1404, 541, 122, 296, 1140, 589, 397, 86, 157, 410, 149, 107, 676, 575, 396, 597, 299, 1444, 804, 814, 865, 1588, 961, 1624, 1009, 858, 1282, 1180, 1313, 1496, 866, 841, 1525, 1182, 1047, 1495, 1090, 1151, 1447, 1087, 1614, 806, 1443, 876, 1242, 925, 1315, 1156, 1665, 805, 1243, 1477, 775, 960, 1008, 1622, 1403, 379, 778, 43, 375, 431]. Trindade [1595]. Trio [488, 1612]. TRIVELLORE [1144]. Trosset [83]. trouble [1346]. Truncated [304, 1660, 262, 1043, 1272, 1370, 1395, 982, 1320]. Truncation [180, 1565, 1471]. Tseng [1120]. tuberculosis [802]. Tuberculous [545]. Tumor [1204, 65, 1063, 1343]. tumors [1342]. Turing [916, 914, 915, 765, 972, 913]. Turing-computerization [916, 914, 915, 913]. Turner [1033]. Tutz [1348, 1188]. Tweedie [1261]. Twin [351, 1055, 941].

Twins [515, 4, 194]. Two [125, 287, 550, 42, 549, 618, 491, 205, 358, 543, 612, 687, 509, 74, 484, 575, 196, 1397, 1241, 1000, 827, 1625, 973, 1466, 1102, 860, 1084, 1443, 1392, 982, 893, 983, 1116]. Two-Arm [575]. Two-Component [509, 484]. Two-Dimensional [491]. two-group [860]. Two-Level [42]. Two-Part [205]. Two-Phase [612, 1084]. Two-Sample [550, 549, 827]. Two-Stage [125, 687, 196, 1397, 1241, 1625, 1466, 1443]. Two-Step [74, 973, 983]. Two-Way [287, 543]. Type [462, 250, 293, 100, 9, 853, 1565, 1049, 1433, 1425, 249]. Tyregod [430].

U.S. [317]. UK [1590]. Ullman [1503]. Ultrahigh [731, 1364, 1095, 1639]. ultrahigh-dimensional [1095, 1639]. Umbach [810]. Unbiased [1061, 216, 1112]. uncertain [900]. Uncertainty [608, 478, 475, 479, 1541, 928, 1247, 1543, 888, 875, 1212]. Unconditional [323]. unconstrained [1520]. Uncovering [25, 76]. undercoverage [933]. Underreporting [364]. Understanding [743]. Unequal [367, 269]. ungrouped [1512]. unification [873]. Unified [139, 1029, 999, 1022]. Uniform [501]. Unimodal [782]. United [1409]. units [1165, 1604]. univariate [1317]. University [1120, 1188, 1409, 1404, 1503, 1119, 1298, 1351]. Unobserved [146, 1572]. unordered [1539]. unrelated [1326]. Unspecified [528]. Untreated [596]. unwanted [1585]. Updating [88]. Upper [369]. Uri [1184]. Usage [243]. Use [543, 367, 569]. Used [101]. useful [1524]. User [1409, 624]. Using [125, 380, 621, 810, 82, 559, 404, 307, 643, 892, 644, 370, 482, 104, 522, 137, 1208, 384, 99, 199, 26, 407, 573, 563, 345, 182, 374, 33, 118, 69, 29, 588, 316, 68, 288, 300, 176, 238, 75, 283, 200, 639, 541, 454, 352, 1653, 1594, 150, 1005, 173, 589, 229, 406, 389, 580, 640, 1213, 636, 188, 2, 728, 427, 9, 451, 1471, 94, 18, 141, 266, 918, 308, 542, 108, 1390, 1391, 861, 1502, 1053, 850, 615, 1205, 865, 741, 735, 1440, 702, 1368, 717, 1191, 1197, 777, 838, 745, 579, 480, 1094, 1023, 1376, 1651, 1650, 1338]. using [967, 1013, 782, 1003, 1400, 871, 832, 530, 732, 1327, 1069, 985, 1105, 1268, 1638, 1341, 1254, 749, 1384, 794, 708, 682, 943, 1636, 1175, 1618, 1617, 1335, 941, 1609, 925, 1315, 902, 1343, 1018, 1262, 19, 900, 1399, 1599, 1170, 975]. Utero [451]. Utilities [640]. Utility [69, 1447]. Utility-Based [69, 1447]. Utilizing [102, 1524].

V [1488, 1487, 1489, 89]. V. [214, 658]. vaccination [774, 816]. Vaccine [211, 595, 598, 149, 858, 1313, 1070, 908, 755]. Vaccines [583]. Valid [588, 473, 589, 1665]. Validated [423]. Validation [88, 813, 394, 1519, 1618, 755, 900, 1564]. Validity [677, 397]. Value [1141, 142, 217, 538, 194, 942]. valued [1659, 698, 1581]. Values [559, 272, 238, 295, 11, 1030, 724, 1330, 1248]. Vance [1507]. Vandamme [46]. variability [1431, 1333]. Variable [245, 610, 140, 1561, 1097, 16, 344, 105, 189, 233, 14, 708, 103, 450, 1265, 390, 342, 1574, 896, 1522, 1073, 1485, 1172, 1096, 817, 830, 1479, 1573, 1013, 1386, 1201, 1562, 791, 824, 1560, 1563, 1439, 1367, 1209, 1266, 829, 1463, 1130, 790]. Variables [613, 137, 637, 238, 401, 859, 1601, 969, 1268, 1643, 749, 704, 1286]. Variance [313, 116, 309, 345, 263, 1648, 670, 528, 1540, 751, 1496, 1247, 1290]. Variances [347, 868]. Variants [220, 1213, 974, 939]. Variation [292, 267, 1054, 1585, 1210]. variation-testing [1585]. variational [706]. Variations [368, 1482]. Varying [20, 161, 526, 1492, 600, 181, 617, 103, 525, 445, 223, 1391, 776, 1224, 1222, 1226, 1108, 1223, 1277, 1225, 1157, 963, 760, 1246, 1611]. Varying-coefficient [1492, 1157, 1611]. Vector [606]. vectors [1536]. Verbeke [130]. Verification [145, 343]. Verlag [1632]. Versus [115, 986, 908]. Verzani [624]. Via [536, 1279, 1672, 1211, 1444, 398, 765, 945, 972, 817, 1359, 1362, 919, 793, 1442, 1479, 1585, 1535, 1252, 868, 1360, 142, 1478, 341, 1361, 1340, 897, 668, 1048, 390, 1392, 886, 1058, 932, 789, 437, 785, 1130, 792]. View [513, 514, 1642, 1295, 1580, 1042]. vine [1566, 852, 1437]. Vining [657]. Violations [683, 1180]. Viral [739, 316, 48, 707]. virus [858, 857, 1626]. visit [1112]. visual [1006]. Visualization [1032]. Visualizations [1254]. Vitro [162]. Volume [555, 89]. vQTL [1169].

W [382, 1674, 84, 85, 374, 1302, 557, 660, 44, 132]. W. [42, 254, 88, 173, 625, 83]. Wai [1034]. Wai-Yuan [1034]. Wainwright [1401]. wait [1339]. wait-listed [1339]. Walker [434]. Wallace [1672]. Wallis [608, 198, 875]. Walter [1594]. Wang [1673, 383, 551, 1630, 691]. Ware [1077]. Warped [751]. Wavelet [562, 578, 1627]. Wavelet-Based [562, 578]. Waves [317]. Way [287, 543, 1346]. weak [1131]. Weaker [518]. wearable [1397]. Web [1304]. wedge [1525, 1495]. Wei [691]. Weight [427]. Weighted [104, 336, 148, 179, 1275, 601, 1672, 750, 882, 1650, 1370, 1192, 886, 887, 1152]. Weighting [402, 1202, 1175]. Weights [9, 983]. Weili [992]. Weinberg [810, 34]. Weisberg [473]. Westfall [384, 743]. While [632]. who [1651]. whole [1102, 1005]. Wickham [259]. Wide [1408, 1327, 1253, 1638]. Wiechung [1505]. Wiegand [1187]. Wienke [470]. Wikle [552, 1667]. Wilcoxon [248, 448]. wild [1435]. Wiley [1139, 1590, 1301, 877, 1077, 1035, 1502, 1215, 1302, 1145, 1304]. Wilks [280]. William [1139]. win [826, 1546, 1545]. windows [1571, 532]. Wishart [1262]. within [1010, 937, 123, 1431, 466]. within-cluster [1431]. Without [546, 383, 1540, 503, 1426, 194, 1325, 873, 273, 1665, 1471]. Women [1122]. Woncheol [690]. World [1504, 1034]. worst [1547, 1294]. Wrapper [642]. Writing [1142].

X. [661, 332]. Xianghon [1120]. Xiaofeng [1630]. Xiaonan [1561, 1562]. Xue [1673]. XXVI [593].

Y. [551, 330]. Yan [1141]. Yang [1561, 1562]. Yanjun [1488, 1487, 1489]. Year [226, 1271]. Yi [1632]. Yichen [1561, 1562]. Yin [386, 1300]. Ying [771]. York [1035, 1184, 1632, 1120, 1215, 951, 1592, 1119, 992, 878, 1351, 1354]. Youden [714, 1281]. Young [76]. Yu [1630]. Yuan [1034]. Yue [1630]. Yuetian [1561, 1562]. Yulia [1122].

Zadeh [1488, 1487, 1489]. Zammit [1667]. Zammit-Mangion [1667]. Zelterman [1216]. Zero [499, 304, 520, 1043, 1027, 1646, 1368, 1293, 979, 809, 1196, 1071]. Zero-Effect [520]. Zero-Inflated [499, 1027, 1646, 1368, 1293, 979, 1196, 1071]. Zero-Truncated [304, 1043]. Zhang [1673, 1349, 840]. Zhao [1670, 168]. Zhou [1120]. ZIG [499]. Zucchini [374, 1594].

References

Follmann:1999:BME

[1] Dean A. Follmann and Paul S. Albert. Bayesian monitoring of event rates with censored data. Biometrics, 55(2):603–607, June 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [622].

Weinberg:1999:UPE

[2] Clarice R. Weinberg and David M. Umbach. Using pooled exposure assessment to improve efficiency in case-control studies. Biometrics, 55(3): 718–726, September 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [810].

Cai:2003:HRI

[3] Tianxi Cai and Rebecca A. Betensky. Hazard regression for intervalcensored data with penalized spline. Biometrics, 59(3):570–579, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comments [653].

Baker:2005:GSP

[4] Stuart G. Baker, Paul Lichtenstein, Jaakko Kaprio, and Niels Holm. Genetic susceptibility to prostate, breast, and colorectal cancer among Nordic twins. Biometrics, 61(1):55–63, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [515].

Louis:2010:RE

[5] Thomas A. Louis, Geert Molenberghs, Geert Verbeke, David Zucker, and Marie Davidian. Report of the Editors — 2009. Biometrics, 66(1):vii– ix, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balasubramanian:2010:EHI

[6] Raji Balasubramanian and Stephen W. Lagakos. Estimating HIV incidence based on combined prevalence testing. Biometrics, 66(1):1–10, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [47].

Jacqmin-Gadda:2010:STC

[7] Hélène Jacqmin-Gadda, Cécile Proust-Lima, Jeremy M. G. Taylor, and Daniel Commenges. Score test for conditional independence between longitudinal outcome and time to event given the classes in the joint latent class model. Biometrics, 66(1):11–19, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rizopoulos:2010:MIB

[8] Dimitris Rizopoulos, Geert Verbeke, and Geert Molenberghs. Multipleimputation-based residuals and diagnostic plots for joint models of longitudinal and survival outcomes. Biometrics, 66(1):20–29, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2010:ILT

[9] Song Yang and Ross Prentice. Improved logrank-type tests for survival data using adaptive weights. Biometrics, 66(1):30–38, March 2010. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2010:SAR

[10] C.-Y. Huang, J. Qin, and M.-C. Wang. Semiparametric analysis for recurrent event data with time-dependent covariates and informative censoring. Biometrics, 66(1):39–49, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2010:SMT

[11] Yingye Zheng, Tianxi Cai, Janet L. Stanford, and Ziding Feng. Semiparametric models of time-dependent predictive values of prognostic biomarkers. Biometrics, 66(1):50–60, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reiss:2010:FGL

[12] Philip T. Reiss and R. Todd Ogden. Functional generalized linear models with images as predictors. Biometrics, 66(1):61–69, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2010:BIS

[13] Yisheng Li, Xihong Lin, and Peter Müller. Bayesian inference in semiparametric mixed models for longitudinal data. Biometrics, 66(1):70–78, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ni:2010:VSS

[14] Xiao Ni, Daowen Zhang, and Hao Helen Zhang. Variable selection for semiparametric mixed models in longitudinal studies. Biometrics, 66(1): 79–88, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2010:PLT

[15] Kai Yu, William Wheeler, Qizhai Li, Andrew W. Bergen, Neil Caporaso, Nilanjan Chatterjee, and Jinbo Chen. A partially linear tree-based regression model for multivariate outcomes. Biometrics, 66(1):89–96, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Garcia:2010:VSC

[16] Ramon I. Garcia, Joseph G. Ibrahim, and Hongtu Zhu. Variable selection in the Cox regression model with covariates missing at random. Biometrics, 66(1):97–104, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2010:BQR

[17] Ying Yuan and Guosheng Yin. Bayesian quantile regression for longitudinal studies with nonignorable missing data. Biometrics, 66(1):105–114, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeng:2010:AMU

[18] Donglin Zeng and Qingxia Chen. Adjustment for missingness using auxiliary information in semiparametric regression. Biometrics, 66(1):115– 122, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2010:AID

[19] Jihnhee Yu, Albert Vexler, and Lili Tian. Analyzing incomplete data subject to a threshold using empirical likelihood methods: an application to a pneumonia risk study in an ICU setting. Biometrics, 66(1):123–130, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Almirall:2010:SNM

[20] Daniel Almirall, Thomas Ten Have, and Susan A. Murphy. Structural nested mean models for assessing time-varying effect moderation. Biometrics, 66(1):131–139, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:2010:RBM

[21] Yuanyuan Liang and Keumhee Chough Carriere. On the role of baseline measurements for crossover designs under the self and mixed carryover effects model. Biometrics, 66(1):140–148, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2010:SRS

[22] Ying Qing Chen. Semiparametric regression in size-biased sampling. Biometrics, 66(1):149–158, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wulfsohn:2010:NSE

[23] Dvoralai Wulfsohn, Marco Sciortino, Jesper M. Aaslyng, and Marta García-Fiñana. Nondestructive, stereological estimation of canopy surface area. Biometrics, 66(1):159–168, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buckland:2010:DOL

[24] Stephen T. Buckland, Jeffrey L. Laake, and David L. Borchers. Doubleobserver line transect methods: Levels of independence. Biometrics, 66 (1):169–177, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Link:2010:ULM

[25] William A. Link, Jun Yoshizaki, Larissa L. Bailey, and Kenneth H. Pollock. Uncovering a latent multinomial: Analysis of mark-recapture data with misidentification. Biometrics, 66(1):178–185, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Foster:2010:ABU

[26] Scott D. Foster and Piers K. Dunstan. The analysis of biodiversity using rank abundance distributions. Biometrics, 66(1):186–195, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2010:IPS

[27] Gang Zheng, Zhaohai Li, Mitchell H. Gail, and Joseph L. Gastwirth. Impact of population substructure on trend tests for genetic case-control association studies. Biometrics, 66(1):196–204, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Basu:2010:LBT

[28] S. Basu, M. Stephens, J. S. Pankow, and E. A. Thompson. A likelihoodbased trait-model-free approach for linkage detection of binary trait. Biometrics, 66(1):205–213, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Javaras:2010:EDP

[29] Kristin N. Javaras, Nan M. Laird, James I. Hudson, and Brian D. Ripley. Estimating disease prevalence using relatives of case and control probands. Biometrics, 66(1):214–221, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2010:GPB

[30] Hanwen Huang, Haibo Zhou, Fuxia Cheng, Ina Hoeschele, and Fei Zou. Gaussian process based Bayesian semiparametric quantitative trait loci interval mapping. Biometrics, 66(1):222–232, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chung:2010:EDM

[31] Yuk-Ka Chung, Yue-Qing Hu, and Wing K. Fung. Evaluation of DNA mixtures from database search. Biometrics, 66(1):233–238, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kodell:2010:DSS

[32] Ralph L. Kodell, Shelly Y. Lensing, Reid D. Landes, K. Sree Kumar, and Martin Hauer-Jensen. Determination of sample sizes for demonstrating efficacy of radiation countermeasures. Biometrics, 66(1):239–248, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Henderson:2010:BCS

[33] D. A. Henderson, R. J. Boys, and D. J. Wilkinson. Bayesian calibration of a stochastic kinetic computer model using multiple data sources. Biometrics, 66(1):249–256, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wakefield:2010:BME

[34] Jon Wakefield. Bayesian methods for examining Hardy-Weinberg equilibrium. Biometrics, 66(1):257–265, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joo:2010:IPT

[35] Jungnam Joo, Minjung Kwak, and Gang Zheng. Improving power for testing genetic association in case-control studies by reducing the alternative space. Biometrics, 66(1):266–276, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kottas:2010:BNM

[36] Athanasios Kottas and Sam Behseta. Bayesian nonparametric modeling for comparison of single-neuron firing intensities. Biometrics, 66(1):277– 286, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2010:AMC

[37] Lanjia Lin, Dipankar Bandyopadhyay, Stuart R. Lipsitz, and Debajyoti Sinha. Association models for clustered data with binary and continuous responses. Biometrics, 66(1):287–293, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2010:JMC

[38] Binbing Yu and Pulak Ghosh. Joint modeling for cognitive trajectory and risk of dementia in the presence of death. Biometrics, 66(1):294– 300, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ahmed:2010:FDR

[39] I. Ahmed, C. Dalmasso, F. Haramburu, F. Thiessard, P. Broët, and P. Tubert-Bitter. False discovery rate estimation for frequentist pharmacovigilance signal detection methods. Biometrics, 66(1):301–309, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2010:MBA

[40] Devin S. Johnson, Jeffrey L. Laake, and Jay M. Ver Hoef. A modelbased approach for making ecological inference from distance sampling data. Biometrics, 66(1):310–318, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berger:2010:BRP

[41] Vance W. Berger. Book review: Permutation Tests for Stochastic Ordering and ANOVA: Theory and Applications with R, by Basso, D., Pesarin, F., Salmaso, L., and Solari, A. Biometrics, 66(1):319–321, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilmour:2010:BRC

[42] Steven G. Gilmour. Book review: A Comprehensive Guide to Factorial Two-Level Experimentation, by Mee, R. W. Biometrics, 66(1):321–322, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liddle:2010:BRD

[43] Roger F. Liddle. Book review: Data and Safety Monitoring Committees in Clinical Trials, by Herson, J. Biometrics, 66(1):322–323, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2010:BRM

[44] Ivy Liu. Book review: Marginal Models for Dependent, Clustered, and Longitudinal Categorical Data, by Bergsma, W., Croon, M., and Hagenaars, J. A. Biometrics, 66(1):323, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCandless:2010:BRA

[45] Lawrence McCandless. Book review: Applying Quantitative Bias Analysis to Epidemiologic Data, by Lash, T. L., Fox, M. P., and Fink, A. K. Biometrics, 66(1):323–324, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steel:2010:BRP

[46] Mike Steel. Book review: The Phylogenetic Handbook: a Practical Approach to Phylogenetic Analysis and Hypothesis Testing, edited by Lemey, P., Salemi, M., and Vandamme, A.-M. Biometrics, 66(1):324–325, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balasubramanian:2010:CAE

[47] Raji Balasubramanian and Stephen W. Lagakos. Correction to article “Estimating HIV Incidence Based on Combined Prevalence Testing”. Biometrics, 66(1):326, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [6].

Wu:2010:JIH

[48] L. Wu, W. Liu, and X. J. Hu. Joint inference on HIV viral dynamics and immune suppression in presence of measurement errors. Biometrics, 66(2):327–335, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nathoo:2010:JSM

[49] F. S. Nathoo. Joint spatial modeling of recurrent infection and growth with processes under intermittent observation. Biometrics, 66(2):336– 346, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diggle:2010:PLA

[50] Peter J. Diggle, Irene Kaimi, and Rosa Abellana. Partial-likelihood analysis of spatio-temporal point-process data. Biometrics, 66(2):347–354, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2010:HJS

[51] Haijun Ma, Bradley P. Carlin, and Sudipto Banerjee. Hierarchical and joint site-edge methods for Medicare hospice service region boundary analysis. Biometrics, 66(2):355–364, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schildcrout:2010:LSB

[52] Jonathan S. Schildcrout and Paul J. Rathouz. Longitudinal studies of binary response data following case-control and stratified case-control sampling: Design and analysis. Biometrics, 66(2):365–373, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2010:CRN

[53] Mengling Liu, Wenbin Lu, and Chi hong Tseng. Cox regression in nested case-control studies with auxiliary covariates. Biometrics, 66(2):374– 381, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2010:SMA

[54] Jing Qin and Yu Shen. Statistical methods for analyzing right-censored length-biased data under Cox model. Biometrics, 66(2):382–392, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xue:2010:ICM

[55] L. Xue, L. Wang, and A. Qu. Incorporating correlation for multivariate failure time data when cluster size is large. Biometrics, 66(2):393–404, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2010:RCS

[56] Menggang Yu and Bin Nan. Regression calibration in semiparametric accelerated failure time models. Biometrics, 66(2):405–414, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeng:2010:RAM

[57] Leilei Zeng, Richard J. Cook, and Theodore E. Warkentin. Regression analysis with a misclassified covariate from a current status observation scheme. Biometrics, 66(2):415–425, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2010:BCS

[58] Jing Cao, Ann Moosman, and Valen E. Johnson. A Bayesian chi-squared goodness-of-fit test for censored data models. Biometrics, 66(2):426– 434, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2010:BSS

[59] Song Zhang, Peter Müller, and Kim-Anh Do. A Bayesian semiparametric survival model with longitudinal markers. Biometrics, 66(2):435–443, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:2010:SBA

[60] Samiran Sinha, Bani K. Mallick, Victor Kipnis, and Raymond J. Carroll. Semiparametric Bayesian analysis of nutritional epidemiology data in the presence of measurement error. Biometrics, 66(2):444–454, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

MacLehose:2010:BSM

[61] Richard F. MacLehose and David B. Dunson. Bayesian semiparametric multiple shrinkage. Biometrics, 66(2):455–462, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2010:BHM

[62] Hongxiao Zhu, Marina Vannucci, and Dennis D. Cox. A Bayesian hierarchical model for classification with selection of functional predictors. Biometrics, 66(2):463–473, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2010:IPN

[63] Wei Pan, Benhuai Xie, and Xiaotong Shen. Incorporating predictor network in penalized regression with application to microarray data. Biometrics, 66(2):474–484, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2010:CFD

[64] Wenge Guo, Sanat K. Sarkar, and Shyamal D. Peddada. Controlling false discoveries in multidimensional directional decisions, with applications to gene expression data on ordered categories. Biometrics, 66(2):485– 492, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2010:SBM

[65] Lianming Wang and David B. Dunson. Semiparametric Bayes multiple testing: Applications to tumor data. Biometrics, 66(2):493–501, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2010:DIC

[66] Xiaofei Wang and Haibo Zhou. Design and inference for cancer biomarker study with an outcome and auxiliary-dependent subsampling. Biometrics, 66(2):502–511, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brinkley:2010:GEA

[67] Jason Brinkley, Anastasios Tsiatis, and Kevin J. Anstrom. A generalized estimator of the attributable benefit of an optimal treatment regime. Biometrics, 66(2):512–522, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2010:BAS

[68] Yun Li, Jeremy M. G. Taylor, and Michael R. Elliott. A Bayesian approach to surrogacy assessment using principal stratification in clinical trials. Biometrics, 66(2):523–531, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Houede:2010:UBO

[69] Nadine Houede, Peter F. Thall, Hoang Nguyen, Xavier Paoletti, and Andrew Kramar. Utility-based optimization of combination therapy using ordinal toxicity and efficacy in Phase I/II trials. Biometrics, 66(2):532– 540, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bekele:2010:RGS

[70] B. Nebiyou Bekele, Yisheng Li, and Yuan Ji. Risk-group-specific dose finding based on an average toxicity score. Biometrics, 66(2):541–548, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davidov:2010:ORS

[71] Ori Davidov, Konstantinos Fokianos, and George Iliopoulos. Orderrestricted semiparametric inference for the power bias model. Biometrics, 66(2):549–557, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Todem:2010:GST

[72] D. Todem, J. Fine, and L. Peng. A global sensitivity test for evaluating statistical hypotheses with nonidentifiable models. Biometrics, 66(2): 558–566, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hyrien:2010:SAM

[73] O. Hyrien, R. Chen, M. Mayer-Pröschel, and M. Noble. Saddlepoint approximations to the moments of multitype age-dependent branching processes, with applications. Biometrics, 66(2):567–577, June 2010. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Valderrama:2010:FPC

[74] Mariano J. Valderrama, Francisco A. Ocaña, Ana M. Aguilera, and Francisco M. Ocaña-Peinado. Forecasting pollen concentration by a twostep functional model. Biometrics, 66(2):578–585, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2010:UOR

[75] Qing Lu, Nancy Obuchowski, Sungho Won, Xiaofeng Zhu, and Robert C. Elston. Using the optimal robust receiver operating characteristic (ROC) curve for predictive genetic tests. Biometrics, 66(2):586–593, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yan:2010:USP

[76] Jun Yan, Yu Cheng, Jason P. Fine, and HuiChuan J. Lai. Uncovering symptom progression history from disease registry data with application to Young cystic fibrosis patients. Biometrics, 66(2):594–602, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weichenthal:2010:BEP

[77] Scott Weichenthal, Lawrence Joseph, Patrick Bélisle, and André Dufresne. Bayesian estimation of the probability of asbestos exposure from lung fiber counts. Biometrics, 66(2):603–612, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2010:CAR

[78] Huilin Li, Barry I. Graubard, and Mitchell H. Gail. Covariate adjustment and ranking methods to identify regions with high and low mortality rates. Biometrics, 66(2):613–620, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Demirkale:2010:LMM

[79] Cumhur Yusuf Demirkale, Dan Nettleton, and Tapabrata Maiti. Linear mixed model selection for false discovery rate control in microarray data analysis. Biometrics, 66(2):621–629, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2010:SMQ

[80] Hui Tang and Terry M. Therneau. Statistical metrics for quality assessment of high-density tiling array data. Biometrics, 66(2):630–635, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reiss:2010:DBP

[81] Philip T. Reiss, M. Henry H. Stevens, Zarrar Shehzad, Eva Petkova, and Michael P. Milham. On distance-based permutation tests for betweengroup comparisons. Biometrics, 66(2):636–643, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Arnold:2010:CRE

[82] Richard Arnold, Yu Hayakawa, and Paul Yip. Capture–recapture estimation using finite mixtures of arbitrary dimension. Biometrics, 66 (2):644–655, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Venables:2010:BRI

[83] W. N. Venables. Book review: An Introduction to Statistical Inference and Its Applications with R, by Trosset, M. W. Biometrics, 66(2):656– 657, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dodd:2010:BRR

[84] Lori E. Dodd. Book review: ROC Curves for Continuous Data, by Krzanowski, W. J. and Hand, D. J. Biometrics, 66(2):657–658, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Doerge:2010:BRG

[85] R. W. Doerge. Book review: A Guide to QTL Mapping with R/qtl, by Broman, K. W. and Sen, S. Biometrics, 66(2):658–659, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shyr:2010:BRD

[86] Yu Shyr. Book review: Design and Analysis of Clinical Trials with Timeto-Event Endpoint, edited by Peace, K. E. Biometrics, 66(2):659–660, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shao:2010:BRT

[87] Yongzhao Shao. Book review: Translational Medicine: Strategies and Statistical Methods, edited by Cosmatos, D. and Chow, S.-C. Biometrics, 66(2):660–661, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Omar:2010:BRC

[88] Rumana Omar. Book review: Clinical Prediction Models: a Practical Approach to Development, Validation and Updating, by Steyerberg, E. W. Biometrics, 66(2):661–662, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2010:BRM

[89] Hao Zhang. Book review: Maximum Penalized Likelihood Estimation: Volume II: Regression, by Eggermont, P. P. and Laricca, V. N. Biometrics, 66(2):662, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kume:2010:BRM

[90] Alfred Kume. Book review: Morphometrics with R, by Claude, J. Biometrics, 66(2):662–663, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Linkletter:2010:BRS

[91] Crystal Linkletter. Book review: Statistical Analysis of Network Data: Methods and Models, by Kolaczyk, E. D. Biometrics, 66(2):663–664, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bravo:2010:MBQ

[92] Héctor Corrada Bravo and Rafael A. Irizarry. Model-based quality assessment and base-calling for second-generation sequencing data. Biometrics, 66(3):665–674, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tai:2010:SES

[93] Yu Chuan Tai, Mark N. Kvale, and John S. Witte. Segmentation and estimation for SNP microarrays: a Bayesian multiple change-point approach. Biometrics, 66(3):675–683, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2010:DGA

[94] Xuesong Yu, Timothy W. Randolph, Hua Tang, and Li Hsu. Detecting genomic aberrations using products in a multiscale analysis. Biometrics, 66(3):684–693, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2010:HDM

[95] Gong Chen and Qing Zhou. Heterogeneity in DNA multiple alignments: Modeling, inference, and applications in Motif finding. Biometrics, 66(3): 694–704, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scheike:2010:EHE

[96] Thomas H. Scheike, Torben Martinussen, and Jeremy D. Silver. Estimating haplotype effects for survival data. Biometrics, 66(3):705–715, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2010:SAI

[97] Jinfeng Xu, John D. Kalbfleisch, and Beechoo Tai. Statistical analysis of illness-death processes and semicompeting risks data. Biometrics, 66(3): 716–725, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finkelstein:2010:STA

[98] Dianne M. Finkelstein, Rui Wang, Linda H. Ficociello, and David A. Schoenfeld. A score test for association of a longitudinal marker and an event with missing data. Biometrics, 66(3):726–732, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donnet:2010:BAG

[99] Sophie Donnet, Jean-Louis Foulley, and Adeline Samson. Bayesian analysis of growth curves using mixed models defined by stochastic differential equations. Biometrics, 66(3):733–741, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Titman:2010:SMM

[100] Andrew C. Titman and Linda D. Sharples. Semi-Markov models with phase-type sojourn distributions. Biometrics, 66(3):742–752, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chaubert-Pereira:2010:MSM

[101] Florence Chaubert-Pereira, Yann Guédon, Christian Lavergne, and Catherine Trottier. Markov and semi-Markov switching linear mixed models used to identify forest tree growth components. Biometrics, 66(3): 753–762, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steinsland:2010:UGM

[102] Ingelin Steinsland and Henrik Jensen. Utilizing Gaussian Markov random field properties of Bayesian animal models. Biometrics, 66(3):763–771, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2010:BVS

[103] Brian J. Reich, Montserrat Fuentes, Amy H. Herring, and Kelly R. Evenson. Bayesian variable selection for multivariate spatially varying coefficient regression. Biometrics, 66(3):772–782, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2010:SCD

[104] Andrea J. Cook, Yi Li, David Arterburn, and Ram C. Tiwari. Spatial cluster detection for weighted outcomes using cumulative geographic residuals. Biometrics, 66(3):783–792, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2010:PVS

[105] Jian Guo, Elizaveta Levina, George Michailidis, and Ji Zhu. Pairwise variable selection for high-dimensional model-based clustering. Biometrics, 66(3):793–804, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Braun:2010:HBD

[106] Thomas M. Braun and Shufang Wang. A hierarchical Bayesian design for Phase I trials of novel combinations of cancer therapeutic agents. Biometrics, 66(3):805–812, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2010:PSM

[107] Zhenzhen Xu and John D. Kalbfleisch. Propensity score matching in randomized clinical trials. Biometrics, 66(3):813–823, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Achy-Brou:2010:ETE

[108] Aristide C. Achy-Brou, Constantine E. Frangakis, and Michael Griswold. Estimating treatment effects of longitudinal designs using regression models on propensity scores. Biometrics, 66(3):824–833, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsonaka:2010:NMI

[109] Roula Tsonaka, Dimitris Rizopoulos, Geert Verbeke, and Emmanuel Lesaffre. Nonignorable models for intermittently missing categorical longitudinal responses. Biometrics, 66(3):834–844, September 2010. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gebregziabher:2010:SMD

[110] Mulugeta Gebregziabher and Bryan Langholz. A semiparametric missing-data-induced intensity method for missing covariate data in individually matched case-control studies. Biometrics, 66(3):845–854, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jones:2010:IMM

[111] Geoffrey Jones, Wesley O. Johnson, Timothy E. Hanson, and Ronald Christensen. Identifiability of models for multiple diagnostic testing in the absence of a gold standard. Biometrics, 66(3):855–863, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2010:ACS

[112] Rui Wang and Stephen W. Lagakos. Augmented cross-sectional prevalence testing for estimating HIV incidence. Biometrics, 66(3):864–874, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

David:2010:EPD

[113] Olivier David, Aurélie Garnier, Catherine Larédo, and Jane Lecomte. Estimation of plant demographic parameters from stage-structured censuses. Biometrics, 66(3):875–882, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pledger:2010:OCR

[114] Shirley Pledger, Kenneth H. Pollock, and James L. Norris. Open capture– recapture models with heterogeneity: II. jolly-seber model. Biometrics, 66(3):883–890, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2010:ECL

[115] Kaifeng Lu. On efficiency of constrained longitudinal data analysis versus longitudinal analysis of covariance. Biometrics, 66(3):891–896, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carrasco:2010:GCC

[116] Josep L. Carrasco. A generalized concordance correlation coefficient based on the variance components generalized linear mixed models for overdispersed count data. Biometrics, 66(3):897–904, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2010:DMM

[117] Bo Cai, David B. Dunson, and Joseph B. Stanford. Dynamic model for multivariate markers of fecundability. Biometrics, 66(3):905–913, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holan:2010:MCP

[118] Scott H. Holan, Christopher K. Wikle, Laura E. Sullivan-Beckers, and Reginald B. Cocroft. Modeling complex phenotypes: Generalized linear models using spectrogram predictors of animal communication signals. Biometrics, 66(3):914–924, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2010:ATC

[119] Meijuan Li, Cavan Reilly, and Tim Hanson. Association tests for a censored quantitative trait and candidate genes in structured populations with multilevel genetic relatedness. Biometrics, 66(3):925–933, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mukherjee:2010:CCS

[120] Bhramar Mukherjee, Jaeil Ahn, Stephen B. Gruber, Malay Ghosh, and Nilanjan Chatterjee. Case-control studies of gene-environment interaction: Bayesian design and analysis. Biometrics, 66(3):934–948, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Derado:2010:MST

[121] Gordana Derado, F. DuBois Bowman, and Clinton D. Kilts. Modeling the spatial and temporal dependence in fMRI data. Biometrics, 66(3): 949–957, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nason:2010:DAC

[122] Martha Nason and Dean Follmann. Design and analysis of crossover trials for absorbing binary endpoints. Biometrics, 66(3):958–965, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leon-Novelo:2010:ATC

[123] L. G. Leon-Novelo, X. Zhou, B. Nebiyou Bekele, and P. Müller. Assessing toxicities in a clinical trial: Bayesian inference for ordinal data nested within categories. Biometrics, 66(3):966–974, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lloyd:2010:BSO

[124] Chris J. Lloyd. Bootstrap and second-order tests of risk difference. Biometrics, 66(3):975–982, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2010:ERB

[125] Paul S. Albert and Joanna H. Shih. On estimating the relationship between longitudinal measurements and time-to-event data using a simple two-stage procedure. Biometrics, 66(3):983–987, September 2010. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ye:2010:ARF

[126] Wen Ye, Jeremy M. G. Taylor, and Xihong Lin. The authors replied as follows. Biometrics, 66(3):987–991, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Koziol:2010:LEC

[127] J. A. Koziol. Letter to the Editor: “Comment on Proschan and Nason (2009), Biometrics 65, 316–322”. Biometrics, 66(3):992–993, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2010:ARFa

[128] Anonymous. The authors replied as follows. Biometrics, 66(3):993, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:2010:BRD

[129] Peter B. Gilbert. Book review: Design of Observational Studies, edited by P. R. Rosenbaum. Biometrics, 66(3):994–995, September 2010. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hogan:2010:BRL

[130] Joseph W. Hogan. Book review: Longitudinal Data Analysis, edited by G. Fitzmaurice, M. Davidian, G. Verbeke, and G. Molenberghs. Biometrics, 66(3):995–996, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stover:2010:BRM

[131] Jason Stover. Book review: Multiple Testing Problems in Pharmaceutical Statistics, edited by A. Dmitrienko, A. C. Tamhane, and F. Bretz. Biometrics, 66(3):996–997, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rinott:2010:BRL

[132] Yosef Rinott. Book review: Life Distributions: Structure of Nonparametric, Semiparametric, and Parametric Families, by A. W. Marshall and I.

Olkin. Biometrics, 66(3):997–998, September 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Saha:2010:TDP

[133] P. Saha and P. J. Heagerty. Time-dependent predictive accuracy in the presence of competing risks. Biometrics, 66(4):999–1011, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:2010:MFA

[134] Joanna H. Shih and Paul S. Albert. Modeling familial association of ages at onset of disease in the presence of competing risk. Biometrics, 66(4): 1012–1023, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2010:CFB

[135] Yingye Zheng, Patrick J. Heagerty, Li Hsu, and Polly A. Newcomb. On combining family-based and population-based case-control data in association studies. Biometrics, 66(4):1024–1033, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2010:SAL

[136] Eunsik Park and Yuan chin I. Chang. Sequential analysis of longitudinal data in a prospective nested case-control study. Biometrics, 66(4):1034– 1042, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daudin:2010:MHR

[137] Jean-Jacques Daudin, Laurent Pierre, and Corinne Vacher. Model for heterogeneous random networks using continuous latent variables and an application to a tree-fungus network. Biometrics, 66(4):1043–1051, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hwang:2010:SSE

[138] Wen-Han Hwang and Tsung-Jen Shen. Small-sample estimation of species richness applied to forest communities. Biometrics, 66(4):1052– 1060, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2010:UAM

[139] Ariel Alonso, Annouschka Laenen, Geert Molenberghs, Helena Geys, and Tony Vangeneugden. A unified approach to multi-item reliability. Biometrics, 66(4):1061–1068, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bondell:2010:JVS

[140] Howard D. Bondell, Arun Krishna, and Sujit K. Ghosh. Joint variable selection for fixed and random effects in linear mixed-effects models. Biometrics, 66(4):1069–1077, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2010:RCU

[141] Jian Zhang and Faming Liang. Robust clustering using exponential power mixtures. Biometrics, 66(4):1078–1086, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2010:BSS

[142] Mihee Lee, Haipeng Shen, Jianhua Z. Huang, and J. S. Marron. Biclustering via sparse singular value decomposition. Biometrics, 66(4): 1087–1095, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2010:BCD

[143] Song Huang, Tiejun Tong, and Hongyu Zhao. Bias-corrected diagonal discriminant rules for high-dimensional classification. Biometrics, 66(4): 1096–1106, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Iaci:2010:MAD

[144] Ross Iaci, T. N. Sriram, and Xiangrong Yin. Multivariate association and dimension reduction: a generalization of canonical correlation analysis. Biometrics, 66(4):1107–1118, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2010:MAN

[145] Danping Liu and Xiao-Hua Zhou. A model for adjusting for nonignorable verification bias in estimation of the ROC curve and its area with likelihood-based approach. Biometrics, 66(4):1119–1128, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gustafson:2010:SBS

[146] P. Gustafson, L. C. McCandless, A. R. Levy, and S. Richardson. Simplified Bayesian sensitivity analysis for mismeasured and unobserved confounders. Biometrics, 66(4):1129–1137, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tchetgen:2010:SCO

[147] Eric J. Tchetgen Tchetgen and James Robins. The semiparametric caseonly estimator. Biometrics, 66(4):1138–1144, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lakhal-Chaieb:2010:IPC

[148] Lajmi Lakhal-Chaieb, Richard J. Cook, and Xihong Lin. Inverse probability of censoring weighted estimates of Kendall’s τ for gap time analyses. Biometrics, 66(4):1145–1152, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wolfson:2010:SIS

[149] Julian Wolfson and Peter Gilbert. Statistical identifiability and the surrogate endpoint problem, with application to vaccine trials. Biometrics, 66(4):1153–1161, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Raftery:2010:EPT

[150] Adrian E. Raftery and Le Bao. Estimating and projecting trends in HIV/ AIDS generalized epidemics using incremental mixture importance sampling. Biometrics, 66(4):1162–1173, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tamhane:2010:TPS

[151] Ajit C. Tamhane, Cyrus R. Mehta, and Lingyun Liu. Testing a primary and a secondary endpoint in a group sequential design. Biometrics, 66(4): 1174–1184, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Westfall:2010:MMT

[152] Peter H. Westfall, James F. Troendle, and Gene Pennello. Multiple Mc-Nemar tests. Biometrics, 66(4):1185–1191, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Henderson:2010:RRO

[153] Robin Henderson, Phil Ansell, and Deyadeen Alshibani. Regretregression for optimal dynamic treatment regimes. Biometrics, 66(4): 1192–1201, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2010:MIA

[154] Kaifeng Lu, Liqiu Jiang, and Anastasios A. Tsiatis. Multiple imputation approaches for the analysis of dichotomized responses in longitudinal studies with missing data. Biometrics, 66(4):1202–1208, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rummel:2010:CCM

[155] D. Rummel, T. Augustin, and H. Küchenhoff. Correction for covariate measurement error in nonparametric longitudinal regression. Biometrics, 66(4):1209–1219, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2010:DAM

[156] Wenguang Sun, Marshall M. Joffe, Jinbo Chen, and Steven M. Brunelli. Design and analysis of multiple events case-control studies. Biometrics, 66(4):1220–1229, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Teerenstra:2010:SSC

[157] Steven Teerenstra, Bing Lu, John S. Preisser, Theo van Achterberg, and George F. Borm. Sample size considerations for GEE analyses of threelevel cluster randomized trials. Biometrics, 66(4):1230–1237, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2010:SMA

[158] Duncan Lee and Gavin Shaddick. Spatial modeling of air pollution in studies of its short-term health effects. Biometrics, 66(4):1238–1246, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marques:2010:PTS

[159] T. A. Marques, S. T. Buckland, D. L. Borchers, D. Tosh, and R. A. McDonald. Point transect sampling along linear features. Biometrics, 66(4):1247–1255, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonner:2010:CCM

[160] Simon J. Bonner, Byron J. T. Morgan, and Ruth King. Continuous covariates in mark-recapture-recovery analysis: a comparison of methods. Biometrics, 66(4):1256–1265, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:2010:BET

[161] Elizabeth R. Brown. Bayesian estimation of the time-varying sensitivity of a diagnostic test with application to mother-to-child transmission of HIV. Biometrics, 66(4):1266–1274, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hennessey:2010:BAD

[162] Violeta G. Hennessey, Gary L. Rosner, Robert C. Bast, Jr., and Min-Yu Chen. A Bayesian approach to dose-response assessment and synergy and its application to in vitro dose-response studies. Biometrics, 66(4): 1275–1283, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mo:2010:BMC

[163] Qianxing Mo and Faming Liang. Bayesian modeling of ChIP-chip data through a high-order Ising model. Biometrics, 66(4):1284–1294, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Held:2010:SRA

[164] L. Held, K. Rufibach, and F. Balabdaoui. A score regression approach to assess calibration of continuous probabilistic predictions. Biometrics, 66(4):1295–1305, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandel:2010:AFT

[165] Micha Mandel and Ya’akov Ritov. The accelerated failure time model under biased sampling. Biometrics, 66(4):1306–1308, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2010:ARFb

[166] Anonymous. The author replied as follows. Biometrics, 66(4):1308, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Solymos:2010:BRP

[167] Péter Sólymos. Book review: A Primer of Ecology with R, by Stevens, M. H. H. Biometrics, 66(4):1309–1310, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Visscher:2010:BRH

[168] Peter M. Visscher. Book review: Handbook on Analyzing Human Genetic Data: Computational Approaches and Software, edited by Lin, S. and Zhao, H. Biometrics, 66(4):1310, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lovric:2010:BRF

[169] Miodrag Lovric. Book review: From Finite Sample to Asymptotic Methods in Statistics, by Sen, P. K., Singer, J. M., and Pedroso de Lima, A. C. Biometrics, 66(4):1310–1311, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Migon:2010:BRD

[170] Helio S. Migon. Book review: Dynamic Linear Models with R, by Petris, G., Petrone, S. and Campagnoli, P. Biometrics, 66(4):1311–1312, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cardot:2010:BRF

[171] Hervé Cardot. Book review: Functional Data Analysis with R and Matlab, by Ramsay, J. O., Hooker, G., and Graves, S. Biometrics, 66(4): 1312–1313, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Branco:2010:BRC

[172] Márcia D’elia Branco. Book review: Continuous Bivariate Distributions, by Balakrishnan, N. and Lai, C.-D. Biometrics, 66(4):1313–1314, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosa:2010:BRG

[173] Guilherme J. M. Rosa. Book review: Gene Expression Studies Using Affymetrix Microarrays, by Göhlmann, H. and Talloen, W. Biometrics, 66(4):1314–1315, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosa:2010:BRE

[174] Guilherme J. M. Rosa. Book review: The Elements of Statistical Learning: Data Mining, Inference, and Prediction, by Hastie, T., Tibshirani, R., and Friedman, J. Biometrics, 66(4):1315, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Louis:2011:RE

[175] Thomas A. Louis, Russell Millar, Geert Verbeke, David Zucker, and Marie Davidian. Report of the Editors — 2010. Biometrics, 67(1):vii– ix, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Logan:2011:MMC

[176] Brent R. Logan, Mei-Jie Zhang, and John P. Klein. Marginal models for clustered time-to-event data with competing risks using pseudovalues. Biometrics, 67(1):1–7, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2011:PSF

[177] Dandan Liu, John D. Kalbfleisch, and Douglas E. Schaubel. A positive stable frailty model for clustered failure time data with covariatedependent frailty. Biometrics, 67(1):8–17, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2011:PHR

[178] Hui Zhang, Douglas E. Schaubel, and John D. Kalbfleisch. Proportional hazards regression for the analysis of clustered survival data from casecohort studies. Biometrics, 67(1):18–28, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schaubel:2011:DIW

[179] Douglas E. Schaubel and Guanghui Wei. Double inverse-weighted estimation of cumulative treatment effects under nonproportional hazards and dependent censoring. Biometrics, 67(1):29–38, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Geskus:2011:CSC

[180] Ronald B. Geskus. Cause-specific cumulative incidence estimation and the fine and Gray model under both left truncation and right censoring. Biometrics, 67(1):39–49, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liao:2011:SAE

[181] Xiaomei Liao, David M. Zucker, Yi Li, and Donna Spiegelman. Survival analysis with error-prone time-varying covariates: a risk set calibration approach. Biometrics, 67(1):50–58, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guedj:2011:JMC

[182] Jeremie Guedj, Rodolphe Thiébaut, and Daniel Commenges. Joint modeling of the clinical progression and of the biomarkers’ dynamics using a mechanistic model. Biometrics, 67(1):59–66, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yi:2011:SIB

[183] G. Y. Yi, W. Liu, and Lang Wu. Simultaneous inference and bias analysis for longitudinal data with covariate measurement error and missing responses. Biometrics, 67(1):67–75, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hillary:2011:NME

[184] R. M. Hillary. A new method for estimating growth transition matrices. Biometrics, 67(1):76–85, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Di:2011:MLC

[185] Chong-Zhi Di and Karen Bandeen-Roche. Multilevel latent class models with Dirichlet mixing distribution. Biometrics, 67(1):86–96, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chagneau:2011:HBM

[186] Pierrette Chagneau, Frédéric Mortier, Nicolas Picard, and Jean-Noël Bacro. A hierarchical Bayesian model for spatial prediction of multivariate non-Gaussian random fields. Biometrics, 67(1):97–105, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tango:2011:STS

[187] Toshiro Tango, Kunihiko Takahashi, and Kazuaki Kohriyama. A spacetime scan statistic for detecting emerging outbreaks. Biometrics, 67(1): 106–115, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Warton:2011:RSE

[188] David I. Warton. Regularized sandwich estimators for analysis of highdimensional data using generalized estimating equations. Biometrics, 67 (1):116–123, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jin:2011:VSC

[189] Man Jin and Yixin Fang. Variable selection in canonical discriminant analysis for family studies. Biometrics, 67(1):124–132, March 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ho:2011:MLA

[190] Yen-Yi Ho, Giovanni Parmigiani, Thomas A. Louis, and Leslie M. Cope. Modeling liquid association. Biometrics, 67(1):133–141, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2011:BHM

[191] Fang Yu, Ming-Hui Chen, Lynn Kuo, Peng Huang, and Wanling Yang. Bayesian hierarchical modeling and selection of differentially expressed genes for the EST data. Biometrics, 67(1):142–150, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2011:PPI

[192] Xuekui Zhang, Gordon Robertson, Martin Krzywinski, Kaida Ning, Arnaud Droit, Steven Jones, and Raphael Gottardo. PICS: Probabilistic inference for ChIP-seq. Biometrics, 67(1):151–163, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:LOP

[193] Pei Wang, Dennis L. Chao, and Li Hsu. Learning oncogenic pathways from binary genomic instability data. Biometrics, 67(1):164–173, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finos:2011:KFC

[194] L. Finos and A. Farcomeni. k -FWER control without p-value adjustment, with application to detection of genetic determinants of multiple sclerosis in Italian twins. Biometrics, 67(1):174–181, March 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2011:HTM

[195] Jing Qin and Kung-Yee Liang. Hypothesis testing in a mixture casecontrol model. Biometrics, 67(1):182–193, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2011:SIT

[196] Haibo Zhou, Rui Song, Yuanshan Wu, and Jing Qin. Statistical inference for a two-stage outcome-dependent sampling design with a continuous outcome. Biometrics, 67(1):194–202, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kulinskaya:2011:THM

[197] Elena Kulinskaya, Michael B. Dollinger, and Kirsten Bjørkestøl. Testing for homogeneity in meta-analysis I. The one-parameter case: Standardized mean difference. Biometrics, 67(1):203–212, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fan:2011:SSK

[198] Chunpeng Fan, Donghui Zhang, and Cun-Hui Zhang. On sample size of the Kruskal-Wallis test with application to a mouse peritoneal cavity study. Biometrics, 67(1):213–224, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drovandi:2011:EPM

[199] C. C. Drovandi and A. N. Pettitt. Estimation of parameters for macroparasite population evolution using approximate Bayesian computation. Biometrics, 67(1):225–233, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCrea:2011:MMR

[200] Rachel S. McCrea and Byron J. T. Morgan. Multistate mark-recapture model selection using score tests. Biometrics, 67(1):234–241, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:BLQ

[201] Chenxi Li, Ying Wei, Rick Chappell, and Xuming He. Bent line quantile regression with application to an allometric study of land mammals’ speed and mass. Biometrics, 67(1):242–249, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lavielle:2011:MLE

[202] Marc Lavielle, Adeline Samson, Ana Karina Fermin, and France Mentré. Maximum likelihood estimation of long-term HIV dynamic models and antiviral response. Biometrics, 67(1):250–259, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2011:BAJ

[203] Yangxin Huang and Getachew Dagne. A Bayesian approach to joint mixed-effects models with a skew-normal distribution and measurement errors in covariates. Biometrics, 67(1):260–269, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCulloch:2011:PRE

[204] Charles E. McCulloch and John M. Neuhaus. Prediction of random effects in linear and generalized linear models under model misspecification. Biometrics, 67(1):270–279, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neelon:2011:BTP

[205] Brian Neelon, A. James O’Malley, and Sharon-Lise T. Normand. A Bayesian two-part latent class model for longitudinal medical expenditure data: Assessing the impact of mental health and substance abuse parity. Biometrics, 67(1):280–289, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dupuis:2011:EOR

[206] Jérôme A. Dupuis, Florent Bled, and Jean Joachim. Estimating the occupancy rate of spatially rare or hard to detect species: a conditional approach. Biometrics, 67(1):290–298, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2011:PLA

[207] Jun Chen, Jichun Xie, and Hongzhe Li. A penalized likelihood approach for bivariate conditional normal models for dynamic co-expression analysis. Biometrics, 67(1):299–308, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haneuse:2011:MDS

[208] S. Haneuse and J. Chen. A multiphase design strategy for dealing with participation bias. Biometrics, 67(1):309–318, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2011:EIC

[209] Stuart G. Baker. Estimation and inference for the causal effect of receiving treatment on a multinomial outcome: an alternative approach. Biometrics, 67(1):319–323, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2011:ARF

[210] Jing Cheng. The authors replied as follows. Biometrics, 67(1):323–325, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alexander:2011:BRD

[211] Neal Alexander. Book review: Design and Analysis of Vaccine Studies, by Halloran, M. E., Longini, Jr., I. M., and Struchiner, C. J. Biometrics, 67(1):326–327, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gentle:2011:BRS

[212] James E. Gentle. Book review: A SAS/IML Companion for Linear Models, by Perrett, J. J. Biometrics, 67(1):327–328, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ripley:2011:BRI

[213] Brian D. Ripley. Book review: Introducing Monte Carlo Methods with R, by Robert, C. P. and Casella, G. Biometrics, 67(1):328, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:BRI

[214] W. K. Li. Book review: Introductory Time Series with R, by Cowpertwait, P. S. P. and Metcalfe, A. V. Biometrics, 67(1):328–329, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hocking:2011:BRL

[215] Ronald R. Hocking. Book review: Linear Model Methodology, by Khuri, A. I. Biometrics, 67(1):329–330, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:ULE

[216] Yuanjia Wang, Qiong Yang, and Daniel Rabinowitz. Unbiased and locally efficient estimation of genetic effect on quantitative trait in the presence of population admixture. Biometrics, 67(2):331–343, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leek:2011:ACS

[217] Jeffrey T. Leek. Asymptotic conditional singular value decomposition for high-dimensional genomic data. Biometrics, 67(2):344–352, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:IDA

[218] Huixia Judy Wang and Jianhua Hu. Identification of differential aberrations in multiple-sample array CGH studies. Biometrics, 67(2):353–362, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bickel:2011:END

[219] David R. Bickel. Estimating the null distribution to adjust observed confidence levels for genome-scale screening. Biometrics, 67(2):363–370, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Capanu:2011:HME

[220] Marinela Capanu and Colin B. Begg. Hierarchical modeling for estimating relative risks of rare genetic variants: Properties of the pseudolikelihood method. Biometrics, 67(2):371–380, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2011:SDP

[221] Brian J. Reich and Howard D. Bondell. A spatial Dirichlet process mixture model for clustering population genetics data. Biometrics, 67(2): 381–390, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2011:SDP

[222] Luping Zhao and Timothy E. Hanson. Spatially dependent Polya tree modeling for survival data. Biometrics, 67(2):391–403, June 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2011:STM

[223] Xingqiu Zhao, Jie Zhou, and Liuquan Sun. Semiparametric transformation models with time-varying coefficients for recurrent and terminal events. Biometrics, 67(2):404–414, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gorfine:2011:FBC

[224] Malka Gorfine and Li Hsu. Frailty-based competing risks model for multivariate survival data. Biometrics, 67(2):415–426, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:ESS

[225] Yi Li, Lu Tian, and Lee-Jen Wei. Estimating subject-specific dependent competing risk profile with censored event time observations. Biometrics, 67(2):427–435, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2011:RPY

[226] Tianxi Cai, Thomas A. Gerds, Yingye Zheng, and Jinbo Chen. Robust prediction of t-year survival with data from multiple studies. Biometrics, 67(2):436–444, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Acar:2011:DCC

[227] Elif F. Acar, Radu V. Craiu, and Fang Yao. Dependence calibration in conditional copulas: a nonparametric approach. Biometrics, 67(2):445– 453, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ryu:2011:BNR

[228] Duchwan Ryu, Erning Li, and Bani K. Mallick. Bayesian nonparametric regression analysis of data with random effects covariates from longitudinal measurements. Biometrics, 67(2):454–466, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sommen:2011:EDI

[229] C. Sommen, D. Commenges, S. Le Vu, L. Meyer, and A. Alioum. Estimation of the distribution of infection times using longitudinal serological markers of HIV: Implications for the estimation of HIV incidence. Biometrics, 67(2):467–475, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2011:MES

[230] Dacheng Liu, Tao Lu, Xu-Feng Niu, and Hulin Wu. Mixed-effects statespace models for analysis of longitudinal dynamic systems. Biometrics, 67(2):476–485, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coull:2011:RIF

[231] Brent A. Coull. A random intercepts-functional slopes model for flexible assessment of susceptibility in longitudinal designs. Biometrics, 67(2): 486–494, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ibrahim:2011:FRE

[232] Joseph G. Ibrahim, Hongtu Zhu, Ramon I. Garcia, and Ruixin Guo. Fixed and random effects selection in mixed effects models. Biometrics, 67(2):495–503, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2011:HDV

[233] Fei Liu, David Dunson, and Fei Zou. High-dimensional variable selection in meta-analysis for censored data. Biometrics, 67(2):504–512, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2011:SDR

[234] Wenbin Lu and Lexin Li. Sufficient dimension reduction for censored regressions. Biometrics, 67(2):513–523, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schmid:2011:RAS

[235] Matthias Schmid, Thomas Hielscher, Thomas Augustin, and Olaf Gefeller. A robust alternative to the schemper-Henderson estimator of prediction error. Biometrics, 67(2):524–535, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsiatis:2011:IDR

[236] Anastasios A. Tsiatis, Marie Davidian, and Weihua Cao. Improved doubly robust estimation when data are monotonely coarsened, with application to longitudinal studies with dropout. Biometrics, 67(2):536–545, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ahn:2011:MED

[237] Jaeil Ahn, Bhramar Mukherjee, Stephen B. Gruber, and Samiran Sinha. Missing exposure data in stereotype regression model: Application to matched case-control study with disease subclassification. Biometrics, 67(2):546–558, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Long:2011:REA

[238] Qi Long, Xiaoxi Zhang, and Brent A. Johnson. Robust estimation of area under ROC curve using auxiliary variables in the presence of missing biomarker values. Biometrics, 67(2):559–567, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2011:OCD

[239] Xin Huang, Gengsheng Qin, and Yixin Fang. Optimal combinations of diagnostic tests based on AUC. Biometrics, 67(2):568–576, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dail:2011:MEA

[240] D. Dail and L. Madsen. Models for estimating abundance from repeated counts of an open metapopulation. Biometrics, 67(2):577–587, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hughes:2011:MMQ

[241] John Hughes and John Fricks. A mixture model for quantum dot images of kinesin motor assays. Biometrics, 67(2):588–595, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartroff:2011:IIC

[242] Jay Bartroff and Tze Leung Lai. Incorporating individual and collective ethics into Phase I cancer trial designs. Biometrics, 67(2):596–603, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tian:2011:HCD

[243] Lu Tian, Rui Wang, Tianxi Cai, and Lee-Jen Wei. The highest confidence density region and its usage for joint inferences about constrained parameters. Biometrics, 67(2):604–610, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mardia:2011:HBM

[244] Kanti V. Mardia, Vysaul B. Nyirongo, Christopher J. Fallaize, Stuart Barber, and Richard M. Jackson. Hierarchical Bayesian modeling of pharmacophores in bioinformatics. Biometrics, 67(2):611–619, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Adebayo:2011:GLV

[245] Samson B. Adebayo, Ludwig Fahrmeir, Christian Seiler, and Christian Heumann. Geoadditive latent variable modeling of count data on multiple sexual partnering in Nigeria. Biometrics, 67(2):620–628, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:LMB

[246] Xiaoyun Li, Dipankar Bandyopadhyay, Stuart Lipsitz, and Debajyoti Sinha. Likelihood methods for binary responses of present components in a cluster. Biometrics, 67(2):629–635, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2011:BRAa

[247] Zhiwei Zhang and Paul S. Albert. Binary regression analysis with pooled exposure measurements: a regression calibration approach. Biometrics, 67(2):636–645, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosner:2011:PSS

[248] Bernard Rosner and Robert J. Glynn. Power and sample size estimation for the clustered Wilcoxon test. Biometrics, 67(2):646–653, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neuhaus:2011:NTI

[249] John M. Neuhaus, Charles E. McCulloch, and Ross Boylan. A note on Type II error under random effects misspecification in generalized linear mixed models. Biometrics, 67(2):654–656, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [250].

Litiere:2011:RNT

[250] Saskia Litière, Ariel Alonso, and Geert Molenberghs. Rejoinder to “A Note on Type II Error Under Random Effects Misspecification in Generalized Linear Mixed Models”. Biometrics, 67(2):656–660, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [249].

Zhou:2011:CRR

[251] Bingqing Zhou, Aurelien Latouche, Vanderson Rocha, and Jason Fine. Competing risks regression for stratified data. Biometrics, 67(2):661– 670, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Commenges:2011:BRH

[252] Daniel Commenges. Book review: Handbook of Spatial Statistics, edited by Gelfand, A. E., Diggle, P. J., Fuentes, M. and Guttorp, P. Biometrics, 67(2):671–672, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lumley:2011:BRC

[253] Thomas Lumley. Book review: Computational Statistics, by Gentle, J. E. Biometrics, 67(2):672, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mattei:2011:BRB

[254] Alessandra Mattei. Book review: Bayesian Missing Data Problems: EM, Data Augmentation and Noniterative Computation, by Tan, M. T., Tian, G.-L., and Ng, K. W. Biometrics, 67(2):672–674, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stangl:2011:BRF

[255] Dalene Stangl. Book review: A First Course in Bayesian Statistical Methods, by Hoff, P. D. Biometrics, 67(2):674, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finley:2011:BRE

[256] Andrew Finley. Book review: Environmental and Ecological Statistics with R, by Qian, S. S. Biometrics, 67(2):674–675, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marasinghe:2011:BRM

[257] Mervyn G. Marasinghe. Book review: A Modern Approach to Regression with R, by Sheather, S. J. Biometrics, 67(2):675–677, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Emerson:2011:BRE

[258] Scott S. Emerson. Book review: Exercises and Solutions in Biostatistical Theory, by Kupper, L. L., Neelon, B. H., and O’Brien, S. M. Biometrics, 67(2):677–678, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilkinson:2011:BRG

[259] Leland Wilkinson. Book review: ggplot2: Elegant Graphics for Data Analysis, by Wickham, H. Biometrics, 67(2):678–679, June 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Horrocks:2011:LAE

[260] Julie Horrocks, David C. Hamilton, and Hal Whitehead. A likelihood approach to estimating animal density from binary acoustic transects. Biometrics, 67(3):681–690, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2011:ASC

[261] Zhi Wen, Kenneth Pollock, James Nichols, and Peter Waser. Augmenting superpopulation capture–recapture models with population assignment data. Biometrics, 67(3):691–700, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:QRL

[262] Ruosha Li and Limin Peng. Quantile regression for left-truncated semicompeting risks data. Biometrics, 67(3):701–710, September 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guan:2011:VES

[263] Yongtao Guan, Jun Yan, and Rajita Sinha. Variance estimation for statistics computed from single recurrent event processes. Biometrics, 67(3):711–718, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsieh:2011:RAR

[264] Jin-Jian Hsieh, A. Adam Ding, and Weijing Wang. Regression analysis for recurrent events data under dependent censoring. Biometrics, 67(3): 719–729, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guan:2011:SOA

[265] Yongtao Guan. Second-order analysis of semiparametric recurrent event processes. Biometrics, 67(3):730–739, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2011:EDR

[266] Min Zhang and Douglas E. Schaubel. Estimating differences in restricted mean lifetime using observational data subject to dependent censoring. Biometrics, 67(3):740–749, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stare:2011:MEV

[267] Janez Stare, Maja Pohar Perme, and Robin Henderson. A measure of explained variation for event history data. Biometrics, 67(3):750–759, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2011:ACS

[268] Chi-Chung Wen and Chien-Tai Lin. Analysis of current status data with missing covariates. Biometrics, 67(3):760–769, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2011:NCP

[269] Xingqiu Zhao and Jianguo Sun. Nonparametric comparison for panel count data with unequal observation processes. Biometrics, 67(3):770– 779, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Titman:2011:FNM

[270] Andrew C. Titman. Flexible nonhomogeneous Markov models for panel observed data. Biometrics, 67(3):780–787, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2011:SMP

[271] Shen-Ming Lee, Mei-Jih Gee, and Shu-Hui Hsieh. Semiparametric methods in the proportional odds model for ordinal response data with missing covariates. Biometrics, 67(3):788–798, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2011:MIM

[272] Hua Yun Chen, Hui Xie, and Yi Qian. Multiple imputation for missing values through conditional semiparametric odds ratio models. Biometrics, 67(3):799–809, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:NMI

[273] Chenguang Wang and Michael J. Daniels. A note on MAR, identifying restrictions, model comparison, and sensitivity analysis in pattern mixture models with and without covariates for incomplete data. Biometrics, 67(3):810–818, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [383].

Basford:2011:ITO

[274] Kaye E. Basford. IBS: Transforming our governance. Biometrics, 67(4): 1185–1188, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buzas:2011:SPD

[275] Jeffrey S. Buzas, Carrie G. Wager, and David M. Lansky. Split-plot designs for robotic serial dilution assays. Biometrics, 67(4):1189–1196, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2011:MPL

[276] Ji-Yeon Yang and Xuming He. A multistep protein lysate array quantification method and its statistical properties. Biometrics, 67(4):1197– 1205, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2011:DLM

[277] W. Evan Johnson, Noah C. Welker, and Brenda L. Bass. Dynamic linear model for the identification of miRNAs in next-generation sequencing data. Biometrics, 67(4):1206–1214, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ji:2011:BMB

[278] Yuan Ji, Yanxun Xu, Qiong Zhang, Kam-Wah Tsui, Yuan Yuan, Clift Norris, Jr., Shoudan Liang, and Han Liang. BM-Map: Bayesian mapping of multireads for next-generation sequencing data. Biometrics, 67(4): 1215–1224, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsui:2011:EES

[279] Shigeyuki Matsui and Hisashi Noma. Estimating effect sizes of differentially expressed genes for power and sample-size assessments in microarray experiments. Biometrics, 67(4):1225–1235, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roverato:2011:WDM

[280] Alberto Roverato and F. Marta L. Di Lascio. Wilks’ Λ dissimilarity measures for gene clustering: an approach based on the identification of transcription modules. Biometrics, 67(4):1236–1248, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Di:2011:LRT

[281] Chong-Zhi Di and Kung-Yee Liang. Likelihood ratio testing for admixture models with application to genetic linkage analysis. Biometrics, 67(4):1249–1259, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoffmann:2011:CDM

[282] Thomas J. Hoffmann, Stijn Vansteelandt, Christoph Lange, Edwin K. Silverman, Dawn L. DeMeo, and Nan M. Laird. Combining disease models to test for gene-environment interaction in nuclear families. Biometrics, 67(4):1260–1270, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maity:2011:PTD

[283] Arnab Maity and Xihong Lin. Powerful tests for detecting a gene effect in the presence of possible gene-gene interactions using garrote kernel machines. Biometrics, 67(4):1271–1284, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yi:2011:PGP

[284] G. Yi, J. Q. Shi, and T. Choi. Penalized Gaussian process regression and classification for high-dimensional nonlinear data. Biometrics, 67 (4):1285–1294, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2011:SFD

[285] Bin Zhu, Peter X.-K. Song, and Jeremy M. G. Taylor. Stochastic functional data analysis: a diffusion model-based approach. Biometrics, 67 (4):1295–1304, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2011:REO

[286] J. Cao, L. Wang, and J. Xu. Robust estimation for ordinary differential equation models. Biometrics, 67(4):1305–1313, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gaugler:2011:TIT

[287] Trent Gaugler and Michael G. Akritas. Testing for interaction in two-way random and mixed effects models: The fully nonparametric approach. Biometrics, 67(4):1314–1320, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:PIL

[288] Yimei Li, E. Paul Wileyto, and Daniel F. Heitjan. Prediction of individual long-term outcomes in smoking cessation trials using frailty models. Biometrics, 67(4):1321–1329, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Du:2011:SSA

[289] Pang Du, Yihua Jiang, and Yuedong Wang. Smoothing spline ANOVA frailty model for recurrent event data. Biometrics, 67(4):1330–1339, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2011:AME

[290] Jianwen Cai and Donglin Zeng. Additive mixed effect model for clustered failure time data. Biometrics, 67(4):1340–1351, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2011:NSE

[291] Jiajia Zhang, Yingwei Peng, and Ou Zhao. A new semiparametric estimation method for accelerated hazard model. Biometrics, 67(4):1352–1360, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pipper:2011:EVF

[292] C. B. Pipper, C. Ritz, and T. H. Scheike. Explained variation in a fully specified model for data-grouped survival data. Biometrics, 67(4): 1361–1368, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ning:2011:BJT

[293] Jing Ning, Jing Qin, and Yu Shen. Buckley–James-type estimator with right-censored and length-biased data. Biometrics, 67(4):1369–1378, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2011:PRC

[294] Brent A. Johnson, Qi Long, and Matthias Chung. On path restoration for censored outcomes. Biometrics, 67(4):1379–1388, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Uno:2011:GPE

[295] Hajime Uno, Tianxi Cai, Lu Tian, and L. J. Wei. Graphical procedures for evaluating overall and subject-specific incremental values from new predictors with censored event time data. Biometrics, 67(4):1389–1396, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nie:2011:IET

[296] Hui Nie, Jing Cheng, and Dylan S. Small. Inference for the effect of treatment on survival probability in randomized trials with noncompliance and administrative censoring. Biometrics, 67(4):1397–1405, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanderWeele:2011:NCC

[297] Tyler J. VanderWeele and Ilya Shpitser. A new criterion for confounder selection. Biometrics, 67(4):1406–1413, December 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanderWeele:2011:ICI

[298] Tyler J. VanderWeele, Yu Chen, and Habibul Ahsan. Inference for causal interactions for continuous exposures under dichotomization. Biometrics, 67(4):1414–1421, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2011:RLS

[299] Yufan Zhao, Donglin Zeng, Mark A. Socinski, and Michael R. Kosorok. Reinforcement learning strategies for clinical trials in nonsmall cell lung cancer. Biometrics, 67(4):1422–1433, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:SAE

[300] Yun Li, Jeremy M. G. Taylor, and Roderick J. A. Little. A shrinkage approach for estimating a treatment effect using intermediate biomarker data in clinical trials. Biometrics, 67(4):1434–1441, December 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2011:CBP

[301] Ying Huang and Peter B. Gilbert. Comparing biomarkers as principal surrogate endpoints. Biometrics, 67(4):1442–1451, December 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:MIM

[302] Tao Wang and Lang Wu. Multiple imputation methods for multivariate one-sided tests with missing data. Biometrics, 67(4):1452–1460, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thomas:2011:MAI

[303] Laine Thomas, Leonard Stefanski, and Marie Davidian. A momentadjusted imputation method for measurement error models. Biometrics, 67(4):1461–1470, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2011:DME

[304] Yih-Huei Huang, Wen-Han Hwang, and Fei-Yin Chen. Differential measurement errors in zero-truncated regression models for count data. Biometrics, 67(4):1471–1480, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:NTH

[305] Jun Li, Jifei Ban, and Louis S. Santiago. Nonparametric tests for homogeneity of species assemblages: a data depth approach. Biometrics, 67(4):1481–1488, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dupuis:2011:ESR

[306] Jérôme A. Dupuis and Michel Goulard. Estimating species richness from quadrat sampling data: a general approach. Biometrics, 67(4):1489– 1497, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonner:2011:SPS

[307] Simon J. Bonner and Carl J. Schwarz. Smoothing population size estimates for time-stratified mark-recapture experiments using Bayesian P-splines. Biometrics, 67(4):1498–1507, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2011:DDO

[308] Yingqi Zhao, Donglin Zeng, Amy H. Herring, Amy Ising, Anna Waller, David Richardson, and Michael R. Kosorok. Detecting disease outbreaks using local spatiotemporal methods. Biometrics, 67(4):1508–1517, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fewster:2011:VES

[309] R. M. Fewster. Variance estimation for systematic designs in spatial surveys. Biometrics, 67(4):1518–1531, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2011:GPM

[310] Ying Guo. A general probabilistic model for group independent component analysis and its estimation methods. Biometrics, 67(4):1532–1542, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2011:DRC

[311] Ying Yuan and Guosheng Yin. Dose-response curve estimation: a semiparametric mixture approach. Biometrics, 67(4):1543–1554, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wages:2011:CRM

[312] Nolan A. Wages, Mark R. Conaway, and John O’Quigley. Continual reassessment method for partial ordering. Biometrics, 67(4):1555–1563, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Al-Khalidi:2011:IRV

[313] Hussein R. Al-Khalidi, Yili Hong, Thomas R. Fleming, and Terry M. Therneau. Insights on the robust variance estimator under recurrentevents model. Biometrics, 67(4):1564–1572, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:HDM

[314] Chia-Ning Wang, Roderick Little, Bin Nan, and Siobán D. Harlow. A hot-deck multiple imputation procedure for gaps in longitudinal recurrent event histories. Biometrics, 67(4):1573–1582, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schildcrout:2011:ODS

[315] Jonathan S. Schildcrout and Patrick J. Heagerty. Outcome-dependent sampling from existing cohorts with longitudinal binary response data: Study planning and analysis. Biometrics, 67(4):1583–1593, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lachos:2011:LNM

[316] Victor H. Lachos, Dipankar Bandyopadhyay, and Dipak K. Dey. Linear and nonlinear mixed-effects models for censored HIV viral loads using normal/independent distributions. Biometrics, 67(4):1594–1604, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bobb:2011:BMA

[317] Jennifer F. Bobb, Francesca Dominici, and Roger D. Peng. A Bayesian model averaging approach for estimating the relative risk of mortality associated with heat waves in 105 U.S. cities. Biometrics, 67(4):1605– 1616, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ruan:2011:EBA

[318] Lingyan Ruan and Ming Yuan. An empirical Bayes’ approach to joint analysis of multiple microarray gene expression studies. Biometrics, 67 (4):1617–1626, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boonstra:2011:BMG

[319] Philip S. Boonstra, Bhramar Mukherjee, Jeremy M. G. Taylor, Mef Nilbert, Victor Moreno, and Stephen B. Gruber. Bayesian modeling for genetic anticipation in presence of mutational heterogeneity: a case study in Lynch syndrome. Biometrics, 67(4):1627–1637, December 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2011:OCB

[320] Peter F. Thall, Aniko Szabo, Hoang Q. Nguyen, Catherine M. Amlie-Lefond, and Osama O. Zaidat. Optimizing the concentration and bolus of a drug delivered by continuous infusion. Biometrics, 67(4):1638–1646, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zetlaoui:2011:EFC

[321] Mélanie Zetlaoui, Max Feinberg, Philippe Verger, and Stephan Clémençon. Extraction of food consumption systems by nonnegative matrix factorization (NMF) for the assessment of food choices. Biometrics, 67(4):1647–1658, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stoklosa:2011:HCR

[322] Jakub Stoklosa, Wen-Han Hwang, Sheng-Hai Wu, and Richard Huggins. Heterogeneous capture–recapture models with covariates: a partial likelihood approach for closed populations. Biometrics, 67(4):1659–1665, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2011:NMA

[323] Xiaochun Li, Mengling Liu, and Judith D. Goldberg. A note on monotonicity assumptions for exact unconditional tests in binary matchedpairs designs. Biometrics, 67(4):1666–1668, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sabo:2011:LEB

[324] Roy T. Sabo and N. Rao Chaganty. Letter to the Editor of Biometrics on “Joint Regression Analysis for Discrete Longitudinal Data” by Madsen and Fang. Biometrics, 67(4):1669–1670, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [371].

Madsen:2011:ARF

[325] L. Madsen and Y. Fang. The authors replied as follows. Biometrics, 67(4):1670–1671, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lou:2011:BRS

[326] Xiang-Yang Lou and David B. Allison. Book review: Statistics in Human Genetics and Molecular Biology, by Reilly, C. Biometrics, 67(4):1672– 1673, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Westfall:2011:BRM

[327] Peter H. Westfall. Book review: Meta-Analysis and Combining Information in Genetics and Genomics, by Guerra, R. and Goldstein, D. R. Biometrics, 67(4):1673–1674, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hubert:2011:BRM

[328] Mia Hubert. Book review: Multivariate Nonparametric Methods with R. An Approach Based on Spatial Signs and Ranks, by Oja, H. Biometrics, 67(4):1674–1675, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonner:2011:BRB

[329] Simon Bonner. Book review: Bayesian Analysis for Population Ecology, by King, R., Morgan, B. J. T., Gimenez, O., and Brooks, S. P. Biometrics, 67(4):1675–1676, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eilers:2011:BRO

[330] Paul H. C. Eilers. Book review: The Oxford Handbook of Functional Data Analysis, edited by Ferraty, F. and Romain, Y. Biometrics, 67 (4):1676, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giesbrecht:2011:BRD

[331] Francis Giesbrecht. Book review: Design and Analysis of Experiments with SAS, by Lawson, J. Biometrics, 67(4):1676–1677, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ribeiro:2011:BRS

[332] Paulo Justiniano Ribeiro, Jr. Book review: Spatial Statistics and Modeling, by Gaetan, C. and Guyon, X. Biometrics, 67(4):1677–1678, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2011:BRAb

[333] Kui Zhang. Book review: Applied Statistical Genetics with R: For Population-based Association Studies, by Foulkes, A. S. Biometrics, 67 (4):1678–1679, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rizopoulos:2011:DPP

[334] Dimitris Rizopoulos. Dynamic predictions and prospective accuracy in joint models for longitudinal and time-to-event data. Biometrics, 68(3): 819–829, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2011:DRE

[335] Baojiang Chen and Xiao-Hua Zhou. Doubly robust estimates for binary longitudinal data analysis with missing response and missing covariates. Biometrics, 68(3):830–842, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2011:ICS

[336] Ying Huang and Brian Leroux. Informative cluster sizes for subclusterlevel covariates and weighted generalized estimating equations. Biometrics, 68(3):843–851, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2011:RAR

[337] Shuang Wu and Hans-Georg Müller. Response-adaptive regression for longitudinal data. Biometrics, 68(3):852–860, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2011:PSA

[338] Huaihou Chen and Yuanjia Wang. A penalized spline approach to functional mixed effects model analysis. Biometrics, 68(3):861–870, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nummi:2011:TCS

[339] Tapio Nummi, Jianxin Pan, Tarja Siren, and Kun Liu. Testing for cubic smoothing splines under dependent data. Biometrics, 68(3):871–875, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2011:PLM

[340] Haibo Zhou, Guoyou Qin, and Matthew P. Longnecker. A partial linear model in the outcome-dependent sampling setting to evaluate the effect of prenatal PCB exposure on cognitive function in children. Biometrics, 68(3):876–885, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2011:SDR

[341] Brian J. Reich, Howard D. Bondell, and Lexin Li. Sufficient dimension reduction via Bayesian mixture modeling. Biometrics, 68(3):886–895, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:PBS

[342] Yuanjia Wang, Huaihou Chen, Runze Li, Naihua Duan, and Roberto Lewis-Fernández. Prediction-based structured variable selection through the receiver operating characteristic curves. Biometrics, 68(3):896–905, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2011:SEC

[343] Danping Liu and Xiao-Hua Zhou. Semiparametric estimation of the covariate-specific ROC curve in presence of ignorable verification bias. Biometrics, 68(3):906–916, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2011:BVS

[344] Joyee Ghosh, Amy H. Herring, and Anna Maria Siega-Riz. Bayesian variable selection for latent class models. Biometrics, 68(3):917–925, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guan:2011:BCV

[345] Yongtao Guan. Bias-corrected variance estimation and hypothesis testing for spatial point and marked point processes using subsampling. Biometrics, 68(3):926–936, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yue:2011:BSI

[346] Yu Ryan Yue and Ji Meng Loh. Bayesian semiparametric intensity estimation for inhomogeneous spatial point processes. Biometrics, 68(3): 937–946, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christensen:2011:FKS

[347] William F. Christensen. Filtered kriging for spatial data with heterogeneous measurement error variances. Biometrics, 68(3):947–957, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2011:NBA

[348] Xiaoyu Jiang, David Gold, and Eric D. Kolaczyk. Network-based autoprobit modeling for protein function prediction. Biometrics, 68(3):958– 966, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2011:CPS

[349] Seunggeun Lee, Fred A. Wright, and Fei Zou. Control of population stratification by correlation-selected principal components. Biometrics, 68(3):967–974, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2011:KMA

[350] Tianxi Cai, Giulia Tonini, and Xihong Lin. Kernel machine approach to testing the significance of multiple genetic markers for risk prediction. Biometrics, 68(3):975–986, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:SIM

[351] Xueqin Wang, Xiaobo Guo, Mingguang He, and Heping Zhang. Statistical inference in mixed models and analysis of twin and family data. Biometrics, 68(3):987–995, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2011:EMC

[352] Wenqin Pan and Donglin Zeng. Estimating mean cost using auxiliary covariates. Biometrics, 68(3):996–1006, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hiriote:2011:MBC

[353] Sasiprapa Hiriote and Vernon M. Chinchilli. Matrix-based concordance correlation coefficient for repeated measures. Biometrics, 68(3):1007– 1016, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2011:NSS

[354] Paul R. Rosenbaum. A new u-statistic with superior design sensitivity in matched observational studies. Biometrics, 68(3):1017–1027, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2011:GCM

[355] Jeffrey M. Albert and Suchitra Nelson. Generalized causal mediation analysis. Biometrics, 68(3):1028–1038, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gutjahr:2011:ACE

[356] Georg Gutjahr, Werner Brannath, and Peter Bauer. An approach to the conditional error rate principle with nuisance parameters. Biometrics, 68(3):1039–1046, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hobbs:2011:HCP

[357] Brian P. Hobbs, Bradley P. Carlin, Sumithra J. Mandrekar, and Daniel J. Sargent. Hierarchical commensurate and power prior models for adaptive incorporation of historical information in clinical trials. Biometrics, 68 (3):1047–1056, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pfeiffer:2011:TCE

[358] R. M. Pfeiffer and M. H. Gail. Two criteria for evaluating risk prediction models. Biometrics, 68(3):1057–1065, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carpenter:2011:ASM

[359] James Carpenter, Gerta Rücker, and Guido Schwarzer. Assessing the sensitivity of meta-analysis to selection bias: a multiple imputation approach. Biometrics, 68(3):1066–1072, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fu:2011:BMM

[360] Rongwei Fu, Dipak K. Dey, and Kent E. Holsinger. A beta-mixture model for assessing genetic population structure. Biometrics, 68(3):1073–1082, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shepherd:2011:ADE

[361] Bryan E. Shepherd and Chang Yu. Accounting for data errors discovered from an audit in multiple linear regression. Biometrics, 68(3):1083–1091, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dodd:2011:ASP

[362] Lori E. Dodd, Edward L. Korn, Boris Freidlin, Robert Gray, and Suman Bhattacharya. An audit strategy for progression-free survival. Biometrics, 68(3):1092–1099, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shepherd:2011:SAC

[363] Bryan E. Shepherd, Peter B. Gilbert, and Charles T. Dupont. Sensitivity analyses comparing time-to-event outcomes only existing in a subset selected postrandomization and relaxing monotonicity. Biometrics, 68 (3):1100–1110, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:SBP

[364] Lianming Wang and David B. Dunson. Semiparametric Bayes’ proportional odds models for current status data with underreporting. Biometrics, 68(3):1111–1118, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:2011:BPI

[365] Sanjoy K. Sinha, Andrea B. Troxel, Stuart R. Lipsitz, Debajyoti Sinha, Garrett M. Fitzmaurice, Geert Molenberghs, and Joseph G. Ibrahim. A bivariate pseudolikelihood for incomplete longitudinal binary data with nonignorable nonmonotone missingness. Biometrics, 68(3):1119–1126, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:2011:APM

[366] Dean Follmann and Martha Nason. An augmented probit model for missing predictable covariates in quantal bioassay with small sample size.

Biometrics, 68(3):1127–1134, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proschan:2011:MUM

[367] Michael Proschan, Erica Brittain, and Lisa Kammerman. Minimize the use of minimization with unequal allocation. Biometrics, 68(3):1135– 1141, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2011:EST

[368] Cindy Xin Feng, Jiguo Cao, and Leah Bendell. Exploring spatial and temporal variations of cadmium concentrations in Pacific oysters from British Columbia. Biometrics, 68(3):1142–1152, September 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hatfield:2011:MEB

[369] Laura A. Hatfield, Steve Gutreuter, Michael A. Boogaard, and Bradley P. Carlin. Multilevel empirical Bayes modeling for improved estimation of toxicant formulations to suppress parasitic sea lamprey in the upper great lakes. Biometrics, 68(3):1153–1162, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2011:BDN

[370] Ming-Hui Chen, Joseph G. Ibrahim, Peter Lam, Alan Yu, and Yuanye Zhang. Bayesian design of noninferiority trials for medical devices using historical data. Biometrics, 68(3):1163–1170, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Madsen:2011:JRA

[371] L. Madsen and Y. Fang. Joint regression analysis for discrete longitudinal data. Biometrics, 68(3):1171–1175, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [372] and letter [324].

Song:2011:RJR

[372] Peter X.-K. Song, Mingyao Li, and Ying Yuan. Rejoinder to “Joint Regression Analysis for Discrete Longitudinal Data” by Madsen and Fang. Biometrics, 68(3):1175–1176, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [371].

Friendly:2011:BRG

[373] Michael Friendly. Book review: Graphics for Statistics and Data Analysis with R, by Keen, K. J. Biometrics, 68(3):1177, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guttorp:2011:BRH

[374] Peter Guttorp. Book review: Hidden Markov Models for Time Series: an Introduction Using R, by Zucchini, W. and MacDonald, I. L. Biometrics, 68(3):1178, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tan:2011:BRB

[375] Say Beng Tan. Book review: Bayesian Adaptive Methods for Clinical Trials, by Berry, S. M., Carlin, B. P., Lee, J. J., and Muller, P. Biometrics, 68(3):1178–1179, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steibel:2011:BRS

[376] Juan P. Steibel. Book review: SAS and R Data Management, Statistical Analysis, and Graphics, by Kleinman, K. and Horton, N. Biometrics, 68(3):1179–1180, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rizzo:2011:BRN

[377] Maria L. Rizzo. Book review: Numerical Analysis for Statisticians, by Lange, K. Biometrics, 68(3):1180–1181, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Waller:2011:BRS

[378] Lance A. Waller. Book review: Statistical Methods for Disease Clustering, by Tango, T. Biometrics, 68(3):1181–1182, September 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2011:BRD

[379] Anonymous. Book review: Design and Analysis of Quality of Life Studies in Clinical Trials, by Fairclough, D. L. Biometrics, 68(3):1182, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2011:BRSa

[380] Anonymous. Book review: Survival Analysis Using SAS: a Practical Guide, by Allison, P. D. Biometrics, 68(3):1182, September 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2011:BRN

[381] Anonymous. Book review: Nonparametric Statistical Inference, by Gibbons, J. D. and Chakraborti, S. Biometrics, 68(3):1182–1183, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2011:BRSb

[382] Anonymous. Book review: Stochastic Processes: an Introduction, by Jones, P. W and Smith, P. Biometrics, 68(3):1183, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2012:CNM

[383] Stuart G. Baker, Paul Lichtenstein, Jaakko Kaprio, and Niels Holm. Correction to “A Note on MAR, Identifying Restrictions, Model Comparison, and Sensitivity Analysis in Pattern Mixture Models with and without Covariates for Incomplete Data,” by Chenguang Wang and Michael J. Daniels; 67, 810–818, September 2011. Biometrics, 61(1):994, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [273].

Dickhaus:2012:BRM

[384] Thorsten Dickhaus. Book review: Multiple Comparisons Using R, by Bretz, F., Hothorn, T., and Westfall, P. Biometrics, 67(3):995, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mitra:2012:BRF

[385] Nandita Mitra. Book review: The Fundamentals of Modern Statistical Genetics, by Laird, N. M. and Lange, C. Biometrics, 67(3):995–996, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ding:2012:BRC

[386] Meichun Ding. Book review: Clinical Trial Design: Bayesian and Frequentist Adaptive Methods, by Yin, G. Biometrics, 67(3):996–997, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pettit:2012:BRB

[387] Lawrence Pettit. Book review: Bayesian Model Selection and Statistical Modeling, by Ando, T. Biometrics, 67(3):997–998, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:RE

[388] Anonymous. Report of the Editors — 2011. Biometrics, 68(1):vii–x, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2012:SFE

[389] Wei Sun. A statistical framework for eQTL mapping using RNA-seq data. Biometrics, 68(1):1–11, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2012:MLM

[390] Wei Sun and Lexin Li. Multiple loci mapping via model-free variable selection. Biometrics, 68(1):12–22, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCormick:2012:DLR

[391] Tyler H. McCormick, Adrian E. Raftery, David Madigan, and Randall S. Burd. Dynamic logistic regression and dynamic model averaging for binary classification. Biometrics, 68(1):23–30, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Croux:2012:REM

[392] Christophe Croux, Irène Gijbels, and Ilaria Prosdocimi. Robust estimation of mean and dispersion functions in extended generalized additive models. Biometrics, 68(1):31–44, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Malinovsky:2012:PDO

[393] Yaakov Malinovsky, Paul S. Albert, and Enrique F. Schisterman. Pooling designs for outcomes under a Gaussian random effects model. Biometrics, 68(1):45–52, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Braun:2012:PCV

[394] Julia Braun, Leonhard Held, and Bruno Ledergerber. Predictive crossvalidation for the choice of linear mixed-effects models with application to data from the Swiss HIV cohort study. Biometrics, 68(1):53–61, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Claggett:2012:ACS

[395] B. Claggett, S. W. Lagakos, and R. Wang. Augmented cross-sectional studies with abbreviated follow-up for estimating HIV incidence. Biometrics, 68(1):62–74, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2012:PRT

[396] Kai Zhang, Mikhail Traskin, and Dylan S. Small. A powerful and robust test statistic for randomization inference in group-randomized trials with matched pairs of groups. Biometrics, 68(1):75–84, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shinohara:2012:BSC

[397] Russell T. Shinohara, Constantine E. Frangakis, and Constantine G. Lyketsos. A broad symmetry criterion for nonparametric validity of parametrically based tests in randomized trials. Biometrics, 68(1):85– 91, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burgette:2012:MAB

[398] Lane F. Burgette and Jerome P. Reiter. Modeling adverse birth outcomes via confirmatory factor quantile regression. Biometrics, 68(1):92–100, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ji:2012:QRD

[399] Shuang Ji, Limin Peng, Yu Cheng, and HuiChuan Lai. Quantile regression for doubly censored data. Biometrics, 68(1):101–112, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perme:2012:ERS

[400] Maja Pohar Perme, Janez Stare, and Jacques Estève. On estimation in relative survival. Biometrics, 68(1):113–120, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yan:2012:ICE

[401] Wei Yan, Yaqin Hu, and Zhi Geng. Identifiability of causal effects for binary variables with baseline data missing due to death. Biometrics, 68 (1):121–128, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seaman:2012:CMI

[402] Shaun R. Seaman, Ian R. White, Andrew J. Copas, and Leah Li. Combining multiple imputation and inverse-probability weighting. Biometrics, 68(1):129–137, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holland-Letz:2012:EAO

[403] Tim Holland-Letz, Holger Dette, and Didier Renard. Efficient algorithms for optimal designs with correlated observations in pharmacokinetics and dose-finding studies. Biometrics, 68(1):138–145, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bertrand:2012:SAA

[404] Julie Bertrand, Emmanuelle Comets, Marylore Chenel, and France Mentré. Some alternatives to asymptotic tests for the analysis of pharmacogenetic data using nonlinear mixed effects models. Biometrics, 68 (1):146–155, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2012:GFD

[405] Ying Yuan and Valen E. Johnson. Goodness-of-fit diagnostics for Bayesian hierarchical models. Biometrics, 68(1):156–164, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

StClair:2012:BME

[406] Katherine St. Clair and Daniel O’Connell. A Bayesian model for estimating population means using a link-tracing sampling design. Biometrics, 68(1):165–173, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gangnon:2012:LMA

[407] Ronald E. Gangnon. Local multiplicity adjustment for the spatial scan statistic using the Gumbel distribution. Biometrics, 68(1):174–182, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jones:2012:IBD

[408] M. C. Jones and Arthur Pewsey. Inverse Batschelet distributions for circular data. Biometrics, 68(1):183–193, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:ICE

[409] Xinlei Wang, Ke Wang, and Johan Lim. Isotonized CDF estimation from judgment poststratification data with empty strata. Biometrics, 68 (1):194–202, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Trippa:2012:BES

[410] Lorenzo Trippa, Gary L. Rosner, and Peter Müller. Bayesian enrichment strategies for randomized discontinuation trials. Biometrics, 68(1):203– 211, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schmidli:2012:D

[411] Heinz Schmidli and Beat Neuenschwander. Discussions. Biometrics, 68 (1):212–214, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gonen:2012:D

[412] Mithat Gönen. Discussions. Biometrics, 68(1):215–216, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perevozskaya:2012:D

[413] Inna Perevozskaya. Discussions. Biometrics, 68(1):217–218, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grieve:2012:D

[414] Andrew P. Grieve. Discussions. Biometrics, 68(1):219–223, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Trippa:2012:R

[415] Lorenzo Trippa, Gary L. Rosner, and Peter Müller. Rejoinder. Biometrics, 68(1):224–225, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2012:MAS

[416] Debashis Ghosh, Jeremy M. G. Taylor, and Daniel J. Sargent. Metaanalysis for surrogacy: Accelerated failure time models and semicompeting risks modeling. Biometrics, 68(1):226–232, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [417, 418] and rejoinder [421].

Molenberghs:2012:DCG

[417] Geert Molenberghs. Discussion contribution to 091037PR4 (Ghosh, Taylor, and Sargent). Biometrics, 68(1):233–235, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [416, 421].

Korn:2012:DPG

[418] Edward L. Korn. Discussion of the paper of Ghosh, Taylor, and Sargent. Biometrics, 68(1):236–238, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [416, 421].

Berger:2012:D

[419] Vance W. Berger, Grant Izmirlian, and Diana Knoll. Discussions. Biometrics, 68(1):239–241, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2012:DOF

[420] John O’Quigley and Philippe Flandre. Discussion by O’Quigley and Flandre. Biometrics, 68(1):242–244, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2012:RMA

[421] Debashis Ghosh, Jeremy M. G. Taylor, and Daniel J. Sargent. Rejoinder for “Meta-analysis for Surrogacy: Accelerated Failure Time Models and Semicompeting Risks Modeling”. Biometrics, 68(1):245–247, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [417].

Baker:2012:PTE

[422] Stuart G. Baker, Daniel J. Sargent, Marc Buyse, and Tomasz Burzykowski. Predicting treatment effect from surrogate endpoints and historical trials: an extrapolation involving probabilities of a binary outcome or survival to a specific time. Biometrics, 68(1):248–257, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Renfro:2012:BAT

[423] Lindsay A. Renfro, Bradley P. Carlin, and Daniel J. Sargent. Bayesian adaptive trial design for a newly validated surrogate endpoint. Biometrics, 68(1):258–267, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2012:BAM

[424] Peng Zhang, Juxin Liu, Jianghu Dong, Jelena L. Holovati, Brenda Letcher, and Locksley E. McGann. A Bayesian adjustment for multiplicative measurement errors for a calibration problem with application to a stem cell study. Biometrics, 68(1):268–274, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joffe:2012:GEA

[425] Marshall M. Joffe, Wei Peter Yang, and Harold Feldman. G-estimation and artificial censoring: Problems, challenges, and applications. Biometrics, 68(1):275–286, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McMahan:2012:IDS

[426] Christopher S. McMahan, Joshua M. Tebbs, and Christopher R. Bilder. Informative Dorfman screening. Biometrics, 68(1):287–296, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Whitmore:2012:CND

[427] G. A. Whitmore, Guangyu Zhang, and Mei-Ling Ting Lee. Constructing normalcy and discrepancy indexes for birth weight and gestational age using a threshold regression mixture model. Biometrics, 68(1):297–306, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2012:CLA

[428] Fang Han and Wei Pan. A composite likelihood approach to latent multivariate Gaussian modeling of SNP data with application to genetic association testing. Biometrics, 68(1):307–315, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daye:2012:HDH

[429] Z. John Daye, Jinbo Chen, and Hongzhe Li. High-dimensional heteroscedastic regression with an application to eQTL data analysis. Biometrics, 68(1):316–326, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Demetrio:2012:BRI

[430] Clarice G. B. Demétrio. Book review: Introduction to General and Generalized Linear Models, by Madsen, H. and Tyregod, P. Biometrics, 68 (1):327, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wittes:2012:BRS

[431] Janet T. Wittes. Book review: Sample Sizes for Clinical Trials, by Julious, S. A. Biometrics, 68(1):327–328, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wit:2012:BRL

[432] Ernst Wit. Book review: Learning and Inference in Computational Systems Biology, by Lawrence, N. D., Girolami, M., Rattray, M., and Sanguinetti, G. Biometrics, 68(1):328–329, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:BRA

[433] Anonymous. Book review: Applied Probability, by Lange, K. Biometrics, 68(1):329–330, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:BRC

[434] Anonymous. Book review: Common Statistical Methods for Clinical Research with SAS Examples, by Walker, G. A. and Shostak, J. Biometrics, 68(1):330, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:BRR

[435] Anonymous. Book review: Regression Estimators: a Comparative Study, by Gruber, M. H. J. Biometrics, 68(1):330, March 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staicu:2012:MFD

[436] Ana-Maria Staicu, Ciprian M. Crainiceanu, Daniel S. Reich, and David Ruppert. Modeling functional data with spatially heterogeneous shape characteristics. Biometrics, 68(2):331–343, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2012:NDB

[437] Hulin Wu, Hongqi Xue, and Arun Kumar. Numerical discretizationbased estimation methods for ordinary differential equation models via penalized spline smoothing with applications in biomedical research. Biometrics, 68(2):344–352, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:PGE

[438] Lan Wang, Jianhui Zhou, and Annie Qu. Penalized generalized estimating equations for high-dimensional longitudinal data analysis. Biometrics, 68(2):353–360, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhadra:2012:BSA

[439] Dhiman Bhadra, Michael J. Daniels, Sungduk Kim, Malay Ghosh, and Bhramar Mukherjee. A Bayesian semiparametric approach for incorporating longitudinal information on exposure history for inference in case-control studies. Biometrics, 68(2):361–370, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jones:2012:FJM

[440] G. Jones, W. O. Johnson, W. D. Vink, and N. French. A framework for the joint modeling of longitudinal diagnostic outcome data and latent infection status: Application to investigating the temporal relationship between infection and disease. Biometrics, 68(2):371–379, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Commenges:2012:CPE

[441] Daniel Commenges, Benoit Liquet, and Cécile Proust-Lima. Choice of prognostic estimators in joint models by estimating differences of expected conditional Kullback–Leibler risks. Biometrics, 68(2):380–387, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2012:EPA

[442] Yingye Zheng, Tianxi Cai, Yuying Jin, and Ziding Feng. Evaluating prognostic accuracy of biomarkers under competing risk. Biometrics, 68 (2):388–396, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shaw:2012:HRE

[443] Pamela A. Shaw and Ross L. Prentice. Hazard ratio estimation for biomarker-calibrated dietary exposures. Biometrics, 68(2):397–407, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dong:2012:ELC

[444] Bin Dong and David E. Matthews. Empirical likelihood for cumulative hazard ratio estimation with covariate adjustment. Biometrics, 68(2):

408–418, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yan:2012:MSC

[445] Jun Yan and Jian Huang. Model selection for Cox models with timevarying coefficients. Biometrics, 68(2):419–428, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2012:SFM

[446] Zhangsheng Yu, Xihong Lin, and Wanzhu Tu. Semiparametric frailty models for clustered failure time data. Biometrics, 68(2):429–436, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lian:2012:IPL

[447] Heng Lian, Xin Chen, and Jian-Yi Yang. Identification of partially linear structure in additive models with an application to gene expression prediction from sequences. Biometrics, 68(2):437–445, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thas:2012:OCA

[448] Olivier Thas, Lieven Clement, John C. W. Rayner, Beatriz Carvalho, and Wim Van Criekinge. An omnibus consistent adaptive percentile modified Wilcoxon rank sum test with applications in gene expression studies. Biometrics, 68(2):446–454, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dawson:2012:EBA

[449] John A. Dawson and Christina Kendziorski. An empirical Bayesian approach for identifying differential coexpression in high-throughput experiments. Biometrics, 68(2):455–465, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sanchez:2012:LVA

[450] Brisa N. Sánchez, Shan Kang, and Bhramar Mukherjee. A latent variable approach to study gene-environment interactions in the presence of multiple correlated exposures. Biometrics, 68(2):466–476, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2012:LAD

[451] Jingyuan Yang and Shili Lin. Likelihood approach for detecting imprinting and in utero maternal effects using general pedigrees from prospective family-based association studies. Biometrics, 68(2):477–485, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2012:PTR

[452] Oliver E. Lee and Thomas M. Braun. Permutation tests for random effects in linear mixed models. Biometrics, 68(2):486–493, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lebreton:2012:SET

[453] J. D. Lebreton, R. Choquet, and O. Gimenez. Simple estimation and test procedures in capture–mark–recapture mixed models. Biometrics, 68(2):494–503, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Okamura:2012:AEL

[454] Hiroshi Okamura, Shingo Minamikawa, Hans J. Skaug, and Toshiya Kishiro. Abundance estimation of long-diving animals using line transect methods. Biometrics, 68(2):504–513, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [621].

Grafstrom:2012:SBS

[455] Anton Grafström, Niklas L. P. Lundström, and Lina Schelin. Spatially balanced sampling through the pivotal method. Biometrics, 68(2):514– 520, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chan:2012:EIP

[456] Kwun Chuen Gary Chan and Mei-Cheng Wang. Estimating incident population distribution from prevalent data. Biometrics, 68(2):521–531, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lendle:2012:GTC

[457] Samuel D. Lendle, Michael G. Hudgens, and Bahjat F. Qaqish. Group testing for case identification with correlated responses. Biometrics, 68 (2):532–540, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Munoz:2012:PIC

[458] Iván Díaz Muñoz and Mark van der Laan. Population intervention causal effects based on stochastic interventions. Biometrics, 68(2):541–549, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leon-Novelo:2012:BSN

[459] L. G. Leon-Novelo, B. Nebiyou Bekele, P. Müller, F. Quintana, and K. Wathen. Borrowing strength with nonexchangeable priors over subpopulations. Biometrics, 68(2):550–558, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Spencer:2012:WDL

[460] Bruce D. Spencer. When do latent class models overstate accuracy for diagnostic and other classifiers in the absence of a gold standard? Biometrics, 68(2):559–566, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2012:BLS

[461] Ruixin Guo, Hongtu Zhu, Sy-Miin Chow, and Joseph G. Ibrahim. Bayesian lasso for semiparametric structural equation models. Biometrics, 68(2):567–577, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ibrahim:2012:BME

[462] Joseph G. Ibrahim, Ming-Hui Chen, H. Amy Xia, and Thomas Liu. Bayesian meta-experimental design: Evaluating cardiovascular risk in new antidiabetic therapies to treat type 2 diabetes. Biometrics, 68(2): 578–586, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Biswas:2012:LBL

[463] Swati Biswas and Shili Lin. Logistic Bayesian LASSO for identifying association with rare haplotypes and application to age-related macular degeneration. Biometrics, 68(2):587–597, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2012:EAR

[464] Nicholas G. Reich, Justin Lessler, Derek A. T. Cummings, and Ron Brookmeyer. Estimating absolute and relative case fatality ratios from infectious disease surveillance data. Biometrics, 68(2):598–606, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meyer:2012:STC

[465] Sebastian Meyer, Johannes Elias, and Michael Höhle. A space-time conditional intensity model for invasive meningococcal disease occurrence. Biometrics, 68(2):607–616, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:ECB

[466] Yue Wang, Robin Mogg, and Jared Lunceford. Evaluating correlationbased metric for surrogate marker qualification within a causal correlation framework. Biometrics, 68(2):617–627, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2012:OMM

[467] Dan Yang, Dylan S. Small, Jeffrey H. Silber, and Paul R. Rosenbaum. Optimal matching with minimal deviation from fine balance in a study of obesity and surgical outcomes. Biometrics, 68(2):628–636, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2012:CEM

[468] Dandan Liu, Douglas E. Schaubel, and John D. Kalbfleisch. Computationally efficient marginal models for clustered recurrent event data. Biometrics, 68(2):637–647, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLain:2012:JME

[469] Alexander C. McLain, Kirsten J. Lum, and Rajeshwari Sundaram. A joint mixed effects dispersion model for menstrual cycle length and timeto-pregnancy. Biometrics, 68(2):648–656, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oakes:2012:BRF

[470] David Oakes. Book review: Frailty Models in Survival Analysis, by Wienke, A. Biometrics, 68(2):657–658, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dobson:2012:BRL

[471] Annette J. Dobson. Book review: Logistic Regression Models, by Hilbe, J. M. Biometrics, 68(2):658, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oehlert:2012:BRD

[472] Gary W. Oehlert. Book review: Design of Experiments: an Introduction Based on Linear Models, by Morris, M. D. Biometrics, 68(2):658–659, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pearl:2012:BRB

[473] Judea Pearl. Book review: Bias and Causation, Models and Judgment for Valid Comparisons, by Weisberg, H. I. Biometrics, 68(2):659–660, June 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:CTC

[474] Anonymous. Correction to the table of contents for the September 2012 issue. Biometrics, 68(3):1, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:BEE

[475] Chi Wang, Giovanni Parmigiani, and Francesca Dominici. Bayesian effect estimation accounting for adjustment uncertainty. Biometrics, 68(3): 661–671, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gutman:2012:AIP

[476] R. Gutman and D. B. Rubin. Analyses that inform policy decisions. Biometrics, 68(3):671–675, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vansteelandt:2012:D

[477] Stijn Vansteelandt. Discussions. Biometrics, 68(3):675–678, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCandless:2012:DAU

[478] Lawrence C. McCandless. Discussion of adjustment uncertainty and propensity scores. Biometrics, 68(3):678–680, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:RBE

[479] Chi Wang, Giovanni Parmigiani, and Francesca Dominici. Rejoinder: Bayesian effect estimation accounting for adjustment uncertainty. Biometrics, 68(3):680–686, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2012:ATS

[480] Ying Huang, Peter B. Gilbert, and Holly Janes. Assessing treatmentselection markers using a potential outcomes framework. Biometrics, 68(3):687–696, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2012:CIQ

[481] Zhiwei Zhang, Zhen Chen, James F. Troendle, and Jun Zhang. Causal inference on quantiles with an obstetric application. Biometrics, 68(3): 697–706, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2012:EPS

[482] Yu-Jen Cheng and Mei-Cheng Wang. Estimating propensity scores and causal survival functions using prevalent survival data. Biometrics, 68(3): 707–716, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2012:NIM

[483] Hongwei Zhao, Chen Zuo, Shuai Chen, and Heejung Bang. Nonparametric inference for median costs with censored data. Biometrics, 68(3): 717–725, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:TCM

[484] Lu Wang, Pang Du, and Hua Liang. Two-component mixture cure rate model with spline estimated nonparametric components. Biometrics, 68(3):726–735, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schrodle:2012:AIM

[485] Birgit Schrödle, Leonhard Held, and Håvard Rue. Assessing the impact of a movement network on the spatiotemporal spread of infectious diseases. Biometrics, 68(3):736–744, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lok:2012:ITS

[486] Judith J. Lok and Victor DeGruttola. Impact of time to start treatment following infection with application to initiating HAART in HIV-positive patients. Biometrics, 68(3):745–754, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Groendyke:2012:NBA

[487] Chris Groendyke, David Welch, and David R. Hunter. A network-based analysis of the 1861 hagelloch measles data. Biometrics, 68(3):755–765, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwender:2012:RTS

[488] Holger Schwender, Margaret A. Taub, Terri H. Beaty, Mary L. Marazita, and Ingo Ruczinski. Rapid testing of SNPs and gene-environment interactions in case-parent trio data based on exact analytic parameter estimation. Biometrics, 68(3):766–773, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kuan:2012:IPK

[489] Pei Fen Kuan and Derek Y. Chiang. Integrating prior knowledge in multiple testing under dependence with applications to detecting differential DNA methylation. Biometrics, 68(3):774–783, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maadooliat:2012:AMP

[490] Mehdi Maadooliat, Jianhua Z. Huang, and Jianhua Hu. Analyzing multiple-probe microarray: Estimation and application of gene expression indexes. Biometrics, 68(3):784–792, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McMahan:2012:TDI

[491] Christopher S. McMahan, Joshua M. Tebbs, and Christopher R. Bilder. Two-dimensional informative array testing. Biometrics, 68(3):793–804, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Serban:2012:MFC

[492] Nicoleta Serban and Huijing Jiang. Multilevel functional clustering analysis. Biometrics, 68(3):805–814, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weiss:2012:CBM

[493] Christian H. Weiß and Philip K. Pollett. Chain binomial models and binomial autoregressive processes. Biometrics, 68(3):815–824, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Motta:2012:EFA

[494] Giovanni Motta and Hernando Ombao. Evolutionary factor analysis of replicated time series. Biometrics, 68(3):825–836, September 2012. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berrocal:2012:STD

[495] Veronica J. Berrocal, Alan E. Gelfand, and David M. Holland. Spacetime data fusion under error in computer model output: an application to modeling air quality. Biometrics, 68(3):837–848, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hund:2012:GAL

[496] Lauren Hund, Jarvis T. Chen, Nancy Krieger, and Brent A. Coull. A geostatistical approach to large-scale disease mapping with temporal misalignment. Biometrics, 68(3):849–858, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nieto-Barajas:2012:TSD

[497] Luis E. Nieto-Barajas, Peter Müller, Yuan Ji, Yiling Lu, and Gordon B. Mills. A time-series DDP for functional proteomics profiles. Biometrics, 68(3):859–868, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2012:SML

[498] Jinbo Chen, Dongyu Lin, and Hagit Hochner. Semiparametric maximum likelihood methods for analyzing genetic and environmental effects with case-control mother-child pair data. Biometrics, 68(3):869–877, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2012:ZFM

[499] Souparno Ghosh, Alan E. Gelfand, Kai Zhu, and James S. Clark. The k-ZIG: Flexible modeling for zero-inflated counts. Biometrics, 68(3): 878–885, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Englert:2012:IFE

[500] Stefan Englert and Meinhard Kieser. Improving the flexibility and efficiency of Phase II designs for oncology trials. Biometrics, 68(3):886–892, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bornkamp:2012:FUP

[501] Björn Bornkamp. Functional uniform priors for nonlinear modeling. Biometrics, 68(3):893–901, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prague:2012:TMH

[502] Mélanie Prague, Daniel Commenges, Julia Drylewicz, and Rodolphe Thiébaut. Treatment monitoring of HIV-infected patients based on mechanistic models. Biometrics, 68(3):902–911, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chao:2012:NLB

[503] Anne Chao and Chih-Wei Lin. Nonparametric lower bounds for species richness and shared species richness under sampling without replacement. Biometrics, 68(3):912–921, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zigler:2012:BAI

[504] Corwin M. Zigler and Thomas R. Belin. A Bayesian approach to improved estimation of causal effect predictiveness for a principal surrogate endpoint. Biometrics, 68(3):922–932, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2012:PBS

[505] Nanhua Zhang and Roderick J. Little. A pseudo-Bayesian shrinkage approach to regression with missing covariates. Biometrics, 68(3):933– 942, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2012:BSN

[506] Yangxin Huang and Getachew Dagne. Bayesian semiparametric nonlinear mixed-effects joint models for data with skewness, missing responses, and measurement errors in covariates. Biometrics, 68(3):943– 953, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2012:BIM

[507] Hongtu Zhu, Joseph G. Ibrahim, Yueh-Yun Chi, and Niansheng Tang. Bayesian influence measures for joint models for longitudinal and survival data. Biometrics, 68(3):954–964, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Higgs:2012:DAM

[508] Megan D. Higgs and Jay M. Ver Hoef. Discretized and aggregated: Modeling dive depth of harbor seals from ordered categorical data with temporal autocorrelation. Biometrics, 68(3):965–974, September 2012. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Todem:2012:EST

[509] David Todem, Wei-Wen Hsu, and KyungMann Kim. On the efficiency of score tests for homogeneity in two-component parametric models for discrete data. Biometrics, 68(3):975–982, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2012:MTC

[510] Shaowu Tang and Jong-Hyeon Jeong. Median tests for censored survival data; a contingency table approach. Biometrics, 68(3):983–989, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berger:2012:FPF

[511] Vance W. Berger. Out of the frying pan and in to the fire? Biometrics, 68(3):990, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proschan:2012:RBL

[512] Michael Proschan, Erica Brittain, and Lisa Kammerman. Reply to Berger’s “Letter to the Editor ‘Comment on Proschan et al. (2011), Out of the Frying Pan and in to the Fire?’ ”. Biometrics, 68(3):990–991, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2012:CNA

[513] Stuart G. Baker. Comment on Nie et al. (2011), Biometrics, early view. Biometrics, 68(3):992, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [514].

Nie:2012:RBL

[514] Hui Nie, Jing Cheng, and Dylan S. Small. Reply to Baker’s “Letter to the Editor ‘Comments on Nie et al. (2011), Biometrics, Early View’ ”. Biometrics, 68(3):992, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [513].

Anonymous:2012:CGS

[515] Anonymous. Correction to “Genetic Susceptibility to Prostate, Breast and Colorectal Cancer among Nordic Twins,” by Stuart G. Baker, Paul Lichtenstein, Jaakko Kaprio and Niels Holm; 61, 55–63, December 2005. Biometrics, 68(3):993, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [4].

Zhang:2012:DRS

[516] Min Zhang and Douglas E. Schaubel. Double-robust semiparametric estimator for differences in restricted mean lifetimes in observational studies. Biometrics, 68(4):999–1009, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2012:RME

[517] Baqun Zhang, Anastasios A. Tsiatis, Eric B. Laber, and Marie Davidian. A robust method for estimating optimal treatment regimes. Biometrics, 68(4):1010–1018, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comments [840].

Vansteelandt:2012:NDI

[518] Stijn Vansteelandt and Tyler J. VanderWeele. Natural direct and indirect effects on the exposed: Effect decomposition under weaker assumptions. Biometrics, 68(4):1019–1027, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daniels:2012:BIC

[519] Michael J. Daniels, Jason A. Roy, Chanmin Kim, Joseph W. Hogan, and Michael G. Perri. Bayesian inference for the causal effect of mediation. Biometrics, 68(4):1028–1036, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sobel:2012:CMM

[520] Michael E. Sobel and Bengt Muthén. Compliance mixture modelling with a zero-effect complier class and missing data. Biometrics, 68(4): 1037–1045, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2012:MSG

[521] Chung-Wei Shen and Yi-Hau Chen. Model selection for generalized estimating equations accommodating dropout missingness. Biometrics, 68 (4):1046–1054, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daniels:2012:BMS

[522] Michael J. Daniels, Arkendu S. Chatterjee, and Chenguang Wang. Bayesian model selection for incomplete data using the posterior predictive distribution. Biometrics, 68(4):1055–1063, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Montagna:2012:BLF

[523] Silvia Montagna, Surya T. Tokdar, Brian Neelon, and David B. Dunson. Bayesian latent factor regression for functional and longitudinal data. Biometrics, 68(4):1064–1073, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fu:2012:EER

[524] Liya Fu and You-Gan Wang. Efficient estimation for rank-based regression with clustered data. Biometrics, 68(4):1074–1082, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Skup:2012:MAM

[525] Martha Skup, Hongtu Zhu, and Heping Zhang. Multiscale adaptive marginal analysis of longitudinal neuroimaging data with time-varying covariates. Biometrics, 68(4):1083–1092, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2012:TVL

[526] Na Cai, Wenbin Lu, and Hao Helen Zhang. Time-varying latent effect model for longitudinal data with informative observation times. Biometrics, 68(4):1093–1102, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fong:2012:RBR

[527] Y. Fong, J. Wakefield, S. De Rosa, and N. Frahm. A robust Bayesian random effects model for nonlinear calibration problems. Biometrics, 68(4):1103–1112, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:TUF

[528] Yuanjia Wang and Huaihou Chen. On testing an unspecified function through a linear mixed effects model with multiple variance components. Biometrics, 68(4):1113–1125, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choi:2012:GCS

[529] Sangbum Choi and Xuelin Huang. A general class of semiparametric transformation frailty models for nonproportional hazards survival data. Biometrics, 68(4):1126–1135, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2012:SBS

[530] Jianchang Lin, Debajyoti Sinha, Stuart Lipsitz, and Adriano Polpo. Semiparametric Bayesian survival analysis using models with log-linear median. Biometrics, 68(4):1136–1145, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lostritto:2012:PDS

[531] Karen Lostritto, Robert L. Strawderman, and Annette M. Molinaro. A partitioning deletion/substitution/addition algorithm for creating survival risk groups. Biometrics, 68(4):1146–1156, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Warren:2012:STM

[532] Joshua Warren, Montserrat Fuentes, Amy Herring, and Peter Langlois. Spatial-temporal modeling of the association between air pollution exposure and preterm birth: Identifying critical Windows of exposure. Biometrics, 68(4):1157–1167, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2012:HSM

[533] Long Qu, Dan Nettleton, and Jack C. M. Dekkers. A hierarchical semiparametric model for incorporating intergene information for analysis of genomic data. Biometrics, 68(4):1168–1177, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2012:IEN

[534] Long Qu, Dan Nettleton, and Jack C. M. Dekkers. Improved estimation of the noncentrality parameter distribution from a large number of t-statistics, with applications to false discovery rate estimation in microarray data analysis. Biometrics, 68(4):1178–1187, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Favaro:2012:NED

[535] Stefano Favaro, Antonio Lijoi, and Igor Prünster. A new estimator of the discovery probability. Biometrics, 68(4):1188–1196, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [655].

Sun:2012:RGG

[536] Hokeun Sun and Hongzhe Li. Robust Gaussian graphical modeling via l_{1} penalization. Biometrics, 69(3):1197–1206, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2012:BMM

[537] Mei-Cheng Wang and Shanshan Li. Bivariate marker measurements and ROC analysis. Biometrics, 69(3):1207–1218, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2012:EPV

[538] Dandan Liu, Tianxi Cai, and Yingye Zheng. Evaluating the predictive value of biomarkers with stratified case-cohort design. Biometrics, 69(3): 1219–1227, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nguyen:2012:MCR

[539] P. Nguyen, P. E. Brown, and J. Stafford. Mapping cancer risk in Southwestern Ontario with changing census boundaries. Biometrics, 69(3): 1228–1237, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2012:HEM

[540] Yang Yang, Ira M. Longini, Jr., M. Elizabeth Halloran, and Valerie Obenchain. A hybrid EM and Monte Carlo EM algorithm and its application to analysis of transmission of infectious diseases. Biometrics, 69(3): 1238–1249, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mehrotra:2012:ALC

[541] Devan V. Mehrotra, Xiaoming Li, Jiajun Liu, and Kaifeng Lu. Analysis of longitudinal clinical trials with missing data using multiple imputation in conjunction with robust regression. Biometrics, 69(3):1250–1259, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2012:RCF

[542] Hongxiao Zhu, Philip J. Brown, and Jeffrey S. Morris. Robust classification of functional and quantitative image data using functional mixed models. Biometrics, 69(3):1260–1268, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piepho:2012:UTW

[543] H. P. Piepho, E. R. Williams, and L. V. Madden. The use of two-way linear mixed models in multitreatment meta-analysis. Biometrics, 69(3): 1269–1277, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2012:MME

[544] Han Chen, Alisa K. Manning, and Josée Dupuis. A method of moments estimator for random effect multivariate meta-analysis. Biometrics, 69 (3):1278–1284, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dendukuri:2012:BMA

[545] Nandini Dendukuri, Ian Schiller, Lawrence Joseph, and Madhukar Pai. Bayesian meta-analysis of the accuracy of a test for tuberculous pleuritis in the absence of a gold standard reference. Biometrics, 69(3):1285–1293, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2012:EDA

[546] Bo Zhang, Zhen Chen, and Paul S. Albert. Estimating diagnostic accuracy of raters without a gold standard by exploiting a group of experts. Biometrics, 69(3):1294–1302, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dorazio:2012:PGD

[547] Robert M. Dorazio. Predicting the geographic distribution of a species from presence-only data subject to detection errors. Biometrics, 69(3): 1303–1312, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deutsch:2012:BDP

[548] Roland C. Deutsch and Walter W. Piegorsch. Benchmark dose profiles for joint-action quantal data in quantitative risk assessment. Biometrics, 69(3):1313–1322, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giorgi:2012:NTS

[549] E. E. Giorgi and T. Bhattacharya. A note on two-sample tests for comparing intra-individual genetic sequence diversity between populations. Biometrics, 69(3):1323–1326, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [550].

Gilbert:2012:RNT

[550] Peter B. Gilbert, A. J. Rossini, and Raj Shankarappa. Rejoinder to “A Note on Two-Sample Tests for Comparing Intra-Individual Genetic Sequence Diversity Between Populations”. Biometrics, 69(3):1326, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [549].

Du:2012:BRS

[551] Pang Du. Book review: Smoothing Splines: Methods and Applications, by Wang, Y. Biometrics, 69(3):1327–1328, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christensen:2012:BRS

[552] Ole F. Christensen. Book review: Statistics for Spatio-Temporal Data, by Cressie, N. and Wikle, C. K. Biometrics, 69(3):1328–1329, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Noma:2012:BRB

[553] Hisashi Noma. Book review: Bayesian Modeling in Bioinformatics, edited by Dey, D. K., Ghosh, S. and Mallick, B. K. Biometrics, 69 (3):1329, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Espejo:2012:BRR

[554] Mariano Ruiz Espejo. Book review: Randomized Response and Indirect Questioning Techniques in Surveys, by Chaudhuri, A. Biometrics, 69(3): 1329–1330, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Trinca:2012:BRD

[555] Luzia Trinca. Book review: Designs and Analysis of Experiments, Volume 3: Special Designs and Applications, edited by Hinkelmann, K. Biometrics, 69(3):1330–1331, December 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2013:BRC

[556] Jaeyong Lee. Book review: Case Studies in Bayesian Statistical Modelling and Analysis, by Alston, C. L., Mengersen, K. L. and Pettitt, A. N. Biometrics, 68(4):798, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jang:2013:BRG

[557] Woncheol Jang. Book review: Generalized Estimating Equations, by Hardin, J. W. and Hilbe, J. M. Biometrics, 68(4):799, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [690].

Anonymous:2013:RE

[558] Anonymous. Report of the Editors — 2012. Biometrics, 69(1):vii–x, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bancroft:2013:EFD

[559] Tim Bancroft, Chuanlong Du, and Dan Nettleton. Estimation of false discovery rate using sequential permutation p-values. Biometrics, 69(1): 1–7, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Palacios:2013:GPB

[560] Julia A. Palacios and Vladimir N. Minin. Gaussian process-based Bayesian nonparametric inference of population size trajectories from gene genealogies. Biometrics, 69(1):8–18, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ren:2013:HFM

[561] Qian Ren and Sudipto Banerjee. Hierarchical factor models for large spatially misaligned data: a low-rank predictive process approach. Biometrics, 69(1):19–30, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giacofci:2013:WBC

[562] M. Giacofci, S. Lambert-Lacroix, G. Marot, and F. Picard. Waveletbased clustering for mixed-effects functional models in high dimension. Biometrics, 69(1):31–40, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goldsmith:2013:CCB

[563] J. Goldsmith, S. Greven, and C. Crainiceanu. Corrected confidence bands for functional data using principal components. Biometrics, 69(1):41–51, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Altstein:2013:LSA

[564] L. Altstein and G. Li. Latent subgroup analysis of a randomized clinical trial through a semiparametric accelerated failure time mixture model. Biometrics, 69(1):52–61, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2013:RAC

[565] Rena Jie Sun and John D. Kalbfleisch. A risk-adjusted O − E CUSUM with monitoring bands for monitoring medical outcomes. Biometrics, 69(1):62–69, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Post:2013:FSS

[566] Justin B. Post and Howard D. Bondell. Factor selection and structural identification in the interaction ANOVA model. Biometrics, 69(1):70– 79, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zucker:2013:RCS

[567] David M. Zucker, Malka Gorfine, Yi Li, Mahlet G. Tadesse, and Donna Spiegelman. A regularization corrected score method for nonlinear regression models with covariate error. Biometrics, 69(1):80–90, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2013:CAE

[568] Danping Liu and Xiao-Hua Zhou. Covariate adjustment in estimating the area under ROC curve with partially missing gold standard. Biometrics, 69(1):91–100, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Porter:2013:PSS

[569] Aaron T. Porter and Jacob J. Oleson. A path-specific SEIR model for use with general latent and infectious time distributions. Biometrics, 69 (1):101–108, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joly:2013:PPC

[570] Pierre Joly, Célia Touraine, Aurore Georget, Jean-François Dartigues, Daniel Commenges, and Hélène Jacqmin-Gadda. Prevalence projections of chronic diseases and impact of public health intervention. Biometrics, 69(1):109–117, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2013:IMM

[571] Paul R. Rosenbaum. Impact of multiple matched controls on design sensitivity in observational studies. Biometrics, 69(1):118–127, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Simpson:2013:PED

[572] Shawn E. Simpson. A positive event dependence model for self-controlled case series with applications in postmarketing surveillance. Biometrics, 69(1):128–136, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gijbels:2013:THM

[573] Irène Gijbels and Marek Omelka. Testing for homogeneity of multivariate dispersions using dissimilarity measures. Biometrics, 69(1):137–145, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dail:2013:EOP

[574] D. Dail and L. Madsen. Estimating open population site occupancy from presence-absence data lacking the robust design. Biometrics, 69(1):146– 156, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zaslavsky:2013:BHT

[575] Boris G. Zaslavsky. Bayesian hypothesis testing in two-arm trials with dichotomous outcomes. Biometrics, 69(1):157–163, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Majumdar:2013:NBS

[576] Arunabha Majumdar, Sourabh Bhattacharya, Analabha Basu, and Saurabh Ghosh. A novel Bayesian semiparametric algorithm for inferring population structure and adjusting for case-control association tests. Biometrics, 69(1):164–173, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leon-Novelo:2013:SBI

[577] Luis G. León-Novelo, Peter Müller, Wadih Arap, Mikhail Kolonin, Jessica Sun, Renata Pasqualini, and Kim-Anh Do. Semiparametric Bayesian inference for phage display data. Biometrics, 69(1):174–183, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jeong:2013:WBB

[578] Jaesik Jeong, Marina Vannucci, and Kyungduk Ko. A wavelet-based Bayesian approach to regression models with long memory errors and its application to fMRI data. Biometrics, 69(1):184–196, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heagerty:2013:ELA

[579] Patrick J. Heagerty and Bryan A. Comstock. Exploration of lagged associations using longitudinal data. Biometrics, 69(1):197–205, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:2013:RTI

[580] Jeremy M. G. Taylor, Yongseok Park, Donna P. Ankerst, Cecile Proust-Lima, Scott Williams, Larry Kestin, Kyoungwha Bae, Tom Pickles, and Howard Sandler. Real-time individual predictions of prostate cancer recurrence using joint models. Biometrics, 69(1):206–213, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2013:MSS

[581] Yuda Zhu and Robert E. Weiss. Modeling seroadaptation and sexual behavior among HIV^{+} study participants with a simultaneously multilevel and multivariate longitudinal count model. Biometrics, 69(1):214–224, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandel:2013:ETD

[582] Micha Mandel, Francois Mercier, Benjamin Eckert, Peter Chin, and Rebecca A. Betensky. Estimating time to disease progression comparing transition models and survival methods — an analysis of multiple sclerosis data. Biometrics, 69(1):225–234, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Auranen:2013:ESS

[583] Kari Auranen, Hanna Rinta-Kokko, and M. Elizabeth Halloran. Estimating strain-specific and overall efficacy of polyvalent vaccines against recurrent pathogens from a cross-sectional study. Biometrics, 69(1):235– 244, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hubbard:2013:SCB

[584] Rebecca A. Hubbard and Diana L. Miglioretti. A semiparametric censoring bias model for estimating the cumulative risk of a false-positive screening test under dependent censoring. Biometrics, 69(1):245–253, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See letter [689].

Gruber:2013:ATM

[585] Susan Gruber and Mark J. van der Laan. An application of targeted maximum likelihood estimation to the meta-analysis of safety data. Biometrics, 69(1):254–262, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zigler:2013:MFB

[586] Corwin M. Zigler, Krista Watts, Robert W. Yeh, Yun Wang, Brent A. Coull, and Francesca Dominici. Model feedback in Bayesian propensity score estimation. Biometrics, 69(1):263–273, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Renner:2013:EMP

[587] Ian W. Renner and David I. Warton. Equivalence of MAXENT and Poisson point process models for species distribution modeling in ecology. Biometrics, 69(1):274–281, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2013:NUR

[588] Jane Paik Kim. A note on using regression models to analyze randomized trials: Asymptotically valid hypothesis tests despite incorrectly specified models. Biometrics, 69(1):282–289, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [589].

Rosenblum:2013:RNU

[589] Michael Rosenblum and Mark J. van der Laan. Rejoinder to “A Note on Using Regression Models to Analyze Randomized Trials: Asymptotically Valid Hypothesis Tests Despite Incorrectly Specified Models”. Biometrics, 69(1):290, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [588].

Zhu:2013:BRSa

[590] Hongtu Zhu. Book review: Statistics of Medical Imaging, by Lei, T. Biometrics, 69(1):291–292, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Assuncao:2013:BRB

[591] Renato Assunção. Book review: Bayesian Smoothing and Regression for Longitudinal, Spatial and Event History Data, by Fahrmeir, L. and Kneib, T. Biometrics, 69(1):292, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rotnitzky:2013:BRT

[592] Andrea Rotnitzky. Book review: Targeted Learning, by van der Laan, M. and Rose, S. Biometrics, 69(1):293, March 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Demetrio:2013:PAX

[593] C. G. B. Demétrio. Presidential address: XXVI international biometric conference, kobe, Japan, August 2012 education in developing countries ’ the role of IBS. Biometrics, 69(2):295–299, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Basford:2013:ITO

[594] Kaye E. Basford. IBS: Transformation of our governance. Biometrics, 69(2):300, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2013:DEE

[595] Ying Huang, Peter B. Gilbert, and Julian Wolfson. Design and estimation for evaluating principal surrogate markers in vaccine trials. Biometrics, 69(2):301–309, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lendle:2013:IEE

[596] Samuel D. Lendle, Meenakshi S. Subbaraman, and Mark J. van der Laan. Identification and efficient estimation of the natural direct effect among the untreated. Biometrics, 69(2):310–317, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2013:CMJ

[597] Zhiwei Zhang, Richard M. Kotz, Chenguang Wang, Shiling Ruan, and Martin Ho. A causal model for joint evaluation of placebo and treatmentspecific effects in clinical trials. Biometrics, 69(2):318–327, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Juraska:2013:MSH

[598] M. Juraska and P. B. Gilbert. Mark-specific hazard ratio model with multivariate continuous marks: an application to vaccine efficacy. Biometrics, 69(2):328–337, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gong:2013:PCE

[599] Qi Gong and Douglas E. Schaubel. Partly conditional estimation of the effect of a time-dependent factor in the presence of dependent censoring. Biometrics, 69(2):338–347, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lian:2013:PLS

[600] Heng Lian, Peng Lai, and Hua Liang. Partially linear structure selection in Cox models with varying coefficients. Biometrics, 69(2):348–357, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2013:WLS

[601] Lili Yu, Liang Liu, and Ding-Geng (Din) Chen. Weighted least-squares method for right-censored data in accelerated failure time model. Biometrics, 69(2):358–365, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kalbfleisch:2013:EFA

[602] John D. Kalbfleisch, Douglas E. Schaubel, Yining Ye, and Qi Gong. An estimating function approach to the analysis of recurrent and terminal events. Biometrics, 69(2):366–374, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2013:QRR

[603] Xianghua Luo, Chiung-Yu Huang, and Lan Wang. Quantile regression for recurrent gap time data. Biometrics, 69(2):375–385, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2013:GPL

[604] Baojiang Chen and Xiao-Hua Zhou. Generalized partially linear models for incomplete longitudinal data in the presence of population-level information. Biometrics, 69(2):386–395, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lopez-Pintado:2013:RGS

[605] Sara López-Pintado and Ian W. McKeague. Recovering gradients from sparsely observed functional data. Biometrics, 69(2):396–404, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schildcrout:2013:OVD

[606] Jonathan S. Schildcrout, Shawn P. Garbett, and Patrick J. Heagerty. Outcome vector dependent sampling with longitudinal continuous response data: Stratified sampling based on summary statistics. Biometrics, 69(2):405–416, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2013:JRA

[607] Colin O. Wu, Gang Zheng, and Minjung Kwak. A joint regression analysis for genetic association studies with outcome stratified samples. Biometrics, 69(2):417–426, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Acar:2013:GKW

[608] Elif F. Acar and Lei Sun. A generalized Kruskal-Wallis test incorporating group uncertainty with application to genetic association studies. Biometrics, 69(2):427–435, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Khalili:2013:RFM

[609] Abbas Khalili and Shili Lin. Regularization in finite mixture of regression models with diverging number of parameters. Biometrics, 69(2):436– 446, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhadra:2013:JHD

[610] Anindya Bhadra and Bani K. Mallick. Joint high-dimensional Bayesian variable and covariance selection with an application to eQTL analysis. Biometrics, 69(2):447–457, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Overstall:2013:SBI

[611] Antony M. Overstall and David C. Woods. A strategy for Bayesian inference for computationally expensive models with application to the estimation of stem cell properties. Biometrics, 69(2):458–468, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ross:2013:BIT

[612] Michelle Ross and Jon Wakefield. Bayesian inference for two-phase studies with categorical covariates. Biometrics, 69(2):469–477, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Altomare:2013:OBS

[613] Davide Altomare, Guido Consonni, and Luca La Rocca. Objective Bayesian search of Gaussian directed acyclic graphical models for ordered variables with non-local priors. Biometrics, 69(2):478–487, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Besag:2013:EGF

[614] J. Besag and D. Mondal. Exact goodness-of-fit tests for Markov chains. Biometrics, 69(2):488–496, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chiu:2013:STS

[615] Sung Nok Chiu and Kwong Ip Liu. Stationarity tests for spatial point processes using discrepancies. Biometrics, 69(2):497–507, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schorgendorfer:2013:BGF

[616] Angela Schörgendorfer, Adam J. Branscum, and Timothy E. Hanson. A Bayesian goodness of fit test and semiparametric generalization of logistic regression with measurement data. Biometrics, 69(2):508–519, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mohammed:2013:NHT

[617] Sandra M. Mohammed, Lorien S. Dalrymple, Damla Sentürk, and Danh V. Nguyen. Naive hypothesis testing for case series analysis with time-varying exposure onset measurement error: Inference for infectioncardiovascular risk in patients on dialysis. Biometrics, 69(2):520–529, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2013:RSS

[618] Fang Liu. A revisit to sample size and power calculations for testing odds ratio in two independent binomials. Biometrics, 69(2):530–536, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rushworth:2013:DLM

[619] Alastair M. Rushworth, Adrian W. Bowman, Mark J. Brewer, and Simon J. Langan. Distributed lag models for hydrological data. Biometrics, 69(2):537–544, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boehm:2013:BCM

[620] Laura Boehm, Brian J. Reich, and Dipankar Bandyopadhyay. Bridging conditional and marginal inference for spatially referenced binary data. Biometrics, 69(2):545–554, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2013:CAE

[621] Anonymous. Correction to “Abundance Estimation of Long-Diving Animals Using Line Transect Methods,” by Hiroshi Okamura, Shingo Minamikawa, Hans J. Skaug and Toshiya Kishiro; 68, 504–513, DOI: 10.1111/j.1541-0420.2011.01689.x June 2012. Biometrics, 69(2):555, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [454].

Anonymous:2013:CBM

[622] Anonymous. Correction to “Bayesian Monitoring of Event Rates with Censored Data,” by D. A. Follmann and P. S. Albert; 55, 603–607, DOI: 10.1111/j.0006-341X.1999.00603.x June 1999. Biometrics, 69(2): 556, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1].

Zhu:2013:BRSb

[623] Jun Zhu. Book review: Spatio-Temporal Heterogeneity: Concepts and Analyses, by Dutilleul, P. R. L. Biometrics, 69(2):557–558, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tusell:2013:BRP

[624] Fernando Tusell. Book review: Programming Graphical User Interfaces, by Lawrence, M. L. and Verzani, J. Biometrics, 69(2):558–559, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Solari:2013:BRS

[625] Aldo Solari. Book review: Simultaneous Inference in Regression, by Liu, W. Biometrics, 69(2):559–560, June 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanderWeele:2013:SMC

[626] Tyler J. VanderWeele. Surrogate measures and consistent surrogates. Biometrics, 69(3):561–565, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [627, 628, 629].

Elliott:2013:DSM

[627] Michael R. Elliott, Anna Conlon, and Yun Li. Discussion on “Surrogate Measures and Consistent Surrogates”. Biometrics, 69(3):565–569, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [626].

Joffe:2013:DSM

[628] Marshall M. Joffe. Discussion on “Surrogate Measures and Consistent Surrogates”. Biometrics, 69(3):569–573, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [626].

Pearl:2013:DSM

[629] Judea Pearl. Discussion on “Surrogate Measures and Consistent Surrogates”. Biometrics, 69(3):573–577, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [626].

VanderWeele:2013:R

[630] Tyler J. VanderWeele. Rejoinder. Biometrics, 69(3):577–581, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2013:NBP

[631] Sunkyung Kim, Wei Pan, and Xiaotong Shen. Network-based penalized regression with application to genomic data. Biometrics, 69(3):582–593, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Si:2013:OTM

[632] Yaqing Si and Peng Liu. An optimal test with maximum average power while controlling FDR with application to RNA-Seq data. Biometrics, 69(3):594–605, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Farcomeni:2013:FCP

[633] A. Farcomeni and L. Finos. FDR control with pseudo-gatekeeping based on a possibly data driven order of the hypotheses. Biometrics, 69(3): 606–613, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boca:2013:DTA

[634] Simina M. Boca, Héctor Corrada Bravo, Brian Caffo, Jeffrey T. Leek, and Giovanni Parmigiani. A decision-theory approach to interpretable set analysis for high-dimensional data. Biometrics, 69(3):614–623, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leung:2013:SEL

[635] Denis Heng-Yan Leung, Dylan S. Small, Jing Qin, and Min Zhu. Shrinkage empirical likelihood estimator in longitudinal analysis with timedependent covariates-application to modeling the health of Filipino children. Biometrics, 69(3):624–632, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Touloumis:2013:GMR

[636] Anestis Touloumis, Alan Agresti, and Maria Kateri. GEE for multinomial responses using a local odds ratios parameterization. Biometrics, 69(3):633–640, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Freue:2013:NRO

[637] Gabriela V. Cohen Freue, Hernan Ortiz-Molina, and Ruben H. Zamar. A natural robustification of the ordinary instrumental variables estimator. Biometrics, 69(3):641–650, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2013:BQR

[638] Brian J. Reich and Luke B. Smith. Bayesian quantile regression for censored data. Biometrics, 69(3):651–660, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McEvoy:2013:BAC

[639] Bradley W. McEvoy, Rajesh R. Nandy, and Ram C. Tiwari. Bayesian approach for clinical trial safety data using an Ising prior. Biometrics, 69(3):661–672, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2013:UJU

[640] Peter F. Thall, Hoang Q. Nguyen, Thomas M. Braun, and Muzaffar H. Qazilbash. Using joint utilities of the times to response and toxicity to adaptively optimize schedule-dose regimes. Biometrics, 69(3):673–682, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cui:2013:MPI

[641] Na Cui, Yuguo Chen, and Dylan S. Small. Modeling parasite infection dynamics when there is heterogeneity and imperfect detectability. Biometrics, 69(3):683–692, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bernau:2013:COR

[642] Christoph Bernau, Thomas Augustin, and Anne-Laure Boulesteix. Correcting the optimal resampling-based error rate by estimating the error rate of wrapper algorithms. Biometrics, 69(3):693–702, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borchers:2013:UHM

[643] D. L. Borchers, W. Zucchini, M. P. Heide-Jørgensen, A. Cañadas, and R. Langrock. Using hidden Markov models to deal with availability bias on line transect surveys. Biometrics, 69(3):703–713, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chakraborty:2013:IOD

[644] Bibhas Chakraborty, Eric B. Laber, and Yingqi Zhao. Inference for optimal dynamic treatment regimes using an adaptive m-out-of-n bootstrap scheme. Biometrics, 69(3):714–723, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2013:TBT

[645] Changyu Shen, Jaesik Jeong, Xiaochun Li, Peng-Sheng Chen, and Alfred Buxton. Treatment benefit and treatment harm rate to characterize heterogeneity in treatment effect. Biometrics, 69(3):724–731, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2013:MCE

[646] Brent A. Johnson, Heather Ribaudo, Roy M. Gulick, and Joseph J. Eron, Jr. Modeling clinical endpoints as a function of time of switch to secondline ART with incomplete data on switching times. Biometrics, 69(3): 732–740, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shotwell:2013:MAC

[647] M. S. Shotwell, K. J. Drake, V. Y. Sidorov, and J. P. Wikswo. Mechanistic analysis of challenge-response experiments. Biometrics, 69(3): 741–747, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Numminen:2013:ETD

[648] Elina Numminen, Lu Cheng, Mats Gyllenberg, and Jukka Corander. Estimating the transmission dynamics of Streptococcus pneumoniae from strain prevalence data. Biometrics, 69(3):748–757, September 2013. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2013:PSE

[649] Chang Xuan Mao, Nan Yang, and Jinhua Zhong. On population size estimators in the Poisson mixture model. Biometrics, 69(3):758–765, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonner:2013:MRM

[650] Simon J. Bonner and Jason Holmberg. Mark-recapture with multiple, non-invasive marks. Biometrics, 69(3):766–775, September 2013. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robertson:2013:BBA

[651] B. L. Robertson, J. A. Brown, T. McDonald, and P. Jaksons. BAS: Balanced acceptance sampling of natural resources. Biometrics, 69(3): 776–784, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Woodard:2013:LFR

[652] D. B. Woodard, T. M. T. Love, S. W. Thurston, D. Ruppert, S. Sathyanarayana, and S. H. Swan. Latent factor regression models for grouped outcomes. Biometrics, 69(3):785–794, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bove:2013:CCB

[653] Daniel Sabanés Bové and Leonhard Held. Comment on Cai and Betensky (2003), On the Poisson Approximation for Hazard Regression. Biometrics, 69(3):795, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [3, 654].

Cai:2013:RSB

[654] Tianxi Cai and Rebecca Betensky. Reply to Sabanés Bové and Held’s “Comment on Cai and Betensky (2003), On the Poisson Approximation for Hazard Regression”. Biometrics, 69(3):796, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [653].

Anonymous:2013:CNE

[655] Anonymous. Correction to “A New Estimator of the Discovery Probability” by Stefano Favaro, Antonio Lijoi, and Igor Prünster; 68, 1188– 1196, December 2012. Biometrics, 69(3):797, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [535].

Park:2013:BRM

[656] Mira Park. Book review: Medical Biostatistics, 3rd edition by Indrayan, Abhaya. Biometrics, 69(3):1086–1087, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oh:2013:BRI

[657] Hee-Seok Oh. Book review: Introduction to Linear Regression Analysis, 5th edition by Montgomery, Douglas C., Peck, Elizabeth A., and Vining, G. Geoffrey. Biometrics, 69(3):1087, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shin:2013:BRS

[658] Dong Wan Shin. Book review: Statistical Methods with Applications to Demography and Life Insurance, by Khmaladze, Estate V. Biometrics, 69(3):1087–1088, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2013:BRC

[659] Donguk Kim. Book review: Categorical Data Analysis, 3rd edition, by Agresti, Alan. Biometrics, 69(3):1088–1089, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lim:2013:BRS

[660] Johan Lim. Book review: Spatio-Temporal Design: Advances in Efficient Data Acquisition, by Mateu, J. and Müller, W. G. Biometrics, 69(4):800, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lopes:2013:BRH

[661] Hedibert Freitas Lopes. Book review: Handbook of Markov Chain Monte Carlo, by Brooks, S., Gelman, A., Jones, G. L. and Meng, X. Biometrics, 69(4):801, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsu:2013:CSA

[662] Jesse Y. Hsu and Dylan S. Small. Calibrating sensitivity analyses to observed covariates in observational studies. Biometrics, 69(4):803–811, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Long:2013:SBP

[663] Dustin M. Long and Michael G. Hudgens. Sharpening bounds on principal effects with covariates. Biometrics, 69(4):812–819, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vock:2013:ACE

[664] David M. Vock, Anastasios A. Tsiatis, Marie Davidian, Eric B. Laber, Wayne M. Tsuang, C. Ashley Finlen Copeland, and Scott M. Palmer. Assessing the causal effect of organ transplantation on the distribution of residual lifetime. Biometrics, 69(4):820–829, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bai:2013:DRE

[665] Xiaofei Bai, Anastasios A. Tsiatis, and Sean M. O’Brien. Doubly-robust estimators of treatment-specific survival distributions in observational studies with stratified sampling. Biometrics, 69(4):830–839, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diao:2013:ESE

[666] Guoqing Diao, Donglin Zeng, and Song Yang. Efficient semiparametric estimation of short-term and long-term hazard ratios with right-censored data. Biometrics, 69(4):840–849, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2013:BSA

[667] Nan Xuan Lin, Stuart Logan, and William Edward Henley. Bias and sensitivity analysis when estimating treatment effects from the Cox model with omitted covariates. Biometrics, 69(4):850–860, December 2013. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinnott:2013:ORA

[668] Jennifer A. Sinnott and Tianxi Cai. Omnibus risk assessment via accelerated failure time kernel machine modeling. Biometrics, 69(4):861–873, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2013:NAM

[669] Ying Guo, Ruosha Li, Limin Peng, and Amita K. Manatunga. New agreement measures based on survival processes. Biometrics, 69(4):874– 882, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2013:LST

[670] Long Qu, Tobias Guennel, and Scott L. Marshall. Linear score tests for variance components in linear mixed models and applications to genetic association studies. Biometrics, 69(4):883–892, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Simpson:2013:MSC

[671] Shawn E. Simpson, David Madigan, Ivan Zorych, Martijn J. Schuemie, Patrick B. Ryan, and Marc A. Suchard. Multiple self-controlled case series for large-scale longitudinal observational databases. Biometrics, 69(4):893–902, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Serban:2013:MCD

[672] Nicoleta Serban, Ana-Maria Staicu, and Raymond J. Carroll. Multilevel cross-dependent binary longitudinal data. Biometrics, 69(4):903–913, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ahn:2013:BLC

[673] Jaeil Ahn, Suyu Liu, Wenyi Wang, and Ying Yuan. Bayesian latentclass mixed-effect hybrid models for dyadic longitudinal data with nonignorable dropouts. Biometrics, 69(4):914–924, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2013:BAT

[674] Shi Li, Bhramar Mukherjee, Stuart Batterman, and Malay Ghosh. Bayesian analysis of time-series data under case-crossover designs: Posterior equivalence and inference. Biometrics, 69(4):925–936, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drovandi:2013:BED

[675] Christopher C. Drovandi and Anthony N. Pettitt. Bayesian experimental design for models with intractable likelihoods. Biometrics, 69(4):937– 948, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2013:RRM

[676] Zhenzhen Xu and John D. Kalbfleisch. Repeated randomization and matching in multi-arm trials. Biometrics, 69(4):949–959, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shao:2013:VTU

[677] Jun Shao and Xinxin Yu. Validity of tests under covariate-adaptive biased coin randomization and generalized linear models. Biometrics, 69(4):960–969, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2013:HMP

[678] Ying Guo and Li Tang. A hierarchical model for probabilistic independent component analysis of multi-subject fMRI studies. Biometrics, 69(4):970–981, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davidov:2013:TMS

[679] Ori Davidov and Shyamal Peddada. Testing for the multivariate stochastic order among ordered experimental groups with application to doseresponse studies. Biometrics, 69(4):982–990, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pledger:2013:BRT

[680] Shirley Pledger, Edward Baker, and Kim Scribner. Breeding return times and abundance in capture–recapture models. Biometrics, 69(4): 991–1001, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cattelan:2013:HPL

[681] Manuela Cattelan and Cristiano Varin. Hybrid pairwise likelihood analysis of animal behavior experiments. Biometrics, 69(4):1002–1011, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schofield:2013:MIS

[682] Matthew R. Schofield, Richard J. Barker, and Peter Taylor. Modeling individual specific fish length from capture–recapture data using the von Bertalanffy growth curve. Biometrics, 69(4):1012–1021, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mitchell:2013:CMC

[683] Shira Mitchell, Al Ozonoff, Alan M. Zaslavsky, Bethany Hedt-Gauthier, Kristian Lum, and Brent A. Coull. A comparison of marginal and conditional models for capture–recapture data with application to human rights violations data. Biometrics, 69(4):1022–1032, December 2013. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohning:2013:GCE

[684] Dankmar Böhning, Alberto Vidal-Diez, Rattana Lerdsuwansri, Chukiat Viwatwongkasem, and Mark Arnold. A generalization of Chao’s estimator for covariate information. Biometrics, 69(4):1033–1042, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nicolaie:2013:DPO

[685] M. A. Nicolaie, J. C. van Houwelingen, T. M. de Witte, and H. Putter. Dynamic pseudo-observations: a robust approach to dynamic prediction in competing risks. Biometrics, 69(4):1043–1052, December 2013. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xia:2013:LNM

[686] Fan Xia, Jun Chen, Wing Kam Fung, and Hongzhe Li. A logistic normal multinomial regression model for microbiome compositional data analysis. Biometrics, 69(4):1053–1063, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tebbs:2013:TSH

[687] Joshua M. Tebbs, Christopher S. McMahan, and Christopher R. Bilder. Two-stage hierarchical group testing for multiple infections with application to the infertility prevention project. Biometrics, 69(4):1064–1073, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miranda:2013:BST

[688] Michelle F. Miranda, Hongtu Zhu, and Joseph G. Ibrahim. Bayesian spatial transformation models with applications in neuroimaging data. Biometrics, 69(4):1074–1083, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2013:LEC

[689] Stuart G. Baker. Letter to the Editor: “Comment on Hubbard and Miglioretti (2013), Consider Also a Selection Model for the Cumulative Risk of False Positive Screening Tests”. Biometrics, 69(4):1084, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [584].

Anonymous:2013:CBR

[690] Anonymous. Correction to “Book Review of ‘Generalized Estimating Equations”’ by Woncheol Jang; 69, 799, September 2013. Biometrics, 69(4):1085, December 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [557].

Weinberg:2014:CLM

[691] Clarice R. Weinberg and David M. Umbach. Correction to “Linear mixed function-on-function regression models,” by Wei Wang; 70, 794– 801, December 2014. Biometrics, 55(3):1062, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [783].

Anonymous:2014:RE

[692] Anonymous. Report of the Editors — 2013. Biometrics, 70(1):vii–ix, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hudgens:2014:PLI

[693] Michael G. Hudgens, Chenxi Li, and Jason P. Fine. Parametric likelihood inference for interval censored competing risks data. Biometrics, 70(1): 1–9, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ning:2014:ETD

[694] Jing Ning and Karen Bandeen-Roche. Estimation of time-dependent association for bivariate failure times in the presence of a competing risk. Biometrics, 70(1):10–20, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:2014:SAL

[695] Samiran Sinha and Yanyuan Ma. Semiparametric analysis of linear transformation models with covariate measurement errors. Biometrics, 70(1): 21–32, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Das:2014:SAS

[696] Kiranmoy Das and Michael J. Daniels. A semiparametric approach to simultaneous covariance estimation for bivariate sparse longitudinal data. Biometrics, 70(1):33–43, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neuhaus:2014:LBA

[697] John M. Neuhaus, Alastair J. Scott, Christopher J. Wild, Yannan Jiang, Charles E. McCulloch, and Ross Boylan. Likelihood-based analysis of longitudinal data from outcome-related sampling designs. Biometrics, 70(1):44–52, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laber:2014:SVD

[698] Eric B. Laber, Daniel J. Lizotte, and Bradley Ferguson. Set-valued dynamic treatment regimes for competing outcomes. Biometrics, 70(1):

53–61, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daniels:2014:FBI

[699] M. J. Daniels, C. Wang, and B. H. Marcus. Fully Bayesian inference under ignorable missingness in the presence of auxiliary covariates. Biometrics, 70(1):62–72, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2014:BMS

[700] Xiaoquan Wen. Bayesian model selection in complex linear systems, as illustrated in genetic association studies. Biometrics, 70(1):73–83, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2014:BEH

[701] Jing Cao and Song Zhang. A Bayesian extension of the hypergeometric test for functional enrichment analysis. Biometrics, 70(1):84–94, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fronczyk:2014:BAA

[702] Kassandra Fronczyk and Athanasios Kottas. A Bayesian approach to the analysis of quantal bioassay studies using nonparametric mixture models. Biometrics, 70(1):95–102, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wolfsegger:2014:HME

[703] Martin J. Wolfsegger, Georg Gutjahr, Werner Engl, and Thomas Jaki. A hybrid method to estimate the minimum effective dose for monotone and non-monotone dose-response relationships. Biometrics, 70(1):103– 109, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stoklosa:2014:FFS

[704] Jakub Stoklosa, Heloise Gibb, and David I. Warton. Fast forward selection for generalized estimating equations with a large number of predictor variables. Biometrics, 70(1):110–120, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2014:SLM

[705] Jian Zhang, Chao Liu, and Gary Green. Source localization with MEG data: a beamforming approach based on covariance thresholding. Biometrics, 70(1):121–131, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nathoo:2014:VBS

[706] F. S. Nathoo, A. Babul, A. Moiseev, N. Virji-Babul, and M. F. Beg. A variational Bayes spatiotemporal model for electromagnetic brain mapping. Biometrics, 70(1):132–143, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schnitzer:2014:MIH

[707] Mireille E. Schnitzer, Erica E. M. Moodie, Mark J. van der Laan, Robert W. Platt, and Marina B. Klein. Modeling the impact of hepatitis C viral clearance on end-stage liver disease in an HIV co-infected cohort with targeted maximum likelihood estimation. Biometrics, 70 (1):144–152, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qian:2014:VSP

[708] Jing Qian, Seyedmehdi Payabvash, André Kemmling, Michael H. Lev, Lee H. Schwamm, and Rebecca A. Betensky. Variable selection and prediction using a nested, matched case-control study: Application to hospital acquired pneumonia in stroke patients. Biometrics, 70(1):153– 163, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Szpiro:2014:EAA

[709] Adam A. Szpiro, Lianne Sheppard, Sara D. Adar, and Joel D. Kaufman. Estimating acute air pollution health effects from cohort study data. Biometrics, 70(1):164–174, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2014:ELR

[710] Paul S. Albert, Aiyi Liu, and Tonja Nansel. Efficient logistic regression designs under an imperfect population identifier. Biometrics, 70(1):175– 184, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murray:2014:SBC

[711] Thomas A. Murray, Brian P. Hobbs, Theodore C. Lystig, and Bradley P. Carlin. Semiparametric Bayesian commensurate survival model for post-market medical device surveillance with non-exchangeable historical data. Biometrics, 70(1):185–191, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2014:AHM

[712] Yuhui Chen, Timothy Hanson, and Jiajia Zhang. Accelerated hazards model based on parametric families generalized with Bernstein polynomials. Biometrics, 70(1):192–201, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mitchell:2014:RSB

[713] Emily M. Mitchell, Robert H. Lyles, Amita K. Manatunga, Michelle Danaher, Neil J. Perkins, and Enrique F. Schisterman. Regression for skewed biomarker outcomes subject to pooling. Biometrics, 70(1):202– 211, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bantis:2014:CCR

[714] Leonidas E. Bantis, Christos T. Nakas, and Benjamin Reiser. Construction of confidence regions in the ROC space after the estimation of the optimal Youden index-based cut-off point. Biometrics, 70(1):212–223, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Risk:2014:EIC

[715] Benjamin B. Risk, David S. Matteson, David Ruppert, Ani Eloyan, and Brian S. Caffo. An evaluation of independent component analyses with an application to resting-state fMRI. Biometrics, 70(1):224–236, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilson:2014:HDR

[716] Ander Wilson, David M. Reif, and Brian J. Reich. Hierarchical doseresponse modeling for high-throughput toxicity screening of environmental chemicals. Biometrics, 70(1):237–246, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerhard:2014:EMP

[717] Daniel Gerhard, Melanie Bremer, and Christian Ritz. Estimating marginal properties of quantitative real-time PCR data using nonlinear mixed models. Biometrics, 70(1):247–254, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2014:BRS

[718] Seungyeoun Lee. Book review: Survival Analysis in Medicine and Genetics, by Li, Jialiang and Ma, Shuangge. Biometrics, 70(1):255–256, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oh:2014:BRB

[719] Man-Suk Oh. Book review: Bayesian Methods in Health Economics, by Baio, Gianluca. Biometrics, 70(1):256, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2014:BRM

[720] Jae kwang Kim. Book review: Methods of Statistical Model Estimation, by Hilbe, Joseph M. and Robinson, Andrew P. Biometrics, 70(1):257, March 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Billard:2014:SRF

[721] Lynne Billard. Sir Ronald A. Fisher and the International Biometric Society. Biometrics, 70(2):259–265, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mayo:2014:FA

[722] Oliver Mayo. Fisher in Adelaide. Biometrics, 70(2):266–269, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hattori:2014:SDR

[723] Satoshi Hattori and Masayuki Henmi. Stratified doubly robust estimators for the average causal effect. Biometrics, 70(2):270–277, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mattei:2014:ICE

[724] Alessandra Mattei, Fabrizia Mealli, and Barbara Pacini. Identification of causal effects in the presence of nonignorable missing outcome values. Biometrics, 70(2):278–288, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2014:TRD

[725] Baojiang Chen and Jing Qin. Test the reliability of doubly robust estimation with missing response data. Biometrics, 70(2):289–298, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:2014:SAN

[726] Samiran Sinha, Krishna K. Saha, and Suojin Wang. Semiparametric approach for non-monotone missing covariates in a parametric regression model. Biometrics, 70(2):299–311, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2014:PEA

[727] Lin S. Chen, Ross L. Prentice, and Pei Wang. A penalized EM algorithm incorporating missing data mechanism for Gaussian parameter estimation. Biometrics, 70(2):312–322, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2014:UIM

[728] Zhi Wen, Kenneth H. Pollock, James D. Nichols, Peter M. Waser, and Weihua Cao. Using imputation and mixture model approaches to integrate multi-state capture-recapture models with assignment information. Biometrics, 70(2):323–334, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2014:SAI

[729] Chi-Chung Wen and Yi-Hau Chen. Semiparametric analysis of incomplete current status outcome data under transformation models. Biometrics, 70(2):335–345, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Knape:2014:ESD

[730] Jonas Knape, Kent M. Daane, and Perry de Valpine. Estimation of stage duration distributions and mortality under repeated cohort censuses. Biometrics, 70(2):346–355, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2014:UDT

[731] Peirong Xu, Lixing Zhu, and Yi Li. Ultrahigh dimensional time course feature selection. Biometrics, 70(2):356–365, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2014:EIR

[732] Suyu Liu, Ying Yuan, Richard Castillo, Thomas Guerrero, and Valen E. Johnson. Evaluation of image registration spatial accuracy using a Bayesian hierarchical model. Biometrics, 70(2):366–377, June 2014. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kapelner:2014:MFS

[733] Adam Kapelner and Abba Krieger. Matching on-the-fly: Sequential allocation with higher power and efficiency. Biometrics, 70(2):378–388, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2014:SBC

[734] Ying Kuen Cheung. Simple benchmark for complex dose finding studies. Biometrics, 70(2):389–397, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2014:CUC

[735] Xingdong Feng, Nell Sedransk, and Jessie Q. Xia. Calibration using constrained smoothing with applications to mass spectrometry data. Biometrics, 70(2):398–408, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Viladomat:2014:ASG

[736] Júlia Viladomat, Rahul Mazumder, Alex McInturff, Douglas J. Mc-Cauley, and Trevor Hastie. Assessing the significance of global and local correlations under spatial autocorrelation: a nonparametric approach. Biometrics, 70(2):409–418, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2014:BLC

[737] Duncan Lee, Alastair Rushworth, and Sujit K. Sahu. A Bayesian localized conditional autoregressive model for estimating the health effects of air pollution. Biometrics, 70(2):419–429, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yun:2014:BHM

[738] Jonghyun Yun, Tao Wang, and Guanghua Xiao. Bayesian hidden Markov models to identify RNA–protein interaction sites in PAR–CLIP. Biometrics, 70(2):430–440, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gouskova:2014:VSH

[739] Natalia A. Gouskova, Stephen R. Cole, Joseph J. Eron, and Jason P. Fine. Viral suppression in HIV studies: Combining times to suppression and rebound. Biometrics, 70(2):441–448, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seaman:2014:MOC

[740] Shaun R. Seaman, Menelaos Pavlou, and Andrew J. Copas. Methods for observed-cluster inference when cluster size is informative: a review and clarifications. Biometrics, 70(2):449–456, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fearnhead:2014:IRN

[741] Paul Fearnhead, Vasilieos Giagos, and Chris Sherlock. Inference for reaction networks using the linear noise approximation. Biometrics, 70 (2):457–466, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nam:2014:BRT

[742] Byung-Ho Nam. Book review: Theory of Drug Development, by Holmgren, E. B. Biometrics, 70(2):467, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Won:2014:BRU

[743] Sungho Won. Book review: Understanding Advanced Statistical Methods, by Westfall, P. H. and Henning, K. S. S. Biometrics, 70(2):467–468, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2014:BRA

[744] Taesung Park. Book review: Analysis of Mixed Data, by de Leon, A. R. and Chough, K. C. Biometrics, 70(2):468–469, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2014:MTJ

[745] Zihuai He, Min Zhang, Xiaowei Zhan, and Qing Lu. Modeling and testing for joint association using a genetic random field model. Biometrics, 70(3):471–479, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2014:IAP

[746] Jin Liu, Jian Huang, Yawei Zhang, Qing Lan, Nathaniel Rothman, Tongzhang Zheng, and Shuangge Ma. Integrative analysis of prognosis data on multiple cancer subtypes. Biometrics, 70(3):480–488, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsouaka:2014:EMG

[747] Roland A. Matsouaka, Junlong Li, and Tianxi Cai. Evaluating marker-guided treatment selection strategies. Biometrics, 70(3):489–499, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sjolander:2014:BCI

[748] A. Sjölander, W. Lee, H. Källberg, and Y. Pawitan. Bounds on causal interactions for binary outcomes. Biometrics, 70(3):500–505, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OMalley:2014:EPE

[749] A. James O’Malley, Felix Elwert, J. Niels Rosenquist, Alan M. Zaslavsky, and Nicholas A. Christakis. Estimating peer effects in longitudinal dyadic data using instrumental variables. Biometrics, 70(3):506–515, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2014:OWD

[750] Huaihou Chen, Philip T. Reiss, and Thaddeus Tarpey. Optimally weighted L^{2} distance for functional data. Biometrics, 70(3):516–525, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gervini:2014:WFA

[751] Daniel Gervini and Patrick A. Carter. Warped functional analysis of variance. Biometrics, 70(3):526–535, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jung:2014:SAS

[752] Sungkyu Jung and Xingye Qiao. A statistical approach to set classification by feature selection with applications to classification of histopathology images. Biometrics, 70(3):536–545, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shin:2014:PES

[753] Seung Jun Shin, Yichao Wu, Hao Helen Zhang, and Yufeng Liu. Probability-enhanced sufficient dimension reduction for binary classification. Biometrics, 70(3):546–555, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Levin:2014:EIP

[754] Gregory P. Levin, Sarah C. Emerson, and Scott S. Emerson. An evaluation of inferential procedures for adaptive clinical trial designs with prespecified rules for modifying the sample size. Biometrics, 70(3):556–567, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2014:PEA

[755] Yang Yang, M. Elizabeth Halloran, Yanjun Chen, and Eben Kenah. A pathway EM-algorithm for estimating vaccine efficacy with a nonmonotone validation set. Biometrics, 70(3):568–578, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2014:AIF

[756] Bo Liu, Wenbin Lu, and Jiajia Zhang. Accelerated intensity frailty model for recurrent events data. Biometrics, 70(3):579–587, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choi:2014:MLE

[757] Sangbum Choi and Xuelin Huang. Maximum likelihood estimation of semiparametric mixture component models for competing risks data. Biometrics, 70(3):588–598, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2014:STM

[758] Huazhen Lin, Ling Zhou, Chunhong Li, and Yi Li. Semiparametric transformation models for semicompeting survival data. Biometrics, 70(3): 599–607, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2014:CEE

[759] Yu-Jen Cheng and Chiung-Yu Huang. Combined estimating equation approaches for semiparametric transformation models with length-biased survival data. Biometrics, 70(3):608–618, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Saegusa:2014:HTE

[760] Takumi Saegusa, Chongzhi Di, and Ying Qing Chen. Hypothesis testing for an extended Cox model with time-varying coefficients. Biometrics, 70(3):619–628, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Concordet:2014:IPR

[761] D. Concordet and R. Servien. Individual prediction regions for multivariate longitudinal data with small samples. Biometrics, 70(3):629–638, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Forkman:2014:PBM

[762] Johannes Forkman and Hans-Peter Piepho. Parametric bootstrap methods for testing multiplicative terms in GGE and AMMI models. Biometrics, 70(3):639–647, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lopiano:2014:PPQ

[763] Kenneth K. Lopiano, Linda J. Young, and Carol A. Gotway. A pseudopenalized quasi-likelihood approach to the spatial misalignment problem with non-normal data. Biometrics, 70(3):648–660, September 2014. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bai:2014:EPC

[764] Yun Bai, Jian Kang, and Peter X.-K. Song. Efficient pairwise composite likelihood estimation for spatial-clustered data. Biometrics, 70(3):661– 670, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chiu:2014:INL

[765] Chun-Huo Chiu, Yi-Ting Wang, Bruno A. Walther, and Anne Chao. An improved nonparametric lower bound of species richness via a modified good-turing frequency formula. Biometrics, 70(3):671–682, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2014:BEP

[766] Joungyoun Kim and Bret Larget. Bayesian estimation of the phylogeography of African gorillas with genome-differentiated population trees. Biometrics, 70(3):683–694, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kang:2014:CBO

[767] Chaeryon Kang, Holly Janes, and Ying Huang. Combining biomarkers to optimize patient treatment recommendations. Biometrics, 70(3):695– 707, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [768, 770, 771] and rejoinder [772].

Laber:2014:DCB

[768] Eric B. Laber, Anastasios A. Tsiatis, Marie Davidian, and Shannon T. Holloway. Discussion of “Combining biomarkers to optimize patient treatment recommendation”. Biometrics, 70(3):707–710, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [767, 772].

Tian:2014:D

[769] Lu Tian. Discussion. Biometrics, 70(3):710–713, September 2014. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2014:DCB

[770] Ying-Qi Zhao and Michael R. Kosorok. Discussion of combining biomarkers to optimize patient treatment recommendations. Biometrics, 70(3): 713–716, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [767].

Yu:2014:DCB

[771] Menggang Yu and Quefeng Li. Discussion of “Combining biomarkers to optimize patient treatment recommendations” by Chaeryon Kang, Holly Janes, and Ying Huang. Biometrics, 70(3):716–719, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [767, 772].

Kang:2014:RCB

[772] Chaeryon Kang, Holly Janes, and Ying Huang. Rejoinder: Combining biomarkers to optimize patient treatment recommendations. Biometrics, 70(3):719–720, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [767].

Taguri:2014:MSC

[773] Masataka Taguri, Yutaka Matsuyama, and Yasuo Ohashi. Model selection criterion for causal parameters in structural mean models based on a quasi-likelihood. Biometrics, 70(3):721–730, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perez-Heydrich:2014:AEC

[774] Carolina Perez-Heydrich, Michael G. Hudgens, M. Elizabeth Halloran, John D. Clemens, Mohammad Ali, and Michael E. Emch. Assessing effects of cholera vaccination in the presence of interference. Biometrics, 70(3):731–741, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2014:EEH

[775] Jingyang Zhang and Elizabeth R. Brown. Estimating the effectiveness in HIV prevention trials by incorporating the exposure process: Application to HPTN 035 data. Biometrics, 70(3):742–750, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Estes:2014:CER

[776] Jason P. Estes, Danh V. Nguyen, Lorien S. Dalrymple, Yi Mu, and Damla Sentürk. Cardiovascular event risk dynamics over time in older patients on dialysis: a generalized multiple-index varying coefficient model approach. Biometrics, 70(3):751–761, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ha:2014:PCM

[777] Min Jin Ha and Wei Sun. Partial correlation matrix estimation using ridge penalty followed by thresholding and re-estimation. Biometrics, 70(3):762–770, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2014:BRR

[778] Jae Won Lee. Book review: Randomized Phase II Cancer Clinical Trials, by Jung, Sin-Ho. Biometrics, 70(3):771–772, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Paik:2014:BRS

[779] Myunghee Cho Paik. Book review: Statistical Methods for Handling Incomplete Data, by Kim, Jae Kwang and Shao, Jun. Biometrics, 70(3): 772–773, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

An:2014:BRB

[780] Hyonggin An. Book review: Biosimilars: Design and Analysis of Follow-on Biologics, by Chow, Shein-Chung. Biometrics, 70(3):773–774, September 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barker:2014:CPC

[781] Richard J. Barker, Matthew R. Schofield, Janine A. Wright, Alain C. Frantz, and Chris Stevens. Closed-population capture-recapture modeling of samples drawn one at a time. Biometrics, 70(4):775–782, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kollmann:2014:URU

[782] Claudia Köllmann, Björn Bornkamp, and Katja Ickstadt. Unimodal regression using Bernstein-Schoenberg splines and penalties. Biometrics, 70(4):783–793, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2014:LMF

[783] Wei Wang. Linear mixed function-on-function regression models. Biometrics, 70(4):794–801, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [691].

Li:2014:HFD

[784] Haocheng Li, John Staudenmayer, and Raymond J. Carroll. Hierarchical functional data with mixed continuous and binary measurements. Biometrics, 70(4):802–811, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xue:2014:IFC

[785] Wenqiong Xue, Jian Kang, F. DuBois Bowman, Tor D. Wager, and Jian Guo. Identifying functional co-activation patterns in neuroimaging studies via Poisson graphical models. Biometrics, 70(4):812–822, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sarkar:2014:BSR

[786] Abhra Sarkar, Bani K. Mallick, and Raymond J. Carroll. Bayesian semiparametric regression in the presence of conditionally heteroscedastic measurement and regression errors. Biometrics, 70(4):823–834, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2014:CME

[787] Shanshan Zhao and Ross L. Prentice. Covariate measurement error correction methods in mediation analysis with failure time data. Biometrics, 70(4):835–844, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2014:TGF

[788] Wenbin Lu, Mengling Liu, and Yi-Hau Chen. Testing goodness-of-fit for the proportional hazards model based on nested case-control data. Biometrics, 70(4):845–851, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilson:2014:CSP

[789] Ander Wilson and Brian J. Reich. Confounder selection via penalized credible regions. Biometrics, 70(4):852–861, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2014:STV

[790] Sihai Dave Zhao and Yi Li. Score test variable screening. Biometrics, 70(4):862–871, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2014:MAB

[791] Quefeng Li, Sijian Wang, Chiang-Ching Huang, Menggang Yu, and Jun Shao. Meta-analysis based variable selection for gene expression data. Biometrics, 70(4):872–880, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2014:MPG

[792] Sihai D. Zhao, T. Tony Cai, and Hongzhe Li. More powerful genetic association testing via a new statistical framework for integrative genomics. Biometrics, 70(4):881–890, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2014:IOB

[793] Ying Huang and Youyi Fong. Identifying optimal biomarker combinations for treatment selection via a robust kernel method. Biometrics, 70(4):891–901, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perin:2014:RAC

[794] J. Perin, J. S. Preisser, C. Phillips, and B. Qaqish. Regression analysis of correlated ordinal data using orthogonalized residuals. Biometrics, 70(4): 902–909, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2014:SEG

[795] Li Liu and Liming Xiang. Semiparametric estimation in generalized linear mixed models with auxiliary covariates: a pairwise likelihood approach. Biometrics, 70(4):910–919, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aeberhard:2014:RIN

[796] William H. Aeberhard, Eva Cantoni, and Stephane Heritier. Robust inference in the negative binomial regression model with an application to falls data. Biometrics, 70(4):920–931, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2014:SMS

[797] Brian J. Reich, Howard H. Chang, and Kristen M. Foley. A spectral method for spatial downscaling. Biometrics, 70(4):932–942, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2014:SIM

[798] Jie Cheng, Elizaveta Levina, Pei Wang, and Ji Zhu. A sparse Ising model with covariates. Biometrics, 70(4):943–953, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Solari:2014:RBM

[799] Aldo Solari, Livio Finos, and Jelle J. Goeman. Rotation-based multiple testing in the multivariate linear model. Biometrics, 70(4):954–961, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vale:2014:MLE

[800] R. T. R. Vale, R. M. Fewster, E. L. Carroll, and N. J. Patenaude. Maximum likelihood estimation for model M_{t,α} for capture-recapture data with misidentification. Biometrics, 70(4):962–971, December 2014. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Millar:2014:OSA

[801] R. B. Millar and S. McKechnie. A one-step-ahead pseudo-DIC for comparison of Bayesian state-space models. Biometrics, 70(4):972–980, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kang:2014:BNM

[802] Jian Kang, Nanhua Zhang, and Ran Shi. A Bayesian nonparametric model for spatially distributed multivariate binary data with application to a multidrug-resistant tuberculosis (MDR-TB) study. Biometrics, 70(4):981–992, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hohle:2014:BND

[803] Michael Höhle and Matthias an der Heiden. Bayesian nowcasting during the STEC O104:H4 outbreak in Germany, 2011. Biometrics, 70(4):993– 1002, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2014:BDS

[804] Ming-Hui Chen, Joseph G. Ibrahim, Donglin Zeng, Kuolung Hu, and Catherine Jia. Bayesian design of superiority clinical trials for recurrent events data with applications to bleeding and transfusion events in myelodyplastic syndrome. Biometrics, 70(4):1003–1013, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2014:ETE

[805] Zhenke Wu, Constantine E. Frangakis, Thomas A. Louis, and Daniel O. Scharfstein. Estimation of treatment effects in matched-pair cluster randomized trials by calibrating covariate imbalance between clusters. Biometrics, 70(4):1014–1022, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schmidli:2014:RMA

[806] Heinz Schmidli, Sandro Gsteiger, Satrajit Roychoudhury, Anthony O’Hagan, David Spiegelhalter, and Beat Neuenschwander. Robust metaanalytic-predictive priors in clinical trials with historical control information. Biometrics, 70(4):1023–1032, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2014:PCC

[807] Eunhee Kim, Zheng Zhang, Youdan Wang, and Donglin Zeng. Power calculation for comparing diagnostic accuracies in a multi-reader, multi-test design. Biometrics, 70(4):1033–1041, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huynh:2014:SRS

[808] Dat Huynh, Oliver Laeyendecker, and Ron Brookmeyer. A serial risk score approach to disease classification that accounts for accuracy and cost. Biometrics, 70(4):1042–1051, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLain:2014:MLD

[809] Alexander C. McLain and Paul S. Albert. Modeling longitudinal data with a random change point and no time-zero: Applications to inference and prediction of the labor curve. Biometrics, 70(4):1052–1060, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2014:CUP

[810] Anonymous. Correction to “Using Pooled Exposure Assessment to Improve Efficiency in Case-Control Studies,” by Clarice R. Weinberg and David M. Umbach; 55, 718–726, September 1999. Biometrics, 70(4): 1061, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [2].

Yoo:2014:BRC

[811] Yun Joo Yoo. Book review: Chen, Jehua. Statistical Methods for QTL Mapping. Chapman and Hall/CRC, Boca Raton, FL. Biometrics, 70 (4):1063, December 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daniel:2015:CMA

[812] R. M. Daniel, B. L. De Stavola, S. N. Cousens, and S. Vansteelandt. Causal mediation analysis with multiple mediators. Biometrics, 71(1):1– 14, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2015:RBC

[813] Ariel Alonso, Wim Van der Elst, Geert Molenberghs, Marc Buyse, and Tomasz Burzykowski. On the relationship between the causal-inference and meta-analytic paradigms for the validation of surrogate endpoints. Biometrics, 71(1):15–24, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choai:2015:ETE

[814] Yuki Choai and Shigeyuki Matsui. Estimation of treatment effects in allcomers randomized clinical trials with a predictive marker. Biometrics, 71(1):25–32, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hedayat:2015:MCI

[815] A. S. Hedayat, Junhui Wang, and Tu Xu. Minimum clinically important difference in medical studies. Biometrics, 71(1):33–41, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Saarela:2015:CBM

[816] Olli Saarela and James A. Hanley. Case-base methods for studying vaccination safety. Biometrics, 71(1):42–52, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Geng:2015:GVS

[817] Zhigeng Geng, Sijian Wang, Menggang Yu, Patrick O. Monahan, Victoria Champion, and Grace Wahba. Group variable selection via convex log-exp-sum penalty with application to a breast cancer survivor study. Biometrics, 71(1):53–62, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2015:SKM

[818] Yuanyuan Shen, Katherine P. Liao, and Tianxi Cai. Sparse kernel machine regression for ordinal outcomes. Biometrics, 71(1):63–70, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2015:RAM

[819] Liang Zhu, Hui Zhao, Jianguo Sun, Wendy Leisenring, and Leslie L. Robison. Regression analysis of mixed recurrent-event and panel-count data with additive rate models. Biometrics, 71(1):71–79, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartolucci:2015:DTE

[820] Francesco Bartolucci and Alessio Farcomeni. A discrete time eventhistory approach to informative drop-out in mixed latent Markov models with covariates. Biometrics, 71(1):80–89, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lange:2015:JMM

[821] Jane M. Lange, Rebecca A. Hubbard, Lurdes Y. T. Inoue, and Vladimir N. Minin. A joint model for multistate disease processes and random informative observation times, with applications to electronic medical records data. Biometrics, 71(1):90–101, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Blanche:2015:QCD

[822] Paul Blanche, Cécile Proust-Lima, Lucie Loubère, Claudine Berr, Jean-François Dartigues, and Hélène Jacqmin-Gadda. Quantifying and comparing dynamic predictive accuracy of joint models for longitudinal marker and time-to-event in presence of censoring and competing risks. Biometrics, 71(1):102–113, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:2015:DRE

[823] Xiaohui Chang, Rasmus Waagepetersen, Herbert Yu, Xiaomei Ma, Theodore R. Holford, Rong Wang, and Yongtao Guan. Disease risk estimation by combining case-control data with aggregated information on the population at risk. Biometrics, 71(1):114–121, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2015:IVA

[824] Jialiang Li, Jason Fine, and Alan Brookhart. Instrumental variable additive hazards models. Biometrics, 71(1):122–130, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2015:SOD

[825] Xinyu Zhang, Jiguo Cao, and Raymond J. Carroll. On the selection of ordinary differential equation models with application to predator-prey dynamical models. Biometrics, 71(1):131–138, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2015:AAC

[826] Xiaodong Luo, Hong Tian, Surya Mohanty, and Wei Yann Tsai. An alternative approach to confidence interval estimation for the win ratio statistic. Biometrics, 71(1):139–145, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fay:2015:COS

[827] Michael P. Fay, Michael A. Proschan, and Erica Brittain. Combining one-sample confidence procedures for inference in the two-sample case. Biometrics, 71(1):146–156, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Touloumis:2015:TMM

[828] Anestis Touloumis, Simon Tavaré, and John C. Marioni. Testing the mean matrix in high-dimensional transposable data. Biometrics, 71(1): 157–166, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vock:2015:SVS

[829] Laura F. Boehm Vock, Brian J. Reich, Montserrat Fuentes, and Francesca Dominici. Spatial variable selection methods for investigating acute health effects of fine particulate matter components. Biometrics, 71(1):167–177, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2015:SVS

[830] Zangdong He, Wanzhu Tu, Sijian Wang, Haoda Fu, and Zhangsheng Yu. Simultaneous variable selection for joint models of longitudinal and survival outcomes. Biometrics, 71(1):178–187, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graziani:2015:BNE

[831] Rebecca Graziani, Michele Guindani, and Peter F. Thall. Bayesian nonparametric estimation of targeted agent effects on biomarker change to predict clinical outcome. Biometrics, 71(1):188–197, March 2015. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2015:MCP

[832] Xing Ju Lee, Christopher C. Drovandi, and Anthony N. Pettitt. Model choice problems using approximate Bayesian computation with applications to pathogen transmission data sets. Biometrics, 71(1):198–207, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quiroz:2015:BAE

[833] Zaida C. Quiroz, Marcos O. Prates, and Håvard Rue. A Bayesian approach to estimate the biomass of anchovies off the coast of Perú. Biometrics, 71(1):208–217, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ventz:2015:BDC

[834] Steffen Ventz and Lorenzo Trippa. Bayesian designs and the control of frequentist characteristics: a practical solution. Biometrics, 71(1):218– 226, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smoot:2015:AHD

[835] E. Smoot and S. Haneuse. On the analysis of hybrid designs that combine group- and individual-level data. Biometrics, 71(1):227–236, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dennis:2015:CAM

[836] Emily B. Dennis, Byron J. T. Morgan, and Martin S. Ridout. Computational aspects of N-mixture models. Biometrics, 71(1):237–246, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shou:2015:SFP

[837] Haochang Shou, Vadim Zipunnikov, Ciprian M. Crainiceanu, and Sonja Greven. Structured functional principal component analysis. Biometrics, 71(1):247–257, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Handcock:2015:ESP

[838] Mark S. Handcock, Krista J. Gile, and Corinne M. Mar. Estimating the size of populations at high risk for HIV using respondent-driven sampling data. Biometrics, 71(1):258–266, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2015:RRR

[839] Anonymous. Response to reader reaction. Biometrics, 71(1):267–273, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:2015:RRR

[840] Jeremy M. G. Taylor, Wenting Cheng, and Jared C. Foster. Reader reaction to “A robust method for estimating optimal treatment regimes” by Zhang et al. (2012). Biometrics, 71(1):267–273, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [517].

Kang:2015:RRA

[841] Seung-Ho Kang. Randomized response-adaptive designs in clinical trials. Biometrics, 71(1):274–277, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ugarte:2015:BRB

[842] Maria Dolores Ugarte. Book review: Banerjee, S. Carlin, B. P., and Gelfand, A. E., Hierarchical Modeling and Analysis for Spatial Data. Second Edition. CRC Press/Chapman and Hall. Monographs on Statistics and Applied Probability 135, Boca Raton, Florida, 2015. 562 pp. $99.95. ISBN-13: 978-1-4398-1917-3 (Hardcover). Biometrics, 71(1):274– 277, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2015:RE

[843] Anonymous. Report of the Editors — 2014. Biometrics, 71(1):fmvii– fmix, March 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Saarela:2015:BEM

[844] Olli Saarela, David A. Stephens, Erica E. M. Moodie, and Marina B. Klein. On Bayesian estimation of marginal structural models. Biometrics, 71(2):279–288, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [845, 846, 847, 848] and rejoinder [849].

Elliott:2015:DBE

[845] Michael R. Elliott and Roderick J. Little. Discussion of “On Bayesian Estimation of Marginal Structural Models”. Biometrics, 71(2):288–291, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [844, 849].

Gustafson:2015:DBE

[846] Paul Gustafson. Discussion of “On Bayesian Estimation of Marginal Structural Models”. Biometrics, 71(2):291–293, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [844, 849].

Mattei:2015:DBE

[847] Alessandra Mattei and Fabrizia Mealli. Discussion of “On Bayesian Estimation of Marginal Structural Models”. Biometrics, 71(2):293–296, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [844, 849].

Robins:2015:DBE

[848] James M. Robins, Miguel A. Hernán, and Larry Wasserman. Discussion of “On Bayesian estimation of marginal structural models”. Biometrics, 71(2):296–299, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [844, 849].

Saarela:2015:RBE

[849] Olli Saarela, David A. Stephens, Erica E. M. Moodie, and Marina B. Klein. Rejoinder “On Bayesian estimation of marginal structural models”. Biometrics, 71(2):299–301, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [844].

Cheng:2015:CEU

[850] Yu-Jen Cheng and Mei-Cheng Wang. Causal estimation using semiparametric transformation models under prevalent sampling. Biometrics, 71 (2):302–312, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2015:SEH

[851] Li Li, Timothy Hanson, and Jiajia Zhang. Spatial extended hazard model with application to prostate cancer survival. Biometrics, 71(2):313–322, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Erhardt:2015:RVM

[852] Tobias Michael Erhardt, Claudia Czado, and Ulf Schepsmeier. R-vine models for spatial time series with an application to daily mean temperature. Biometrics, 71(2):323–332, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Arribas-Gil:2015:CLD

[853] Ana Arribas-Gil, Rolando De la Cruz, Emilie Lebarbier, and Cristian Meza. Classification of longitudinal data through a semiparametric mixed-effects model based on lasso-type estimators. Biometrics, 71(2): 333–343, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goldsmith:2015:GMF

[854] Jeff Goldsmith, Vadim Zipunnikov, and Jennifer Schrack. Generalized multilevel function-on-scalar regression and principal component analysis. Biometrics, 71(2):344–353, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2015:MSG

[855] Yanming Li, Bin Nan, and Ji Zhu. Multivariate sparse group lasso for the multivariate multiple linear regression with an arbitrary group structure. Biometrics, 71(2):354–363, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

deUna-Alvarez:2015:NET

[856] Jacobo de Uña-Álvarez and Luís Meira-Machado. Nonparametric estimation of transition probabilities in the non-Markov illness-death model: a comparative study. Biometrics, 71(2):364–375, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goldstein:2015:ARP

[857] Joshua Goldstein, Murali Haran, Ivan Simeonov, John Fricks, and Francesca Chiaromonte. An attraction-repulsion point process model for respiratory syncytial virus infections. Biometrics, 71(2):376–385, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:2015:IFV

[858] Dean Follmann and Chiung-Yu Huang. Incorporating founder virus information in vaccine field trials. Biometrics, 71(2):386–396, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2015:IHH

[859] Xin Cheng, Wenbin Lu, and Mengling Liu. Identification of homogeneous and heterogeneous variables in pooled cohort studies. Biometrics, 71(2): 397–403, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Komori:2015:GST

[860] Osamu Komori, Shinto Eguchi, and John B. Copas. Generalized tstatistic for two-group classification. Biometrics, 71(2):404–416, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baayen:2015:TED

[861] C. Baayen, P. Hougaard, and C. B. Pipper. Testing effect of a drug using multiple nested models for the dose-response. Biometrics, 71(2): 417–427, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chekouo:2015:MTG

[862] Thierry Chekouo, Francesco C. Stingo, James D. Doecke, and Kim-Anh Do. miRNA-target gene regulatory networks: a Bayesian integrative approach to biomarker selection with application to kidney cancer. Biometrics, 71(2):428–438, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2015:DME

[863] Weining Shen, Jing Ning, and Ying Yuan. A direct method to evaluate the time-dependent predictive accuracy for biomarkers. Biometrics, 71 (2):439–449, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2015:SMA

[864] Ying Kuen Cheung, Bibhas Chakraborty, and Karina W. Davidson. Sequential multiple assignment randomized trial (SMART) with adaptive randomization for quality improvement in depression treatment program. Biometrics, 71(2):450–459, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conlon:2015:IEC

[865] A. S. C. Conlon, J. M. G. Taylor, and D. J. Sargent. Improving efficiency in clinical trials using auxiliary information: Application of a multi-state cure model. Biometrics, 71(2):460–468, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jeffries:2015:LCT

[866] Neal Jeffries and Nancy L. Geller. Longitudinal clinical trials with adaptive choice of follow-up time. Biometrics, 71(2):469–477, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goldstein:2015:DCM

[867] Benjamin A. Goldstein, Themistocles Assimes, Wolfgang C. Winkelmayer, and Trevor Hastie. Detecting clinically meaningful biomarkers with repeated measurements: an illustration with electronic health records. Biometrics, 71(2):478–486, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2015:JMC

[868] Bei Jiang, Michael R. Elliott, Mary D. Sammel, and Naisyin Wang. Joint modeling of cross-sectional health outcomes and longitudinal predictors via mixtures of means and variances. Biometrics, 71(2):487–497, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moreno-Betancur:2015:DLI

[869] Margarita Moreno-Betancur, Grégoire Rey, and Aurélien Latouche. Direct likelihood inference and sensitivity analysis for competing risks regression with missing causes of failure. Biometrics, 71(2):498–507, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:2015:MQF

[870] Luke B. Smith, Brian J. Reich, Amy H. Herring, Peter H. Langlois, and Montserrat Fuentes. Multilevel quantile function modeling with application to birth outcomes. Biometrics, 71(2):508–519, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Langrock:2015:NIH

[871] Roland Langrock, Thomas Kneib, Alexander Sohn, and Stacy L. DeRuiter. Nonparametric inference in hidden Markov models using Psplines. Biometrics, 71(2):520–528, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neely:2015:PLA

[872] Megan L. Neely, Howard D. Bondell, and Jung-Ying Tzeng. A penalized likelihood approach for investigating gene-drug interactions in pharmacogenetic studies. Biometrics, 71(2):529–537, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2015:UMM

[873] Yulun Liu, Yong Chen, and Haitao Chu. A unification of models for meta-analysis of diagnostic accuracy studies without a gold standard. Biometrics, 71(2):538–547, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Demetrashvili:2015:CII

[874] Nino Demetrashvili and Edwin R. Van den Heuvel. Confidence intervals for intraclass correlation coefficients in a nonlinear dose-response metaanalysis. Biometrics, 71(2):548–555, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2015:RRG

[875] Baolin Wu and Weihua Guan. Reader reaction on the generalized Kruskal-Wallis test for genetic association studies incorporating group uncertainty. Biometrics, 71(2):556–557, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2015:CTM

[876] Juwon Song. Clinical trials with missing data. Biometrics, 71(2):558– 559, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2015:BRC

[877] Sohee Park. Book review: Carpenter, J. R., and Kenward, M. G., Multiple Imputation and its Application. Wiley, Hoboken, New Jersey. Biometrics, 71(2):559–560, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2015:BRJ

[878] Seungyeoun Lee. Book review: Jeong, J.-H. Statistical Inference on Residual Life. Springer, New York. Biometrics, 71(2):560–561, June 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meyer:2015:BFF

[879] Mark J. Meyer, Brent A. Coull, Francesco Versace, Paul Cinciripini, and Jeffrey S. Morris. Bayesian function-on-function regression for multilevel functional data. Biometrics, 71(3):563–574, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quick:2015:BMA

[880] Harrison Quick, Sudipto Banerjee, and Bradley P. Carlin. Bayesian modeling and analysis for gradients in spatiotemporal processes. Biometrics, 71(3):575–584, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ni:2015:BNM

[881] Yang Ni, Francesco C. Stingo, and Veerabhadran Baladandayuthapani. Bayesian nonlinear model selection for gene regulatory networks. Biometrics, 71(3):585–595, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2015:EWS

[882] Summer S. Han, Philip S. Rosenberg, Arpita Ghosh, Maria Teresa Landi, Neil E. Caporaso, and Nilanjan Chatterjee. An exposure-weighted score test for genetic associations integrating environmental risk factors. Biometrics, 71(3):596–605, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2015:SBT

[883] Zihuai He, Min Zhang, Seunggeun Lee, Jennifer A. Smith, Xiuqing Guo, Walter Palmas, Sharon L. R. Kardia, Ana V. Diez Roux, and Bhramar Mukherjee. Set-based tests for genetic association in longitudinal studies. Biometrics, 71(3):606–615, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qiu:2015:MBE

[884] Jin Qiu and Lang Wu. A moving blocks empirical likelihood method for longitudinal data. Biometrics, 71(3):616–624, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2015:QRC

[885] Chenxi Li, N. Maritza Dowling, and Rick Chappell. Quantile regression with a change-point model for longitudinal data: an application to the study of cognitive changes in preclinical alzheimer’s disease. Biometrics, 71(3):625–635, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wallace:2015:DRD

[886] Michael P. Wallace and Erica E. M. Moodie. Doubly-robust dynamic treatment regimen estimation via weighted least squares. Biometrics, 71(3):636–644, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correlation [1672].

Xu:2015:ROW

[887] Yaoyao Xu, Menggang Yu, Ying-Qi Zhao, Quefeng Li, Sijian Wang, and Jun Shao. Regularized outcome weighted subgroup identification for differential treatment effects. Biometrics, 71(3):645–653, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2015:AUC

[888] Chi Wang, Francesca Dominici, Giovanni Parmigiani, and Corwin Matthew Zigler. Accounting for uncertainty in confounder and effect modifier selection when estimating average causal effects in generalized linear models. Biometrics, 71(3):654–665, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2015:ECS

[889] Shanshan Li and Yang Ning. Estimation of covariate-specific timedependent ROC curves in the presence of missing biomarkers. Biometrics, 71(3):666–676, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eriksson:2015:AGF

[890] Frank Eriksson and Thomas Scheike. Additive gamma frailty models with applications to competing risks in related individuals. Biometrics, 71(3):677–686, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eriksson:2015:POC

[891] Frank Eriksson, Jianing Li, Thomas Scheike, and Mei-Jie Zhang. The proportional odds cumulative incidence model for competing risks. Biometrics, 71(3):687–695, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borgan:2015:UCS

[892] Ørnulf Borgan and Ying Zhang. Using cumulative sums of martingale residuals for model checking in nested case-control studies. Biometrics, 71(3):696–703, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2015:DIM

[893] Yanqing Wang, Suojin Wang, and Raymond J. Carroll. The direct integral method for confidence intervals for the ratio of two location parameters. Biometrics, 71(3):704–713, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schmertmann:2015:APS

[894] Carl P. Schmertmann. Adjusting for population shifts and covariates in space-time interaction tests. Biometrics, 71(3):714–720, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tounkara:2015:MRM

[895] Fodé Tounkara and Louis-Paul Rivest. Mixture regression models for closed population capture-recapture data. Biometrics, 71(3):721–730, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Breheny:2015:GEL

[896] Patrick Breheny. The group exponential lasso for bi-level variable selection. Biometrics, 71(3):731–740, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shu:2015:MTN

[897] Hai Shu, Bin Nan, and Robert Koeppe. Multiple testing for neuroimaging via hidden Markov random field. Biometrics, 71(3):741–750, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fogarty:2015:BHR

[898] Colin B. Fogarty, Michael P. Fay, Jennifer A. Flegg, Kasia Stepniewska, Rick M. Fairhurst, and Dylan S. Small. Bayesian hierarchical regression on clearance rates in the presence of “lag” and “tail” phases with an application to malaria parasites. Biometrics, 71(3):751–759, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

deCastro:2015:BPS

[899] Mário de Castro, Ming-Hui Chen, and Yuanye Zhang. Bayesian path specific frailty models for multi-state survival data with applications. Biometrics, 71(3):760–771, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zee:2015:NDS

[900] Jarcy Zee and Sharon X. Xie. Nonparametric discrete survival function estimation with uncertain endpoints using an internal validation subsample. Biometrics, 71(3):772–781, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2015:PRI

[901] Ying Wu and Richard J. Cook. Penalized regression for interval-censored times of disease progression: Selection of HLA markers in psoriatic arthritis. Biometrics, 71(3):782–791, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2015:PCC

[902] Yuan Wang, Brian P. Hobbs, Jianhua Hu, Chaan S. Ng, and Kim-Anh Do. Predictive classification of correlated targets with application to detection of metastatic cancer using functional CT imaging. Biometrics, 71(3):792–802, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cassese:2015:BMI

[903] Alberto Cassese, Michele Guindani, Philipp Antczak, Francesco Falciani, and Marina Vannucci. A Bayesian model for the identification of differentially expressed genes in Daphnia magna exposed to munition pollutants. Biometrics, 71(3):803–811, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hua:2015:EKM

[904] Wen-Yu Hua and Debashis Ghosh. Equivalence of kernel machine regression and kernel distance covariance for multidimensional phenotype association studies. Biometrics, 71(3):812–820, September 2015. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Raffa:2015:MLD

[905] Jesse D. Raffa and Joel A. Dubin. Multivariate longitudinal data analysis with mixed effects hidden Markov models. Biometrics, 71(3):821–831, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lipsitz:2015:TIC

[906] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Debajyoti Sinha, Nathanael Hevelone, Edward Giovannucci, and Jim C. Hu. Testing for independence in J × K contingency tables with complex sample survey data. Biometrics, 71(3):832–840, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2015:SMH

[907] Jingnan Zhang, Yicheng Kang, Yang Yang, and Peihua Qiu. Statistical monitoring of the hand, foot and mouth disease in China. Biometrics, 71(3):841–850, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Silva:2015:CVG

[908] I. R. Silva and M. Kulldorff. Continuous versus group sequential analysis for post-market drug and vaccine safety surveillance. Biometrics, 71(3): 851–858, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dong:2015:RNE

[909] Bin Dong and David Matthews. Rejoinder to “A note on the empirical likelihood confidence band for hazards ratio with covariate adjustment”. Biometrics, 71(3):859–563, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [910].

Zhu:2015:NEL

[910] Shihong Zhu, Yifan Yang, and Mai Zhou. A note on the empirical likelihood confidence band for hazards ratio with covariate adjustment. Biometrics, 71(3):859–563, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [909].

Choi:2015:BRB

[911] Taeryon Choi. Book review: Bayesian networks with examples in R. Biometrics, 71(3):864–865, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2015:BRD

[912] Mira Park. Book review: Ding-Geng, Chen, and Karl, E. Peace, Applied meta-analysis with R. CRC Press. Biometrics, 71(3):864–865, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frangakis:2015:DDT

[913] Constantine E. Frangakis, Tianchen Qian, Zhenke Wu, and Ivan Diaz. Deductive derivation and Turing-computerization of semiparametric efficient estimation. Biometrics, 71(4):867–874, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [914, 915] and rejoinder [916].

Luedtke:2015:DDD

[914] Alexander R. Luedtke, Marco Carone, and Mark J. van der Laan. Discussion of “Deductive derivation and Turing-computerization of semiparametric efficient estimation” by Frangakis et al. Biometrics, 71(4):

875–879, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [913, 913].

Stephens:2015:DDD

[915] David A. Stephens. Discussion of “Deductive derivation and Turingcomputerization of semiparametric efficient estimation” by Frangakis et al. Biometrics, 71(4):880, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [913, 913].

Frangakis:2015:RDD

[916] Constantine E. Frangakis, Tianchen Qian, Zhenke Wu, and Iván Díaz. Rejoinder to discussions on: “Deductive derivation and Turingcomputerization of semiparametric efficient estimation”. Biometrics, 71(4):881–883, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [913].

Ciarleglio:2015:TDB

[917] Adam Ciarleglio, Eva Petkova, R. Todd Ogden, and Thaddeus Tarpey. Treatment decisions based on scalar and functional baseline covariates. Biometrics, 71(4):884–894, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2015:UDL

[918] Yichi Zhang, Eric B. Laber, Anastasios Tsiatis, and Marie Davidian. Using decision lists to construct interpretable and parsimonious treatment regimes. Biometrics, 71(4):895–904, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2015:LAM

[919] Yue Hu and Genevera I. Allen. Local-aggregate modeling for big data via distributed optimization: Applications to neuroimaging. Biometrics, 71(4):905–917, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2015:MKL

[920] Tianle Chen, Donglin Zeng, and Yuanjia Wang. Multiple kernel learning with random effects for predicting longitudinal outcomes and data integration. Biometrics, 71(4):918–928, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2015:MML

[921] Fei Wang, Peter X.-K. Song, and Lu Wang. Merging multiple longitudinal studies with study-specific missing covariates: a joint estimating function approach. Biometrics, 71(4):929–940, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ha:2015:SEP

[922] Jinkyung Ha and Alexander Tsodikov. Semiparametric estimation in the proportional hazard model accounting for a misclassified cause of failure. Biometrics, 71(4):941–949, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:2015:TDP

[923] Abigail R. Smith and Douglas E. Schaubel. Time-dependent prognostic score matching for recurrent event analysis to evaluate a treatment assigned during follow-up. Biometrics, 71(4):950–959, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2015:RLB

[924] Yuanyuan Shen, Tianxi Cai, Yu Chen, Ying Yang, and Jinbo Chen. Retrospective likelihood-based methods for analyzing case-cohort genetic association studies. Biometrics, 71(4):960–968, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Villar:2015:RAR

[925] Sofía S. Villar, James Wason, and Jack Bowden. Response-adaptive randomization for multi-arm clinical trials using the forward looking Gittins index rule. Biometrics, 71(4):969–978, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kennes:2015:IBR

[926] Lieven N. Kennes, William F. Rosenberger, and Ralf-Dieter Hilgers. Inference for blocked randomization under a selection bias model. Biometrics, 71(4):979–984, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2015:ECI

[927] Weizhen Wang and Guogen Shan. Exact confidence intervals for the relative risk and the odds ratio. Biometrics, 71(4):985–995, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dette:2015:DRS

[928] Holger Dette, Stefanie Titoff, Stanislav Volgushev, and Frank Bretz. Dose response signal detection under model uncertainty. Biometrics, 71(4):996–1008, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2015:LBI

[929] Jason Xu, Peter Guttorp, Midori Kato-Maeda, and Vladimir N. Minin. Likelihood-based inference for discretely observed birth-death-shift processes, with applications to evolution of mobile genetic elements. Biometrics, 71(4):1009–1021, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jalilian:2015:MPS

[930] Abdollah Jalilian, Yongtao Guan, Jorge Mateu, and Rasmus Waagepetersen. Multivariate product-shot-noise Cox point process models. Biometrics, 71(4):1022–1033, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Titman:2015:TPE

[931] Andrew C. Titman. Transition probability estimates for non-Markov multi-state models. Biometrics, 71(4):1034–1041, December 2015. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Willis:2015:EDF

[932] Amy Willis and John Bunge. Estimating diversity via frequency ratios. Biometrics, 71(4):1042–1049, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beliveau:2015:AUA

[933] Audrey Béliveau, Richard A. Lockhart, Carl J. Schwarz, and Steven K. Arndt. Adjusting for undercoverage of access-points in creel surveys with fewer overflights. Biometrics, 71(4):1050–1059, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borchers:2015:DOL

[934] D. L. Borchers and R. Langrock. Double-observer line transect surveys with Markov-modulated Poisson process models for animal availability. Biometrics, 71(4):1060–1069, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schofield:2015:CLM

[935] Matthew R. Schofield and Simon J. Bonner. Connecting the latent multinomial. Biometrics, 71(4):1070–1080, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dang:2015:MMP

[936] Utkarsh J. Dang, Ryan P. Browne, and Paul D. McNicholas. Mixtures of multivariate power exponential distributions. Biometrics, 71(4):1081– 1089, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hasenstab:2015:ILT

[937] Kyle Hasenstab, Catherine A. Sugar, Donatello Telesca, Kevin McEvoy, Shafali Jeste, and Damla Sentürk. Identifying longitudinal trends within EEG experiments. Biometrics, 71(4):1090–1100, December 2015. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2015:MBN

[938] Haiming Zhou, Timothy Hanson, and Roland Knapp. Marginal Bayesian nonparametric model for time to disease arrival of threatened amphibian populations. Biometrics, 71(4):1101–1110, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCallum:2015:EBS

[939] Kenneth J. McCallum and Iuliana Ionita-Laza. Empirical Bayes scan statistics for detecting clusters of disease risk variants in genetic studies. Biometrics, 71(4):1111–1120, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Konikoff:2015:SSM

[940] Jacob Konikoff and Ron Brookmeyer. Sample size methods for estimating HIV incidence from cross-sectional surveys. Biometrics, 71(4):1121–1129, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsonaka:2015:MGE

[941] Roula Tsonaka, Diane van der Woude, and Jeanine Houwing-Duistermaat. Marginal genetic effects estimation in family and twin studies using random-effects models. Biometrics, 71(4):1130–1138, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2015:AIV

[942] Qian M. Zhou, Yingye Zheng, Lori B. Chibnik, Elizabeth W. Karlson, and Tianxi Cai. Assessing incremental value of biomarkers with multiphase nested case-control studies. Biometrics, 71(4):1139–1149, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seaman:2015:HMD

[943] Shaun R. Seaman and Ruth H. Keogh. Handling missing data in matched case-control studies using multiple imputation. Biometrics, 71(4):1150– 1159, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daniels:2015:PMM

[944] Michael J. Daniels, Dan Jackson, Wei Feng, and Ian R. White. Pattern mixture models for the analysis of repeated attempt designs. Biometrics, 71(4):1160–1167, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fang:2015:BMA

[945] Q. Fang, W. W. Piegorsch, S. J. Simmons, X. Li, C. Chen, and Y. Wang. Bayesian model-averaged benchmark dose analysis via reparameterized quantal-response models. Biometrics, 71(4):1168–1175, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dattner:2015:MBI

[946] Itai Dattner. A model-based initial guess for estimating parameters in systems of ordinary differential equations. Biometrics, 71(4):1176–1184, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sabourin:2015:PAS

[947] Jeremy A. Sabourin, William Valdar, and Andrew B. Nobel. A permutation approach for selecting the penalty parameter in penalized model selection. Biometrics, 71(4):1185–1194, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2015:MTE

[948] Hongyue Wang, Yun Zhang, Changyong Feng, and Xin M. Tu. On the misuse of Taylor expansion. Biometrics, 71(4):1195, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [910].

Zhao:2015:RMT

[949] Shanshan Zhao and Ross L. Prentice. Rejoinder to “On the Misuse of Taylor Expansion”. Biometrics, 71(4):1195, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [948].

Carriere:2015:BRA

[950] K. C. Carriere. Book review: Antognini, Alessandro B. and Giovagnoli, Alessandra, Adaptive Designs for Sequential Treatment Allocation. Boca Raton, FL: Chapman and Hall/CRC. Biometrics, 71(4):1196–1197, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2015:BRG

[951] Jinheum Kim. Book review: Geskus, Ronald B., Data Analysis with Competing Risks and Intermediate States. New York: Chapman and Hall/CRC. Biometrics, 71(4):1197–1198, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2015:BRBa

[952] Jin Kyung Park. Book review: Becker, Niels G., Modeling to Inform Infectious Disease Control. Boca Raton, FL: Chapman and Hall/CRC. Biometrics, 71(4):1198, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2015:BRBb

[953] Taesung Park. Book review: Bilder, Christopher R. and Loughin, Thomas M., Analysis of Categorical Data with R. Boca Raton, FL: Chapman and Hall/CRC. Biometrics, 71(4):1198–1199, December 2015. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IITa

[954] Anonymous. Issue information — TOC. Biometrics, 72(1):1–2, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIMa

[955] Anonymous. Issue information — masthead. Biometrics, 72(1):3, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIIa

[956] Anonymous. Issue information — info for authors. Biometrics, 72(1): 4, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIEa

[957] Anonymous. Issue information — Ed board. Biometrics, 72(1):5–6, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:RE

[958] Anonymous. Report of the Editors — 2015. Biometrics, 72(1):7–9, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2016:CTP

[959] Xi Lu, Kevin G. Lynch, David W. Oslin, and Susan Murphy. Comparing treatment policies with assistance from the structural nested mean model. Biometrics, 72(1):10–19, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2016:NMT

[960] Zhiwei Zhang, Lei Nie, Guoxing Soon, and Zonghui Hu. New methods for treatment effect calibration, with applications to non-inferiority trials. Biometrics, 72(1):20–29, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2016:ACO

[961] James Y. Dai, Xinyi Cindy Zhang, Ching-Yun Wang, and Charles Kooperberg. Augmented case-only designs for randomized clinical trials with failure time endpoints. Biometrics, 72(1):30–38, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2016:CMH

[962] Zhongxue Chen, Hanwen Huang, and Peihua Qiu. Comparison of multiple hazard rate functions. Biometrics, 72(1):39–45, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rathbun:2016:MEM

[963] Stephen L. Rathbun and Saul Shiffman. Mixed effects models for recurrent events data with partially observed time-varying covariates: Ecological momentary assessment of smoking. Biometrics, 72(1):46–55, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bilgili:2016:BFP

[964] D. Bilgili, D. Ryu, Ö. Ergönül, and N. Ebrahimi. Bayesian framework for parametric bivariate accelerated lifetime modeling and its application to hospital acquired infections. Biometrics, 72(1):56–63, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2016:QRA

[965] Xiaoyan Sun, Limin Peng, Amita Manatunga, and Michele Marcus. Quantile regression analysis of censored longitudinal data with irregular outcome-dependent follow-up. Biometrics, 72(1):64–73, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frumento:2016:PMQ

[966] Paolo Frumento and Matteo Bottai. Parametric modeling of quantile regression coefficient functions. Biometrics, 72(1):74–84, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:2016:DGG

[967] Hung Hung, Yu-Ting Lin, Penweng Chen, Chen-Chien Wang, Su-Yun Huang, and Jung-Ying Tzeng. Detection of gene-gene interactions using multistage sparse and low-rank regression. Biometrics, 72(1):85–94, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2016:OSD

[968] Fangyuan Zhang, Abbas Khalili, and Shili Lin. Optimum study design for detecting imprinting and maternal effects based on partial likelihood. Biometrics, 72(1):95–105, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kipnis:2016:BME

[969] Victor Kipnis, Laurence S. Freedman, Raymond J. Carroll, and Douglas Midthune. A bivariate measurement error model for semicontinuous and continuous variables: Application to nutritional epidemiology. Biometrics, 72(1):106–115, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Farcomeni:2016:GCR

[970] A. Farcomeni. A general class of recapture models based on the conditional capture probabilities. Biometrics, 72(1):116–124, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fegatelli:2016:FBC

[971] Danilo Alunni Fegatelli and Luca Tardella. Flexible behavioral capturerecapture modeling. Biometrics, 72(1):125–135, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Favaro:2016:RGT

[972] Stefano Favaro, Bernardo Nipoti, and Yee Whye Teh. Rediscovery of good-Turing estimators via Bayesian nonparametrics. Biometrics, 72 (1):136–145, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ha:2016:PTS

[973] Min Jin Ha, Wei Sun, and Jichun Xie. PenPC: A two-step approach to estimate the skeletons of high-dimensional directed acyclic graphs. Biometrics, 72(1):146–155, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2016:TRV

[974] Xinyi Lin, Seunggeun Lee, Michael C. Wu, Chaolong Wang, Han Chen, Zilin Li, and Xihong Lin. Test for rare variants by environment interactions in sequencing association studies. Biometrics, 72(1):156–164, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2016:PAR

[975] Yi-Hui Zhou. Pathway analysis for RNA-Seq data using a score-based approach. Biometrics, 72(1):165–174, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2016:CCS

[976] Paul R. Rosenbaum. The cross-cut statistic and its sensitivity to bias in observational studies with ordered doses of treatment. Biometrics, 72 (1):175–183, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2016:NBM

[977] Jing Zhou, Amy H. Herring, Anirban Bhattacharya, Andrew F. Olshan, David B. Dunson, and The National Birth Defects Prevention Study. Nonparametric Bayes modeling for case control studies with many predictors. Biometrics, 72(1):184–192, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lum:2016:BJM

[978] Kirsten J. Lum, Rajeshwari Sundaram, Germaine M. Buck Louis, and Thomas A. Louis. A Bayesian joint model of menstrual cycle length and fecundity. Biometrics, 72(1):193–203, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2016:JFM

[979] Lei Liu, Xuelin Huang, Alex Yaroshinsky, and Janice N. Cormier. Joint frailty models for zero-inflated recurrent events in the presence of a terminal event. Biometrics, 72(1):204–214, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2016:RMS

[980] Lihui Zhao, Brian Claggett, Lu Tian, Hajime Uno, Marc A. Pfeffer, Scott D. Solomon, Lorenzo Trippa, and L. J. Wei. On the restricted mean survival time curve in survival analysis. Biometrics, 72(1):215– 221, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2016:FCE

[981] Lianming Wang, Christopher S. McMahan, Michael G. Hudgens, and Zaina P. Qureshi. A flexible, computationally efficient method for fitting the proportional hazards model to interval-censored data. Biometrics, 72(1):222–231, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsai:2016:CIR

[982] Tsung-Hsien Tsai, Wei-Yann Tsai, Yunchan Chi, and Sheng-Mao Chang. Confidence intervals for the ratio of two median residual lifetimes with left-truncated and right-censored data. Biometrics, 72(1):232–241, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2016:TSS

[983] Hanzhi Zhou, Michael R. Elliott, and Trviellore E. Raghunathan. A two-step semiparametric method to accommodate sampling weights in multiple imputation. Biometrics, 72(1):242–252, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nooraee:2016:AML

[984] Nazanin Nooraee, Fentaw Abegaz, Johan Ormel, Ernst Wit, and Edwin R. van den Heuvel. An approximate marginal logistic distribution for the analysis of longitudinal ordinal data. Biometrics, 72(1):253–261, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lyons:2016:PSS

[985] James E. Lyons, William L. Kendall, J. Andrew Royle, Sarah J. Converse, Brad A. Andres, and Joseph B. Buchanan. Population size and stopover duration estimation using mark-resight data and Bayesian analysis of a superpopulation model. Biometrics, 72(1):262–271, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeBin:2016:SVB

[986] Riccardo De Bin, Silke Janitza, Willi Sauerbrei, and Anne-Laure Boulesteix. Subsampling versus bootstrapping in resampling-based model selection for multivariable regression. Biometrics, 72(1):272–280, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alexeeff:2016:BMQ

[987] Stacey E. Alexeeff, Gabriele G. Pfister, and Doug Nychka. A Bayesian model for quantifying the change in mortality associated with future ozone exposures under climate change. Biometrics, 72(1):281–288, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:2016:EEP

[988] Vivekananda Roy, Evangelos Evangelou, and Zhengyuan Zhu. Efficient estimation and prediction for the Bayesian binary spatial model with flexible link functions. Biometrics, 72(1):289–298, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Malinovsky:2016:RRN

[989] Yaakov Malinovsky, Paul S. Albert, and Anindya Roy. Reader reaction: a note on the evaluation of group testing algorithms in the presence of misclassification. Biometrics, 72(1):299–302, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinders [990, 991].

McMahan:2016:RRR

[990] Christopher S. McMahan, Joshua M. Tebbs, and Christopher R. Bilder. Rejoinder to “Reader reaction: a note on the evaluation of group testing algorithms in the presence of misclassification”. Biometrics, 72(1):303– 304, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [989].

Hudgens:2016:RRR

[991] Michael G. Hudgens. Rejoinder to “Reader reaction: a note on the evaluation of group testing algorithms in the presence of misclassification”. Biometrics, 72(1):304, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [990].

Kim:2016:BRW

[992] Seonwoo Kim. Book review: He, Weili, Pinheiro, José, and Kuznetsova, Olga M., Practical Considerations for Adaptive Trial Design and Implementation. Springer, New York. Biometrics, 72(1):305–306, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joo:2016:BRE

[993] Jungnam Joo. Book review: Edited by Sverdlov O. Modern. Adaptive Randomized Clinical Trials: Statistical and Practical Aspects. Boca Raton, FL. Chapman and Hall/CRC. Biometrics, 72(1):306–307, March 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IITb

[994] Anonymous. Issue information — TOC. Biometrics, 72(2):309–310, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIMb

[995] Anonymous. Issue information — masthead. Biometrics, 72(2):311, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIIb

[996] Anonymous. Issue information — info for authors. Biometrics, 72(2): 312, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIEb

[997] Anonymous. Issue information — Ed board. Biometrics, 72(2):313– 314, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Niu:2016:MIA

[998] Mu Niu, Paul G. Blackwell, and Anna Skarin. Modeling interdependent animal movement in continuous time. Biometrics, 72(2):315–324, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mukhopadhyay:2016:LSS

[999] Subhadeep Mukhopadhyay. Large-scale signal detection: a unified perspective. Biometrics, 72(2):325–334, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:2016:EAA

[1000] Grant D. Brown, Jacob J. Oleson, and Aaron T. Porter. An empirically adjusted approach to reproductive number estimation for stochastic compartmental models: a case study of two Ebola outbreaks. Biometrics, 72(2):335–343, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drovandi:2016:ASB

[1001] Christopher C. Drovandi and Roy A. McCutchan. Alive SMC^{2}: Bayesian model selection for low-count time series models with intractable likelihoods. Biometrics, 72(2):344–353, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2016:EDM

[1002] Tao Wang, Mengjie Chen, and Hongyu Zhao. Estimating DNA methylation levels by joint modeling of multiple methylation profiles from microarray data. Biometrics, 72(2):354–363, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kong:2016:TEM

[1003] Dehan Kong, Arnab Maity, Fang-Chi Hsu, and Jung-Ying Tzeng. Testing and estimation in marker-set association study using semiparametric quantile regression kernel machine. Biometrics, 72(2):364–371, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Payne:2016:KMT

[1004] Rebecca Payne, Matey Neykov, Majken Karoline Jensen, and Tianxi Cai. Kernel machine testing for risk prediction with stratified case cohort studies. Biometrics, 72(2):372–381, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rivera:2016:UWC

[1005] C. Rivera and T. Lumley. Using the whole cohort in the analysis of countermatched samples. Biometrics, 72(2):382–391, June 2016. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Casalicchio:2016:RBP

[1006] Giuseppe Casalicchio, Bernd Bischl, Anne-Laure Boulesteix, and Matthias Schmid. The residual-based predictiveness curve: a visual tool to assess the performance of prediction models. Biometrics, 72(2):392– 401, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2016:HTM

[1007] Yen-Tsung Huang and Wen-Chi Pan. Hypothesis test of mediation effect in causal mediation model with high-dimensional continuous mediators. Biometrics, 72(2):402–413, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2016:OBA

[1008] Yifan Zhang, Lorenzo Trippa, and Giovanni Parmigiani. Optimal Bayesian adaptive trials when treatment efficacy depends on biomarkers. Biometrics, 72(2):414–421, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diaz:2016:EPA

[1009] Iván Díaz, Elizabeth Colantuoni, and Michael Rosenblum. Enhanced precision in the analysis of randomized trials with ordinal outcomes. Biometrics, 72(2):422–431, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dutta:2016:RST

[1010] Sandipan Dutta and Somnath Datta. A rank-sum test for clustered data when the number of subjects in a group within a cluster is informative. Biometrics, 72(2):432–440, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bible:2016:CAR

[1011] Joe Bible, James D. Beck, and Somnath Datta. Cluster adjusted regression for displaced subject data (CARDS): Marginal inference under potentially informative temporal cluster size profiles. Biometrics, 72(2): 441–451, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DiRienzo:2016:PCS

[1012] Albert Gregory DiRienzo. Parsimonious covariate selection with censored outcomes. Biometrics, 72(2):452–462, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kjaersgaard:2016:IVM

[1013] Maiken I. S. Kjaersgaard and Erik T. Parner. Instrumental variable method for time-to-event data using a pseudo-observation approach. Biometrics, 72(2):463–472, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Garcia-Zattera:2016:FAM

[1014] María José García-Zattera, Alejandro Jara, and Arnost Komárek. A flexible AFT model for misclassified clustered interval-censored data. Biometrics, 72(2):473–483, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2016:STG

[1015] Sen Zhao and Ali Shojaie. A significance test for graph-constrained estimation. Biometrics, 72(2):484–493, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bansal:2016:TMH

[1016] Naveen K. Bansal, Gholamhossein G. Hamedani, and Mehdi Maadooliat. Testing multiple hypotheses with skewed alternatives. Biometrics, 72(2): 494–502, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rathnayake:2016:TBF

[1017] Lasitha N. Rathnayake and Pankaj K. Choudhary. Tolerance bands for functional data. Biometrics, 72(2):503–512, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xiao:2016:MPU

[1018] Luo Xiao, Bing He, Annemarie Koster, Paolo Caserotti, Brittney Lange-Maia, Nancy W. Glynn, Tamara B. Harris, and Ciprian M. Crainiceanu. Movement prediction using accelerometers in a human population. Biometrics, 72(2):513–524, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shu:2016:SMC

[1019] Xu Shu and Douglas E. Schaubel. Semiparametric methods to contrast gap time survival functions: Application to repeat kidney transplantation. Biometrics, 72(2):525–534, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2016:NMA

[1020] Chi Hyun Lee, Xianghua Luo, Chiung-Yu Huang, Todd E. DeFor, Claudio G. Brunstein, and Daniel J. Weisdorf. Nonparametric methods for analyzing recurrent gap time data with application to infections after hematopoietic cell transplant. Biometrics, 72(2):535–545, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2016:MPE

[1021] Yan Li, Orestis A. Panagiotou, Amanda Black, Dandan Liao, and Sholom Wacholder. Multivariate piecewise exponential survival modeling. Biometrics, 72(2):546–553, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2016:UBS

[1022] Lili Zhao, Dai Feng, Guoan Chen, and Jeremy M. G. Taylor. A unified Bayesian semiparametric approach to assess discrimination ability in survival analysis. Biometrics, 72(2):554–562, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2016:MAS

[1023] Yen-Tsung Huang and Tianxi Cai. Mediation analysis for survival data using semiparametric probit models. Biometrics, 72(2):563–574, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2016:BPM

[1024] Junsheng Ma, Francesco C. Stingo, and Brian P. Hobbs. Bayesian predictive modeling for genomic based personalized treatment selection. Biometrics, 72(2):575–583, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Satagopan:2016:BEB

[1025] Jaya M. Satagopan, Ananda Sen, Qin Zhou, Qing Lan, Nathaniel Rothman, Hilde Langseth, and Lawrence S. Engel. Bayes and empirical Bayes methods for reduced rank regression models in matched case-control studies. Biometrics, 72(2):584–595, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2016:BHF

[1026] Shuo Chen, F. DuBois Bowman, and Helen S. Mayberg. A Bayesian hierarchical framework for modeling brain connectivity for neuroimaging data. Biometrics, 72(2):596–605, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choo-Wosoba:2016:MRM

[1027] Hyoyoung Choo-Wosoba, Steven M. Levy, and Somnath Datta. Marginal regression models for clustered count data based on zero-inflated Conway–Maxwell–Poisson distribution with applications. Biometrics, 72 (2):606–618, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2016:SMA

[1028] Ruth King and Roland Langrock. Semi-Markov arnason-Schwarz models. Biometrics, 72(2):619–628, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2016:USB

[1029] Qianying Liu, Lin S. Chen, Dan L. Nicolae, and Brandon L. Pierce. A unified set-based test with adaptive filtering for gene-environment interaction analyses. Biometrics, 72(2):629–638, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:2016:FDR

[1030] Kun Liang. False discovery rate estimation for large-scale homogeneous discrete p-values. Biometrics, 72(2):639–648, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giraud:2016:COD

[1031] Christophe Giraud, Clément Calenge, Camille Coron, and Romain Julliard. Capitalizing on opportunistic data for monitoring relative abundances of species. Biometrics, 72(2):649–658, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:2016:BRF

[1032] Alan Agresti. Book review: Friendly, Michael and Meyer, David. Discrete Data Analysis with R: Visualization and Modeling Techniques for Categorical and Count Data. Boca Raton, FL: CRC Press. Biometrics, 72(2):659–660, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

vanLieshout:2016:BRB

[1033] M. N. M. van Lieshout. Book review: Baddeley, Adrian, Rubak, Ege, and Turner, Rolf, Spatial Point Patterns: Methodology and Applications with R. Boca Raton, FL: CRC Press. Biometrics, 72(2):660–661, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rolles:2016:BRW

[1034] Silke Rolles. Book review: Wai-Yuan, Tan, Stochastic Models with Applications to Genetics, Cancers, AIDS, and Other Biomedical Systems, 2nd ed. Singapore: World Scientific Publishing Co. Biometrics, 72(2): 661, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ankerst:2016:BRH

[1035] Donna Pauler Ankerst. Book review: Harron, Katie, Goldstein, Harvey, and Dibben, Chris, eds., Methodological Developments in Data Linkage. New York: Wiley. Biometrics, 72(2):661–662, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IITc

[1036] Anonymous. Issue information — TOC. Biometrics, 72(3):663–664, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIMc

[1037] Anonymous. Issue information — masthead. Biometrics, 72(3):665, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIIc

[1038] Anonymous. Issue information — info for authors. Biometrics, 72(3):666, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIEc

[1039] Anonymous. Issue information — Ed board. Biometrics, 72(3):667–668, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2016:ITA

[1040] Ariel Alonso, Wim Van der Elst, Geert Molenberghs, Marc Buyse, and Tomasz Burzykowski. An information-theoretic approach for the evaluation of surrogate endpoints based on causal inference. Biometrics, 72(3): 669–677, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huque:2016:SRC

[1041] Md Hamidul Huque, Howard D. Bondell, Raymond J. Carroll, and Louise M. Ryan. Spatial regression with covariate measurement error: a semiparametric approach. Biometrics, 72(3):678–686, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moller:2016:MST

[1042] Jesper Møller, Mohammad Ghorbani, and Ege Rubak. Mechanistic spatio-temporal point process models for marked point processes, with a view to forest stand data. Biometrics, 72(3):687–696, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohning:2016:FRR

[1043] Dankmar Böhning, Irene Rocchetti, Marco Alfó, and Heinz Holling. A flexible ratio regression approach for zero-truncated capture-recapture counts. Biometrics, 72(3):697–706, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2016:FLF

[1044] Dan Li, Xia Wang, Lizhen Lin, and Dipak K. Dey. Flexible link functions in nonparametric binary regression with Gaussian process priors. Biometrics, 72(3):707–719, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2016:RNE

[1045] Ying Zhang, Gang Cheng, and Wanzhu Tu. Robust nonparametric estimation of monotone regression functions with interval-censored observations. Biometrics, 72(3):720–730, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2016:MSI

[1046] Brent A. Johnson, Qi Long, Yijian Huang, Kari Chansky, and Mary Redman. Model selection and inference for censored lifetime medical expenditures. Biometrics, 72(3):731–741, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2016:ETE

[1047] Shuli Li and Robert J. Gray. Estimating treatment effect in a proportional hazards model in randomized clinical trials with all-or-nothing compliance. Biometrics, 72(3):742–750, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Su:2016:SEC

[1048] Xiaogang Su, Chalani S. Wijayasinghe, Juanjuan Fan, and Ying Zhang. Sparse estimation of Cox proportional hazards models via approximated information criteria. Biometrics, 72(3):751–759, September 2016. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graffeo:2016:LRT

[1049] Nathalie Grafféo, Fabienne Castell, Aurélien Belot, and Roch Giorgi. A log-rank-type test to compare net survival distributions. Biometrics, 72(3):760–769, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2016:NFD

[1050] Jing Yang and Limin Peng. A new flexible dependence measure for semicompeting risks. Biometrics, 72(3):770–779, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2016:NAB

[1051] Chiung-Yu Huang, Chenguang Wang, and Mei-Cheng Wang. Nonparametric analysis of bivariate gap time with competing risks. Biometrics, 72(3):780–790, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2016:ESD

[1052] Sung Won Han and Hua Zhong. Estimation of sparse directed acyclic graphs for multivariate counts data. Biometrics, 72(3):791–803, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonafede:2016:MOH

[1053] Elisabetta Bonafede, Franck Picard, Stéphane Robin, and Cinzia Viroli. Modeling overdispersion heterogeneity in differential expression analysis using mixtures. Biometrics, 72(3):804–814, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2016:SMS

[1054] Jianhua Hu, Liwen Zhang, and Huixia Judy Wang. Sequential model selection-based segmentation to detect DNA copy number variation. Biometrics, 72(3):815–826, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2016:LCT

[1055] Stuart G. Baker. The latent class twin method. Biometrics, 72(3):827– 834, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pini:2016:ITP

[1056] Alessia Pini and Simone Vantini. The interval testing procedure: a general framework for inference in functional data analysis. Biometrics, 72(3):835–845, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2016:IFP

[1057] Zhenhua Lin, Liangliang Wang, and Jiguo Cao. Interpretable functional principal component analysis. Biometrics, 72(3):846–854, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wallace:2016:MAD

[1058] Michael P. Wallace, Erica E. M. Moodie, and David A. Stephens. Model assessment in dynamic treatment regimen estimation via double robustness. Biometrics, 72(3):855–864, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chakraborty:2016:EOS

[1059] Bibhas Chakraborty, Palash Ghosh, Erica E. M. Moodie, and A. John Rush. Estimating optimal shared-parameter dynamic regimens with application to a multistage depression clinical trial. Biometrics, 72(3):865– 876, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2016:PEP

[1060] Junlong Li, Lihui Zhao, Lu Tian, Tianxi Cai, Brian Claggett, Andrea Callegaro, Benjamin Dizier, Bart Spiessens, Fernando Ulloa-Montoya, and Lee-Jen Wei. A predictive enrichment procedure to identify potential responders to a new therapy for randomized, comparative controlled clinical studies. Biometrics, 72(3):877–887, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tayob:2016:UEB

[1061] Nabihah Tayob, Kim-Anh Do, and Ziding Feng. Unbiased estimation of biomarker panel performance when combining training and testing data in a group sequential design. Biometrics, 72(3):888–896, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2016:AIA

[1062] Ming Wang and Qi Long. Addressing issues associated with evaluating prediction models for survival endpoints based on the concordance statistic. Biometrics, 72(3):897–906, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krol:2016:JML

[1063] Agnieszka Król, Loïc Ferrer, Jean-Pierre Pignon, Cécile Proust-Lima, Michel Ducreux, Olivier Bouché, Stefan Michiels, and Virginie Rondeau. Joint model for left-censored longitudinal data, recurrent events and terminal event: Predictive abilities of tumor burden for cancer evolution with application to the FFCD 2000–05 trial. Biometrics, 72(3):907–916, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2016:CJM

[1064] Sungduk Kim and Paul S. Albert. A class of joint models for multivariate longitudinal measurements and a binary event. Biometrics, 72(3):917– 925, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ramchandani:2016:GRT

[1065] Ritesh Ramchandani, David A. Schoenfeld, and Dianne M. Finkelstein. Global rank tests for multiple, possibly censored, outcomes. Biometrics, 72(3):926–935, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shaby:2016:TGM

[1066] Benjamin A. Shaby, Gaia Skibinski, Michael Ando, Eva S. LaDow, and Steven Finkbeiner. A three-groups model for high-throughput survival screens. Biometrics, 72(3):936–944, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shi:2016:NRE

[1067] Jingchunzi Shi and Seunggeun Lee. A novel random effect model for GWAS meta-analysis and its application to trans-ethnic meta-analysis. Biometrics, 72(3):945–954, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2016:BIL

[1068] Yanxun Xu, Peter Müller, and Donatello Telesca. Bayesian inference for latent biologic structure with determinantal point processes (DPP). Biometrics, 72(3):955–964, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lyles:2016:EDS

[1069] Robert H. Lyles, Emily M. Mitchell, Clarice R. Weinberg, David M. Umbach, and Enrique F. Schisterman. An efficient design strategy for logistic regression using outcome- and covariate-dependent pooling of biospecimens prior to assay. Biometrics, 72(3):965–975, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mehtala:2016:EIH

[1070] Juha Mehtälä, Ron Dagan, and Kari Auranen. Estimation and interpretation of heterogeneous vaccine efficacy against recurrent infections. Biometrics, 72(3):976–985, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Todem:2016:MMM

[1071] David Todem, KyungMann Kim, and Wei-Wen Hsu. Marginal mean models for zero-inflated count data. Biometrics, 72(3):986–994, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Devlin:2016:BMR

[1072] Sean M. Devlin, Irina Ostrovnaya, and Mithat Gönen. Boomerang: a method for recursive reclassification. Biometrics, 72(3):995–1002, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chan:2016:RRI

[1073] Kwun Chuen Gary Chan. Reader reaction: Instrumental variable additive hazards models with exposure-dependent censoring. Biometrics, 72(3):1003–1005, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steyerberg:2016:BRH

[1074] E. W. Steyerberg. Book review: Harrell, Frank E., Regression Modeling Strategies: With Applications, to Linear Models, Logistic and Ordinal Regression, and Survival Analysis, 2nd ed. Heidelberg: Springer. Biometrics, 72(3):1006–1007, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Paulus:2016:BRB

[1075] Julie Paulus. Book review: Broemeling, Lyle D., Bayesian Methods in Epidemiology. Boca Raton, FL: CRC Press. Biometrics, 72(3):1006– 1008, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tellier:2016:BRD

[1076] Aurélien Tellier. Book review: Datta, Somnath and Nettleton, Dan, Statistical Analysis of Next-Generation Sequencing Data. Heidelberg: Springer. Biometrics, 72(3):1008, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ankerst:2016:BRF

[1077] Donna Pauler Ankerst. Book review: Fitzmaurice, Garrett M., Laird, Nan M., and Ware, James H., Applied Longitudinal Analysis. Hoboken: Wiley. Biometrics, 72(3):1009, September 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IITd

[1078] Anonymous. Issue information — TOC. Biometrics, 72(4):1011–1012, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIMd

[1079] Anonymous. Issue information — masthead. Biometrics, 72(4):1013, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIId

[1080] Anonymous. Issue information — info for authors. Biometrics, 72(4): 1014, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:IIEd

[1081] Anonymous. Issue information — Ed board. Biometrics, 72(4):1015– 1016, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2016:IPM

[1082] Yuanyuan Shen and Tianxi Cai. Identifying predictive markers for personalized treatment selection. Biometrics, 72(4):1017–1025, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schnell:2016:BCS

[1083] Patrick M. Schnell, Qi Tang, Walter W. Offen, and Bradley P. Carlin. A Bayesian credible subgroups approach to identifying patient subgroups with positive treatment effects. Biometrics, 72(4):1026–1036, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Payne:2016:RRP

[1084] Rebecca Payne, Ming Yang, Yingye Zheng, Majken K. Jensen, and Tianxi Cai. Robust risk prediction with biomarkers under two-phase stratified cohort design. Biometrics, 72(4):1037–1045, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gabriel:2016:CBT

[1085] Erin E. Gabriel, Michael J. Daniels, and M. Elizabeth Halloran. Comparing biomarkers as trial level general surrogates. Biometrics, 72(4): 1046–1054, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2016:PSM

[1086] Shu Yang, Guido W. Imbens, Zhanglin Cui, Douglas E. Faries, and Zbigniew Kadziola. Propensity score matching and subclassification in observational studies with multi-level treatments. Biometrics, 72(4):1055– 1065, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prague:2016:AIC

[1087] Melanie Prague, Rui Wang, Alisa Stephens, Eric Tchetgen Tchetgen, and Victor DeGruttola. Accounting for interactions and complex intersubject dependency in estimating treatment effect in cluster-randomized trials with missing outcomes. Biometrics, 72(4):1066–1077, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2016:SIT

[1088] Dan-Yu Lin, Jianjian Gong, Paul Gallo, Paul H. Bunn, and David Couper. Simultaneous inference on treatment effects in survival studies with factorial designs. Biometrics, 72(4):1078–1085, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2016:SEC

[1089] Yongxiu Cao, Jian Huang, Yanyan Liu, and Xingqiu Zhao. Sieve estimation of Cox models with latent structures. Biometrics, 72(4):1086–1097, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2016:CIH

[1090] Dan-Yu Lin, Luyan Dai, Gang Cheng, and Martin Oliver Sailer. On confidence intervals for the hazard ratio in randomized clinical trials. Biometrics, 72(4):1098–1102, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2016:RAC

[1091] Peijie Wang, Hui Zhao, and Jianguo Sun. Regression analysis of case k interval-censored failure time data in the presence of informative censoring. Biometrics, 72(4):1103–1112, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2016:MRA

[1092] X. Joan Hu and Rhonda J. Rosychuk. Marginal regression analysis of recurrent events with coarsened censoring times. Biometrics, 72(4):1113– 1122, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rouanet:2016:JLC

[1093] Anaïs Rouanet, Pierre Joly, Jean-François Dartigues, Cécile Proust-Lima, and Hélène Jacqmin-Gadda. Joint latent class model for longitudinal data and interval-censored semi-competing events: Application to dementia. Biometrics, 72(4):1123–1135, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2016:IEU

[1094] Yih-Huei Huang, Wen-Han Hwang, and Fei-Yin Chen. Improving efficiency using the Rao-Blackwell theorem in corrected and conditional score estimation methods for joint models. Biometrics, 72(4):1136–1144, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2016:SII

[1095] Jialiang Li, Qi Zheng, Limin Peng, and Zhipeng Huang. Survival impact index and ultrahigh-dimensional model-free screening with survival outcomes. Biometrics, 72(4):1145–1154, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fang:2016:FVS

[1096] Zaili Fang, Inyoung Kim, and Patrick Schaumont. Flexible variable selection for recovering sparsity in nonadditive nonparametric models. Biometrics, 72(4):1155–1163, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2016:VSB

[1097] Wenning Feng, Abdhi Sarkar, Chae Young Lim, and Tapabrata Maiti. Variable selection for binary spatial regression: Penalized quasilikelihood approach. Biometrics, 72(4):1164–1172, December 2016. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wong:2016:OEP

[1098] Kin Yau Wong, Yair Goldberg, and Jason P. Fine. Oracle estimation of parametric models under boundary constraints. Biometrics, 72(4): 1173–1183, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2016:FLA

[1099] Fei Wang, Lu Wang, and Peter X.-K. Song. Fused lasso with the adaptation of parameter ordering in combining multiple studies with repeated measurements. Biometrics, 72(4):1184–1193, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agniel:2016:ETM

[1100] Denis Agniel, Katherine P. Liao, and Tianxi Cai. Estimation and testing for multiple regulation of multivariate mixed outcomes. Biometrics, 72 (4):1194–1205, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Imholte:2016:BHM

[1101] Gregory Imholte and Raphael Gottardo. Bayesian hierarchical modeling for subject-level response classification in peptide microarray immunoassays. Biometrics, 72(4):1206–1215, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2016:CWG

[1102] Yiming Hu and Hongyu Zhao. CCor: a whole genome network-based similarity measure between two genes. Biometrics, 72(4):1216–1225, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2016:SSS

[1103] Pei-Sheng Lin, Yi-Hung Kung, and Murray Clayton. Spatial scan statistics for detection of multiple clusters with arbitrary shapes. Biometrics, 72(4):1226–1234, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haines:2016:MLE

[1104] Linda M. Haines. Maximum likelihood estimation for N-mixture models. Biometrics, 72(4):1235–1245, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Manrique-Vallier:2016:BPS

[1105] Daniel Manrique-Vallier. Bayesian population size estimation using Dirichlet process mixtures. Biometrics, 72(4):1246–1254, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lloyd-Jones:2016:MTD

[1106] Luke R. Lloyd-Jones, Hien D. Nguyen, Geoffrey J. McLachlan, Wayne Sumpton, and You-Gan Wang. Mixture of time-dependent growth models with an application to blue swimmer crab length-frequency data. Biometrics, 72(4):1255–1265, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rueda:2016:CPR

[1107] Cristina Rueda, Miguel A. Fernández, Sandra Barragán, Kanti V. Mardia, and Shyamal D. Peddada. Circular piecewise regression with applications to cell-cycle data. Biometrics, 72(4):1266–1274, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2016:SIV

[1108] Xinchao Luo, Lixing Zhu, and Hongtu Zhu. Single-index varying coefficient model for functional responses. Biometrics, 72(4):1275–1284, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wan:2016:CSO

[1109] Fang Wan, Wei Liu, Frank Bretz, and Yang Han. Confidence sets for optimal factor levels of a response surface. Biometrics, 72(4):1285–1293, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2016:ECC

[1110] Shen-Ming Lee, Wen-Han Hwang, and Jean de Dieu Tapsoba. Estimation in closed capture-recapture models when covariates are missing at random. Biometrics, 72(4):1294–1304, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dennis:2016:GAI

[1111] Emily B. Dennis, Byron J. T. Morgan, Stephen N. Freeman, Tom M. Brereton, and David B. Roy. A generalized abundance index for seasonal invertebrates. Biometrics, 72(4):1305–1314, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McCulloch:2016:BUE

[1112] Charles E. McCulloch, John M. Neuhaus, and Rebecca L. Olin. Biased and unbiased estimation in longitudinal studies with informative visit processes. Biometrics, 72(4):1315–1324, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2016:LNR

[1113] Seokho Lee, Hyejin Shin, and Sang Han Lee. Label-noise resistant logistic regression for functional data classification with an application to Alzheimer’s disease study. Biometrics, 72(4):1325–1335, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fraser:2016:AMR

[1114] Raphael André Fraser, Stuart R. Lipsitz, Debajyoti Sinha, Garrett M. Fitzmaurice, and Yi Pan. Approximate median regression for complex survey data with skewed response. Biometrics, 72(4):1336–1347, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsu:2016:SST

[1115] Wei-Wen Hsu, David Todem, and KyungMann Kim. A sup-score test for the cure fraction in mixture models for long-term survivors. Biometrics, 72(4):1348–1357, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zoh:2016:PPC

[1116] Roger S. Zoh, Bani Mallick, Ivan Ivanov, Veera Baladandayuthapani, Ganiraju Manyam, Robert S. Chapkin, Johanna W. Lampe, and Raymond J. Carroll. PCAN: Probabilistic correlation analysis of two nonnormal data sets. Biometrics, 72(4):1358–1368, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Potgieter:2016:MRM

[1117] Cornelis J. Potgieter, Rubin Wei, Victor Kipnis, Laurence S. Freedman, and Raymond J. Carroll. Moment reconstruction and moment-adjusted imputation when exposure is generated by a complex, nonlinear random effects modeling process. Biometrics, 72(4):1369–1377, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2016:MAF

[1118] Guang Yang, Dungang Liu, Junyuan Wang, and Min ge Xie. Metaanalysis framework for exact inferences with application to the analysis of rare events. Biometrics, 72(4):1378–1386, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shepherd:2016:BRI

[1119] Bryan E. Shepherd, Ryan Jarrett, and Lingjun Fu. Book review: Imbens, Guido, Rubin, Donald, Causal Inference for Statistics, Social, and Biomedical Sciences: an Introduction. New York: Cambridge University Press. Biometrics, 72(4):1387–1388, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gelfond:2016:BRG

[1120] Jonathan A. L. Gelfond. Book review: George Tseng, Debashish Ghosh, Xianghon Jasmine Zhou, Eds. Integrating Omics Data. New York: Cambridge University Press. Biometrics, 72(4):1388–1389, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laake:2016:BRB

[1121] Jeff Laake. Book review: Buckland, S. T., Rexstad, E. A., Marques, T. A., Oedekoven, C. S., Distance Sampling: Methods and Applications. Heidelberg: Springer. Biometrics, 72(4):1389–1390, December 2016. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stangl:2016:BRG

[1122] Dalene Stangl. Book review: Geelbek, Amanda L., Olkin, Ingram, Gel, Yulia R., Leadership and Women in Statistics. Boca Raton: CRC Press. Biometrics, 72(4):1390–1391, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2016:A

[1123] Anonymous. Acknowledgements. Biometrics, 72(4):1392–1393, December 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IITa

[1124] Anonymous. Issue information — TOC. Biometrics, 73(1):1–2, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIMa

[1125] Anonymous. Issue information — masthead. Biometrics, 73(1):3, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIIa

[1126] Anonymous. Issue information — info for authors. Biometrics, 73(1): 4, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIEa

[1127] Anonymous. Issue information — Ed board. Biometrics, 73(1):5–6, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:RE

[1128] Anonymous. Report of the Editors — 2016. Biometrics, 73(1):7–9, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chi:2017:CB

[1129] Eric C. Chi, Genevera I. Allen, and Richard G. Baraniuk. Convex biclustering. Biometrics, 73(1):10–19, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2017:TOR

[1130] Wenjing Yang and Yuhong Yang. Toward an objective and reproducible model choice via variable selection deviation. Biometrics, 73(1):20–30, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:2017:CLC

[1131] Jinyuan Chang, Wen Zhou, Wen-Xin Zhou, and Lan Wang. Comparing large covariance matrices under weak conditions on the dependence structure and its application to gene clustering. Biometrics, 73(1):31– 41, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fu:2017:SMD

[1132] Rong Fu, Pei Wang, Weiping Ma, Ayumu Taguchi, Chee-Hong Wong, Qing Zhang, Adi Gazdar, Samir M. Hanash, Qinghua Zhou, Hua Zhong, and Ziding Feng. A statistical method for detecting differentially expressed SNVs based on next-generation RNA-seq data. Biometrics, 73 (1):42–51, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2017:REM

[1133] Jincheol Park and Shili Lin. A random effect model for reconstruction of spatial chromatin structure. Biometrics, 73(1):52–62, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drikvandi:2017:DMR

[1134] Reza Drikvandi, Geert Verbeke, and Geert Molenberghs. Diagnosing misspecification of the random-effects distribution in mixed models. Biometrics, 73(1):63–71, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2017:JPL

[1135] Sehee Kim, Donglin Zeng, and Jeremy M. G. Taylor. Joint partially linear model for longitudinal data with informative drop-outs. Biometrics, 73 (1):72–82, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maziarz:2017:LPT

[1136] Marlena Maziarz, Patrick Heagerty, Tianxi Cai, and Yingye Zheng. On longitudinal prediction with time-to-event outcome: Comparison of modeling options. Biometrics, 73(1):83–93, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2017:MIC

[1137] Yuanshan Wu and Guosheng Yin. Multiple imputation for cure rate quantile regression with censored data. Biometrics, 73(1):94–103, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gouskova:2017:NAC

[1138] Natalia A. Gouskova, Feng-Chang Lin, and Jason P. Fine. Nonparametric analysis of competing risks data with event category missing at random. Biometrics, 73(1):104–113, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1256].

Crowley:2017:BRR

[1139] John Crowley. Book review: Rosenberger, William F., Lachin, John M., Randomization in Clinical Trials: Theory and Practice, 2nd edition. Hoboken: John Wiley and Sons. Biometrics, 73(1):1054–1055, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Panagiotou:2017:BRM

[1140] Orestis A. Panagiotou. Book review: Matsui, Shigeyuki, Buyse, Marc, Simon, Richard, Eds. Design and Analysis of Clinical Trials for Predictive Medicine. Boca Raton: CRC Press. Biometrics, 73(1):1055–1057, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cooley:2017:BRD

[1141] Dan Cooley. Book review: Dey, Dipak K., Yan, Jun, Eds. Extreme Value Modeling and Risk Analysis: Methods and Applications. Boca Raton: CRC Press. Biometrics, 73(1):1057–1058, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finkelstein:2017:BRC

[1142] Dianne Finkelstein. Book review: Chasan-Taber, Lisa. Writing Dissertation and Grant Proposals: Epidemiology, Preventive Medicine and Biostatistics. Boca Raton: CRC Press. Biometrics, 73(1):1058, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hothorn:2017:BRB

[1143] Torsten Hothorn. Book review: Buehlmann, Peter, Drineas, Petros, Kane, Michael, van der Laan, Mark. Handbook of Big Data. Boca Raton: CRC Press. Biometrics, 73(1):1058–1059, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Belin:2017:BRR

[1144] Thomas R. Belin. Book review: RAGHUNATHAN, TRIVELLORE. Missing Data Analysis in Practice. Boca Raton: CRC Press. Biometrics, 73(1):1059–1060, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kauermann:2017:BRA

[1145] Goeran Kauermann. Book review: Agresti, Alan. Foundations of Linear and Generalized Linear Models. Hoboken: Wiley. Biometrics, 73(1): 1060–1061, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IITb

[1146] Anonymous. Issue information — TOC. Biometrics, 73(2):361–362, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIMb

[1147] Anonymous. Issue information — masthead. Biometrics, 73(2):363, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIIb

[1148] Anonymous. Issue information — info for authors. Biometrics, 73(2): 364, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIEb

[1149] Anonymous. Issue information — Ed board. Biometrics, 73(2):365– 366, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2017:SCB

[1150] Wentian Guo, Yuan Ji, and Daniel V. T. Catenacci. A subgroup clusterbased Bayesian adaptive design for precision medicine. Biometrics, 73 (2):367–377, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2017:SMA

[1151] Ying Liu, Yuanjia Wang, and Donglin Zeng. Sequential multiple assignment randomization trials with enrichment design. Biometrics, 73 (2):378–390, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2017:GOW

[1152] Ruoqing Zhu, Ying-Qi Zhao, Guanhua Chen, Shuangge Ma, and Hongyu Zhao. Greedy outcome weighted tree learning of optimal personalized treatment rules. Biometrics, 73(2):391–400, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2017:FBN

[1153] Chanmin Kim, Michael J. Daniels, Bess H. Marcus, and Jason A. Roy. A framework for Bayesian nonparametric inference for causal effects of mediation. Biometrics, 73(2):401–409, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cefalu:2017:MAD

[1154] Matthew Cefalu, Francesca Dominici, Nils Arvold, and Giovanni Parmigiani. Model averaged double robust estimation. Biometrics, 73(2):410– 421, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2017:AMH

[1155] Paul R. Rosenbaum and Dylan S. Small. An adaptive Mantel–Haenszel test for sensitivity analysis in observational studies. Biometrics, 73(2): 422–430, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2017:IRT

[1156] Chenguang Wang, Daniel O. Scharfstein, Elizabeth Colantuoni, Timothy D. Girard, and Ying Yan. Inference in randomized trials with death and missingness. Biometrics, 73(2):431–440, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qi:2017:GSV

[1157] Li Qi, Yanqing Sun, and Peter B. Gilbert. Generalized semiparametric varying-coefficient model for longitudinal data with applications to adaptive treatment randomizations. Biometrics, 73(2):441–451, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fong:2017:MRI

[1158] Youyi Fong, Chongzhi Di, Ying Huang, and Peter B. Gilbert. Modelrobust inference for continuous threshold regression models. Biometrics, 73(2):452–462, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rice:2017:STE

[1159] John D. Rice and Alex Tsodikov. Semiparametric time-to-event modeling in the presence of a latent progression event. Biometrics, 73(2):463– 472, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sasieni:2017:SRS

[1160] Peter Sasieni and Adam R. Brentnall. On standardized relative survival. Biometrics, 73(2):473–482, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chauvel:2017:SMC

[1161] Cécile Chauvel and John O’Quigley. Survival model construction guided by fit and predictive strength. Biometrics, 73(2):483–494, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peng:2017:RBM

[1162] Yingwei Peng and Jeremy M. G. Taylor. Residual-based model diagnosis methods for mixture cure models. Biometrics, 73(2):495–505, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2017:QAB

[1163] Ruosha Li, Yu Cheng, Qingxia Chen, and Jason Fine. Quantile association for bivariate survival data. Biometrics, 73(2):506–516, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2017:AQP

[1164] Ruosha Li and Limin Peng. Assessing quantile prediction with censored quantile regression models. Biometrics, 73(2):517–528, June 2017. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2017:SOL

[1165] Yang Liu, Francesca Chiaromonte, and Bing Li. Structured ordinary least squares: a sufficient dimension reduction approach for regressions with partitioned predictors and heterogeneous units. Biometrics, 73(2): 529–539, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2017:SCS

[1166] Jian Zhang. Screening and clustering of sparse regressions with finite non-Gaussian mixtures. Biometrics, 73(2):540–550, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Su:2017:HTF

[1167] Yu-Ru Su, Chong-Zhi Di, and Li Hsu. Hypothesis testing in functional linear models. Biometrics, 73(2):551–561, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Habiger:2017:MTH

[1168] Joshua Habiger, David Watts, and Michael Anderson. Multiple testing with heterogeneous multinomial distributions. Biometrics, 73(2):562– 570, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hong:2017:SMV

[1169] Chuan Hong, Yang Ning, Peng Wei, Ying Cao, and Yong Chen. A semiparametric model for vQTL mapping. Biometrics, 73(2):571–581, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2017:IGR

[1170] Sihai Dave Zhao. Integrative genetic risk prediction using non-parametric empirical Bayes classification. Biometrics, 73(2):582–592, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borchers:2017:DSD

[1171] David Louis Borchers and Martin James Cox. Distance sampling detection functions: 2D or not 2D? Biometrics, 73(2):593–602, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2017:BVS

[1172] Siying Chen, Sara Nunez, Muredach P. Reilly, and Andrea S. Foulkes. Bayesian variable selection for post-analytic interrogation of susceptibility loci. Biometrics, 73(2):603–614, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chekouo:2017:BIA

[1173] Thierry Chekouo, Francesco C. Stingo, James D. Doecke, and Kim-Anh Do. A Bayesian integrative approach for multi-platform genomic data: a kidney cancer case study. Biometrics, 73(2):615–624, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coley:2017:BHM

[1174] Rebecca Yates Coley, Aaron J. Fisher, Mufaddal Mamawala, Herbert Ballentine Carter, Kenneth J. Pienta, and Scott L. Zeger. A Bayesian hierarchical model for prediction of latent health states from multiple data sources with application to active surveillance of prostate cancer. Biometrics, 73(2):625–634, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2017:EOR

[1175] Jincheng Shen, Lu Wang, and Jeremy M. G. Taylor. Estimation of the optimal regime in treatment of prostate cancer recurrence from observational data using flexible weighting models. Biometrics, 73(2):635–645, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barrado:2017:EDA

[1176] Leandro García Barrado, Els Coart, and Tomasz Burzykowski. Estimation of diagnostic accuracy of a combination of continuous biomarkers allowing for conditional dependence between the biomarkers and the imperfect reference-test. Biometrics, 73(2):646–655, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hou:2017:HGT

[1177] Peijie Hou, Joshua M. Tebbs, Christopher R. Bilder, and Christopher S. McMahan. Hierarchical group testing for multiple infections. Biometrics, 73(2):656–665, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2017:JMA

[1178] Shuling Liu, Amita K. Manatunga, Limin Peng, and Michele Marcus. A joint modeling approach for multivariate survival data with random length. Biometrics, 73(2):666–677, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burke:2017:MPR

[1179] K. Burke and G. MacKenzie. Multi-parameter regression survival modeling: an alternative to proportional hazards. Biometrics, 73(2):678–686, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2017:TVE

[1180] Gang Han, Michael J. Schell, Heping Zhang, Daniel Zelterman, Lajos Pusztai, Kerin Adelson, and Christos Hatzis. Testing violations of the exponential assumption in cancer clinical trials with survival endpoints. Biometrics, 73(2):687–695, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perin:2017:ALR

[1181] Jamie Perin and John S. Preisser. Alternating logistic regressions with improved finite sample properties. Biometrics, 73(2):696–705, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Korn:2017:RRE

[1182] Edward L. Korn and Boris Freidlin. Reader reaction on estimation of treatment effects in all-comers randomized clinical trials with a predictive marker. Biometrics, 73(2):706–708, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Venables:2017:BRC

[1183] Bill Venables. Book review: Chambers, John M., Extending R. Boca Raton: CRC Press. Biometrics, 73(2):709–710, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:2017:BRK

[1184] D. Andrew Brown. Book review: Kass, Robert E., Eden, Uri T., Brown, Emery N., Analysis of Neural Data. New York: Springer Science + Business Media. Biometrics, 73(2):710–713, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Escudero:2017:BRF

[1185] Luis Angel García Escudero. Book review: Farcomeni, Alessio, Greco, Luca. Robust Methods for Data Reduction. Boca Raton: CRC Press. Biometrics, 73(2):713–714, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2017:BRC

[1186] Shuai Chen and Menggang Yu. Book review: Commenges, Daniel, Jacqmin-Gadda, Helene. Dynamical Biostatistical Models. Boca Raton: CRC Press. Biometrics, 73(2):714–715, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matiu:2017:BRW

[1187] Michael Matiu. Book review: Wiegand, Thorsten, Moloney, Kirk A., Handbook of Spatial Point-Pattern Analysis in Ecology. Boca Raton: CRC Press. Biometrics, 73(2):715, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hothorn:2017:BRT

[1188] Torsten Hothorn. Book review: Tutz, Gerhard. Regression for Categorical Data. Cambridge: Cambridge University Press. Biometrics, 73 (2):715–716, June 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kang:2017:EEA

[1189] Suhyun Kang, Wenbin Lu, and Mengling Liu. Efficient estimation for accelerated failure time model under case-cohort and nested case-control sampling. Biometrics, 73(3):114–123, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Komukai:2017:DRE

[1190] Sho Komukai and Satoshi Hattori. Doubly robust estimator for net survival rate in analyses of cancer registry data. Biometrics, 73(3):124–133, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gong:2017:EAT

[1191] Qi Gong and Douglas E. Schaubel. Estimating the average treatment effect on survival based on observational data and using partly conditional modeling. Biometrics, 73(3):134–144, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tao:2017:ACW

[1192] Yebin Tao and Lu Wang. Adaptive contrast weighted learning for multi-stage multi-treatment decision-making. Biometrics, 73(3):145–155, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2017:AMB

[1193] Lifeng Lin, Haitao Chu, and James S. Hodges. Alternative measures of between-study heterogeneity in meta-analysis: Reducing the impact of outlying studies. Biometrics, 73(3):156–166, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2017:PEC

[1194] Chang Xuan Mao, Ruochen Huang, and Sijia Zhang. Petersen estimator, Chapman adjustment, list effects, and heterogeneity. Biometrics, 73(3): 167–173, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Canale:2017:RBN

[1195] Antonio Canale and Igor Prünster. Robustifying Bayesian nonparametric mixtures for count data. Biometrics, 73(3):174–184, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neelon:2017:LBL

[1196] Brian Neelon and Dongjun Chung. The LZIP: a Bayesian latent factor model for correlated zero-inflated counts. Biometrics, 73(3):185–196, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gutjahr:2017:LRT

[1197] Georg Gutjahr and Björn Bornkamp. Likelihood ratio tests for a doseresponse effect using multiple nonlinear regression models. Biometrics, 73(3):197–205, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holland-Letz:2017:COD

[1198] Tim Holland-Letz. On the combination of C- and D-optimal designs: General approaches and applications in dose-response studies. Biometrics, 73(3):206–213, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:2017:OPE

[1199] Alan Agresti and Maria Kateri. Ordinal probability effect measures for group comparisons in multinomial cumulative link models. Biometrics, 73(3):214–219, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tomassi:2017:SDR

[1200] Diego Tomassi, Liliana Forzani, Efstathia Bura, and Ruth Pfeiffer. Sufficient dimension reduction for censored predictors. Biometrics, 73(3): 220–231, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2017:MBV

[1201] Kyu Ha Lee, Mahlet G. Tadesse, Andrea A. Baccarelli, Joel Schwartz, and Brent A. Coull. Multivariate Bayesian variable selection exploiting dependence structure among outcomes: Application to air pollution effects on DNA methylation. Biometrics, 73(3):232–241, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Held:2017:APW

[1202] Leonhard Held and Rafael Sauter. Adaptive prior weighting in generalized regression. Biometrics, 73(3):242–251, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balmert:2017:NIQ

[1203] Lauren Balmert and Jong-Hyeon Jeong. Nonparametric inference on quantile lost lifespan. Biometrics, 73(3):252–259, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rancoita:2017:TDF

[1204] Paola M. V. Rancoita, Morten Valberg, Romano Demicheli, Elia Biganzoli, and Clelia Di Serio. Tumor dormancy and frailty models: a novel approach. Biometrics, 73(3):260–270, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choi:2017:MSC

[1205] Yun-Hee Choi, Laurent Briollais, Aung K. Win, John Hopper, Dan Buchanan, Mark Jenkins, and Lajmi Lakhal-Chaieb. Modeling of successive cancer risks in Lynch syndrome families in the presence of competing risks using copulas. Biometrics, 73(3):271–282, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fisher:2017:TSM

[1206] Leigh Fisher, Jon Wakefield, Cici Bauer, and Steve Self. Time series modeling of pathogen-specific disease probabilities with subsampled data. Biometrics, 73(3):283–293, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prague:2017:DME

[1207] Mélanie Prague, Daniel Commenges, Jon Michael Gran, Bruno Ledergerber, Jim Young, Hansjakob Furrer, and Rodolphe Thiébaut. Dynamic models for estimating the effect of HAART on CD4 in observational studies: Application to the aquitaine cohort and the Swiss HIV cohort study. Biometrics, 73(3):294–304, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Desmee:2017:USA

[1208] Solène Desmée, France Mentré, Christine Veyrat-Follet, Bernard Sébastien, and Jérémie Guedj. Using the SAEM algorithm for mechanistic joint models characterizing the relationship between nonlinear PSA kinetics and survival in prostate cancer patients. Biometrics, 73 (3):305–312, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2017:HML

[1209] Xinyuan Song, Yemao Xia, and Hongtu Zhu. Hidden Markov latent variable models with multivariate longitudinal data. Biometrics, 73(3): 313–323, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2017:CMF

[1210] Juhyun Park and Jeongyoun Ahn. Clustering multivariate functional data with phase variation. Biometrics, 73(3):324–333, March 2017. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bao:2017:ADT

[1211] Junshu Bao, Timothy Hanson, Garnett P. McMillan, and Kristin Knight. Assessment of DPOAE test-retest difference curves via hierarchical Gaussian processes. Biometrics, 73(3):334–343, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2017:DRC

[1212] Yuan Zhang, Shili Lin, and Swati Biswas. Detecting rare and common haplotype-environment interaction under uncertainty of geneenvironment independence assumption. Biometrics, 73(3):344–355, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:2017:BRB

[1213] John Thompson. Book review: Burgess, Stephen, Thompson, Simon G., Mendelian Randomization: Methods for Using Genetic Variants in Causal Estimation. Boca Raton: CRC Press. Biometrics, 73(3):356, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grill:2017:BRN

[1214] Sonja Grill. Book review: Ng, Tie-Hua. Noninferiority Testing in Clinical Trials: Issues and Challenges. Boca Raton: CRC Press. Biometrics, 73 (3):357, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jak:2017:BRC

[1215] Suzanne Jak. Book review: Cheung, MW-L. Meta-analysis: a Structural Equation Modeling Approach. New York: Wiley. Biometrics, 73(3):357– 358, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ankerst:2017:BRZ

[1216] Donna Pauler Ankerst. Book review: Zelterman, Daniel, Applied Multivariate Statistics with R. Heidelberg: Springer. Biometrics, 73(3):358– 359, March 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IITc

[1217] Anonymous. Issue information — TOC. Biometrics, 73(3):717–718, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIMc

[1218] Anonymous. Issue information — masthead. Biometrics, 73(3):719, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIIc

[1219] Anonymous. Issue information — info for authors. Biometrics, 73(3):720, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIEc

[1220] Anonymous. Issue information — Ed board. Biometrics, 73(3):721–722, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:BJI

[1221] Anonymous. Biometrics, JABES and the International Biometric Society. Biometrics, 73(3):723–724, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:2017:RBM

[1222] C. Jason Liang and Patrick J. Heagerty. A risk-based measure of timevarying prognostic discrimination for survival models. Biometrics, 73(3): 725–734, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [1223, 1224, 1225] and rejoinder [1226].

Michael:2017:DRB

[1223] H. Michael and L. Tian. Discussion of “A risk-based measure of timevarying prognostic discrimination for survival models,” by C. Jason Liang and Patrick J. Heagerty. Biometrics, 73(3):735–738, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1222, 1226].

Gerds:2017:DRB

[1224] Thomas Alexander Gerds and Martin Schumacher. Discussion of “A risk-based measure of time-varying prognostic discrimination for survival models,” by C. Jason Liang and Patrick J. Heagerty. Biometrics, 73(3): 739–741, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1222, 1226].

Parast:2017:DRB

[1225] Layla Parast and Carolyn M. Rutter. Discussion of “A risk-based measure of time-varying prognostic discrimination for survival models,” by C. Jason Liang and Patrick J. Heagerty. Biometrics, 73(3):742–744, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1222, 1226].

Liang:2017:RDR

[1226] C. Jason Liang and Patrick J. Heagerty. Rejoinder to discussions on: “A risk-based measure of time-varying prognostic discrimination for survival models”. Biometrics, 73(3):745–748, September 2017. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1222, 1223, 1224, 1225].

Morris:2017:STS

[1227] Samuel A. Morris, Brian J. Reich, Emeric Thibaud, and Daniel Cooley. A space-time skew- t model for threshold exceedances. Biometrics, 73(3): 749–758, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shand:2017:MNS

[1228] Lyndsay Shand and Bo Li. Modeling nonstationarity in space and time. Biometrics, 73(3):759–768, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2017:JEG

[1229] Zhixiang Lin, Tao Wang, Can Yang, and Hongyu Zhao. On joint estimation of Gaussian graphical models for spatial and temporal data. Biometrics, 73(3):769–779, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xia:2017:HTM

[1230] Yin Xia and Lexin Li. Hypothesis testing of matrix graph model with application to brain connectivity analysis. Biometrics, 73(3):780–791, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2017:DTM

[1231] Tao Wang and Hongyu Zhao. A Dirichlet-tree multinomial regression model for associating dietary nutrients with gut microorganisms. Biometrics, 73(3):792–801, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sang:2017:PFP

[1232] Peijun Sang, Liangliang Wang, and Jiguo Cao. Parametric functional principal component analysis. Biometrics, 73(3):802–810, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kimes:2017:SSH

[1233] Patrick K. Kimes, Yufeng Liu, David Neil Hayes, and James Stephen Marron. Statistical significance for hierarchical clustering. Biometrics, 73(3):811–821, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:2017:PCS

[1234] Joanna H. Shih and Michael P. Fay. Pearson’s chi-square test and rank correlation inferences for clustered data. Biometrics, 73(3):822–834, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deng:2017:PHM

[1235] Yu Deng, Donglin Zeng, Jinying Zhao, and Jianwen Cai. Proportional hazards model with a change point for clustered event data. Biometrics, 73(3):835–845, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Groll:2017:SEC

[1236] Andreas Groll, Trevor Hastie, and Gerhard Tutz. Selection of effects in Cox frailty models by regularization methods. Biometrics, 73(3):846–856, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2017:SRA

[1237] Lu Mao, Dan-Yu Lin, and Donglin Zeng. Semiparametric regression analysis of interval-censored competing risks data. Biometrics, 73(3): 857–865, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2017:ETD

[1238] Wei Zhao, Ying Qing Chen, and Li Hsu. On estimation of time-dependent attributable fraction from population-based case-control studies. Biometrics, 73(3):866–875, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2017:SRA

[1239] Wei Dai, Ming Yang, Chaolong Wang, and Tianxi Cai. Sequence robust association test for familial data. Biometrics, 73(3):876–884, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laporte:2017:ERC

[1240] Fabien Laporte, Alain Charcosset, and Tristan Mary-Huard. Estimation of the relatedness coefficients from biallelic markers, application in plant mating designs. Biometrics, 73(3):885–894, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Broberg:2017:CET

[1241] Per Broberg and Frank Miller. Conditional estimation in two-stage adaptive designs. Biometrics, 73(3):895–904, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ventz:2017:BRA

[1242] Steffen Ventz, William T. Barry, Giovanni Parmigiani, and Lorenzo Trippa. Bayesian response-adaptive designs for basket trials. Biometrics, 73(3):905–915, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2017:MOD

[1243] Sheng Wu, Weng Kee Wong, and Catherine M. Crespi. Maximin optimal designs for cluster randomized trials. Biometrics, 73(3):916–926, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bogacka:2017:ODN

[1244] Barbara Bogacka, Mahbub A. H. M. Latif, Steven G. Gilmour, and Kuresh Youdim. Optimum designs for non-linear mixed effects models in the presence of covariates. Biometrics, 73(3):927–937, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gasparrini:2017:PFD

[1245] Antonio Gasparrini, Fabian Scheipl, Ben Armstrong, and Michael G. Kenward. A penalized framework for distributed lag non-linear models. Biometrics, 73(3):938–948, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1675].

Zhang:2017:EVC

[1246] Xinyu Zhang, Jiguo Cao, and Raymond J. Carroll. Estimating varying coefficients for partial differential equation models. Biometrics, 73(3): 949–959, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Soave:2017:GLS

[1247] David Soave and Lei Sun. A generalized Levene’s scale test for variance heterogeneity in the presence of sample correlation and group uncertainty. Biometrics, 73(3):960–971, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2017:ORI

[1248] Heng Wang and Ping-Shou Zhong. Order-restricted inference for means with missing values. Biometrics, 73(3):972–980, September 2017. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2017:DEA

[1249] Xin Lu and Brent A. Johnson. Direct estimation for adaptive treatment length policies: Methods and application to evaluating the effect of delayed PEG insertion. Biometrics, 73(3):981–989, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gu:2017:CIR

[1250] Chenyang Gu and Roee Gutman. Combining item response theory with multiple imputation to equate health assessment questionnaires. Biometrics, 73(3):990–998, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hasenstab:2017:MDF

[1251] Kyle Hasenstab, Aaron Scheffler, Donatello Telesca, Catherine A. Sugar, Shafali Jeste, Charlotte DiStefano, and Damla Sentürk. A multidimensional functional principal components analysis of EEG data. Biometrics, 73(3):999–1009, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hwang:2017:ETP

[1252] Youngdeok Hwang, Samantha Wright, and Bret M. Hanlon. Estimation and testing problems in auditory neuroscience via clustering. Biometrics, 73(3):1010–1017, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lock:2017:BGE

[1253] Eric F. Lock and David B. Dunson. Bayesian genome- and epigenomewide association studies with gene level dependence. Biometrics, 73(3): 1018–1028, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McMillan:2017:VGA

[1254] L. F. McMillan and R. M. Fewster. Visualizations for genetic assignment analyses using the saddlepoint approximation method. Biometrics, 73(3): 1029–1041, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yan:2017:IEP

[1255] Ying Yan, Haibo Zhou, and Jianwen Cai. Improving efficiency of parameter estimation in case-cohort studies with multivariate failure time data. Biometrics, 73(3):1042–1052, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:CNA

[1256] Anonymous. Correction to “Nonparametric analysis of competing risks data with event category missing at random,” by Natalia Gouskova, Feng-Chang Lin, and Jason Fine; 73, 104–113, March 2017. Biometrics, 73(3):1053, September 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1138].

Anonymous:2017:IITd

[1257] Anonymous. Issue information — TOC. Biometrics, 73(4):1065–1066, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIMd

[1258] Anonymous. Issue information — masthead. Biometrics, 73(4):1067, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIId

[1259] Anonymous. Issue information — info for authors. Biometrics, 73(4): 1068, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:IIEd

[1260] Anonymous. Issue information — Ed board. Biometrics, 73(4):1069– 1070, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wood:2017:GFS

[1261] Simon N. Wood and Matteo Fasiolo. A generalized fellner-schall method for smoothing parameter optimization with application to Tweedie location, scale and shape models. Biometrics, 73(4):1071–1081, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2017:EBH

[1262] Jingjing Yang, Dennis D. Cox, Jong Soo Lee, Peng Ren, and Taeryon Choi. Efficient Bayesian hierarchical functional data analysis with basis function approximations using Gaussian-Wishart processes. Biometrics, 73(4):1082–1091, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reiss:2017:PIM

[1263] Philip T. Reiss, Lei Huang, Pei-Shien Wu, Huaihou Chen, and Stan Colcombe. Pointwise influence matrices for functional-response regression. Biometrics, 73(4):1092–1101, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2017:CFD

[1264] Hanwen Huang. Controlling the false discoveries in LASSO. Biometrics, 73(4):1102–1110, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shortreed:2017:OAL

[1265] Susan M. Shortreed and Ashkan Ertefaie. Outcome-adaptive lasso: Variable selection for causal inference. Biometrics, 73(4):1111–1122, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tchetgen:2017:GIV

[1266] Eric J. Tchetgen Tchetgen and Kathleen E. Wirth. A general instrumental variable framework for regression analysis with outcome missing not at random. Biometrics, 73(4):1123–1131, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2017:DRE

[1267] Huiming Lin, Bo Fu, Guoyou Qin, and Zhongyi Zhu. Doubly robust estimation of generalized partial linear models for longitudinal data with dropouts. Biometrics, 73(4):1132–1139, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martinussen:2017:IVE

[1268] Torben Martinussen, Stijn Vansteelandt, Eric J. Tchetgen Tchetgen, and David M. Zucker. Instrumental variables estimation of exposure effects on a time-to-event endpoint using structural cumulative survival models. Biometrics, 73(4):1140–1149, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2017:JML

[1269] Qing Cai, Mei-Cheng Wang, and Kwun Chuen Gary Chan. Joint modeling of longitudinal, recurrent events and failure time data for survivor’s population. Biometrics, 73(4):1150–1160, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2017:SEA

[1270] Fei Gao, Donglin Zeng, and Dan-Yu Lin. Semiparametric estimation of the accelerated failure time model with partly interval-censored data. Biometrics, 73(4):1161–1168, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2017:AEY

[1271] Yu Zheng and Tianxi Cai. Augmented estimation for t-year survival with censored regression models. Biometrics, 73(4):1169–1178, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frumento:2017:PMQ

[1272] Paolo Frumento and Matteo Bottai. Parametric modeling of quantile regression coefficient functions with censored and truncated data. Biometrics, 73(4):1179–1188, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berchenko:2017:MAR

[1273] Yakir Berchenko, Jonathan D. Rosenblatt, and Simon D. W. Frost. Modeling and analyzing respondent-driven sampling as a counting process. Biometrics, 73(4):1189–1198, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2017:GSF

[1274] Shuai Chen, Lu Tian, Tianxi Cai, and Menggang Yu. A general statistical framework for subgroup identification and comparative treatment scoring. Biometrics, 73(4):1199–1209, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sofer:2017:WPS

[1275] Tamar Sofer, Elizabeth D. Schifano, David C. Christiani, and Xihong Lin. Weighted pseudolikelihood for SNP set analysis with multiple secondary outcomes in case-control genetic association studies. Biometrics, 73(4):1210–1220, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jia:2017:LGR

[1276] Bochao Jia, Suwa Xu, Guanghua Xiao, Vishal Lamba, and Faming Liang. Learning gene regulatory networks from next generation sequencing data.

Biometrics, 73(4):1221–1230, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nie:2017:ETV

[1277] Yunlong Nie, LiangLiang Wang, and Jiguo Cao. Estimating time-varying directed gene regulation networks. Biometrics, 73(4):1231–1242, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2017:GEM

[1278] Yeonhee Park, Zhihua Su, and Hongtu Zhu. Groupwise envelope models for imaging genetic analysis. Biometrics, 73(4):1243–1253, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agniel:2017:AMD

[1279] Denis Agniel and Tianxi Cai. Analysis of multiple diverse phenotypes via semiparametric canonical correlation analysis. Biometrics, 73(4): 1254–1265, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shi:2017:MPM

[1280] Pixu Shi and Hongzhe Li. A model for paired-multinomial data and its application to analysis of data on a taxonomic tree. Biometrics, 73(4): 1266–1278, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

deCarvalho:2017:NBC

[1281] Vanda Inácio de Carvalho, Miguel de Carvalho, and Adam J. Branscum. Nonparametric Bayesian covariate-adjusted estimation of the Youden index. Biometrics, 73(4):1279–1288, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2017:DMM

[1282] Pranab Ghosh, Lingyun Liu, P. Senchaudhuri, Ping Gao, and Cyrus Mehta. Design and monitoring of multi-arm multi-stage clinical trials. Biometrics, 73(4):1289–1299, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:2017:SBH

[1283] Jinyuan Chang, Chao Zheng, Wen-Xin Zhou, and Wen Zhou. Simulationbased hypothesis testing of high dimensional means under covariance heterogeneity. Biometrics, 73(4):1300–1310, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deng:2017:SOQ

[1284] Chong Deng, Yongtao Guan, Rasmus P. Waagepetersen, and Jingfei Zhang. Second-order quasi-likelihood for spatial point processes. Biometrics, 73(4):1311–1320, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cowen:2017:HMM

[1285] Laura L. E. Cowen, Panagiotis Besbeas, Byron J. T. Morgan, and Carl J. Schwarz. Hidden Markov models for extended batch data. Biometrics, 73(4):1321–1331, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vaughan:2017:SGE

[1286] Gregory Vaughan, Robert Aseltine, Kun Chen, and Jun Yan. Stagewise generalized estimating equations with grouped variables. Biometrics, 73 (4):1332–1342, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2017:PNM

[1287] Huaihou Chen, Donglin Zeng, and Yuanjia Wang. Penalized nonlinear mixed effects model to identify biomarkers that predict disease progression. Biometrics, 73(4):1343–1354, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morita:2017:BPF

[1288] Satoshi Morita and Peter Müller. Bayesian population finding with biomarkers in a randomized clinical trial. Biometrics, 73(4):1355–1365, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Petroff:2017:SCC

[1289] David Petroff. A simple correction to completer analyses and improvement on baseline observation carried forward. Biometrics, 73(4):1366– 1378, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2017:MIV

[1290] Yongqiang Tang. On the multiple imputation variance estimator for control-based and delta-adjusted pattern mixture models. Biometrics, 73(4):1379–1387, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Abrams:2017:POF

[1291] Steven Abrams, Marc Aerts, Geert Molenberghs, and Niel Hens. Parametric overdispersed frailty models for current status data. Biometrics, 73(4):1388–1400, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2017:AFT

[1292] Kyu Ha Lee, Virginie Rondeau, and Sebastien Haneuse. Accelerated failure time models for semi-competing risks data in the presence of complex censoring. Biometrics, 73(4):1401–1412, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Juarez-Colunga:2017:JMZ

[1293] E. Juarez-Colunga, G. L. Silva, and C. B. Dean. Joint modeling of zeroinflated panel count and severity outcomes. Biometrics, 73(4):1413–1423, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hasegawa:2017:SAM

[1294] Raiden Hasegawa and Dylan Small. Sensitivity analysis for matched pair analysis of binary data: From worst case to average case analysis. Biometrics, 73(4):1424–1432, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1547].

Li:2017:ICI

[1295] Gen Li and Sungkyu Jung. Incorporating covariates into integrated factor analysis of multi-view data. Biometrics, 73(4):1433–1442, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McMahan:2017:BRG

[1296] Christopher S. McMahan, Joshua M. Tebbs, Timothy E. Hanson, and Christopher R. Bilder. Bayesian regression for group testing data. Biometrics, 73(4):1443–1452, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhan:2017:FSS

[1297] Xiang Zhan, Anna Plantinga, Ni Zhao, and Michael C. Wu. A fast small-sample kernel independence test for microbiome community-level association analysis. Biometrics, 73(4):1453–1463, December 2017. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:2017:BRE

[1298] Jeremy M. G. Taylor. Book review: Efron, Bradley and Hastie, Trevor. Computer Age Statistical Inference. Cambridge: Cambridge University Press. Biometrics, 73(4):1464–1465, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Galecki:2017:BRF

[1299] Andrzej Galecki. Book review: Faraway, Julian J., Extending the Linear Model with R: Generalized Linear, Mixed Effects, and Nonparametric Regression Models, 2nd edition. Boca Raton: CRC Press. Biometrics, 73(4):1465–1466, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Straubinger:2017:BRC

[1300] Johanna Straubinger. Book review: Cheung, Yin Bun. Statistical Analysis of Human Growth and Development. Boca Raton: CRC Press. Biometrics, 73(4):1466–1467, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mansmann:2017:BRC

[1301] Ulrich Mansmann. Book reviews: Combs, C. D., Sokolowski, John A. and Banks, Catherine M., The Digital Patient: Advancing Healthcare, Research, and Education. Hoboken: Wiley. Chandan K. Reddy and Charu C. Aggarwal. Healthcare Data Analytics. Boca Raton: CRC Press. Biometrics, 73(4):1467–1468, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Janes:2017:BRL

[1302] Holly Janes. Book review: Looney, Stephen W. and Hagan, Joseph L., Analysis of Biomarker Data: a Practical Guide. Hoboken: Wiley. Biometrics, 73(4):1468, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goros:2017:BRH

[1303] Martin Goros. Book review: Heiberger, Richard M. and Holland, Burt. Statistical Analysis and Data Display. Heidelberg: Springer. Biometrics, 73(4):1468–1469, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Selig:2017:BRM

[1304] Katharina Selig. Book review: Munzert, Simon, Rubba, Christian, Meissner, Peter, and Nyhuis, Dominic. Automated Data Collection with R: a Practical Guide to Web Scraping and Text Mining. Hoboken: Wiley. Biometrics, 73(4):1469, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2017:BRK

[1305] Xulei Liu. Book review: Kleinman, Ken and Horton, Nicholas J., SAS and R: Data Management, Statistical Analysis, and Graphics. Boca Raton: CRC Press. Biometrics, 73(4):1470, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2017:A

[1306] Anonymous. Acknowledgements. Biometrics, 73(4):1471–1472, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IITa

[1307] Anonymous. Issue information — TOC and cover information. Biometrics, 74(1):1–2, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIMa

[1308] Anonymous. Issue information — masthead. Biometrics, 74(1):3, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIEa

[1309] Anonymous. Issue information — Ed board. Biometrics, 74(1):4, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:RE

[1310] Anonymous. Report of the Editors — 2017. Biometrics, 74(1):5–7, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Koch:2018:CSG

[1311] Brandon Koch, David M. Vock, and Julian Wolfson. Covariate selection with group lasso and doubly robust estimation of causal effects. Biometrics, 74(1):8–17, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Butler:2018:IPP

[1312] Emily L. Butler, Eric B. Laber, Sonia M. Davis, and Michael R. Kosorok. Incorporating patient preferences into estimation of optimal individualized treatment rules. Biometrics, 74(1):18–26, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2018:EPS

[1313] Ying Huang. Evaluating principal surrogate markers in vaccine trials in the presence of multiphase sampling. Biometrics, 74(1):27–39, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tamhane:2018:GPT

[1314] Ajit C. Tamhane, Jiangtao Gou, Christopher Jennison, Cyrus R. Mehta, and Teresa Curto. A gatekeeping procedure to test a primary and a secondary endpoint in a group sequential design with multiple interim looks. Biometrics, 74(1):40–48, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Villar:2018:CAR

[1315] Sofía S. Villar and William F. Rosenberger. Covariate-adjusted responseadaptive randomization for multi-arm clinical trials using a modified forward looking Gittins index rule. Biometrics, 74(1):49–57, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2018:ODS

[1316] Qingning Zhou, Jianwen Cai, and Haibo Zhou. Outcome-dependent sampling with interval-censored failure time data. Biometrics, 74(1):58–67, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2018:SPM

[1317] Hao Liu and Jing Qin. Semiparametric probit models with univariate and bivariate current-status data. Biometrics, 74(1):68–76, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2018:SFO

[1318] Yifei Sun, Kwun Chuen Gary Chan, and Jing Qin. Simple and fast overidentified rank estimation for right-censored length-biased data and backward recurrence time. Biometrics, 74(1):77–85, March 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2018:LAP

[1319] Tian Dai, Ying Guo, Limin Peng, and Amita K. Manatunga. A local agreement pattern measure based on hazard functions for survival outcomes. Biometrics, 74(1):86–99, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2018:PLA

[1320] Fan Wu, Sehee Kim, Jing Qin, Rajiv Saran, and Yi Li. A pairwise likelihood augmented Cox estimator for left-truncated data. Biometrics, 74(1):100–108, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kong:2018:FFL

[1321] Dehan Kong, Joseph G. Ibrahim, Eunjee Lee, and Hongtu Zhu. FLCRM: Functional linear Cox regression model. Biometrics, 74(1):109–117, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2018:CRD

[1322] Yijian Huang and Ching-Yun Wang. Cox regression with dependent error in covariates. Biometrics, 74(1):118–126, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tekwe:2018:FMI

[1323] Carmen D. Tekwe, Roger S. Zoh, Fuller W. Bazer, Guoyao Wu, and Raymond J. Carroll. Functional multiple indicators, multiple causes measurement error models. Biometrics, 74(1):127–134, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buonaccorsi:2018:MBB

[1324] John P. Buonaccorsi, Giovanni Romeo, and Magne Thoresen. Modelbased bootstrapping when correcting for measurement error with application to logistic regression. Biometrics, 74(1):135–144, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:2018:RML

[1325] Hung Hung, Zhi-Yu Jou, and Su-Yun Huang. Robust mislabel logistic regression without modeling mislabel probabilities. Biometrics, 74(1): 145–154, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2018:CAS

[1326] Yi-Hui Zhou, James S. Marron, and Fred A. Wright. Computation of ancestry scores with mixed families and unrelated individuals. Biometrics, 74(1):155–164, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2018:MPA

[1327] Zhonghua Liu and Xihong Lin. Multiple phenotype association tests using summary statistics in genome-wide association studies. Biometrics, 74(1):165–175, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2018:MAA

[1328] Qianchuan He, Yang Liu, Ulrike Peters, and Li Hsu. Multivariate association analysis with somatic mutation data. Biometrics, 74(1):176–184, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhadra:2018:INS

[1329] Anindya Bhadra, Arvind Rao, and Veerabhadran Baladandayuthapani. Inferring network structure in non-normal and mixed discretecontinuous genomic data. Biometrics, 74(1):185–195, March 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Segal:2018:FAS

[1330] Brian D. Segal, Thomas Braun, Michael R. Elliott, and Hui Jiang. Fast approximation of small p-values in permutation tests by partitioning the permutations. Biometrics, 74(1):196–206, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scharfstein:2018:GSA

[1331] Daniel Scharfstein, Aidan McDermott, Iván Díaz, Marco Carone, Nicola Lunardon, and Ibrahim Turkoz. Global sensitivity analysis for repeated measures studies with informative drop-out: a semi-parametric approach. Biometrics, 74(1):207–219, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2018:PLA

[1332] Ziqi Chen, Man-Lai Tang, and Wei Gao. A profile likelihood approach for longitudinal data analysis. Biometrics, 74(1):220–228, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2018:FBM

[1333] Meihua Wu, Ana Diez-Roux, Trivellore E. Raghunathan, and Brisa N. Sánchez. FPCA-based method to select optimal sampling schedules that capture between-subject variability in longitudinal studies. Biometrics, 74(1):229–238, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2018:MFS

[1334] Weining Shen, Jing Ning, Ying Yuan, Anna S. Lok, and Ziding Feng. Model-free scoring system for risk prediction with application to hepatocellular carcinoma study. Biometrics, 74(1):239–248, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tayob:2018:BSA

[1335] Nabihah Tayob, Francesco Stingo, Kim-Anh Do, Anna S. F. Lok, and Ziding Feng. A Bayesian screening approach for hepatocellular carcinoma using multiple longitudinal biomarkers. Biometrics, 74(1):249–259, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bruce:2018:CAB

[1336] Scott A. Bruce, Martica H. Hall, Daniel J. Buysse, and Robert T. Krafty. Conditional adaptive Bayesian spectral analysis of nonstationary biomedical time series. Biometrics, 74(1):260–269, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shim:2018:BBC

[1337] Heejung Shim and Bret Larget. BayesCAT: Bayesian co-estimation of alignment and tree. Biometrics, 74(1):270–279, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huggins:2018:ESO

[1338] Richard Huggins, Jakub Stoklosa, Cameron Roach, and Paul Yip. Estimating the size of an open population using sparse capture-recapture data. Biometrics, 74(1):280–288, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dharmarajan:2018:ECP

[1339] Sai H. Dharmarajan, Douglas E. Schaubel, and Rajiv Saran. Evaluating center performance in the competing risks setting: Application to outcomes of wait-listed end-stage renal disease patients. Biometrics, 74 (1):289–299, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Safo:2018:IAT

[1340] Sandra E. Safo, Shuzhao Li, and Qi Long. Integrative analysis of transcriptomic and metabolomic data via sparse canonical correlation analysis with incorporation of biological information. Biometrics, 74(1):300– 312, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsui:2018:MSG

[1341] Shigeyuki Matsui, Hisashi Noma, Pingping Qu, Yoshio Sakai, Kota Matsui, Christoph Heuck, and John Crowley. Multi-subgroup gene screening using semi-parametric hierarchical mixture models and the optimal discovery procedure: Application to a randomized clinical trial in multiple myeloma. Biometrics, 74(1):313–320, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mauguen:2018:EPC

[1342] Audrey Mauguen, Venkatraman E. Seshan, Irina Ostrovnaya, and Colin B. Begg. Estimating the probability of clonal relatedness of pairs of tumors in cancer patients. Biometrics, 74(1):321–330, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2018:PES

[1343] You Wu, Jeremy Gaskins, Maiying Kong, and Susmita Datta. Profiling the effects of short time-course cold ischemia on tumor protein phosphorylation using a Bayesian approach. Biometrics, 74(1):331–341, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Montagna:2018:SBL

[1344] Silvia Montagna, Tor Wager, Lisa Feldman Barrett, Timothy D. Johnson, and Thomas E. Nichols. Spatial Bayesian latent factor regression modeling of coordinate-based meta-analysis data. Biometrics, 74(1):342– 353, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hedeker:2018:NMR

[1345] Donald Hedeker, Stephen H. C. du Toit, Hakan Demirtas, and Robert D. Gibbons. A note on marginalization of regression parameters from mixed models of binary outcomes. Biometrics, 74(1):354–361, March 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Warton:2018:WYC

[1346] David I. Warton. Why you cannot transform your way out of trouble for small counts. Biometrics, 74(1):362–368, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barker:2018:RMM

[1347] Richard J. Barker, Matthew R. Schofield, William A. Link, and John R. Sauer. On the reliability of N-mixture models for count data. Biometrics, 74(1):369–377, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beyersmann:2018:BRT

[1348] Jan Beyersmann. Book review: Tutz, Gerhard and Schmid, Matthias. Modeling Discrete Time-to-Event Data. Heidelberg, Springer. Biometrics, 74(1):378–379, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Michalek:2018:BRA

[1349] Joel Michalek. Book review: Ahn, Chul, Heo, Moonseong and Zhang, Song. Sample Size Calculations for Clustered and Longitudinal Outcomes in Clinical Research. Boca Raton, CRC Press. Biometrics, 74(1):379, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tattar:2018:BRS

[1350] Prabhanjan Tattar. Book review: Scutari, Marco and Denis, Jean-Baptiste. Bayesian Networks: With Examples in R. Boca Raton, CRC Press. Biometrics, 74(1):379–380, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:2018:DJI

[1351] Johannes Müller. Duan, jinqiao. an introduction to stochastic dynamics. new York, Cambridge University Press. Biometrics, 74(1):381, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kurz:2018:BRB

[1352] Christoph F. Kurz. Book review: Bottle, Alex and Aylin, Paul. Statistical Methods for Healthcare Performance Monitoring. Boca Raton, CRC Press. Biometrics, 74(1):381–382, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kuttler:2018:BRB

[1353] Christina Kuttler. Book review: Becker, Niels G., Modeling to Inform Infectious Disease Control. Boca Raton, CRC Press. Biometrics, 74(1): 382–383, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rizopoulos:2018:KMJ

[1354] Dimitris Rizopoulos. Kuhn, max and Johnson, kjell. applied predictive modeling. new York, Springer. Biometrics, 74(1):383, March 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jacqmin-Gadda:2018:BRE

[1355] Hélène Jacqmin-Gadda. Book review: Elashoff, Robert, Li, Gang and Li, Ning. Joint Modeling of Longitudinal and Time-to-Event Data. Boca Raton, CRC Press. Biometrics, 74(1):383–384, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IITb

[1356] Anonymous. Issue information — TOC and cover information. Biometrics, 74(2):385–386, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIMb

[1357] Anonymous. Issue information — masthead. Biometrics, 74(2):387, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIEb

[1358] Anonymous. Issue information — Ed board. Biometrics, 74(2):388, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haggstrom:2018:DDC

[1359] Jenny Häggström. Data-driven confounder selection via Markov and Bayesian networks. Biometrics, 74(2):389–398, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [1360, 1361] and rejoinder [1362].

Kennedy:2018:DDD

[1360] Edward H. Kennedy and Sivaraman Balakrishnan. Discussion of “Datadriven confounder selection via Markov and Bayesian networks” by Jenny Häggström. Biometrics, 74(2):399–402, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1359].

Richardson:2018:DDD

[1361] Thomas S. Richardson, James M. Robins, and Linbo Wang. Discussion of “Data-driven confounder selection via Markov and Bayesian networks” by Häggström. Biometrics, 74(2):403–406, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1359, 1362].

Haggstrom:2018:RDD

[1362] Jenny Häggström. Rejoinder to discussions on: “Data-driven confounder selection via Markov and Bayesian networks”. Biometrics, 74(2):407– 410, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1360, 1361].

Efford:2018:SCM

[1363] Murray G. Efford and Christine M. Hunter. Spatial capture-mark-resight estimation of animal population density. Biometrics, 74(2):411–420, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hong:2018:IPD

[1364] Hyokyoung G. Hong, Xuerong Chen, David C. Christiani, and Yi Li. Integrated powered density: Screening ultrahigh dimensional covariates with survival outcomes. Biometrics, 74(2):421–429, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2018:JPT

[1365] Yuping Zhang and Zhengqing Ouyang. Joint principal trend analysis for longitudinal high-dimensional data. Biometrics, 74(2):430–438, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2018:EST

[1366] Yi-Hui Zhou, J. S. Marron, and Fred A. Wright. Eigenvalue significance testing for genetic association. Biometrics, 74(2):439–447, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sohn:2018:GBL

[1367] Michael B. Sohn and Hongzhe Li. A GLM-based latent variable ordination method for microbiome samples. Biometrics, 74(2):448–457, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gauran:2018:ENE

[1368] Iris Ivy M. Gauran, Junyong Park, Johan Lim, DoHwan Park, John Zylstra, Thomas Peterson, Maricel Kann, and John L. Spouge. Empirical null estimation using zero-inflated discrete mixture distributions and its application to protein domain data. Biometrics, 74(2):458–471, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mertens:2018:ADR

[1369] Karl Mertens and Stijn Vansteelandt. Augmented and doubly robust Gestimation of causal effects under a structural nested failure time model.

Biometrics, 74(2):472–480, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandel:2018:IPW

[1370] Micha Mandel, Jacobo de Uña-Álvarez, David K. Simon, and Rebecca A. Betensky. Inverse probability weighted Cox regression for doubly truncated data. Biometrics, 74(2):481–487, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2018:SLB

[1371] Liang Zhu, Ying Zhang, Yimei Li, Jianguo Sun, and Leslie L. Robison. A semiparametric likelihood-based method for regression analysis of mixed panel-count data. Biometrics, 74(2):488–497, June 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McIntyre:2018:MCM

[1372] Julie McIntyre, Brent A. Johnson, and Stephen M. Rappaport. Monte Carlo methods for nonparametric regression with heteroscedastic measurement error. Biometrics, 74(2):498–505, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lou:2018:OTA

[1373] Zhilan Lou, Jun Shao, and Menggang Yu. Optimal treatment assignment to maximize expected outcome with multiple treatments. Biometrics, 74 (2):506–516, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qiu:2018:EEL

[1374] Xin Qiu, Donglin Zeng, and Yuanjia Wang. Estimation and evaluation of linear individualized treatment rules to guarantee performance. Biometrics, 74(2):517–528, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schmidt:2018:ADO

[1375] Rene Schmidt, Andreas Faldum, and Robert Kwiecien. Adaptive designs for the one-sample log-rank test. Biometrics, 74(2):529–537, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2018:EDM

[1376] Hengzhen Huang, Hong-Bin Fang, and Ming T. Tan. Experimental design for multi-drug combination studies using signaling networks. Biometrics, 74(2):538–547, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jackson:2018:MBM

[1377] Dan Jackson, Sylwia Bujkiewicz, Martin Law, Richard D. Riley, and Ian R. White. A matrix-based method of moments for fitting multivariate network meta-analysis models with multiple outcomes and random inconsistency effects. Biometrics, 74(2):548–556, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huling:2018:RPH

[1378] Jared D. Huling, Menggang Yu, Muxuan Liang, and Maureen Smith. Risk prediction for heterogeneous populations with application to hospital admission prediction. Biometrics, 74(2):557–565, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rice:2018:RRP

[1379] John D. Rice, Robert L. Strawderman, and Brent A. Johnson. Regularity of a renewal process estimated from binary data. Biometrics, 74(2):566– 574, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2018:ARM

[1380] Chi Hyun Lee, Jing Ning, and Yu Shen. Analysis of restricted mean survival time for length-biased data. Biometrics, 74(2):575–583, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zuanetti:2018:CDM

[1381] Daiane Aparecida Zuanetti, Peter Müller, Yitan Zhu, Shengjie Yang, and Yuan Ji. Clustering distributions with the marginalized nested Dirichlet process. Biometrics, 74(2):584–594, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2018:CAS

[1382] Qi Liu, Chun Li, Valentine Wanga, and Bryan E. Shepherd. Covariateadjusted Spearman’s rank correlation with probability-scale residuals. Biometrics, 74(2):595–605, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ni:2018:HRG

[1383] Yang Ni, Peter Müller, Yitan Zhu, and Yuan Ji. Heterogeneous reciprocal graphical models. Biometrics, 74(2):606–615, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Palowitch:2018:ECE

[1384] John Palowitch, Andrey Shabalin, Yi-Hui Zhou, Andrew B. Nobel, and Fred A. Wright. Estimation of cis-eQTL effect sizes using a log of linear model. Biometrics, 74(2):616–625, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schofield:2018:CTC

[1385] Matthew R. Schofield, Richard J. Barker, and Nicholas Gelling. Continuous-time capture-recapture in closed populations. Biometrics, 74(2):626–635, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Koslovsky:2018:BVS

[1386] Matthew D. Koslovsky, Michael D. Swartz, Wenyaw Chan, Luis Leon-Novelo, Anna V. Wilkinson, Darla E. Kendzor, and Michael S. Businelle. Bayesian variable selection for multistate Markov models with intervalcensored data in an ecological momentary assessment study of smoking cessation. Biometrics, 74(2):636–644, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2018:FBS

[1387] Brian J. Reich, Joseph Guinness, Simon N. Vandekar, Russell T. Shinohara, and Ana-Maria Staicu. Fully Bayesian spectral methods for imaging data. Biometrics, 74(2):645–652, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2018:TGE

[1388] Ryan Sun, Raymond J. Carroll, David C. Christiani, and Xihong Lin. Testing for gene-environment interaction under exposure misspecification. Biometrics, 74(2):653–662, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Skrivankova:2018:SIM

[1389] Veronika Skrivankova and Patrick J. Heagerty. Single index methods for evaluation of marker-guided treatment rules based on multivariate marker panels. Biometrics, 74(2):663–672, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Amatya:2018:SSD

[1390] Anup Amatya and Dulal K. Bhaumik. Sample size determination for multilevel hierarchical designs using generalized linear mixed models. Biometrics, 74(2):673–684, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andrinopoulou:2018:IDP

[1391] Eleni-Rosalina Andrinopoulou, Paul H. C. Eilers, Johanna J. M. Takkenberg, and Dimitris Rizopoulos. Improved dynamic predictions from joint models of longitudinal and survival data with time-varying effects using P-splines. Biometrics, 74(2):685–693, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tian:2018:ETS

[1392] Lu Tian, Haoda Fu, Stephen J. Ruberg, Hajime Uno, and Lee-Jen Wei. Efficiency of two sample tests via the restricted mean survival time for analyzing event time observations. Biometrics, 74(2):694–702, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2018:MCE

[1393] Liangyuan Hu, Joseph W. Hogan, Ann W. Mwangi, and Abraham Siika. Modeling the causal effect of treatment initiation time on survival: Application to HIV/TB co-infection. Biometrics, 74(2):703–713, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2018:MAB

[1394] Huayu Liu, Nichole E. Carlson, Gary K. Grunwald, and Alex J. Polotsky. Modeling associations between latent event processes governing time series of pulsing hormones. Biometrics, 74(2):714–724, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rennert:2018:CRM

[1395] Lior Rennert and Sharon X. Xie. Cox regression model with doubly truncated data. Biometrics, 74(2):725–733, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2018:IRE

[1396] Sehee Kim, Douglas E. Schaubel, and Keith P. McCullough. A C-index for recurrent event data: Application to hospitalizations among dialysis patients. Biometrics, 74(2):734–743, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bai:2018:TSM

[1397] Jiawei Bai, Yifei Sun, Jennifer A. Schrack, Ciprian M. Crainiceanu, and Mei-Cheng Wang. A two-stage model for wearable device data. Biometrics, 74(2):744–752, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2018:COA

[1398] James Y. Dai, C. Jason Liang, Michael LeBlanc, Ross L. Prentice, and Holly Janes. Case-only approach to identifying markers predicting treatment effects on the relative risk scale. Biometrics, 74(2):753–763, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhan:2018:RRN

[1399] Xiang Zhan and Michael C. Wu. Reader reaction: a note on testing and estimation in marker-set association study using semiparametric quantile regression kernel machine. Biometrics, 74(2):764–766, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [1400].

Kong:2018:RNT

[1400] Dehan Kong, Arnab Maity, Fang-Chi Hsu, and Jung-Ying Tzeng. Rejoinder to “A note on testing and estimation in marker-set association study using semiparametric quantile regression kernel machine”. Biometrics, 74(2):767–768, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1399].

Kondofersky:2018:BRT

[1401] Ivan Kondofersky and Fabian J. Theis. Book review: Trevor Hastie, Robert Tibshirani, and Martin Wainwright. Statistical Learning with Sparsity: The Lasso and Generalizations. Boca Raton: CRC Press. Biometrics, 74(2):769, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eddelbuettel:2018:BRN

[1402] Dirk Eddelbuettel. Book review: Norman Matloff. Parallel Computing for Data Science: With Examples in R, C++, and CUDA. Boca Raton: CRC Press. Biometrics, 74(2):770, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ankerst:2018:BRJ

[1403] Donna Pauler Ankerst. Book review: Jay Herson. Data and Safety Monitoring Committees in Clinical Trials, 2nd edition. Boca Raton: CRC Press. Biometrics, 74(2):771, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsui:2018:BRR

[1404] Shigeyuki Matsui. Book review: Richard M. Simon. Genomic Clinical Trials and Predictive Medicine. Cambridge: Cambridge University Press. Biometrics, 74(2):772, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jackson:2018:BRV

[1405] Christopher Jackson. Book review: van den Hout, Ardo. Multi-State Survival Models for Interval-Censored Data. Boca Raton: CRC Press. Biometrics, 74(2):773, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Erosheva:2018:BRI

[1406] Elena A. Erosheva. Book review: Ian Foster, Rayid Ghani, Ron S. Jarmin, Frauke Kreuter, Julia Lane. Big Data and Social Science: a Practical Guide to Methods and Tools. Boca Raton: CRC Press. Biometrics, 74(2):774–775, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pfeiffer:2018:BRC

[1407] Ruth Pfeiffer. Book review: Claudio Carini, Sandeep M. Menon, and Mark Chang, eds. Clinical and Statistical Considerations in Personalized Medicine. Boca Raton: CRC Press. Biometrics, 74(2):775–776, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2018:BRD

[1408] Min Zhang. Book review: Daniel O. Stram. Design, Analysis, and Interpretation of Genome-Wide Association Scans. Heidelberg: Springer. Biometrics, 74(2):777–778, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Irvine:2018:BRB

[1409] Kathryn M. Irvine. Book review: Bill Shipley. Cause and Correlation in Biology: a User’s Guide to Path Analysis, Structural Equations and Causal Inference with R, 2nd ed. United Kingdom: Cambridge University Press. Biometrics, 74(2):779–780, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IITc

[1410] Anonymous. Issue information — TOC and cover information. Biometrics, 74(3):781–782, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIMc

[1411] Anonymous. Issue information — masthead. Biometrics, 74(3):783, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIEc

[1412] Anonymous. Issue information — Ed board. Biometrics, 74(3):784, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2018:QPB

[1413] Lifeng Lin and Haitao Chu. Quantifying publication bias in metaanalysis. Biometrics, 74(3):785–794, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [1414, 1415, 1416] and rejoinder [1417].

Jackson:2018:DQP

[1414] Dan Jackson. Discussion on quantifying publication bias in metaanalysis. Biometrics, 74(3):795–796, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1413].

Schmid:2018:DQP

[1415] Christopher H. Schmid. Discussion of “Quantifying publication bias in meta-analysis” by Lin et al. Biometrics, 74(3):797–799, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1413].

Geller:2018:DQP

[1416] Nancy L. Geller. Discussion on “Quantifying Publication Bias in Meta-Analysis” by Lin and Chu. Biometrics, 74(3):800, September 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1413].

Lin:2018:RQP

[1417] Lifeng Lin and Haitao Chu. Rejoinder to “Quantifying publication bias in meta-analysis”. Biometrics, 74(3):801–802, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1414, 1415, 1416, 1413].

Li:2018:RFA

[1418] Qunhua Li and Feipeng Zhang. A regression framework for assessing covariate effects on the reproducibility of high-throughput experiments.

Biometrics, 74(3):803–813, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

You:2018:SCH

[1419] Na You, Shun He, Xueqin Wang, Junxian Zhu, and Heping Zhang. Subtype classification and heterogeneous prognosis model construction in precision medicine. Biometrics, 74(3):814–822, September 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Castruccio:2018:SMR

[1420] Stefano Castruccio, Hernando Ombao, and Marc G. Genton. A scalable multi-resolution spatio-temporal model for brain activation and connectivity in fMRI data. Biometrics, 74(3):823–833, September 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2018:MMI

[1421] Hojin Yang, Hongtu Zhu, and Joseph G. Ibrahim. MILFM: Multiple index latent factor model based on high-dimensional features. Biometrics, 74(3):834–844, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mardia:2018:SMH

[1422] Kanti V. Mardia, Karthik Sriram, and Charlotte M. Deane. A statistical model for helices with applications. Biometrics, 74(3):845–854, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cho:2018:SIG

[1423] Hyunkeun Ryan Cho. Statistical inference in a growth curve quantile regression model for longitudinal data. Biometrics, 74(3):855–862, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bose:2018:TDT

[1424] Maitreyee Bose, James S. Hodges, and Sudipto Banerjee. Toward a diagnostic toolkit for linear models with Gaussian-process distributed random effects. Biometrics, 74(3):863–873, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nikolakopoulos:2018:DBT

[1425] Stavros Nikolakopoulos, Ingeborg van der Tweel, and Kit C. B. Roes. Dynamic borrowing through empirical power priors that control type i error. Biometrics, 74(3):874–880, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2018:GSI

[1426] Shih-Wei Chen and Chin-Tsang Chiang. General single-index survival regression models for incident and prevalent covariate data and prevalent data without follow-up. Biometrics, 74(3):881–890, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2018:CLN

[1427] Baqun Zhang and Min Zhang. C-learning: a new classification framework to estimate optimal dynamic treatment regimes. Biometrics, 74(3):891– 899, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2018:MSD

[1428] Shu Yang, Anastasios A. Tsiatis, and Michael Blazing. Modeling survival distribution as a function of time to treatment discontinuation: a dynamic treatment regime approach. Biometrics, 74(3):900–909, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2018:ARE

[1429] Jianxuan Liu, Yanyuan Ma, and Lan Wang. An alternative robust estimator of average treatment effect in causal inference. Biometrics, 74(3): 910–923, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2018:EIT

[1430] Jingxiang Chen, Haoda Fu, Xuanyao He, Michael R. Kosorok, and Yufeng Liu. Estimating individualized treatment rules for ordinal treatments. Biometrics, 74(3):924–933, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2018:MSS

[1431] Chung-Wei Shen and Yi-Hau Chen. Model selection for semiparametric marginal mean regression accounting for within-cluster subsampling variability and informative cluster size. Biometrics, 74(3):934–943, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chiou:2018:SEA

[1432] Sy Han Chiou, Gongjun Xu, Jun Yan, and Chiung-Yu Huang. Semiparametric estimation of the accelerated mean model with panel count data under informative examination times. Biometrics, 74(3):944–953, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2018:GAR

[1433] Huijuan Ma, Limin Peng, Zhumin Zhang, and HuiChuan J. Lai. Generalized accelerated recurrence time model for multivariate recurrent event data with missing event type. Biometrics, 74(3):954–965, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2018:MMR

[1434] Hsiang Yu, Yu-Jen Cheng, and Ching-Yun Wang. Methods for multivariate recurrent event data with measurement error and informative censoring. Biometrics, 74(3):966–976, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bluhmki:2018:WBA

[1435] Tobias Bluhmki, Claudia Schmoor, Dennis Dobler, Markus Pauly, Juergen Finke, Martin Schumacher, and Jan Beyersmann. A wild bootstrap approach for the Aalen–Johansen estimator. Biometrics, 74(3):977–985, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2018:BNA

[1436] Dandan Xu, Michael J. Daniels, and Almut G. Winterstein. A Bayesian nonparametric approach to causal inference on quantiles. Biometrics, 74(3):986–996, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Killiches:2018:DVC

[1437] Matthias Killiches and Claudia Czado. A D-vine copula-based model for repeated measurements extending linear mixed models with homogeneous correlation structure. Biometrics, 74(3):997–1005, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martens:2018:GST

[1438] Michael J. Martens and Brent R. Logan. A group sequential test for treatment effect based on the fine-Gray model. Biometrics, 74(3):1006– 1013, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2018:RAS

[1439] Yinghao Pan, Jianwen Cai, Sangmi Kim, and Haibo Zhou. Regression analysis for secondary response variable in a case-cohort study. Biometrics, 74(3):1014–1022, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:2018:SAU

[1440] Dean Follmann and Chiung-Yu Huang. Sieve analysis using the number of infecting pathogens. Biometrics, 74(3):1023–1033, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilson:2018:MAC

[1441] Ander Wilson, Corwin M. Zigler, Chirag J. Patel, and Francesca Dominici. Model-averaged confounder adjustment for estimating multivariate exposure effects with linear regression. Biometrics, 74(3):1034–1044, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2018:RCT

[1442] Shuang Huang, Chengcheng Hu, Melanie L. Bell, Dean Billheimer, Stefano Guerra, Denise Roe, Monica M. Vasquez, and Edward J. Bedrick. Regularized continuous-time Markov model via elastic net. Biometrics, 74(3):1045–1054, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shi:2018:BET

[1443] Haolun Shi and Guosheng Yin. Bayesian enhancement two-stage design for single-arm Phase II clinical trials with binary and time-to-event endpoints. Biometrics, 74(3):1055–1064, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Braun:2018:MSS

[1444] Thomas M. Braun. Motivating sample sizes in adaptive Phase I trials via Bayesian posterior credible intervals. Biometrics, 74(3):1065–1071, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jeffries:2018:DTD

[1445] Neal O. Jeffries, James F. Troendle, and Nancy L. Geller. Detecting treatment differences in group sequential longitudinal studies with covariate adjustment. Biometrics, 74(3):1072–1081, September 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kaizer:2018:MSA

[1446] Alexander M. Kaizer, Brian P. Hobbs, and Joseph S. Koopmeiners. A multi-source adaptive platform design for testing sequential combinatorial therapeutic strategies. Biometrics, 74(3):1082–1094, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murray:2018:UBD

[1447] Thomas A. Murray, Ying Yuan, Peter F. Thall, Joan H. Elizondo, and Wayne L. Hofstetter. A utility-based design for randomized comparative trials with ordinal outcomes and prognostic subgroups. Biometrics, 74 (3):1095–1103, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maidman:2018:NSM

[1448] Adam Maidman and Lan Wang. New semiparametric method for predicting high-cost patients. Biometrics, 74(3):1104–1111, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Elmi:2018:AJM

[1449] Angelo F. Elmi, Katherine L. Grantz, and Paul S. Albert. An approximate joint model for multiple paired longitudinal outcomes and timeto-event data. Biometrics, 74(3):1112–1119, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2018:RRF

[1450] Bin Guo and Baolin Wu. Reader reaction on the fast small-sample kernel independence test for microbiome community-level association analysis. Biometrics, 74(3):1120–1124, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Connor:2018:BR

[1451] Jason T. Connor. Book review. Biometrics, 74(3):1125–1126, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fraley:2018:BR

[1452] Chris Fraley. Book review. Biometrics, 74(3):1127–1128, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baran:2018:BR

[1453] Sándor Baran. Book review. Biometrics, 74(3):1129–1130, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoering:2018:BR

[1454] Antje Hoering. Book review. Biometrics, 74(3):1131, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ankerst:2018:BR

[1455] Donna Pauler Ankerst. Book review. Biometrics, 74(3):1132, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2018:BR

[1456] Joon Jin Song. Book review. Biometrics, 74(3):1133, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jo:2018:BR

[1457] Booil Jo. Book review. Biometrics, 74(3):1134–1136, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nelson:2018:BR

[1458] Wayne B. Nelson. Book review. Biometrics, 74(3):1137–1138, September 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IITd

[1459] Anonymous. Issue information — TOC and cover information. Biometrics, 74(4):1137–1138, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIMd

[1460] Anonymous. Issue information — masthead. Biometrics, 74(4):1139, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:IIEd

[1461] Anonymous. Issue information — Ed board. Biometrics, 74(4):1140, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nattino:2018:MAS

[1462] Giovanni Nattino and Bo Lu. Model assisted sensitivity analyses for hidden bias with binary outcomes. Biometrics, 74(4):1141–1149, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2018:SAP

[1463] Xuran Wang, Yang Jiang, Nancy R. Zhang, and Dylan S. Small. Sensitivity analysis and power for instrumental variable studies. Biometrics, 74(4):1150–1160, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2018:PAS

[1464] Kwonsang Lee, Dylan S. Small, and Paul R. Rosenbaum. A powerful approach to the study of moderate effect modification in observational studies. Biometrics, 74(4):1161–1170, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Antonelli:2018:DRM

[1465] Joseph Antonelli, Matthew Cefalu, Nathan Palmer, and Denis Agniel. Doubly robust matching estimators for high dimensional confounding adjustment. Biometrics, 74(4):1171–1179, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hager:2018:OTS

[1466] Rebecca Hager, Anastasios A. Tsiatis, and Marie Davidian. Optimal twostage dynamic treatment regimes from a classification perspective with censored survival data. Biometrics, 74(4):1180–1192, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:2018:BNG

[1467] Jason Roy, Kirsten J. Lum, Bret Zeldow, Jordan D. Dworkin, Vincent Lo Re III, and Michael J. Daniels. Bayesian nonparametric generative models for causal inference with missing at random covariates. Biometrics, 74(4):1193–1202, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

deUna-Alvarez:2018:NET

[1468] Jacobo de Uña-Álvarez and Micha Mandel. Nonparametric estimation of transition probabilities for a general progressive multi-state model under cross-sectional sampling. Biometrics, 74(4):1203–1212, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2018:SRA

[1469] Fei Gao, Donglin Zeng, and Dan-Yu Lin. Semiparametric regression analysis of interval-censored data with informative dropout. Biometrics, 74(4):1213–1222, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2018:PCS

[1470] Chi-Chung Wen and Yi-Hau Chen. Pseudo and conditional score approach to joint analysis of current count and current status data. Biometrics, 74(4):1223–1231, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2018:USI

[1471] Fan Yang and Peng Ding. Using survival information in truncation by death problems without the monotonicity assumption. Biometrics, 74(4): 1232–1239, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2018:PLP

[1472] Minggen Lu and Christopher S. McMahan. A partially linear proportional hazards model for current status data. Biometrics, 74(4):1240– 1249, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2018:AMS

[1473] Soyoung Kim, Donglin Zeng, and Jianwen Cai. Analysis of multiple survival events in generalized case-cohort designs. Biometrics, 74(4): 1250–1260, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qian:2018:TRA

[1474] Jing Qian, Sy Han Chiou, Jacqueline E. Maye, Folefac Atem, Keith A. Johnson, and Rebecca A. Betensky. Threshold regression to accommodate a censored covariate. Biometrics, 74(4):1261–1270, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Price:2018:EOS

[1475] Brenda L. Price, Peter B. Gilbert, and Mark J. van der Laan. Estimation of the optimal surrogate based on a randomized trial. Biometrics, 74(4): 1271–1281, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Castilla:2018:NRS

[1476] Elena Castilla, Abhik Ghosh, Nirian Martin, and Leandro Pardo. New robust statistical procedures for the polytomous logistic regression models. Biometrics, 74(4):1282–1291, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ye:2018:RAS

[1477] Ting Ye and Menggang Yu. A robust approach to sample size calculation in cancer immunotherapy trials with delayed treatment effect. Biometrics, 74(4):1292–1300, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2018:EFF

[1478] Gen Li, Jianhua Z. Huang, and Haipeng Shen. Exponential family functional data analysis via a low-rank model. Biometrics, 74(4):1301–1310, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hui:2018:OSS

[1479] Francis K. C. Hui, Emi Tanaka, and David I. Warton. Order selection and sparsity in latent variable models via the ordered factor LASSO. Biometrics, 74(4):1311–1319, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2018:BOI

[1480] Ruitao Lin. Bayesian optimal interval design with multiple toxicity constraints. Biometrics, 74(4):1320–1330, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Canale:2018:CMR

[1481] Antonio Canale, Daniele Durante, and David B. Dunson. Convex mixture regression for quantitative risk assessment. Biometrics, 74(4):1331–1340, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2018:IDA

[1482] Yichen Cheng, James Y. Dai, Xiaoyu Wang, and Charles Kooperberg. Identifying disease-associated copy number variations by a doubly penalized regression model. Biometrics, 74(4):1341–1350, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Griffin:2018:DMP

[1483] Paula J. Griffin, Yuqing Zhang, William Evan Johnson, and Eric D. Kolaczyk. Detection of multiple perturbations in multi-omics biological networks. Biometrics, 74(4):1351–1361, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Safo:2018:SGE

[1484] Sandra E. Safo, Jeongyoun Ahn, Yongho Jeon, and Sungkyu Jung. Sparse generalized eigenvalue problem with application to canonical correlation analysis for integrative analysis of methylation and gene expression data. Biometrics, 74(4):1362–1371, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:2018:SBV

[1485] Changgee Chang, Suprateek Kundu, and Qi Long. Scalable Bayesian variable selection for structured high-dimensional data. Biometrics, 74 (4):1372–1382, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Estes:2018:TDP

[1486] Jason P. Estes, Danh V. Nguyen, Yanjun Chen, Lorien S. Dalrymple, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Damla Sentürk. Timedynamic profiling with application to hospital readmission among patients on dialysis. Biometrics, 74(4):1383–1394, December 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [1487, 1488, 1489] and rejoinder [1490].

Haneuse:2018:DTD

[1487] Sebastien Haneuse, José Zubizarreta, and Sharon-Lise T. Normand. Discussion on “Time-dynamic profiling with application to hospital readmission among patients on dialysis,” by Jason P. Estes, Danh V. Nguyen, Yanjun Chen, Lorien S. Dalrymple, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Damla Senturk. Biometrics, 74(4):1395–1397, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1486].

Goetghebeur:2018:DTD

[1488] Els Goetghebeur. Discussion on “Time-dynamic profiling with application to hospital readmission among patients on dialysis,” by Jason P. Estes, Danh V. Nguyen, Yanjun Chen, Lorien S. Dalrymple, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Damla Senturk. Biometrics, 74(4):

1398–1400, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1486].

Kalbfleisch:2018:DTD

[1489] John D. Kalbfleisch and Kevin He. Discussion on “Time-dynamic profiling with application to hospital readmission among patients on dialysis,” by Jason P. Estes, Danh V. Nguyen, Yanjun Chen, Lorien S. Dalrymple, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Damla Senturk. Biometrics, 74(4):1401–1403, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1486].

Estes:2018:RTD

[1490] Jason P. Estes, Danh V. Nguyen, Yanjun Chen, Lorien S. Dalrymple, Connie M. Rhee, Kamyar Kalantar-Zadeh, and Damla Sentürk. Rejoinder: Time-dynamic profiling with application to hospital readmission among patients on dialysis. Biometrics, 74(4):1404–1406, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1487, 1488, 1489, 1486].

Boatman:2018:ECE

[1491] Jeffrey A. Boatman and David M. Vock. Estimating the causal effect of treatment regimes for organ transplantation. Biometrics, 74(4):1407– 1416, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2018:VCS

[1492] Jialiang Li, Xiaochao Xia, Weng Kee Wong, and David Nott. Varyingcoefficient semiparametric model averaging prediction. Biometrics, 74 (4):1417–1426, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wen:2018:SPM

[1493] Lan Wen and Shaun R. Seaman. Semi-parametric methods of handling missing data in mortal cohorts under non-ignorable missingness. Biometrics, 74(4):1427–1437, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Keogh:2018:MIM

[1494] Ruth H. Keogh, Shaun R. Seaman, Jonathan W. Bartlett, and Angela M. Wood. Multiple imputation of missing data in nested case-control and case-cohort studies. Biometrics, 74(4):1438–1449, December 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2018:SSD

[1495] Fan Li, Elizabeth L. Turner, and John S. Preisser. Sample size determination for GEE analyses of stepped wedge cluster randomized trials. Biometrics, 74(4):1450–1458, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Igeta:2018:PSS

[1496] Masataka Igeta, Kunihiko Takahashi, and Shigeyuki Matsui. Power and sample size calculation incorporating misspecifications of the variance function in comparative clinical trials with over-dispersed count data. Biometrics, 74(4):1459–1467, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2018:ADT

[1497] Minjung Lee, Eric J. Feuer, and Jason P. Fine. On the analysis of discrete time competing risks data. Biometrics, 74(4):1468–1481, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2018:MRL

[1498] Xiao Lin, Tao Lu, Fangrong Yan, Ruosha Li, and Xuelin Huang. Mean residual life regression with functional principal component analysis on longitudinal data for dynamic prediction. Biometrics, 74(4):1482–1491, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buttarazzi:2018:BCD

[1499] Davide Buttarazzi, Giuseppe Pandolfo, and Giovanni C. Porzio. A boxplot for circular data. Biometrics, 74(4):1492–1501, December 2018. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ray:2018:PAC

[1500] Evan L. Ray, Jeffer E. Sasaki, Patty S. Freedson, and John Staudenmayer. Physical activity classification with dynamic discriminative methods. Biometrics, 74(4):1502–1511, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brintz:2018:AAM

[1501] Ben Brintz, Claudio Fuentes, and Lisa Madsen. An asymptotic approximation to the N-mixture model for the estimation of disease prevalence. Biometrics, 74(4):1512–1518, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Basu:2018:BRC

[1502] Sumanta Basu. Book review: Cichosz, Pawel. Data Mining Algorithms explained using R. Hoboken, Wiley. Biometrics, 74(4):1519–1520, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Richter:2018:BRL

[1503] Lothar Richter. Book review: Leskovec, Jure, Rajaraman, Anand, and Ullman, Jeffrey D., Mining of Massive Datasets. Cambridge, Cambridge University Press. Biometrics, 74(4):1520–1521, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bird:2018:BRA

[1504] Geoff Bird. Book review: Allen, Rory. Statistics and Experimental Design for Psychologists. Singapore, World Scientific Publishing Co Pte Ltd. Biometrics, 74(4):1521–1523, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gu:2018:BRS

[1505] Hui Gu. Book review: Shih, Wiechung Joe and Aisner, Joseph. Statistical Design and Analysis of Clinical Trials: Principles and Methods. Boca Raton, CRC Press. Biometrics, 74(4):1523, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parry:2018:BRG

[1506] Katharina Parry and Matthieu Vignes. Book review: Giraud, Christophe. Introduction to High-Dimensional Statistics. Boca Raton, CRC Press. Biometrics, 74(4):1524–1525, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2018:BRB

[1507] Yiyao Chen. Book review: Berger, Vance. Randomization, Masking, and Allocation Concealment. CRC Press. Biometrics, 74(4):1525–1526, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2018:A

[1508] Anonymous. Acknowledgements. Biometrics, 74(4):1527–1529, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IITa

[1509] Anonymous. Issue information — TOC and cover information. Biometrics, 75(1):1–2, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIMa

[1510] Anonymous. Issue information — masthead. Biometrics, 75(1):3, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIEa

[1511] Anonymous. Issue information — Ed board. Biometrics, 75(1):4, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2019:ABR

[1512] Chunling Lu and Yuhong Yang. On assessing binary regression models based on ungrouped data. Biometrics, 75(1):5–12, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gregory:2019:AEN

[1513] Karl B. Gregory, Dewei Wang, and Christopher S. McMahan. Adaptive elastic net for group testing. Biometrics, 75(1):13–23, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2019:RMA

[1514] Ming Zhou, Rachel S. McCrea, Eleni Matechou, Diana J. Cole, and Richard A. Griffiths. Removal models accounting for temporary emigration. Biometrics, 75(1):24–35, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2019:JSE

[1515] Jianyu Liu, Wei Sun, and Yufeng Liu. Joint skeleton estimation of multiple directed acyclic graphs for heterogeneous population. Biometrics, 75(1):36–47, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wrobel:2019:REF

[1516] Julia Wrobel, Vadim Zipunnikov, Jennifer Schrack, and Jeff Goldsmith. Registration for exponential family functional data. Biometrics, 75(1): 48–57, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thomadakis:2019:LTD

[1517] Christos Thomadakis, Loukia Meligkotsidou, Nikos Pantazis, and Giota Touloumi. Longitudinal and time-to-drop-out joint models can lead to seriously biased estimates when the drop-out mechanism is at random. Biometrics, 75(1):58–68, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2019:RLC

[1518] Jiehuan Sun, Jose D. Herazo-Maya, Philip L. Molyneaux, Toby M. Maher, Naftali Kaminski, and Hongyu Zhao. Regularized latent class model for joint analysis of high-dimensional longitudinal biomarkers and a time-to-event outcome. Biometrics, 75(1):69–77, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hong:2019:SSV

[1519] Chuan Hong, Katherine P. Liao, and Tianxi Cai. Semi-supervised validation of multiple surrogate outcomes with application to electronic medical records phenotyping. Biometrics, 75(1):78–89, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Babino:2019:MRE

[1520] Lucia Babino, Andrea Rotnitzky, and James Robins. Multiple robust estimation of marginal structural mean models for unconstrained outcomes. Biometrics, 75(1):90–99, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dukes:2019:DRE

[1521] Oliver Dukes, Torben Martinussen, Eric J. Tchetgen Tchetgen, and Stijn Vansteelandt. On doubly robust estimation of the hazard difference. Biometrics, 75(1):100–109, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brueckner:2019:IVE

[1522] Matthias Brueckner, Andrew Titman, and Thomas Jaki. Instrumental variable estimation in semi-parametric additive hazards models. Biometrics, 75(1):110–120, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2019:SRA

[1523] Fei Gao and Kwun Chuen Gary Chan. Semiparametric regression analysis of length-biased interval-censored data. Biometrics, 75(1):121–132, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsu:2019:UDU

[1524] Chih-Yuan Hsu, Chen-Hsin Chen, Ken-Ning Hsu, and Ya-Hung Lu. A useful design utilizing the information fraction in a group sequential clinical trial with censored survival data. Biometrics, 75(1):133–143, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kasza:2019:ICC

[1525] Jessica Kasza and Andrew B. Forbes. Information content of clusterperiod cells in stepped wedge trials. Biometrics, 75(1):144–152, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tomer:2019:PSS

[1526] Anirudh Tomer, Daan Nieboer, Monique J. Roobol, Ewout W. Steyerberg, and Dimitris Rizopoulos. Personalized schedules for surveillance of low-risk prostate cancer patients. Biometrics, 75(1):153–162, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McKeague:2019:MST

[1527] Ian W. McKeague and Min Qian. Marginal screening of 2 × 2 tables in large-scale case-control studies. Biometrics, 75(1):163–171, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2019:BIH

[1528] Junghi Kim, Kim-Anh Do, Min Jin Ha, and Christine B. Peterson. Bayesian inference of hub nodes across multiple networks. Biometrics, 75(1):172–182, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2019:BNB

[1529] Qiwei Li, Alberto Cassese, Michele Guindani, and Marina Vannucci. Bayesian negative binomial mixture regression models for the analysis of sequence count and methylation data. Biometrics, 75(1):183–192, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wheeler:2019:BAA

[1530] Matthew W. Wheeler. Bayesian additive adaptive basis tensor product models for modeling high dimensional surfaces: an application to highthroughput toxicity testing. Biometrics, 75(1):193–201, March 2019. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chandra:2019:NBM

[1531] Noirrit Kiran Chandra, Richa Singh, and Sourabh Bhattacharya. A novel Bayesian multiple testing approach to deregulated miRNA discovery harnessing positional clustering. Biometrics, 75(1):202–209, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shokoohi:2019:HMM

[1532] Farhad Shokoohi, David A. Stephens, Guillaume Bourque, Tomi Pastinen, Celia M. T. Greenwood, and Aurélie Labbe. A hidden Markov model for identifying differentially methylated sites in bisulfite sequencing data. Biometrics, 75(1):210–221, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2019:LRA

[1533] Yunpeng Zhao, Qing Pan, and Chengan Du. Logistic regression augmented community detection for network data with application in identifying autism-related gene pathways. Biometrics, 75(1):222–234, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2019:GLM

[1534] Jiarui Lu, Pixu Shi, and Hongzhe Li. Generalized linear models with linear constraints for microbiome compositional data. Biometrics, 75 (1):235–244, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:2019:SDR

[1535] Hung Hung and Su-Yun Huang. Sufficient dimension reduction via random-partitions for the large-p-small-n problem. Biometrics, 75(1): 245–255, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2019:DLR

[1536] Zongliang Hu, Tiejun Tong, and Marc G. Genton. Diagonal likelihood ratio test for equality of mean vectors in high-dimensional data. Biometrics, 75(1):256–267, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gronsbell:2019:AFS

[1537] Jessica Gronsbell, Jessica Minnier, Sheng Yu, Katherine Liao, and Tianxi Cai. Automated feature selection of predictors in electronic medical records data. Biometrics, 75(1):268–277, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bilder:2019:IGT

[1538] Christopher R. Bilder, Joshua M. Tebbs, and Christopher S. McMahan. Informative group testing for multiplex assays. Biometrics, 75(1):278– 288, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rose:2019:DRE

[1539] Sherri Rose and Sharon-Lise Normand. Double robust estimation for multiple unordered treatments and clustered observations: Evaluating drug-eluting coronary artery stents. Biometrics, 75(1):289–296, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bertrand:2019:FPA

[1540] Aurélie Bertrand, Ingrid Van Keilegom, and Catherine Legrand. Flexible parametric approach to classical measurement error variance estimation without auxiliary data. Biometrics, 75(1):297–307, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deng:2019:PSS

[1541] Qiqi Deng, Xiaofei Bai, Dacheng Liu, Dooti Roy, Zhiliang Ying, and Dan-Yu Lin. Power and sample size for dose-finding studies with survival endpoints under model uncertainty. Biometrics, 75(1):308–314, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hwang:2019:BMD

[1542] Beom Seuk Hwang, Zhen Chen, Germaine M. Buck Louis, and Paul S. Albert. A Bayesian multi-dimensional couple-based latent risk model with an application to infertility. Biometrics, 75(1):315–325, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stevenson:2019:CCR

[1543] Ben C. Stevenson, David L. Borchers, and Rachel M. Fewster. Cluster capture-recapture to account for identification uncertainty on aerial surveys of animal populations. Biometrics, 75(1):326–336, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeVito:2019:MSF

[1544] Roberta De Vito, Ruggero Bellio, Lorenzo Trippa, and Giovanni Parmigiani. Multi-study factor analysis. Biometrics, 75(1):337–346, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2019:AHW

[1545] Lu Mao. On the alternative hypotheses for the win ratio. Biometrics, 75 (1):347–351, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [1546].

Luo:2019:RAH

[1546] Xiaodong Luo, Hong Tian, Surya Mohanty, and Wei Yann Tsai. Rejoinder to “On the alternative hypotheses for the win ratio”. Biometrics, 75 (1):352–354, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1545].

Anonymous:2019:CSA

[1547] Anonymous. Correction to “Sensitivity Analysis for matched pair analysis of binary data: From worst case to average case analysis,” by Raiden Hasegawa and Dylan Small; 73, 1424-1432, December 2017. Biometrics, 75(1):355, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1294].

Malinovsky:2019:BR

[1548] Yaakov Malinovsky. Book review. Biometrics, 75(1):356, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bauerle:2019:BR

[1549] Stephen Bäuerle. Book review. Biometrics, 75(1):357, March 2019. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OMalley:2019:BR

[1550] A. James O’Malley. Book review. Biometrics, 75(1):358–360, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shannon:2019:BR

[1551] Bill Shannon. Book review. Biometrics, 75(1):361, March 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IITb

[1552] Anonymous. Issue information — TOC and cover information. Biometrics, 75(2):363–364, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIMb

[1553] Anonymous. Issue information — masthead. Biometrics, 75(2):365, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIEb

[1554] Anonymous. Issue information — Ed board. Biometrics, 75(2):366, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:RE

[1555] Anonymous. Report of the Editors — 2018. Biometrics, 75(2):367– 370, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chapple:2019:HPI

[1556] Andrew G. Chapple and Peter F. Thall. A hybrid Phase I–II/III clinical trial design allowing dose re-optimization in Phase III. Biometrics, 75 (2):371–381, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [1557, 1558] and rejoinder [1559].

Leifer:2019:DHP

[1557] Eric S. Leifer and Nancy L. Geller. Discussion of “A Hybrid Phase I–II/III Clinical Trial Design Allowing Dose Re-Optimization in Phase III” by Andrew G. Chapple and Peter F. Thall. Biometrics, 75(2):382– 384, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1556].

Zhou:2019:DHP

[1558] Tianjian Zhou and Yuan Ji. Discussion of “A hybrid phase I–II/III clinical trial design allowing dose re-optimization in Phase III” by Andrew G. Chapple and Peter F. Thall. Biometrics, 75(2):385–388, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1556].

Chapple:2019:RHP

[1559] Andrew G. Chapple and Peter F. Thall. Rejoinder to “A hybrid phase I–II/III clinical trial design allowing dose reoptimization in phase III”. Biometrics, 75(2):389–391, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1556, 1557, 1558].

Li:2019:MCB

[1560] Yang Li, Yuetian Luo, Davide Ferrari, Xiaonan Hu, and Yichen Qin. Model confidence bounds for variable selection. Biometrics, 75(2):392– 403, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [1559, 1562] and rejoinder [1563].

Claeskens:2019:DMC

[1561] Gerda Claeskens and Maarten Jansen. Discussion on “Model Confidence Bounds for Variable Selection” by Yang Li, Yuetian Luo, Davide Ferrari, Xiaonan Hu, and Yichen Qin. Biometrics, 75(2):404–406, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1560, 1563].

Leeb:2019:DMC

[1562] Hannes Leeb, Benedikt M. Pötscher, and Danijel Kivaranovic. Discussion on “Model confidence bounds for variable selection” by Yang Li, Yuetian Luo, Davide Ferrari, Xiaonan Hu, and Yichen Qin. Biometrics, 75(2): 407–410, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1560, 1563].

Li:2019:RDM

[1563] Yang Li, Yuetian Luo, Davide Ferrari, Xiaonan Hu, and Yichen Qin. Rejoinder to discussions on: “Model confidence bounds for variable selection”. Biometrics, 75(2):411–413, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1560, 1559, 1562].

Zucker:2019:MPL

[1564] David M. Zucker, Xin Zhou, Xiaomei Liao, Yi Li, and Donna Spiegelman. A modified partial likelihood score method for Cox regression with covariate error under the internal validation design. Biometrics, 75(2): 414–427, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2019:EJD

[1565] Yu-Jen Cheng, Mei-Cheng Wang, and Chang-Yu Tsai. Estimations of the joint distribution of failure time and failure type with dependent truncation. Biometrics, 75(2):428–438, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barthel:2019:DMR

[1566] Nicole Barthel, Candida Geerdens, Claudia Czado, and Paul Janssen. Dependence modeling for recurrent event times subject to right-censoring with D-vine copulas. Biometrics, 75(2):439–451, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Amico:2019:SIC

[1567] Maïlis Amico, Ingrid Van Keilegom, and Catherine Legrand. The singleindex/Cox mixture cure model. Biometrics, 75(2):452–462, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chu:2019:DFE

[1568] Chenghao Chu, Ying Zhang, and Wanzhu Tu. Distribution-free estimation of local growth rates around interval censored anchoring events. Biometrics, 75(2):463–474, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Besbeas:2019:EII

[1569] P. Besbeas and B. J. T. Morgan. Exact inference for integrated population modelling. Biometrics, 75(2):475–484, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Michael:2019:EIR

[1570] Haben Michael, Suzanne Thornton, Minge Xie, and Lu Tian. Exact inference on the random-effects model for meta-analyses with few studies. Biometrics, 75(2):485–493, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xia:2019:NGS

[1571] Meng Xia, Susan Murray, and Nabihah Tayob. Nonparametric group sequential methods for evaluating survival benefit from multiple shortterm follow-up windows. Biometrics, 75(2):494–505, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Genback:2019:CIA

[1572] Minna Genbäck and Xavier de Luna. Causal inference accounting for unobserved confounding after outcome regression and doubly robust estimation. Biometrics, 75(2):506–515, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huling:2019:IVB

[1573] Jared D. Huling, Menggang Yu, and A. James O’Malley. Instrumental variable based estimation under the semiparametric accelerated failure time model. Biometrics, 75(2):516–527, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhingare:2019:SBL

[1574] Apurva Bhingare, Debajyoti Sinha, Debdeep Pati, Dipankar Bandyopadhyay, and Stuart R. Lipsitz. Semiparametric Bayesian latent variable regression for skewed multivariate data. Biometrics, 75(2):528–538, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2019:BSA

[1575] Zhen Chen and Beom Seuk Hwang. A Bayesian semiparametric approach to correlated ROC surfaces with stochastic order constraints. Biometrics, 75(2):539–550, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fei:2019:DIH

[1576] Zhe Fei, Ji Zhu, Moulinath Banerjee, and Yi Li. Drawing inferences for high-dimensional linear models: a selection-assisted partial regression and smoothing approach. Biometrics, 75(2):551–561, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2019:SBG

[1577] Stephanie T. Chen, Luo Xiao, and Ana-Maria Staicu. A smoothingbased goodness-of-fit test of covariance for functional data. Biometrics, 75(2):562–571, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duarte:2019:ODM

[1578] Belmiro P. M. Duarte and Weng Kee Wong. Optimal design of multipleobjective lot quality assurance sampling (LQAS) plans. Biometrics, 75 (2):572–581, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OConnell:2019:LMF

[1579] Michael J. O’Connell and Eric F. Lock. Linked matrix factorization. Biometrics, 75(2):582–592, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2019:IMV

[1580] Gen Li, Xiaokang Liu, and Kun Chen. Integrative multi-view regression: Bridging group-sparse and low-rank models. Biometrics, 75(2):593–602, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2019:CCA

[1581] Cheolwoo Park, Hosik Choi, Chris Delcher, Yanning Wang, and Young Joo Yoon. Convex clustering analysis for histogram-valued data. Biometrics, 75(2):603–612, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bates:2019:LRL

[1582] Stephen Bates and Robert Tibshirani. Log-ratio lasso: Scalable, sparse estimation for log-ratio models. Biometrics, 75(2):613–624, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2019:CKM

[1583] Ni Zhao, Haoyu Zhang, Jennifer J. Clark, Arnab Maity, and Michael C. Wu. Composite kernel machine regression based on likelihood ratio test for joint testing of genetic and gene-environment interaction effect. Biometrics, 75(2):625–637, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2019:NRP

[1584] Xiaoyu Dai, Nan Lin, Daofeng Li, and Ting Wang. A non-randomized procedure for large-scale heterogeneous multiple discrete testing based on randomized tests. Biometrics, 75(2):638–649, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:2019:RRU

[1585] Hung Hung. A robust removing unwanted variation-testing procedure via γ-divergence. Biometrics, 75(2):650–662, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ji:2019:BHM

[1586] Tieming Ji. A Bayesian hidden Markov model for detecting differentially methylated regions. Biometrics, 75(2):663–673, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mi:2019:BDL

[1587] Xinlei Mi, Fei Zou, and Ruoqing Zhu. Bagging and deep learning in optimal individualized treatment rules. Biometrics, 75(2):674–684, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dahabreh:2019:GCI

[1588] Issa J. Dahabreh, Sarah E. Robertson, Eric J. Tchetgen, Elizabeth A. Stuart, and Miguel A. Hernán. Generalizing causal inferences from individuals in randomized trials to all trial-eligible individuals. Biometrics, 75(2):685–694, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2019:CCE

[1589] Liangyuan Hu and Joseph W. Hogan. Causal comparative effectiveness analysis of dynamic continuous-time treatment initiation rules with sparsely measured outcomes and death. Biometrics, 75(2):695–707, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hogan:2019:BRC

[1590] Joseph W. Hogan. Book review: Causal inference in statistics: A primer. Judea Pearl, Maria Glymour, and Nicholas Jewell, John Wiley and Sons, Ltd., Chichester, UK. Biometrics, 75(2):708–709, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Comulada:2019:BRQ

[1591] W. Scott Comulada. Book review: Quantitative methods for HIV/AIDS research. Cliburn Chan, Michael G. Hudgens, Shein-Chung Chow, CRC Press. Biometrics, 75(2):710–711, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mahmood:2019:BRB

[1592] Nuha Mahmood. Book review: Biostatistics for clinical and public health research, Melody S. Goodman (Ed.), New York: Routledge. Biometrics, 75(2):712, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2019:BRE

[1593] Yi Xu and David Banks. Book review: Errors, Blunders, and Lies: How to Tell the Difference. Salsburg, David S., Raton, Boca: Chapman and Hall/ASA-CRC Series. Biometrics, 75(2):713–714, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Patterson:2019:BRH

[1594] Toby Patterson. Book review: Hidden Markov Models for Time Series: An Introduction Using R. Zucchini, Walter, McDonald, Iain L., and Langrock, Roland, Boca Raton, CRC Press. Biometrics, 75(2):715– 716, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ravishanker:2019:BRS

[1595] Nalini Ravishanker. Book review: State Space Methods for Time Series Analysis: Theory, Applications And Software. Casals, Jose, Garica-Hiernaux, Alfredo, Jerez, Miguel, Sotoca, Sonia, and Trindade, A. Alexandre, Boca Raton: CRC Press. Biometrics, 75(2):717, June 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IITc

[1596] Anonymous. Issue information — TOC and cover information. Biometrics, 75(3):719–720, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIMc

[1597] Anonymous. Issue information — masthead. Biometrics, 75(3):721, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIEc

[1598] Anonymous. Issue information — Ed board. Biometrics, 75(3):722, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2019:FLB

[1599] W. Zhang, M. V. Bravington, and R. M. Fewster. Fast likelihood-based inference for latent count models using the saddlepoint approximation. Biometrics, 75(3):723–733, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mai:2019:IPL

[1600] Qing Mai and Xin Zhang. An iterative penalized least squares approach to sparse canonical correlation analysis. Biometrics, 75(3):734– 744, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Derkach:2019:HDM

[1601] Andriy Derkach, Ruth M. Pfeiffer, Ting-Huei Chen, and Joshua N. Sampson. High dimensional mediation analysis with latent variables. Biometrics, 75(3):745–756, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2019:HOT

[1602] Yangyi Xu, Inyoung Kim, and Raymond J. Carroll. A hybrid omnibus test for generalized semiparametric single-index models with highdimensional covariate sets. Biometrics, 75(3):757–767, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miles:2019:CIW

[1603] Caleb H. Miles, Maya Petersen, and Mark J. van der Laan. Causal inference when counterfactuals depend on the proportion of all subjects exposed. Biometrics, 75(3):768–777, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Papadogeorgou:2019:CII

[1604] Georgia Papadogeorgou, Fabrizia Mealli, and Corwin M. Zigler. Causal inference with interfering units for cluster and population level treatment allocation programs. Biometrics, 75(3):778–787, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2019:GMA

[1605] Yi Zhao and Xi Luo. Granger mediation analysis of multiple time series with an application to functional magnetic resonance imaging. Biometrics, 75(3):788–798, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McClintock:2019:IME

[1606] Brett T. McClintock, Gary C. White, and Moira A. Pryde. Improved methods for estimating abundance and related demographic parameters from mark-resight data. Biometrics, 75(3):799–809, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scharf:2019:APA

[1607] Henry R. Scharf, Mevin B. Hooten, Ryan R. Wilson, George M. Durner, and Todd C. Atwood. Accounting for phenology in the analysis of animal movement. Biometrics, 75(3):810–820, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2019:CAR

[1608] Wei Zhang, Aiyi Liu, Larry L. Tang, and Qizhai Li. A cluster-adjusted rank-based test for a clinical trial concerning multiple endpoints with application to dietary intervention assessment. Biometrics, 75(3):821– 830, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1674].

Vettori:2019:BMA

[1609] Sabrina Vettori, Raphaël Huser, and Marc G. Genton. Bayesian modeling of air pollution extremes using nested multivariate max-stable processes. Biometrics, 75(3):831–841, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maziarz:2019:ICC

[1610] Marlena Maziarz, Yukun Liu, Jing Qin, and Ruth M. Pfeiffer. Inference for case-control studies with incident and prevalent cases. Biometrics, 75(3):842–852, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2019:VCG

[1611] Jie Zhou, Jiajia Zhang, Alexander C. Mclain, Wenbin Lu, Xuemei Sui, and James W. Hardin. A varying-coefficient generalized odds rate model with time-varying exposure: an application to fitness and cardiovascular disease mortality. Biometrics, 75(3):853–863, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhabotynsky:2019:SMJ

[1612] Vasyl Zhabotynsky, Kaoru Inoue, Terry Magnuson, J. Mauro Calabrese, and Wei Sun. A statistical method for joint estimation of cis -eQTLs and parent-of-origin effects under family trio design. Biometrics, 75(3): 864–874, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2019:PAM

[1613] Tao Wang, Can Yang, and Hongyu Zhao. Prediction analysis for microbiome sequencing data. Biometrics, 75(3):875–884, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robertson:2019:FEC

[1614] D. S. Robertson and J. M. S. Wason. Familywise error control in multiarmed response-adaptive trials. Biometrics, 75(3):885–894, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qiu:2019:TSC

[1615] Yumou Qiu and Janaka S. S. Liyanage. Threshold selection for covariance estimation. Biometrics, 75(3):895–905, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dobler:2019:CBM

[1616] Dennis Dobler, Markus Pauly, and ThomasH. Scheike. Confidence bands for multiplicative hazards models: Flexible resampling approaches. Biometrics, 75(3):906–916, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Su:2019:SAA

[1617] Li Su, Qiuju Li, Jessica K. Barrett, and Michael J. Daniels. A sensitivity analysis approach for informative dropout using shared parameter models. Biometrics, 75(3):917–926, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Siddique:2019:MEC

[1618] Juned Siddique, Michael J. Daniels, Raymond J. Carroll, Trivellore E. Raghunathan, Elizabeth A. Stuart, and Laurence S. Freedman. Measurement error correction and sensitivity analysis in longitudinal dietary intervention studies using an external validation study. Biometrics, 75(3): 927–937, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mitani:2019:MAO

[1619] Aya A. Mitani, Elizabeth K. Kaye, and Kerrie P. Nelson. Marginal analysis of ordinal clustered longitudinal data with informative cluster size. Biometrics, 75(3):938–949, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2019:ELB

[1620] Chixiang Chen, Biyi Shen, Lijun Zhang, Yuan Xue, and Ming Wang. Empirical-likelihood-based criteria for model selection on marginal analysis of longitudinal data with dropout missingness. Biometrics, 75(3): 950–965, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1673].

Tancredi:2019:ABI

[1621] Andrea Tancredi. Approximate Bayesian inference for discretely observed continuous-time multi-state models. Biometrics, 75(3):966–977, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2019:BHM

[1622] Jincheng Zhou, James S. Hodges, M. Fareed K. Suri, and Haitao Chu. A Bayesian hierarchical model estimating CACE in meta-analysis of randomized clinical trials with noncompliance. Biometrics, 75(3):978–987, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zanini:2019:BRP

[1623] Carlos Tadeu Pagani Zanini, Peter Müller, Yuan Ji, and Fernando A. Quintana. A Bayesian random partition model for sequential refinement and coagulation. Biometrics, 75(3):988–999, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diao:2019:EMS

[1624] Guoqing Diao, Guanghan F. Liu, Donglin Zeng, William Wang, Xianming Tan, Joseph F. Heyse, and Joseph G. Ibrahim. Efficient methods for signal detection from correlated adverse events in clinical trials. Biometrics, 75(3):1000–1008, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ferguson:2019:TSE

[1625] Jake M. Ferguson, Tanya A. Miura, and Craig R. Miller. A two-stage experimental design for dilution assays. Biometrics, 75(3):1009–1016, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2019:CSH

[1626] Yuejia Xu, Oliver Laeyendecker, and Rui Wang. Cross-sectional human immunodeficiency virus incidence estimation accounting for heterogeneity across communities. Biometrics, 75(3):1017–1028, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2019:DWT

[1627] Minchun Zhou, David Badre, and Hakmook Kang. Double-wavelet transform for multisubject task-induced functional magnetic resonance imaging data. Biometrics, 75(3):1029–1040, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ankerst:2019:BRC

[1628] Reviewed by Donna P. Ankerst. Book review: Chain event graphs. Collazo, Rodrigo A., Goergen, Christiane, Smith, Jim Q., Boca Raton: CRC Press. Biometrics, 75(3):1041, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerds:2019:BRA

[1629] Reviewed by Thomas A. Gerds. Book review: Absolute Risk. Pfeiffer, Ruth M. and Gail, Mitchell H., Boca Raton: CRC Press. Biometrics, 75 (3):1042, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tabb:2019:BRB

[1630] Loni P. Tabb. Book review: Bayesian regression modeling with INLA. Wang, Xiaofeng, Yue, Yu R. and Faraway, Julian J., Boca Raton: CRC Press. Biometrics, 75(3):1042–1044, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kidwell:2019:BRS

[1631] Kelley M. Kidwell. Book review: Statistical approaches in oncology clinical development: Current paradigm and methodological advancement. Roychoudhury, Satrajit and Lahiri, Soumi. Boca Raton: Chapman and Hall/CRC. Biometrics, 75(3):1044–1045, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2019:BRS

[1632] Li-Pang Chen. Book review: Statistical analysis with measurement error or misclassification: Strategy, method and application. Yi, Grace Y. New York: Springer-Verlag. Biometrics, 75(3):1045–1046, September 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IITd

[1633] Anonymous. Issue information — TOC and cover information. Biometrics, 75(4):1047–1048, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIMd

[1634] Anonymous. Issue information — masthead. Biometrics, 75(4):1049, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:IIEd

[1635] Anonymous. Issue information — Ed board. Biometrics, 75(4):1050, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sekula:2019:DDE

[1636] Michael Sekula, Jeremy Gaskins, and Susmita Datta. Detection of differentially expressed genes in discrete single-cell RNA sequencing data using a hurdle model with correlated random effects. Biometrics, 75(4): 1051–1062, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gao:2019:IAG

[1637] Bin Gao, Xu Liu, Hongzhe Li, and Yuehua Cui. Integrative analysis of genetical genomics data incorporating network structures. Biometrics, 75(4):1063–1075, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Masotti:2019:PIA

[1638] Maria Masotti, Bin Guo, and Baolin Wu. Pleiotropy informed adaptive association test of multiple traits using genome-wide association study summary data. Biometrics, 75(4):1076–1085, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2019:MLD

[1639] Yanming Li, Hyokyoung G. Hong, and Yi Li. Multiclass linear discriminant analysis with ultrahigh-dimensional features. Biometrics, 75(4): 1086–1097, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2019:SER

[1640] Jingru Zhang and Wei Lin. Scalable estimation and regularization for the logistic normal multinomial model. Biometrics, 75(4):1098–1108, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2019:CRS

[1641] Wenjing Wang, Xin Zhang, and Lexin Li. Common reducing subspace model and network alternation analysis. Biometrics, 75(4):1109–1120, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gaynanova:2019:SLI

[1642] Irina Gaynanova and Gen Li. Structural learning and integrative decomposition of multi-view data. Biometrics, 75(4):1121–1132, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nghiem:2019:SSE

[1643] Linh Nghiem and Cornelis Potgieter. Simulation-selection-extrapolation: Estimation in high-dimensional errors-in-variables models. Biometrics, 75(4):1133–1144, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vandekar:2019:RSE

[1644] Simon N. Vandekar, Theodore D. Satterthwaite, Cedric H. Xia, Azeez Adebimpe, Kosha Ruparel, Ruben C. Gur, Raquel E. Gur, and Russell T. Shinohara. Robust spatial extent inference with a semiparametric bootstrap joint inference procedure. Biometrics, 75(4):1145–1155, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2019:CBS

[1645] Yanlin Tang, Huixia J. Wang, Ying Sun, and Amanda S. Hering. Copulabased semiparametric models for spatiotemporal data. Biometrics, 75(4): 1156–1167, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diao:2019:SFM

[1646] Guoqing Diao, Donglin Zeng, Kuolung Hu, and Joseph G. Ibrahim. Semiparametric frailty models for zero-inflated event count data in the presence of informative dropout. Biometrics, 75(4):1168–1178, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2019:PLP

[1647] Wu Wang and Ying Sun. Penalized local polynomial regression for spatial data. Biometrics, 75(4):1179–1190, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2019:VCT

[1648] Yen-Tsung Huang. Variance component tests of multivariate mediation effects under composite null hypotheses. Biometrics, 75(4):1191–1204, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wallace:2019:MSE

[1649] Michael P. Wallace, Erica E. M. Moodie, and David A. Stephens. Model selection for G-estimation of dynamic treatment regimes. Biometrics, 75(4):1205–1215, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2019:MIT

[1650] Xinyang Huang, Yair Goldberg, and Jin Xu. Multicategory individualized treatment regime using outcome weighted learning. Biometrics, 75(4):1216–1227, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2019:CCI

[1651] Emily J. Huang, Ethan X. Fang, Daniel F. Hanley, and Michael Rosenblum. Constructing a confidence interval for the fraction who benefit from treatment, using randomized trial data. Biometrics, 75(4):1228– 1239, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2019:LND

[1652] Ming Sun, Donglin Zeng, and Yuanjia Wang. Leveraging nonlinear dynamic models to predict progression of neuroimaging biomarkers. Biometrics, 75(4):1240–1252, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parast:2019:USM

[1653] Layla Parast, Tianxi Cai, and Lu Tian. Using a surrogate marker for early testing of a treatment effect. Biometrics, 75(4):1253–1263, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Soh:2019:DRR

[1654] J. E. Soh and Yijian Huang. Dynamic regression with recurrent events. Biometrics, 75(4):1264–1275, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stensrud:2019:NHS

[1655] Mats J. Stensrud, Kjetil Røysland, and Pål C. Ryalen. On null hypotheses in survival analysis. Biometrics, 75(4):1276–1287, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leday:2019:FBI

[1656] Gwenaël G. R. Leday and Sylvia Richardson. Fast Bayesian inference in large Gaussian graphical models. Biometrics, 75(4):1288–1298, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beer:2019:IPI

[1657] Joanne C. Beer, Howard J. Aizenstein, Stewart J. Anderson, and Robert T. Krafty. Incorporating prior information with fused sparse group lasso: Application to prediction of clinical measures from neuroimages. Biometrics, 75(4):1299–1309, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shin:2019:ANM

[1658] Yei Eun Shin, Huiyan Sang, Dawei Liu, Toby A. Ferguson, and Peter X. K. Song. Autologistic network model on binary data for disease progression study. Biometrics, 75(4):1310–1320, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kowal:2019:IVF

[1659] Daniel R. Kowal. Integer-valued functional data analysis for measles forecasting. Biometrics, 75(4):1321–1333, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lawrence:2019:TST

[1660] John Lawrence. Truncated Simes tests. Biometrics, 75(4):1334–1344, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glennie:2019:OPM

[1661] Richard Glennie, David L. Borchers, Matthew Murchie, Bart J. Harmsen, and Rebecca J. Foster. Open population maximum likelihood spatial capture-recapture. Biometrics, 75(4):1345–1355, December 2019. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finazzi:2019:QPE

[1662] Francesco Finazzi and Lucia Paci. Quantifying personal exposure to air pollution from smartphone-based location data. Biometrics, 75(4): 1356–1366, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jang:2019:AAB

[1663] Jeong Hoon Jang, Limin Peng, and Amita K. Manatunga. Assessing alignment between functional markers and ordinal outcomes based on broad sense agreement. Biometrics, 75(4):1367–1379, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2019:DPO

[1664] Ruoqi Yu and Paul R. Rosenbaum. Directional penalties for optimal matching in observational studies. Biometrics, 75(4):1380–1390, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2019:ACR

[1665] Bingkai Wang, Elizabeth L. Ogburn, and Michael Rosenblum. Analysis of covariance in randomized trials: More precision and valid confidence intervals, without model assumptions. Biometrics, 75(4):1391–1400, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baldoni:2019:IDE

[1666] Pedro L. Baldoni, Naim U. Rashid, and Joseph G. Ibrahim. Improved detection of epigenomic marks with mixed-effects hidden Markov models. Biometrics, 75(4):1401–1413, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peterson:2019:BRS

[1667] Adam Peterson. Book review: Spatio-temporal statistics with R, Wikle, Christopher K., Zammit-Mangion, Andrew, Cressie, Noel, Boca Raton, FL: CRC Press, 2019. Biometrics, 75(4):1414, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Isheden:2019:BRB

[1668] Gabriel Isheden. Book review: Bayesian inference for stochastic processes, Broemeling, Lyle D, Boca Raton: CRC Press. Biometrics, 75(4): 1414–1415, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sanchez:2019:BRT

[1669] Brisa N. Sánchez. Book review: Textual data science with R, Bécue– Bertaut, Mónica, Boca Raton: CRC Press. Biometrics, 75(4):1415–1416, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2019:BRS

[1670] D. Y. Lin. Book review: The statistical analysis of multivariate failure time data: a marginal modeling approach, Prentice, Ross L., Zhao, Shanshan, Boca Raton, FL: CRC Press. Biometrics, 75(4):1416–1417, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2019:A

[1671] Anonymous. Acknowledgements. Biometrics, 75(4):1418–1420, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2020:CDR

[1672] Anonymous. Correction to “Doubly-robust dynamic treatment regimen estimation via weighted least squares,” by Michael P. Wallace and Erica E. M. Moodie; 71 (3), 636–644, September 2015. Biometrics, 76(2): 670, June 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [886].

Anonymous:2021:CEL

[1673] Anonymous. Correction to “Empirical-likelihood-based criteria for model selection on marginal analysis of longitudinal data with dropout missingness,” by Chen, C., Shen, B., Zhang, L., Xue, Y. and Wang, M.; 75(3), 950–965, 2019. Biometrics, 77(2):779, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1620].

Chen:2021:CCA

[1674] Chixiang Chen, Biyi Shen, Lijun Zhang, Yuan Xue, and Ming Wang. Correction to “A cluster-adjusted rank-based test for a clinical trial concerning multiple endpoints with application to dietary intervention assessment,” by Zhang, W., Liu, A., Tang, L.L. and Li, Q; 75(3), 821–830, 2019. Biometrics, 75(3):780, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gasparrini:2022:CPF

[1675] Antonio Gasparrini, Fabian Scheipl, Ben Armstrong, and Michael G. Kenward. Correction to “A penalized framework for distributed lag nonlinear models” by Antonio Gasparrini, Fabian Scheipl, Ben Armstrong, and Michael G. Kenward; 73, 938–948, September 2017. Biometrics, 78 (2):812, June 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1245].